A Complete Bibliography of The Journal of the American Statistical Association: 2010–2019

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

09 November 2023 Version 1.29

Title word cross-reference

$119.95 [Pea19]. $135.00 [Tri19]. 24 [SJ14]. 2 × 2 [LLgX14]. _{2} [Cre18a]. E [God17]. ell_{1} [CLL11]. F [HD10]. g [LC18, XG17]. h [XG17]. I [GJS16]. K [HDL^{+}16, JYL15, MW17, THY16]. L_{1} [SL15, XMZ12]. L_{q} [QP13]. M [Ros14, SLS18, WF12]. P [LX19, AEI18, Ber17, Bri17, GC17, LS17a, PDF15, RDHL19, ZSS19]. p > 1 [FCSZ16]. p ≥ N [GBW16]. Q [CLXY15, LLS17, ZZS19]. R [HPV14, HM15]. S [BSB15]. t [MG12]. T^{2} [CPPW11]. z [Cai10, Efr10b, Efr10c, Hel10, Sch10, Wes10].

-and- [XG17]. -Estimation [HPV14, HM15]. -Estimators [BSB15, SLS18]. -Hour [SJ14]. -Learning [LLS17, ZZS19]. -Likelihood [QP13]. -Matrix [CLXY15]. -Means [HDL^{+}16]. -Optimal [GJS16]. -Penalized [XMZ12]. -Priors [LC18]. -Regression [WF12]. -Regularization [SL15]. -Sample [JYL15]. -Segment [THY16]. -statistics [Ros14]. -Tests [HD10]. -Value [Ber17, GC17, God17, LX19]. -Values [AEI18, Bri17, Cai10, Efr10b, Efr10c, Hel10, LS17a, RDHL19, Sch10, Wes10, ZSS19].

0 [Pea19].

108 [CDQ14]. 1323 [CDQ14].

2nd [Yan19].

3. [Tri19]. 3rd [Vat19].

4Rs [VHJB13].

978 [Pea19, Tri19]. 978-1-13-857833-3. [Tri19]. 978-1-49-876858-0 [Pea19].

A/H1N1 [FBCA14]. A5095 [LER^{+}12]. Aberrations [BJT^{+}10]. Absence [WCZ^{+}13]. Absolute [CGLY10]. Abundance [LBS13]. Abuse [HLG14]. Accelerated [CKY15, NQS14, PPLS18, SS17]. Accelerating [CMPS17]. Accelerometry [HBI^{+}19]. Accelerometry-Measured [HBI^{+}19]. Accommodate [NTC13]. Accounting [HMM16, LS19a]. Accumulate [CG15]. Accumulation [LB17a]. Accuracy [Cai10, CZ11, CCF18, CTM10, Efr10b, Efr10c, Efr14a, Hel10, LW19a, LCLZ18, Sch10, Wes10]. Accurate [DFRS14, LX19, ZL11]. Acoustic [MBCM13]. Acquisition [KLCT11]. Across [MMW^{+}17, Mar19, PM19, TPAC19a, TPAC19b]. Act [JML^{+}14]. Action [Dav13]. Activation [YRR19, YPOR18]. Active [MZC16]. Activity [HBI^{+}19, JKL11, SFM18]. Acute [LD15]. Acyclic [HCCZ16, KM17]. Adaptation [Efr10a]. Adapted [DEV19]. Adaptive [BGB12, BGSB19, BCS18, Büh11, CL11a, CY12, CLM14, CY16, Düm11, EG19, FSB^{+}19, GK11, HCCZ16, HZH15, HWF15, Hun11, LM11a, LM11b, Lei14, LB18, LK19, LJSL16, Ma13, MHZ15, MQ15a, MAS14, MYLC15, Pol11, RJ10, RWS12, Sam11, SCG13, WN11, YC11, ZB13, ZBM11, ZD13]. AdaptSPEC [RWS12]. Addiction [JLZ14]. Additive [ASB11, CFL18, CB18b, Cul11, FFS11, FKH11, HHY19, KKL15, LCZ14, LL13, LYH13, PHSS15, SFK12, WKKS12, WLZ19, WLSA17, WLXL14, XQZ10, YY16, You19, ZPlW13]. Addressing [PD16]. Adherence [MSK^{+}17]. Adhesion [HWZ^{+}13]. Adjacency [STFP12]. Adjust [KC10]. Adjusted [CLL19, CRZZ16, DZ19, FKSZ19, HZH15, JLZ14, PMZ19, RS10, TGZ15, WFB13, ZJZ12]. Adjusting [LS10]. Adjustment [BGK^{+}18, GK18, MRB12]. Adjustments [GK18]. Administrative [OKEK17]. Admissibility [WO19]. Admission [KHK19]. Adults [IZR^{+}13]. Advanced [WRL^{+}12a]. Adverse [BAN^{+}12]. Aerosol [WJYJ13]. Affine [PBD13]. After [Efr14a, Zub12]. Age [CWC^{+}14, HN18, SYS^{+}19, TEKM12]. Age-at-Onset [SYS^{+}19]. Age-Crime [TEKM12]. Age-Dependent [CWC^{+}14]. Age-Period-Cohort [HN18]. Agent [FTZ15, HW10]. Agent-Based [HW10]. Agents [GY17]. Aggregate [FG13, HCKS18, QLL10]. Aggregated [DZ15, DGH^{+}10, MZ15]. Aggregation [DHLL14, Per12]. Agreement [PLGM11]. Agricultural [RGH13a]. AIDS [LER^{+}12]. Air [GDZ11]. al [BFH12]. al. [Mar19, PM19]. Alastair [Pea19]. Algebra [Tri19]. Algorithm [JNR15, LJSL16, LFU11, QP13, YBT13, ZW14, ZR15]. Algorithms [QNLS11]. All-Encompassing [Gel11]. Allele [HLSZ14, HSTP15]. Allele-Specific [HLSZ14, HSTP15]. Allocation [GBZ17, LHW^{+}13, Sna19, XMY^{+}15]. Alternate [ZNSR12]. Alternative [BOSB16, CS10, HH12]. Alternatives [JTF^{+}19, YP17]. Alzheimer [AVWW14]. Ambiguous [SLW19]. Ambulance [ZMW^{+}15]. American [CDQ14, BWH16]. Among [CMS^{+}18, EUW17, PGY14, RSI14, YL16]. Amounts [FLJL17]. Amplitude [HAME15]. Analyses [Ros15a]. Analysing [Mar19]. Analysis [ASX19, BGH^{+}18, BIL^{+}11, BAN^{+}12, BIZ15, BA16, BSB15, CL11b, CCZF15, CPPW11, CCMW11, CCS14, CWZ15, CL15b, CLXY15, CIB^{+}18, CGO19, CXT14, CMS17, CCL^{+}11a, CCL^{+}11b, CLZ15, DQZ18, Dai19, DRC^{+}12, Dav12, DM11, DDM^{+}10, DPS19, DDH15, EI19, FSL17, FAIW^{+}11, FPSE15b, FCSZ16, FSMS17, FS17, FDPS10, Fuk15, GZCL19, GBDL10, GWZ13, GSH13, GDCL11, HAME15, HM15, Han14, HL18, HG18, HSM^{+}15, HJRS10, HGK12, HTGT13, HLSZ14, HD10, HZZ^{+}17, HBI^{+}19, HJCPV11, Ima11, JPCD19, JCL10, KJNW11, KYBB19, KWG15, KRS^{+}17, KCP^{+}11, LVC14, LNN19, Laz11, LW10a, LST^{+}11, LZP10, LZ11a, LCZ14, LG14, LHTB15, LHH^{+}17b, LK19, LHC19, LCS^{+}13, Lin12, LLgX14, LLX15, LNM11, LMH14, MH17, MS18b, MJ14, MT17, MM13, NC10, NLOP10, NSK16, PHSS15, PWL16, Pap10, PAHJ11, QJWG16, RSI14, RDLC10, RBF^{+}17, RMR19, RG17b, RGH13b]. Analysis [RDHL19, RLP^{+}18, SdCG^{+}15, SLW16, She11, SH15, SY17, SM16, SMQ^{+}13, SKBM17, SSZL12, Tad13a, TLG16, TZF^{+}15, VS15, WSSQ16, WY19, WSL^{+}16, Wu11, XSS11, XLN18, ZLZ13, ZHS17, ZLWT17]. Analysis-With [CCS14]. Analytic [HDL^{+}16]. Analytics [Yu19]. Analyze [GAZ13]. Analyzing [BF18, CSSK16, HLR15, PM19, TPAC19a, TPAC19b, WWL12]. ANCOVA [JTF^{+}19]. Anesthesia [ZNSR12]. Angle [ABvdW19]. Angular [MZN^{+}16]. Angular-Sampling-Based [MZN^{+}16]. Animal [HJ17]. Annealing [LCL14]. Anomalies [ACCTW18, PARS19]. Anomaly [TARS17]. ANOVA [JHH14, MS18a, SLR^{+}15]. Anthropogenic [RPS19]. Antiretroviral [MSK^{+}17]. Any [AGS12]. Aphasic [GDCL11]. Apparent [TR19]. Application [BSMR12, BWH16, BGK^{+}18, CZ19, CWPC13, CCS14, CXCR13, CMER11, DHM11, DP11, DM18, DK18, DV11, DTYG12, DLR11, EI19, FRG^{+}17, GSDR19, Gil11, GK18, GHP13, GBZ17, HS13, HTGT13, Hos13, HZT11, yHQ12, HLG14, HBI^{+}19, HC15b, JCL10, JCRG17, KZCS16, KHK19, KLCT11, KCZ^{+}13, KMO15, KWG15, KXZ19, KKLL17, KRS^{+}17, LVC14, LST^{+}11, LMZJ13, LDSH16, LS18b, LMR^{+}19, LQ11, LCZ12, LCZ14, LCG^{+}15, LS18c, LFL15, Lin12, LD15, LHW^{+}13, LLgX14, LZPH14, MZN^{+}16, MZT^{+}17, MJPW13b, MŞDN12, NSH^{+}19b, PARS19, RPS19, RD12, RLP^{+}18, SF13, SCM^{+}18, SKS^{+}15, SYS^{+}19, SSS18, Tad10, TARS17, VRS19, WBG^{+}18, WDSL10, WGM12, WSSQ16, WFZ18, WGDG14b, WKKS12, WH17, XTR^{+}14, XXL^{+}19, YHD^{+}10, You19, ZC17a, ZSS19]. Applications [ASB11, CCZ16, CWH^{+}15, CMS17, Dai19, DPV11, Dui16, FWYZ13, HJRS10, HD10, HWZ^{+}13, JZT19, KJ10, LZ10b, LG14, LD17, NN17, NCB12, RN16, SW11, SPHL16, TZL17, TPLG10, WY19, XCC18, YRR19, Zha05, Zho11, ZLZ13]. Applied [MHC13, RGH13a]. Appreciating [Utt17]. Approach [AL14, AB17b, BFT18, BGMP12, CLL11, CL11b, CWPC13, CGBY13, CWH18, DR10, FTZ15, FLJL17, GC11, GDZ11, HMQA10, HAME15, HMM16, HB14, HP10a, HY12, HMW^{+}14, HC17, HC15b, JAW16, KY11b, KLSY13, LN10, LD15, LLgX14, LZ18, LS17b, MZ12, MH17, MRB^{+}17, Mar19, MJ14, MYT18, PS10, PM19, PCSK16, PLLG18, QDC19, QLL10, RG14, SBG^{+}16, SLM11, ST15, SHF17, SGVC13, TPAC19a, TPAC19b, VW15, WJYJ13, WGE^{+}18, WMY15, WQxYW19, XKBS17, YPQ^{+}16, YPOR18, Zha13a, ZDP10, ZHS17, Zho19]. Approaches [TWYZ11]. Approximate [Blu10, CWZM18, CZMHS17, Dui16, FKSZ19, KPBSK10, Ros11, Wan17a, Zha05]. Approximated [HB14]. Approximately [BCK19]. Approximation [Kat17, LCS^{+}13, LCL14, LX19, PDF15, ZL11]. Approximations [BG16, CMPS17, HYSM19, PCB11]. Arbitrarily [Wan17a, ZH18]. Arbitrary [FHG12a, MKG13, YT17]. Area [DHM11, DM15, JNR11, LZ10a, TGHS18]. Area-Level [DHM11]. Areas [FBM11]. Arising [KJ10]. ARMA [MS18c, ZL15b]. Armed [BFT18]. Array [BJT^{+}10]. Arrays [QA10, ZX17]. ARV [LER^{+}12]. ASA [LS17a]. Assays [BTI19, SFM18, Zha11]. Assessing [BAWM18, CCS14, KA18, PLGM11, PLLG18]. Assessment [HTGT13, LHW^{+}13, LPH^{+}17, SHW17, WH17]. Assessments [LZS^{+}16, RS10]. Assignment [MLC^{+}16]. Assignments [HRS14, ZB13]. Assimilation [SSL^{+}10]. Assistance [JML^{+}14]. Assisted [DH14, WCZ^{+}13]. Association [ACF17, BML17, GWZ13, HLSZ14, JHH14, LS10, LCF14, LZ11b, LL19b, Ma13, TZL17, ZLW10, ZL11, ZJZ12, CDQ14]. Associative [FO17]. Assumption [DH12, GA12]. Assumptions [Pap10]. Astronomical [CMW^{+}19, FG13]. Asymmetric [DGS19, HM15]. Asymmetry [ZSZ12]. Asymptotic [Dui16, FL13, JZ15, MW17, QZL^{+}10, Zha05, Dai19]. Asymptotically [CRZZ16, CLQY16, CMPS17, DV11]. Asymptotics [CGW17, SM12a]. Asynchronous [ASFX10]. Atlantic [RLGL11]. Atmospheric [Cre18a, FSB^{+}19]. Atrophy [AVWW14]. Attenuation [PSR16]. Attributable [LS19a]. Attributes [FH15]. Attrition [FH14]. Augmentation [FFJT16, GC11, JLTY14, SK12, SS17, Zho14]. Augmented [YYC17]. Autocorrelated [LST^{+}11]. Autocorrelation [ZS15a, Zho13]. Autocorrelation-Based [ZS15a]. Autocorrelations [DV11, MS18c]. Automated [PPD^{+}14]. Automatic [CY17, RG17b, SCK19, ZCL11]. Autopsies [MLC^{+}16]. Autoregression [CP12, WLY^{+}14]. Autoregressive [GKM19, IY14, LLT15, PR14, PY12, Tad10, WX15, WC18, YTL15].

Auxiliary [GJL17, HQT16, WCZ^{+}13]. Availability [LHW^{+}13]. Average [CCJ10, CCJ13a, CY17, HP10a, LZWZ11, OR17, SHM^{+}18, WX15]. Averaged [ZD14]. Averaging [AL14, CLLL18, GC11, KRG11, Ma15, RHCT14, SGR10, ZYZL17, ZWZZ19]. Away [AR15b]. AX [KKLL17].

Back [GK18]. Back-Door [GK18]. Bahadur [Ros15a]. Balance [MR15, PKSR15, Zub15]. Balanced [SL14]. Balancing [HFQ12, LMZ18]. Ball [ACPLRC19, TY17]. Banding [BBX16, Bie19]. Bandit [BFT18]. Bands [CLM14, KKC10]. Bandwidth [QC15]. Bankruptcy [CGL10]. Base [BA16]. Base-Pair [BA16]. Based [ACCTW18, BA16, CMM^{+}13, CLXY15, CF17, DT15, DHLL14, DVV14, DG16a, FMP18, FZW16, FL14, Fuk15, GSSVF13, GDLMVF19, GY17, HS12, HH13, HZL^{+}17, HH10, HNW^{+}17, HW10, HGZ18, HLSZ14, HZH15, HZT11, HSZ15, Hun11, JP15, Ken19, LTJM15, LCAL12, LS18c, LCS^{+}13, LLL^{+}10, LJZ14, MS18b, MZN^{+}16, MTY^{+}17, NEB13, NH11, OP13, PRS10, RK11, Sar12, SBG^{+}16, SPZ12, TGZ15, TMP18, WBG^{+}18, WYS19, WW15, Wol11, WG11, WQxYW19, ZLW10, Zha13b, Zha13a, ZS15a, ZZWJ12, ZX17, ZJZ12, ZL15b, RDHL19]. Bases [SWSK19]. Basis [HJ17, RLM13, ZA16]. Basketball [CDBG16]. Bat [MBCM13]. Batch [CY17, LW19b]. Bayes [Bro17, DDV17a, DDV17b, FF17b, HZ13, KM14, KD14, LB17b, MYLC15, WB19c, WMBZ18, WLNC14, XMY^{+}15, YOD11, ZDD19, ZG18, ZD13]. Bayesian [ACBK10, AVWW14, ADR^{+}12, BJT^{+}10, BAA^{+}15, BCR18, BFT18, Blu10, BR12a, BWH16, BR12b, CKP19, CD11, Cha17, CGL10, CIB^{+}18, CSS18, CWZM18, CMER11, DQZ18, DHLL14, DDM^{+}10, DLR11, FOW11, FBCA14, FTM12, FPSE15b, FDPS10, FMV16, FSB^{+}19, FK14a, GNL19, GS10, GdMb13, GDM17, GC11, GKM19, GPD14, GD15, GY17, GAZ13, HC15a, HMM16, HMW18a, HBHC12, HTFK10, Hos13, HC15b, JPCD19, JS11, JPBM13, JL15, JR12, KC11, KJNW11, KCZ^{+}13, KPC^{+}17, KKL15, KKLL17, KMR17, KD13, LNN19, LMZJ13, LNP14, LTJM15, LMR^{+}19, LZ10b, Li12a, LK19, LJGZH19, LHC19, LSY13, LLZ18, LD15, Lin18, Lin13, LGY18, LJW13, Ma15, MMC^{+}12a, MVR17, MBCM13, MD19, MML^{+}13, MDCL13, MR17, MYT18, NC10, NSB17, NSB19, NSH^{+}19b, OPG16, OCAG19, PBD13, PGW^{+}18, PH18, PSV15, PCSK16, Plu17, PSW13, QJWG16, RS10]. Bayesian [RFD11, RDLC10, RBF^{+}17, RHCT14, RG14, RG17b, RD12, RR18, Sad17, SD17a, SPC^{+}18, SCM^{+}18, SZGM14, SLM11, SJDT19, SYS^{+}19, SGR10, SK12, SHF17, SGVC13, SMQ^{+}13, SLR^{+}15, SJ14, TMP18, VS15, VCBT19, WDSL10, WJYJ13, WBCD15, WGE^{+}18, WH17, XTR^{+}14, XMWT16b, YHD^{+}10, YD16, YKC^{+}15, YPQ^{+}16, YPOR18, Zaj12, ZB13, Zho11, ZBHD15, ZMB^{+}16, ZH18, ZKLI14, ZD14, ZSCM18]. Beamforming [ZS15a]. Bed [FMV16]. Before [CW17]. Behavior [CW17]. Being [LS19a]. Benefit [CYT12, SHW17, WFZ18]. Benefit-Risk [SHW17]. Bernoulli [HB14]. Bernstein [BJQ17]. Best [JNR11]. BET [Zha19a]. Beta [ACB^{+}14].

Better [Mor15]. Between [FH15, God17, JK16, KX17, LS10, LSLR12, PLGM11, ST15, TAGT14, XCC18, Zha13a, ZB13, ZNSR12]. Beyond [RR18, ZSZ12]. Bi [MW19]. Bi-Level [MW19]. Bias [CCF18, Efr11b, GK18, GDLMVF19, Li12a, MKG13, SMRGG18, VV15]. Bias-Corrected [Li12a]. Bias-Reduced [VV15]. Biased [yHQ12, KLSY13, QNLS11, TR19, XSWH17]. Biases [HRS14, PSR16]. Big [CCMC16a, PD16, Rod13, WYS19, WQxYW19, Yu19, ZBHD15]. Binary [BAN^{+}12, CHAP16, CW12, DP11, FSMS17, GC11, KLH11, KLM13, NSK16, ST15, SHM^{+}18, WDSL10]. Binomial [BAA^{+}15, ZPS16]. BioCycle [DRC^{+}12]. Biofilms [PPLS18]. Bioinformatics [BBW14]. Biological [ADR^{+}12, LPH^{+}17, PCSK16, ZCJ^{+}17]. Biology [FCDA15]. Biomarker [CZ11, CZW^{+}11, GY17, JHH14]. Biomarker-Based [GY17]. Biophysics [TP16a]. Bipartite [RK11, Sad17, UL17]. Birth [CMS14]. Bitmap [CCSN11]. Bivariate [BBW14, LCCG14, MZN^{+}16, ZHS17]. Black [WLSA17]. Blackwellization [GC11]. Blei [AIP19]. Blessing [LCFW18]. Blessings [AIP19, OST19, WB19a, WB19b]. Blinding [LZS^{+}16]. Block [DG18, God17, KLN13, NL12, Sar12]. Block-Diagonal [DG18]. Blocking [SFM18]. Blockmodel [HPX12, STFP12]. Blockwise [Sar12]. Blocs [SQ10]. Blood [ASS^{+}15]. BNP [DQZ18]. BNP-Seq [DQZ18]. Board [Ano12f, Ano13e, Ano14f, Ano15e, Ano17i, Ano17j]. Boca [Pea19, Tri19]. Body [NC10]. Bonds [CMER11]. Book [Ano10b, Ano10c, Ano10d, Ano10e, Ano11b, Ano11c, Ano11d, Ano11e, Ano12a, Ano12b, Ano12c, Ano12d, Ano13a, Ano13b, Ano13c, Ano13d, Ano14a, Ano14b, Ano14c, Ano14d, Ano15d, Ano15a, Ano15b, Ano15c, Ano16c, Ano16a, Ano16b, Ano17a, Ano17b, Ano17c, Ano17d, Ano17e, Ano18a, Ano18b, Ano18c, Ano19a, Dai19, Har19, HLR15, Kim19, Pea19, Sna19, Su19, Tri19, Vat19, Yan19, Yu19, Zho19]. Boost [Sob12]. Boosting [FKH11, LB18]. Bootstrap [AT11, FCSZ16, FLL17, Hag17, pKP12, OR17, PD14, SVS16, Sha10]. Bootstrapping [CL11c, DGHM19, FPW10, MML12]. Bootstraps [NL12]. Borgan [Pea19]. Both [ZB13]. Bouncy [BCVD18]. Boundary [BMW19]. Bounded [BJQ17, SLW19]. Bounding [IKP11]. Bovine [DDM^{+}10]. Brain [FO17, LS18c, LSS^{+}19a, Lin12, YPQ^{+}16, YPOR18, ZWL^{+}15, ZDD19]. Branching [BSMR12, HV11]. Break [CYZ14]. Breakdown [Cer10]. Breaking [RDG10, SD14]. Breaks [LQS17, PPD15, SD17b, YTL15]. Breast [CWH^{+}15, GHP13, WDSL10]. Breslow [Pea19]. Bridges [LCM10]. Britain [Mar19, PM19, TPAC19a, TPAC19b]. Broad [PLGM11]. Broken [DBNZ16]. Broken-Stick [DBNZ16]. Brownian [ZA16]. Building [BSLR10, PD16, Rod13, TE11].

Calculation [FSL17, LX19]. Calibrated [KRG11]. Calibrating [CHAP16, CYT12, PJY16]. Calibration [BCR18, CCMC16a, CMW^{+}19, FBCA14, KKLL17, KC10, LHD19, PD16, Plu17, SWSK19, SLR^{+}15]. Calibration-Optimal [SWSK19]. California [ADH10, DK18]. Call [Dav13, LHS18]. Calls [HSM^{+}15]. Canadian [CAJ14, HJCPV11]. Cancer [CWH^{+}15, EUW17, GHP13, LDSH16, MZT^{+}17, MTY^{+}17, NSH^{+}19b, SYS^{+}19, WDSL10, WRL^{+}12a, ZCJ^{+}17]. Cancer-Specific [SYS^{+}19]. Cancers [CYT12]. Canonical [GBW16]. Capacity [GMS14]. Capture [BMMS17, BSK^{+}15, KKLL17, YRR19]. Capture-Recapture [BSK^{+}15]. CAR [ZBZ^{+}16]. Carbon [BMMS17, Cre18a, KKLL17]. Cardiac [PJY16]. Cardiovascular [MŞDN12, XAM^{+}14]. Care [BSLR10, HTGT13, LDSH16]. Careers [TEKM12]. Carlo [Vat19, BCVD18, LCT16, LCM10, LX19, WBCD15, Zho14]. Cascading [CR10]. Case [ADH10, CZ11, CZ13, FAIW^{+}11, GWZ13, HGZ18, LJZ14, LZNS17, MŞDN12, Pea19, Sid13, SCHR13, TP16a, TEKM12, TMP18, VHJB13, WSL^{+}16, XWG19, ZK17]. Case-Control [CZ11, CZ13, GWZ13, HGZ18, LJZ14, Pea19, WSL^{+}16, XWG19]. Categorical [BD12, Cra15, HJCPV11, JSPD19, MVR17, MR17, SLR^{+}15, SLC^{+}15]. Causal [Bla17, BS19, BAWM18, CZ19, DGYZ11, EI19, FMV16, FZ13, GMS14, Hos13, Ken19, LDS10, LD17, Lin12, LH14, LM17, MDR18, PBP19, SLM11, SL14, ZD14]. Cause [MLC^{+}16]. Cause-of-Death [MLC^{+}16]. Causes [AIP19, OST19, SHM19, WB19a, WB19b, IJ19]. Cautionary [LHD19]. Celebration [Dav13]. Cell [BFH12, Ion12, LMH14, MMC^{+}12a, MTY^{+}17, HWZ^{+}13]. Cells [PJY16]. Censored [DEV19, DH13, GLL13, GZCL19, HLSY16, LZ11a, LJGZH19, LW19a, NZK11, WF12, WGM12, WC18, WZL18, WY13, WMY15, ZHS17, ZH18]. Censoring [Fra15, SM16, SDS19]. Census [DR10, CTM10]. Centers [HMQA10, LHS18]. Central [CZ14b, LD17]. Centrality [PV13]. Centric [Pan11]. Cessation [GDM17]. CGH [BJT^{+}10]. Chain [BL16, BCVD18, HV11]. Chains [SD17a]. Challenges [IJ19]. Change [BWH16, CMS^{+}18, FSL17, FRG^{+}17, GSDR19, HLL10, LLSS18, LER^{+}12, MJ14, PARS19, SO19, SZ10, XY12, XCH^{+}17, ZL18, Zho13]. Change-Plane [FSL17]. Change-Point [HLL10, SO19, XY12, ZL18]. Changepoint [FR19, RGL16]. Changepoints [KFE12, RLGL11]. Changes [KMO15, KLL17]. Changing [Dup12, GSDR19]. Channel [PJY16, SO19]. Chapman [Pea19, Tri19]. Characteristic [LHC19, RLP^{+}18]. Characteristics [FSB^{+}19]. Characterizing [ZMB^{+}16]. Charting [TGZ15]. Charts [ZZWJ12]. Chatterjee [Pea19]. Checking [DV11, MS18c]. Chemical [XBZ^{+}14]. Chi [Dui16, Zha05]. Chi-Squared-Type [Dui16, Zha05]. Child [KPGJ12, XTR^{+}14]. Childhood [FKH11, FPSE15b, JCRG17]. Chinese [HAME15]. ChIP [KCP^{+}11, MML^{+}13]. ChIP-Seq [KCP^{+}11, MML^{+}13]. Chirp [MBCM13]. Choice [BM10, GRH10, LZWZ11, PGW^{+}18, SWCB16]. Choose [Mor19]. Christopher [Pea19]. Chromatographic [WGDG14b]. Chromosome [LQ11, XWK11]. Chronic [GDZ11, Pre10]. Circadian [KA18]. Circle [KJ10, PBG16]. Circuit [HH13]. Circulation [HB14]. cis} [HSTP15]. cis}-eQTL [HSTP15]. Cities [Dup12]. Civil [Fuk15]. Claims [HZS18]. Class [AGS12, DTYG12, EUW17, HW19, LN14, MW10a, RW13, WCZ^{+}13, XS18, YY16, YOD11]. Classes [ST15]. Classical [SH13]. Classification [BCT18, Büh11, CWC^{+}14, CLXY15, DH13, Düm11, FFJT16, FTM12, GK11, LM11a, LM11b, Li12a, LCAL12, LB18, LZW11, MYLC15, PBD13, PMZ19, Pol11, Sam11, SWW14, WCZ^{+}13, WSSQ16, WN11, YC11, YD16, ZWQ18]. Classified [JRFN18, JCRG17]. Classifier [LCAL12, SQC16]. Classifiers [SLW19]. Classifying [CDH12]. Climate [FRG^{+}17, GH18, LNA10b, RPS19, SBG^{+}16, SSS18]. Clinical [GSH13, HBI^{+}19, JLTY14, LVC14, LER^{+}12, LD15, LZS^{+}16, MHZ15, MZC16, PTC14]. Cloning [LNS10]. Closed [BB10, GPD14]. Clostridium [CMS^{+}18]. Clouds [GPD14]. Cluster [Hag17, HTGT13, JP15, WLS17, XBZ^{+}14]. Cluster-Based [JP15]. Cluster-Robust [Hag17]. Clustered [CKY15, FPW10, FMV16, HRS14, NLOP10, ZK17]. Clustering [AB17b, BDIK12, BT11a, bCH10, CH17, Cra15, HLG14, HSZ15, LMZJ13, SL19, WBG^{+}18, WT10b, WG11, Zha13b, Zha13a]. Clusters [CXCR13, MML12, RTT18, Zha13a]. cmenet [MW19]. CO [Cre18a]. Coarsening [MD19]. Coating [LHH^{+}17b]. Cocaine [GLS11, HLG14]. Cod [CCS14]. Coefficient [CHZ16, FMD14, HLMH18, HHY19, JWXJ13, KXZ19, LHH17a, LLW14, LJZ14, MS15, ŞM10, ZFK14, ZWZZ19]. Coefficients [DR17, FZ16, JK16, LS19b, LSQ15, ZC11]. Cognition [BAA^{+}15]. Cohort [CAJ14, HN18, yHQ12, KWG15, SZGM14, XAM^{+}14, ZL14]. Cointegration [ZRY19]. coli [BBBH10]. coli-SecYEG-Pore [BBBH10]. Collaborators [Ano10a, Ano11a, Ano12g, Ano13f, Ano14g, Ano15f, Ano17k, Ano17l, Ano19e, Ano19f]. Collapsed [BOSB16]. Collective [MZN^{+}16]. Combination [BCR18, CLQY16, SCG13]. Combinatorial [WBCD15]. Combined [WY19]. Combining [DNFZ10, HMW^{+}14, LN14, RHCT14, TWYZ11]. Comment [ALM12, AIP19, BFH12, BW15, BC15, Ble13, Ble17, Bro17, BM15, BJ16, Büh11, Cai10, CT17, CvdL12, CL15a, Che13, CZ14a, CLZ^{+}16, CH19, CB18a, CR18, CT10, CO10, D’A19, DH12, Det13, Du14, Düm11, Dun14, FY17, FF17b, GV14, GJ18b, GK17, GK11, GS17b, Gri13, GLR16, GL14, HL16, Han17a, HR16, Har18, HM16, Hel10, Her18, Hjo14, Hod15, HP10b, Hud15, IJ19, Ion12, Jin12, Kar18, Kne17, Kou12, Kuh18, Laz11, Lee15, Lem15, LNA10a, LMZ15, LB17b, LNM11, LK16, LvdL17, Mam13, MZW16, Mor16, MQ14, Ogb17, OST19, PD16, PY12, Pol14, Pol11, Qia17, Raf17, RG17a, Ros17, Rot18, Sah10, Sam11, Sch10, Sch12, SS15, She11, She13, Smi10, Spi19, SL11, Tad17, TB17, TP16b, Tu17, WSWT10, WSG16, WMS17, WSL14, Was12, WN11, Wes10, WH15, Woo17, Wu11, YC11, Yee17, ZW13, Zen13]. Comment [ZL15a, ZZ18]. Comments [DVV14]. Common [HPV14, ZC17a]. Commons [SQ10]. Communication [GC17, JLY19]. Communication-Efficient [JLY19]. Communities [CL18, GBP19, RBF^{+}17]. Community [BWH16, GBP19, HPX12]. Compact [MML12]. Comparative [ADH10, JTF^{+}19, ZTC^{+}13]. Compare [GRR^{+}17]. Comparing [ST15]. Comparison [CH19, CSS18, CH17, Hah12, LLL^{+}10, LPH^{+}17, MTY^{+}17, PKM10, PL19, RLP^{+}18, SDT10, SJDT19, Spi19, YZ19, ZEL19a, ZEL19b]. Comparisons [BT16, FS17]. Competing [BKRF19, LY16, MTY^{+}17, SW17c, YHD^{+}10]. Complement [SLW16]. Completely [DD16]. Completion [CCZ16, MCW19]. Complex [AT11, ADR^{+}12, CMS17, DHS11, GDLMVF19, HGK12, MDR18, TE11, YPOR18, Yu19]. Complex-Valued [YPOR18]. Complexity [FSLR19]. Compliance [BG11, Sid13]. Complier [KHK19]. Component [ASX19, BSB15, CL15b, FCSZ16, HM15, HL18, HG18, HJCPV11, LST^{+}11, LG14, LL19b, MT17, RMR19, SdCG^{+}15, SLW16, ZSPL17]. Components [AGS12, BGH^{+}18, HPV14, KS15, LWC13, MT11, MH18, PY16, PSY19, RN16]. Composite [AY17, BGMP12, FSMS17, GS10, GSSVF13, yHQ12, KXZ19, SM12b, ZX17]. Composition [CLL19]. Composition-Adjusted [CLL19]. Compositional [CLL19, HAME15, SdCG^{+}15, SW17a]. Compound [BGR13, KM14]. Comprehension [GDCL11]. Compressed [GD15]. Compression [GH18, LQ17]. Computation [Blu10, CLGS18, CH17, SJDT19]. Computational [KFE12, MCIS19, SLR^{+}15]. Computationally [CMPS17]. Computer [BJ11, FSB^{+}19, Plu17, SWSK19]. Computerized [CLvdL19]. Computing [SMQ^{+}13, WG11]. Concave [MH17]. Concealment [Sna19]. Concentrations [GBDL10]. Concerning [MM13]. Concordance [CMW^{+}19]. Concordant [CMS^{+}18]. Concurrent [DRS18]. Condition [ACPLRC19, CSS18]. Conditional [BFV16, Bla17, DM18, Efr10a, FJM19, GH18, JR14, JT12, KNK^{+}18, KRS^{+}17, LCZ12, Ma13, MW19, RW13, WfTQ12, WLH12, WL13, WPH^{+}15, WT13, YD16, ZC13]. Conditioning [Sch12]. Conditions [BCdB14, LRZ17]. Conduct [JLTY14]. Confidence [BBS17, Büh11, CLM14, Düm11, GK11, HZ13, JZ15, KKC10, LM11a, LM11b, PCB11, Pol11, Sam11, SSS18, Wan15, WN11, WFB13, Wil19, XSS11, YC11, ZYH14]. Confidentiality [HTFK10]. Confirmatory [BGB12]. Confounders [MRB12]. Confounding [GK18]. Conjoint [EI19]. Conjunction [WO19]. Connecting [ZDD19]. Connectivity [HH13, LW10a, Lin12, WGE^{+}18, ZWL^{+}15]. Conservative [ZSS19]. Consideration [WFZ18]. Consistency [Fu16, GKM19, SJDT19, WB19c]. Consistent [BR12a, NSH19a, STFP12, TE11, XQZ10]. Consonant [BB10]. Constants [LJSL16]. Constrained [CLL11, CDH11, CCMC16a, FS17, Fu16, LHW^{+}13, PD14, SB18, SKBM17, TT10]. Constraint [CLZ10]. Constraints [DRC^{+}12, DJP18, FZY12, HRC12, KM14, THY16]. Constructed [PSR16]. Constructing [FSLR19, ST17, WZ17a]. Construction [LS18a, PK12]. Constructions [PCJ12, SY18]. Contain [JK16, ZWMC17]. Contaminated [YY16]. Contested [GRH10]. Contingency [HS13, KD13]. Continual [YY11]. Continuous [ASFL^{+}17, GW15, HJCPV11, KCK^{+}15, LCM10, MR17, PLGM11, SY17, SLC^{+}15]. Contoured [BF15]. Contrasting [ZNSR12]. Contributed [CCS14]. Control [ADH10, BCDS18, CZ11, CZ13, CB18b, EG19, FKSZ19, FG13, GWZ13, HZ10, HGZ18, HZZ10, KLH11, LCT16, LB17a, LJZ14, Pea19, PSR16, RSK17, WSL^{+}16, XWG19]. Controlled [FAIW^{+}11, MSM^{+}15, OPG16, ZCC^{+}17]. Controlling [KN18, SCG13]. Controls [Ros14]. Convergence [LS15, ZR15]. Convex [BBX16, KM14, MCIS19, RDHL19]. Convoluted [XBZ^{+}14]. Cooling [LCL14]. Cooperative [WGM12]. Coordinate [FZ13, MFH11]. Copula [CW14, DVV14, Fuk15, KHG18, LG13, MDCL13, NEB13, OP13, PCJ12, SY18, SMAC10, SK12]. Copula-Based [DVV14, Fuk15, NEB13, OP13]. Copulas [BG11]. Copy [BJT^{+}10, HLSZ14, JCL10]. Corporate [CMER11]. Correct [DO12, NSK16]. Corrected [BTI19, Li12a, WMY15]. Correcting [LZ11b, Per10, SEdS19]. Correction [Ano10f, Ano11f, Ano11g, Ano11h, Ano12e, Ano14e, Ano17f, Ano17g, Ano17h, Ano19b, Ano19c, Dui16, GDLMVF19, LSS17, LW10b, LW19b, ML13a, PKSR17, WT10a]. Correlated [AS15, Cai10, Efr10b, Efr10c, Hel10, LMR^{+}19, LHTB15, MB16, QLL10, RB18, Sch10, SLR^{+}15, Wes10, XQZ10, ZBZ^{+}16, ZDP10, ZHM^{+}10]. Correlation [DR17, GHK10, HR11, LZZ12, LSQ15, MLT17, SZ14, WPH^{+}15, ZSZ12, ZQ12]. Correlations [BDIK12, CL16a, LLT15, ST15]. Corridor [ZYH14]. Corrigendum [Ano18d, Ano19d]. Cortex [SKS^{+}15]. Cost [KFE12, LWN^{+}18]. Costs [GAZ13]. Count [BWH16, CCS18, DQZ18, Ima11, KKL15, MS18b, SEdS19, ZZWJ12, ZPS16]. Count-Valued [BWH16]. Countable [Guh10]. Counterclaims [Ros15b]. Countering [Cha17]. Counting [CCSN11, SPHL16]. Counts [CD11, LNN19, ZLP^{+}14]. Coupled [LLLW11]. Couples [KWG15]. Coupling [MW11]. Course [SW11]. Covariance [ASFX10, AGS12, BGMP12, BBX16, Bie19, CL11a, CLX13, CY16, CLL19, CY17, CZZ10, CL16b, DG18, FHG12a, FCDA15, GKS10, GBZ17, GZZ19, Han17b, HFSZ19, LZP10, MT17, MM13, PBG16, QC15, TLG16, XMZ12, YP17, YKC^{+}15, ZA16, ZLL18, ZLWT17]. Covariate [BCS18, CYC10, CRZZ16, DZ19, FF13, HZH15, LZ10b, MHZ15, MCW19, PMZ19, PKSR15, SW14, SLC^{+}15, YNLC19, ZJZ12]. Covariate-Adaptive [BCS18, MHZ15]. Covariate-Adjusted [CRZZ16, DZ19, HZH15, PMZ19, ZJZ12]. Covariates [CCZF15, CJN18, DZ15, HLMH18, LS10, LMZ18, LLW14, LM17, MR15, NTC13, SC12, SM16, SS17, TAGT14, WF12, WH17, YJFL19, YMSC15, ZWMC17, Zub15]. Coverage [CCF18, CTM10, MN16]. Cox [HQT16, RD12]. CPS [HBZ15]. CRC [Pea19, Tri19]. Credible [BR12a]. Crime [MSB^{+}11, RSK17, Tad10, TEKM12]. Criminal [RD12, TEKM12]. Crisis [WG11]. Criteria [GS10]. Criterion [LNP14, ZLT10]. Criticism [BML17, PS10]. Critics [PS10]. Cross [AGS12, CAJ14, CG15, CL18, DS12, GKS10, HJRS10, JCRG17, PBG16, SHS10, ZSR18, ZMB^{+}16]. Cross-Classified [JCRG17]. Cross-Covariance [AGS12, GKS10, PBG16]. Cross-Dimensional [DS12]. Cross-Match [HJRS10]. Cross-Neuronal [ZMB^{+}16]. Cross-Screening [ZSR18]. Cross-Sectional [CAJ14]. Cross-Validated [CG15]. Cross-Validation [CL18, SHS10]. Crossover [HZ10]. Cubic [SLS18]. Cubic-Rate [SLS18]. Cum [LC14]. Cumulative [PHP^{+}12, TGZ15, XAM^{+}14, ZZF10]. Cure [MW10a, WY13]. Currents [BBBH10]. Curvature [TMP18]. Curve [HC15b, LDGX15, TEKM12]. Curves [CGBY13, DMVB18, GPD14, KSKD12, SKBM17, SCK19, ZDD19]. Cutoff [AR15b]. Cycles [LTJM15]. Cyclone [RLGL11].

DAE [RSI14]. DAE-seq [RSI14]. Daily [CGBY13, WLSA17]. Data [ACBK10, ASFX10, AB17b, ASX19, BJT^{+}10, BBW14, BAA^{+}15, BJQ17, BL16, BMW19, BGMP12, BIL^{+}11, BKD^{+}17, BD12, BAN^{+}12, BA16, BWH16, BGK^{+}18, BR12b, CCZ16, CCT15, CLL19, CAJ14, CDH12, CCJ10, bCH10, CHAP16, CL10, CCMC16a, CYC10, CCMW11, CWPC13, CF17, CL18, CIB^{+}18, CCS18, CGO19, CMS17, Cra15, DQZ18, DYZ12, DBNZ16, DHM11, DP11, DM18, DM11, DH13, DZ15, DHLL14, DZ19, DS12, DDM^{+}10, DPT14, DLR11, DPS19, DDH15, DLP12, FOW11, FLY12, FLJL17, FPW10, FPSE15b, FDPS10, FF17a, FMPR12, GLL13, GS10, GA12, GCNC14, GC11, Gil11, GRH10, GLS11, GZZ19, HMQA10, HAME15, Han14, HL14a, HZ14, HG18, Har19, HLSY16, HCKS18, HZS18, HTGT13, HRZ17, HTFK10, HFSZ19, HFQ12, HSTP15, HD10, HZT11, HLG14, HMW^{+}14, HSZ15, HLR15, HJCPV11, HZ13, IKG^{+}10, JZ15, JNR15, JAW16]. Data [JLTY14, JKL11, JP15, JSPD19, KJNW11, KOL^{+}12, Ken15, KY11a, Kim19, KMO15, KKL15, KWG15, KNK^{+}18, KHG18, KCP^{+}11, LVC14, LDS10, LDSH16, LMR^{+}19, LNS10, LZP10, LZ11a, LWC13, LY16, LHH17a, LQS17, LS18c, LJGZH19, LW19a, LCS^{+}13, LN14, LFU11, LSZD17, LD15, LRZ17, LLgX14, LZPH14, LJZ14, LS17b, LT14, LMH14, LSS^{+}19b, MS18b, MVR17, Mar19, MJ14, MRB12, MZ15, MDG17, MML^{+}13, MDCL13, NZK11, NC10, NN17, NLOP10, NCB12, NSK16, NFG^{+}16, OKEK17, PCJ12, Pan11, PD16, PM19, PBP19, QNLS11, QLL10, QJWG16, QKVT19, RQJ15, RSI14, RBB13, RHCT14, RGH13a, RGL16, RSK17, SD14, SdCG^{+}15, SW17a, SW19, SBG^{+}16, SO19, SVS16, ŞM10, SLS18, SMAC10, SK12, Sob12, SL14, SEdS19, SSL^{+}10, SSZL12, ST18, TLG16, TYY^{+}19, TZF^{+}15, TPAC19a, TPAC19b, TE11, VH15, WDSL10, WGM12, WJYJ13, WM14]. Data [WS14, WG14, WYS19, WGE^{+}18, WSL^{+}16, WRCG13, WZL18, WLZ19, WZL19, WLSA17, WY13, XWK11, XKBS17, XSWH17, XLN18, XWG19, XQZ10, YY16, YRR19, YL13, Yu19, ZC13, ZPlW13, ZS15a, ZH19, ZS15b, ZYH14, ZHM^{+}10, ZQ12, ZZWJ12, ZLZ13, ZHS17, ZLL18, ZH18, ZLLZ11, ZFK14, Zub15, Yan19, Yu19]. Data-Centric [Pan11]. Data-Driven [CCT15, CCJ10]. Database [KN18]. Datasets [DBFG16, DVF11, Guh10, Kat17, LS19b, MML12, ZBZ^{+}16]. David [AIP19, Tri19]. DD} [LCAL12]. DD}-Classifier [LCAL12]. DD}-Plot [LCAL12]. Death [CMS14, DGYZ11, LDS10, MLC^{+}16]. Decision [BGR13, CGL10, ZMB^{+}16]. Decision-Making [ZMB^{+}16]. Decisions [KLM13, KM14, Mor15]. Decomposing [DFM19]. Decomposition [AGT14, CCL^{+}11a, CCL^{+}11b, Laz11, LNM11, She11, SMAC10, Wu11, ZH19]. Deconfounder [D’A19]. Deconvolution [ADR^{+}12, CDH12, DH14, SPC^{+}18]. Deconvolve [JAW16]. Decoupling [HC15a]. Deduplication [SHF17]. Default [DTYG12]. Deficits [GDCL11]. Defining [MZT^{+}17]. Definite [XMZ12]. Definition [SCHR13]. Definitive [ZX17]. Degeneracy [Sch11]. Degeneration [HS12]. Degrees [KS11, LMR14]. Deicing [LHH^{+}17b]. Delayed [JLTY14]. Delays [NFG^{+}16]. Demand [LS17b, ZMW^{+}15]. Dementia [CAJ14, yHQ12]. Demonstrating [BBBH10]. Dense [ZB18]. Densities [WW15]. Density [BG16, CDH11, CCJ10, CLQY16, CGW17, DH14, Efr10a, FPSE15b, Gee14, GMS14, HJCPV11, JT12, LJZ14, LJW13, LT14, MHC11, MZN^{+}16, MMNS11, SPC^{+}18, SHS10, TARS17, TT12, dCD14]. Density-Ratio [LJZ14]. Density-Weighted [CCJ10]. Departures [TEKM12]. Dependence [BBS17, BT16, CZ19, CO10, FHG12a, GR11, GLS11, HP10b, HW19, JK16, KD13, MJPW13b, MR17, Sah10, SD17b, SMAC10, VH15]. Dependencies [FH15, RSI14]. Dependency [MM19]. Dependent [BJQ17, CWC^{+}14, DS12, FAIW^{+}11, HCM13, LCLZ18, MS18c, OCAG19, RPS19, Sha10, SSZL12, TZF^{+}15, XAM^{+}14, ZFW17]. Depends [Fuk15]. Deployment [FRG^{+}17]. Depth [CHSV14, NN17, PV13]. Derivatives [CCJ10, CCJ13a]. Derived [GLS11, QA10, ZPS16]. Descent [FZ13, MFH11]. Description [ACLZ14]. Design [BIZ15, CL10, FMP18, GJ18a, GJS16, HSR13, IS12, JPCD19, KYBB19, LHH^{+}17b, LGY18, Ros10, Ros14, RLP^{+}18, SCG13, SCHR13, TGP11, ZB13, ZDP10]. Design-Based [FMP18]. Designed [MW10b]. Designs [AT11, BJ11, BGB12, DT12, FMV16, GA12, God17, He17, HZ10, HZH15, Hun11, JM14, KW15, LCLZ18, Pap12, Plu14, Qia12, SVTG17, ST17, VCBT19, WLS17, XCQ15, YBT13, ZC11, ZX14, ZX17]. Detect [CXCR13]. Detected [CYT12]. Detecting [CMS^{+}18, FG13, MQ15a, PARS19, TP16a, YP17]. Detection [ASB11, ACCTW18, BTI19, BMW19, Cer10, FSL17, FR19, GSDR19, GBP19, HZT11, JP15, JHH14, KFE12, KMO15, LNN19, NFG^{+}16, PAHJ11, PCSK16, PPD15, RPS19, SO19, SKS^{+}15, SO11, SCK19, TARS17, YPOR18, ZLP^{+}14, ZCC^{+}17, Zho13]. Detections [LBS13]. Detects [SLC^{+}15]. Determination [KXZ19]. Determinations [JML^{+}14]. Determine [Tra18, WFB13]. Determining [CL18]. Deterministic [Plu14]. Development [XTR^{+}14]. Developmental [FK14a]. Developments [Sha15]. Deviation [HS12]. Diabetes [WFZ18]. Diagnosing [BMW19]. Diagnosis [HC15b]. Diagnostic [CLXY15, CLQY16, LHC19, MS18c]. Diagnostics [LZ18]. Diagonal [DG18]. Diagram [Fu16]. Diagrams [CWH^{+}15]. Dialect [Mar19, PM19, TPAC19a, TPAC19b]. Dialysis [MŞDN12]. Dichotomization [LS17a, MG17a, MG17b]. Didanosine [XAM^{+}14]. Difference [LS18b, RRW17, SZ14]. Differences [TP16a]. Different [HG18, LS15, LN14]. Differentiable [LvdL18]. Differential [BIL^{+}11, BCdB14, CSW17, DQZ18, MWX14, SLC^{+}15, WZ10, WLXL14, WQxYW19, XCM^{+}13]. Diffusion [KOLL12, LCM10, PT17, SDT10, YL13]. Diffusion-Weighted [YL13]. Diffusions [LT10, TT12]. Digraphs [KM17].

Dimension [BH13, BSZ12, BB14, CH12, CZ14b, DW11, Efr10a, FWYZ13, FL14, GCBL15, GLLL15, HC17, LS18b, LLZ10, MZ12, QDC19, WWL12, XZX10, YNLC19, ZZF10]. Dimension-Reduced [BB14]. Dimensional [AL14, AB17b, BCK19, BR12a, CLX13, CLvdL19, bCH10, CHAP16, CZZ10, CZW^{+}11, CCMW11, CSW17, CFL18, CLLL18, CW19, CHZ16, CLZ15, DS12, DG18, FFS11, FMD14, FFJT16, FK18, FZW16, GS10, GKM19, GHJZ16, GBZ17, GWCL19, HL14a, HL18, HZ14, HFZ18, HG18, HSZ15, HZ13, IKG^{+}10, JR12, KXZ19, LHL15, LLSS18, LZ10b, Li12a, LSY13, LSQ15, LL13, LFL15, Lin18, LLW14, LLLW11, LC14, MB16, NSB17, PGY14, PWL16, QC15, RGH13a, RSK17, RT17, SZ14, SL15, SZLI15, VRS19, WPL15, XWK11, XC14, YD16, YP17, YYY18, Zha13b, ZC17b, ZH19, ZC11, ZLLZ11, ZSPL17, ZB18, ZZS19]. Dimensionality [HLMH18, LCFW18]. Dimensions [PMZ19, PS10, RJ10, ZSCM18]. Dimming [MMM11a, MMM11b, SL11]. Dioxide [Cre18a]. Direct [AT11, CL11b, WKKS12]. Directed [HCCZ16, VCBT19, YJFL19]. Directional [DFRS14, SW17a, SW19, WG14, ZWL^{+}15]. Dirichlet [BPPD15, HTGT13, ZEMD18]. Disablement [IZR^{+}13]. Discipline [Gel11]. Disclosure [MVR12]. Discontinuities [ZFK14]. Discontinuity [AR15b, CCT15]. Discontinuous [ASFL^{+}17]. Discovering [BH13]. Discovery [BCDS18, BFT18, BH18, FHG12a, FKSZ19, GR11, HZZ10, SKS^{+}15, TKPS18, ZCL11, Zha11]. Discrepancy [BMMS17, QDC19]. Discrete [BBS12, BM10, DTYG12, FMGS16, HW10, KLL17, LLgX14, PCJ12, Sch11, SY17, SK12, VW15]. Discriminant [CL11b, CLZ15, PWL16]. Discrimination [QZL^{+}10]. Discussion [AD19, DPS19, LB17b, Mar19, PL19, PM19, YZ19]. Disease [ASB11, AVWW14, GZJLM12, HDL^{+}16, Ken15, LG14, NFG^{+}16, Pre10, XXL^{+}19, ZCC^{+}17]. Disentangling [FMV16, SMRGG18]. Disparate [FPSE15b]. Dispersed [HN18]. Distance [BSK^{+}15, CZ19, LZZ12, LMR14, MT17, PBG16, QZL^{+}10, WPH^{+}15]. Distinct [CCSN11]. Distinctive [MZT^{+}17]. Distorted [DZ19]. Distress [TNZM14]. Distributed [DH12, JLY19, WG11]. Distribution [ACCTW18, AS15, DDT17, Han17b, Hua14, LRW13, LGR^{+}18, PARS19, SYY15, WfTQ12, ZG18]. Distribution-Free [ACCTW18, LRW13, LGR^{+}18]. Distributional [RW13]. Distributions [BCR18, Dui16, HMW18a, HL18, JZ15, KJ10, LJSL16, MJG18, SW19, Tri19, VS15, WLH12, WTM19, WT13, XSS11, Zha05, ZFW17]. Divergence [LS18b]. Diversity [ACBK10, SPU17]. DNA [KSGB10, PKM10]. Do [KLH11, MMNS11, MSZ10]. Do-Validation [MMNS11]. Does [JK16, Sob12]. Domain [GCNC14, MS18a, Zho11]. Domains [HG18, NC10, XTR^{+}14]. Domestic [Sid13]. Don’t [KBKS18]. Door [GK18]. Dose [CZK17a, GY17, LTJM15, TNZM14, YY11, ZB13]. Dose-Finding [LTJM15]. Double [HFQ12, JLPZ17, SVS16]. Doubly [VV15]. Downstream [SHM19]. Driven [BIL^{+}11, CCT15, CCJ10, CWH18, Gil11]. Driving [KCZ^{+}13]. Dropout [GDM17]. Drosophila [ZC17a]. Drug [HZT11, LLZ18]. Dual [CTM10]. Due [BR12b, LDS10]. Duration [Fuk15]. During [FO17, ZMB^{+}16]. Dynamic [BMMS17, BT11b, CL16b, CIB^{+}18, CLLL18, CGO19, DM18, DK18, FBCA14, FO17, FF17a, GBDL10, GBZ17, Hua17, Ion12, JYL15, JAW16, KLL17, KMR17, Lin13, LLLW11, MMC^{+}12a, MT11, MYT18, PGW^{+}18, PY16, PSY19, RDLC10, RHCT14, SC15, SL19, Tad10, TGP11, WRL^{+}12a, WGE^{+}18, WLXL14, XMWT16b, Zaj12, ZDP11, ZWL^{+}15, ZLDT18, ZZLK15, ZMB^{+}16, ZZS19]. Dynamical [CB18b]. Dynamically [RLM13]. Dynamics [KWG15, PR14, TP16a].

E-mail [FSS^{+}16]. E-MS [JNR15]. Early [EUW17, LER^{+}12]. Early-Stage [EUW17]. EBIC [LC14]. ECA [HL18]. ECoG [ZWL^{+}15]. Economic [RGH13a]. ed. [Vat19, Yan19]. Eddy [MSW^{+}18]. Edge [CCS18, CD18]. Edge-Count [CCS18]. Edit [KCK^{+}15, MVR17]. Edit-Imputation [KCK^{+}15]. Editor [Ano10h, Ano10i, Ano11j, FHL15]. Editorial [Ano10a, Ano11a, Ano12g, Ano13f, Ano14g, Ano15f, Ano17k, Ano17l, Ano19e, Ano19f, Ano12f, Ano13e, Ano14f, Ano15e, Ano17i, Ano17j]. eds [Pea19]. Educational [VHJB13, ZSS19]. EEBoost [Wol11]. EEG [KMO15, LSS^{+}19a, SO19]. Efavirenz [LZNS17]. Effect [ADH10, BAWM18, CCF18, DFM19, FMPR12, HSR13, JTF^{+}19, KX17, KBKS18, LER^{+}12, MSK^{+}17, PTC14, PSSS17, Sch12, SHM^{+}18, WM14, XAM^{+}14]. Effective [HC17, ZWL^{+}15]. Effectively [ZTC^{+}13]. Effectiveness [ZK17]. Effects [ASFL^{+}17, AR15b, BML17, BGC19, Bla17, CR10, Cho17, Dai19, DHM11, DM15, DGYZ11, Fra15, FOvS10, GLL13, GW15, GA12, GMS14, GDZ11, HAME15, HV11, HZ10, KC11, KOL^{+}12, Ken19, KHK19, KY11b, KPGJ12, LVC14, LDS10, LSS11, LHTB15, LQS17, LDL^{+}19, LH14, LZS^{+}16, LLLW11, LW19b, MW19, OR17, PHP^{+}12, RLY14, SCM^{+}18, SHM19, SW17a, SM12a, Sid13, Sob12, SWCB16, TGHS18, TPLG10, VHJB13, WA18, WfTQ12, WLS17, ZYZL17, ZHM^{+}10, ZD14]. Efficacy [GSH13, LTJM15]. Efficiency [FHS13, HZH15, KX17, LLX15, Ros10, Ros15a, SS17, Tad13b, Tsi17, Zaj12]. Efficient [BS13, CLvdL19, CLGS18, CWZ15, CRZZ16, GW15, HPV14, HZ10, HQT16, JZT19, JLY19, LCT16, LYH13, LX19, MKG13, MS18a, MSW^{+}18, MW10a, QLL10, QKVT19, TZL17, VW11, WZL18, YBT13, ZL14, ZLL18]. Efficiently [MSZ10]. Eigenanalysis [ZRY19]. Eigenvalues [SDT10, YZBE18]. Eigenvectors [SDT10]. Elastic [Han11, PPD^{+}14, SKBM17, SCK19]. Elderly [CAJ14, HH10]. Election [SMRGG18]. Elections [GRH10, Lin13]. Electricity [CGBY13, LS17b]. Electrocorticographic [ZWL^{+}15]. Elimination [LQ11]. Elliott [YZ19, Spi19]. Elliptical [HL14a, HL18]. Elliptically [BF15]. Embedded [NQS14]. Embedding [STFP12]. Embracing [LCFW18]. Emergence [XXL^{+}19]. Empirical [AO17, AS15, BS13, HLSY16, HZ13, KLN13, KM14, KWX^{+}19, LJZ14, RW10, WMBZ18]. Emulation [FBCA14, GH18, MSW^{+}18]. Emulators [CMM^{+}13]. EMVS [RG14]. Encompassing [Gel11]. Encouragement [FMV16]. End [GAZ13, LDSH16]. End-of-Life [GAZ13, LDSH16]. Endogeneity [Fra15, FH14]. Endogenous [KPGJ12]. Endometriosis [HC15b]. Enriched [SD14]. Enrichment [WY19]. Ensemble [CDQ13, CDQ14, DO12, SSL^{+}10]. Ensembles [RGH13b, SGR10, SDS19]. Entrant [HJCPV11]. Envelope [CZ15, GLLL15]. Environment [HRC12, HZL^{+}17]. Environmental [HD10, LDGX15]. EOV [Ano12f, Ano13e, Ano14f, Ano15e, Ano17i, Ano17j]. Epidemic [FBCA14]. Epidemics [DLP12]. Epidemiological [HMW^{+}14]. Epidemiology [DGH^{+}10, MZT^{+}17]. Epileptic [SO19]. Epistatic [Zha13a]. eQTL [HSTP15]. Equality [GCBL15, GZZ19, ZSCM18]. Equation [BIL^{+}11, LSS^{+}19b, MRB^{+}17, MWX14, WZL18, XCM^{+}13]. Equations [BCdB14, CSW17, NQS14, WRL10, Wol11, WLXL14, WQxYW19]. Equipment [OCAG19]. Equity [Zaj12]. Equivalence [DMVB18, FL13]. Equivalent [LSQ15]. ERIC [HWF15]. Erratum [Ano12h, Ano12i, Ano18e]. Error [ABvdW19, BGK^{+}18, Büh11, CDH11, CMM^{+}13, CG15, CGLY10, CFL18, DDH15, Düm11, FG13, GHK10, GK11, GLS11, Hua14, KLH11, LM11a, LM11b, LS19a, LM17, MR11, MŞDN12, NSK16, PY12, Pol11, SLW19, Sam11, SH13, SW14, SWW14, TR19, WN11, WSW17, YY16, YC11, YMSC15]. Error-Contaminated [YY16]. Error-Prone [DDH15, LM17]. Errors [DH12, MS18c, SM16, SL14, SMQ^{+}13, WMY15, ZWMC17]. Errors-in-Covariates [SM16]. Essential [Nus18]. Estimability [LNS10]. Estimands [BS19]. Estimate [PHP^{+}12]. Estimates [ASFX10, Cai10, CG15, Efr10b, Efr10c, FPW10, Hel10, MTC11, Sch10, WMBZ18, Wes10, XKB12]. Estimating [ADH10, AB13, BGC19, BGMP12, CAJ14, CDH11, CYC10, CHC^{+}12, CB18b, CMER11, DK18, DGH^{+}10, FHG12a, FCDA15, HB14, JTF^{+}19, JKL11, LS19a, LZPH14, MVR12, MRB^{+}17, MSZ10, NQS14, QLL10, RN16, TAGT14, VT17, WRL10, WGE^{+}18, Wol11, XAM^{+}14, YZBE18, ZZRK12, ZZLK15, ZMHKK17]. Estimation [ASFL^{+}17, ABvdW19, AR15b, ACPLRC19, ASS^{+}15, BOSB16, BT11b, BCdB14, CL11a, CLL11, CL11b, CY16, CCF18, CLL19, CLvdL19, CW17, CCMC16a, CGLY10, CRZZ16, CFL18, CSS18, Cho17, CMS14, CWH18, DYZ12, DBNZ16, DM15, DH14, DGYZ11, DJW18a, Efr10a, Efr14a, FLY12, FK18, FZ16, FFvSK17, FPSE15b, FHS13, FMV16, FOvS10, FLL17, GLL13, GW15, GBW16, Gee14, GPPVW12, GRR^{+}17, Gil11, GDZ11, GSSVF13, GWCL19, HPV14, HM15, Han14, HCCZ16, HLL10, Har19, HFSZ19, Hos13, HGZ18, HFQ12, HHY19, yHQ12, Hua14, HMW^{+}14, HQT16, HC17, HMW18b, HC15b, IR15, JT12, JNR11, JLPZ17, JCRG17, KX17, KZCS16, KLCT11, KY11a, KOLL12, LDGX15, LLSS18, LZ10a, LCZ12, LWN^{+}18, LYH13, LZS^{+}16, LZ10c, LJW13, LT14, MHC11, MKG13, MZN^{+}16, MMNS11, MW10a, MM15, MM13, NZK11, NEB13, OCAG19, OP13, PTC14, PDF15, PD14, QP13]. Estimation [QC15, RTT18, RRW17, RD12, RWS12, RT17, Sad17, SHS10, SH13, SYY15, SWJ18, SPZ12, SYS^{+}19, SK12, SWCB16, SS17, TGHS18, TT12, Tsi17, VV15, WA18, WLH12, WGM12, WL13, WM14, WX15, WC18, WFDS19, WZL18, WZL10, WT13, WQxYW19, XZX10, XMWT16b, XSWH17, XMZ12, XCM^{+}13, YT17, YTL15, YLRH19, You19, YL13, ZL14, ZYZL17, ZEMD18, ZQ12, ZW14, ZLL18, ZZF10, ZD14, Zub15]. Estimations [QNLS11]. Estimator [GMS14, Sar12, TR19, WW15, Zho14, ZWZZ19]. Estimators [AI12, AY17, ACF17, BY11, BSB15, CCJ13a, Cer10, CY17, CL11c, CW12, CZ14b, DZ15, Fu16, KKC10, LS15, LZWZ11, LCLZ18, OR17, SLS18]. Ethics [HZH15]. Etiologic [MZT^{+}17]. European [RWF^{+}13]. Evaluating [AB17a, FMPR12, Gne11, LER^{+}12, OKEK17, RSK17, SW17c]. Evaluation [CZ11, CTM10, DIvdB13, ESR18, FL10, FH14, OPG16, WRL^{+}12a]. Evaluative [SL14]. Event [BAN^{+}12, BGK^{+}18, FAIW^{+}11, Fuk15, GSH13, IS12, KNK^{+}18, LW19a, LZPH14, PCC12, RHCT14, SSZL12, SHW17]. Event-Dependent [FAIW^{+}11]. Events [CW17, GZCL19, KCZ^{+}13, LLgX14, RPS19, SPHL16, SW17c, XCH^{+}17]. Ever [Nus18]. Every [XCQ15]. Evidence [bCH10, JPCD19, LS17a, MMM11a, MMM11b, MG17a, MG17b, Mor10, OPG16, Ros11, SL11, ZSL^{+}11, ZNSR12]. Evolution [FO17]. Evolutionary [JPBM13]. EWMA [ZZWJ12]. Exact [AEI18, BS19, CMPS17, FCSZ16, LLgX14, MM19, Mar15, TTLT16a, Wan15]. Exam [AR15b]. Examples [Har19]. Exceedances [You19]. Excess [TR19]. Exchange [LJSL16]. Exchangeability [RGH13b]. Exchangeable [CD18, ZFW17]. Excited [WLY^{+}14]. Exciting [FSS^{+}16, MSB^{+}11]. Excursion [SSS18]. Exogenous [WC18]. Expansion [BSZ12]. Expansions [GMS14]. Expectation [BC14, QP13]. Expectation-Maximization [QP13]. Expected [Zha11]. Expenditure [SW17a]. Experience [SY18]. Experiment [FO17, FSB^{+}19, LHH^{+}17b, LSLR12]. Experimental [FZ13, Plu14]. Experimentation [MW10b, RSA16, Wu15]. Experiments [ABvdW19, BJ11, BF18, Bla17, BS19, CMM^{+}13, CSSK16, Cho17, DD16, EI19, GJS16, HWZ^{+}13, LN14, MP13, MDR18, PLLG18, SW11]. Exploiting [FL10]. Exploration [HPX12]. Exponential [BDF16, DFRS14, DHS11, HNW^{+}17, Sch11, WJHZ13, XMY^{+}15]. Exposure [FZY12]. Expression [DQZ18, DDM^{+}10, HH12, HSTP15, JL15, LNN19, PCSK16, SLC^{+}15, TMK^{+}12, XWK11]. Expression/Usage [SLC^{+}15]. Extended [CHC^{+}12, Lu16]. Extending [NTC13, Raf17]. Extensions [GDCL11]. External [CCMC16a, MRB12]. Extraction [DKK16, MSW^{+}18, MM15]. Extremal [HW19, NN17]. Extreme [Dup12, EEdH19, RPS19, WL13]. Extremes [DRS18, dCD14]. Extremiles [DGS19]. Extrinsic [LSZD17].

Facilitate [JLTY14]. Factor [AB17b, BD12, BT11b, CLGS18, FK18, FKSZ19, HCM13, KWX^{+}19, KHG18, LCFW18, MDCL13, RG17b, TT10, ZSY16]. Factor-Adjusted [FKSZ19]. Factorial [EI19, ZC11, ZX14]. Factorizations [YD16, ZBHD15]. Factors [FKH11, LDGX15, Ros11, Tra18, ZSL^{+}11, ZG18]. Factory [HB14]. Failure [CW17, CKY15, LER^{+}12, LHW^{+}13, NQS14, PHP^{+}12, SS17, XCH^{+}17, ZHS17, Zub12]. Faint [YP17]. False [BCDS18, FHG12a, FKSZ19, GR11, HZZ10, SKS^{+}15, TKPS18]. Familial [GHP13, KSGB10]. Families [DFRS14, HJCPV11, Sch11]. Family [BH18, BDF16, HZH15, KJ10, XMY^{+}15, ZPS16]. FANS [FFJT16]. Farms [LDGX15]. FarmTest [FKSZ19]. Fast [DBNZ16, DVF11, FCSZ16, LFU11, Plu14, RG17b, Wan17a, ZL11, ZEMD18]. FDA [HZT11]. FDP [Sch12]. FDR [KN18, LB17a, SCG13, Sch12]. Feature [CLZ15, FFJT16, LS18a, LZZ12, LLW14, LC14, WT10b, XC14, XMY^{+}15, ZCJ^{+}17, ZLLZ11, ZSP13]. Features [KLH11, KSKD12, Li12a, RTT18, ZCJ^{+}17]. Feedback [CGL10]. Feeding [SFM18]. Fellegi [SF13]. Fertility [SZGM14]. Fever [LS19a]. Fiber [ZDD19]. Fiducial [HILL16, LHL15]. Field [ASS^{+}15, BMW19, MSM^{+}15, SSS18]. Fields [AGS12, COS^{+}10, FPLM18, FJM19, FSLR19, GKS10, MMC^{+}12a, XG17]. Filament [GPPVW12]. File [GAZ13]. Filling [ST17, ZX14]. Filter [GJL17, RLM13, SSL^{+}10]. Filtering [CCR15, HL14b]. Filters [YYC17]. Finance [ZDP11]. Financial [ASFX10, AB17b]. Find [CLQY16]. Finding [CZK17a, GY17, LTJM15, YY11, ZC17a]. Findings [BH13]. Fine [BOSB16]. Fingerprints [KSGB10, WGDG14b]. Finite [AT11, FTZ15, FMP18, LC10, LD17, MDR18, PPLS18]. Firm [CGL10]. Firms [HJCPV11]. Fisher [MM19, SW19]. Fisherian [Wu15]. Fit [ASB11, CS10, CRVF10, FG12, LMR14]. Fitting [ZH18]. Fixed [AIZ14, CXT14, GLL13, LQS17, MW17, NL12, Sar12, Sob12]. Fixed- [MW17]. Fjord [CCS14]. FL [Pea19, Tri19]. Fleeting [SY17]. Flexible [LD15, XAM^{+}14, Yan19]. Flow [ASS^{+}15, BG16, CIB^{+}18, MM15]. Flows [PPY14]. Flu [DLP12]. Fluids [LPH^{+}17]. fMRI [LS18c, LSLR12, SL14, WGE^{+}18, YPOR18]. Focus [PV13]. Follow [BH13]. Follow-Up [BH13]. Following [HCKS18]. Food [KPGJ12]. Forecast [WAH19]. Forecasting [BRGS10, CGBY13, DHS11, HG10, JT12, KRG11, Lin13, LS17b, PSY19, SZGM14, SGR10, SJ14]. Forecasts [Gne11, SPU17]. Forest [FBM11]. Forests [WA18]. Forms [LD17, ZNSR12]. Formula [Efr11b, LS15]. Formulas [GK18]. Forward [CHZ16]. Foundation [BGB12]. Foundations [CZ15, Yu19]. Four [Dup12]. Fractile [SC12]. Fraction [LS19a, WY13]. Fractional [ZX14]. Fragmentary [WRCG13]. Frailty [GHP13]. Frame [RW10]. Framework [BG11, CCL^{+}11a, CCL^{+}11b, DT15, FK14a, HLSZ14, HDL^{+}16, KCP^{+}11, Laz11, LNM11, ML13b, MCIS19, NSB17, PLGM11, RTT18, RSK17, SF13, She11, SLS18, SEdS19, SDWM18, TYY^{+}19, WZ10, WT10b, Wu11, XSS11, YDZ16, ZH18, ZBM11]. Fraumeni [SYS^{+}19]. Free [ACCTW18, BC14, BCVD18, CLZ15, LRW13, LGR^{+}18, ML13b, ML15, MBCM13, RKFL19, WZL10, YDZ16, ZLLZ11]. Free-Tailed [MBCM13]. Freedom [KS11, LMR14]. Frequencies [MM15, ZFW17]. Frequency [ASFX10, ASX19, DGHM19, FLY12, FK18, JKL11, LK19, SO19, SJ14, TWYZ11, WM14, ZFW17]. Frequency-Specific [SO19]. Frequentist [LZWZ11, Mar15, RS10, WB19c]. FreSpeD [SO19]. Front [GK18]. Front-Door [GK18]. Frontier [Tsi17]. FSEM [LSS^{+}19b]. Fuel [AIC14].

Fully [BJQ17, LNN19]. Function [DEV19, HJ17, HGZ18, LQ17, Pap12, RLM13, SFK12, SDWM18, YPQ^{+}16, ZTS11, ZZS19]. Function-on-Function [LQ17, SDWM18]. Functional [ANH15, BGH^{+}18, BCT18, BSB15, CY12, CCMW11, CM12, CWPC13, CL15b, CGO19, CHSV14, DH13, GHK10, GCNC14, GZZ19, HAME15, HG18, HH10, HJCPV11, JAW16, KJNW11, KOL^{+}12, KMR17, LS18a, LST^{+}11, LMR^{+}19, Lei14, LWC10, LWC13, LG14, LHH17a, LS18c, LHS18, Lin12, LS17b, LSS^{+}19b, MS18a, MBCM13, NN17, PSSS17, PJY16, QGJ19, RQJ15, RN16, SD14, ŞM10, TP16a, WS14, WGE^{+}18, WLZ19, XKBS17, XLN18, YY16, YMSC15, ZBZ^{+}16, ZYH14, ZHM^{+}10, ZLL18, ZBM11]. Functionals [KY11a]. Functions [AGS12, BBS17, BGMP12, CLM14, CYC10, GKS10, GZZ19, LB18, LM17, LT14, MZN^{+}16, Mar15, NQS14, PKM10, PH18, Plu14, PBG16, QLL10, UL17, VW15, WfTQ12, ZA16, ZLL18]. Fused [CZ14b]. Fusion [DW11, MH17, NCB12, Kim19]. Fusion-Refinement [DW11]. Future [ZTC^{+}13].

Gail [Pea19]. Gamma [PSW13]. Gap [MS18b]. Gaussian [BG16, CRZZ16, CH17, DBFG16, DG18, DLR11, FJM19, FZ13, FSLR19, HL18, LSQ15, LX19, MDCL13, PKM10, PPLS18, PSV15, RMR19, RD12, WS14, WG14, WDP^{+}14, ZA16, ZD13]. Gene [DDM^{+}10, HRC12, HCCZ16, HZL^{+}17, LNN19, LCZ12, LN10, LLLW11, PCSK16, WY19, WLXL14, XWK11, ZC17a]. Gene-Environment [HRC12, HZL^{+}17]. Gene-Gene [HRC12]. General [BT16, CMS14, DLR11, HS13, KLSY13, LD17, LL19a, RTT18, RGL16, SM12a, Wol11, WPS17b, YBT13, YDZ16]. Generalized [ACB^{+}14, ASB11, AGT14, BML17, BY11, BS13, CCJ13a, CYC10, CRVF10, HS13, HILL16, HNW^{+}17, Hua14, HLG14, JLZ14, JHZ16, KKL15, LHL15, LNS10, LWC10, LC18, LSY13, LYH13, LS17b, MW10a, MWX14, OKEK17, PY16, SF13, TLG16, WS14, WZ17b, WLSA17, XQZ10, You19, ZLW10, ZLT10, ZYZL17, ZS15b, ZEMD18, ZSZ12, ZJZ12, ZKLI14]. Generalizing [SPHL16]. Generating [LCM10]. Generation [BA16]. Generic [PWXZ19, ZW14]. Genes [LLZ18, ZCC^{+}17]. Genetic [BFWE10, BML17, GWCL19, JLZ14, LCG^{+}15, LLQ17, RN16, TZL17, ZKLI14]. Genetical [LFL15]. Genetically [ZCC^{+}17]. Genetics [MZW16, SGVC13]. Genome [BA16, ZL11]. Genome-Wide [BA16, ZL11]. Genomewide [LZ11b]. Genomic [CCZ16, FLJL17, HCKS18, JZT19, ZCJ^{+}17]. Genomics [HJRS10, LZ10b, LFL15, WZL19]. Geometric [yCtW13, LL19b, LMWA17, XKBS17]. Geometry [GPPVW12, PKM10, Yu19]. Geostatistical [DBFG16, KRG11, LCS^{+}13]. Geostatistics [DG16a]. Geriatric [HTGT13]. Get [AR15b]. Getting [WSG16]. Gibbs [NSH19a]. Gigadata [WLSA17]. Gini [FMC11]. Glaucoma [BMW19, LMR^{+}19]. Global [CGSW10, FPSE15b, Lei14, LCL14, MMM11a, MMM11b, MAS14, SL11, TGHS18]. Global-Local [TGHS18]. GLS [BT11b]. Gold [LHW^{+}13]. Goodness [CS10, CRVF10, FG12, LMR14].

Goodness-of-Fit [CS10, LMR14]. Google [DLP12]. Gradient [FL14, Sar12]. Gradient-Based [FL14]. Graph [Bie19, CY16, CF17, DPS19, LN10, ZSP13]. Graph-Based [CF17]. Graph-Guided [Bie19]. Graphical [BT16, CRZZ16, DG18, DLR11, FJM19, GNL19, LT12, LCZ12, LS18c, LSQ15, MML^{+}13, NSB17, NSB19, PSV15, QGJ19, SHF17, YL16]. Graphs [HCCZ16, STFP12, WY19, ZSP14]. Great [Mar19, PM19, PBG16, TPAC19a, TPAC19b]. Grid [Plu14]. Gross [FZY12]. Gross-Exposure [FZY12]. Group [BQWS17, BGSB19, CYZ14, DM11, DZ15, LER^{+}12, SDT10, SWJ18, VHJB13, WMBZ18]. Group-Linear [WMBZ18]. Group-Randomized [VHJB13]. Group-Specific [BQWS17]. Grouped [WAH19]. Grouping [SH10, ZSP13]. Groups [BGSB19, HZZ10, PSR16]. Groupwise [GLLL15, LLZ10]. Groves [MS18a]. Guided [Bie19]. Guiding [MMC^{+}12a]. Gusts [You19]. Gyration [ABvdW19].

H [Pea19, Tri19]. H1N1 [FBCA14]. Halfspace [CHSV14]. Hall [Pea19, Tri19]. Hall/CRC [Pea19, Tri19]. Hamming [TY17]. Handbook [Pea19]. Happens [Fuk15]. Hard [LZW11]. Hardware [SMQ^{+}13]. Harville [Tri19]. Hazard [CGL10, HGZ18, LN14, XAM^{+}14]. Hazards [CCZF15, CGSW10, LHS18, LL13, PHSS15, SW14, YY16]. Health [BAWM18, HD10, KPGJ12, LZPH14, PKSR15, Pre10]. Healthcare [HZS18]. Heat [HSM^{+}15]. Heat-Related [HSM^{+}15]. Heavy [HL14b, WLH12, Zho19, ZL15b]. Heavy-Tailed [WLH12, ZL15b]. Heckman [MG12]. Height [COS^{+}10]. Help [CB18b]. Heredity [CLZ10]. Heritability [RN16]. Heterogeneity [HZS18, JS11, WWL12, XMY^{+}15]. Heterogeneity-Feature [XMY^{+}15]. Heterogeneous [EUW17, FBM11, LLX15, RKFL19, WA18]. Heteroscedastic [BGH^{+}18, SM12b, WMBZ18, XKB12, ZL15b]. Heteroscedasticity [CJN18, Zho13]. Heterosis [LNN19]. Hidden [BL16, CSSK16, CWH18, DGM12, FTM12, FFvSK17, HWZ^{+}13, IZR^{+}13, LN10, THY16]. Hierarchic [Yu19]. Hierarchical [BFWE10, BT11a, CSSK16, CT10, CMER11, DBFG16, FBM11, GS17a, JPBM13, KCZ^{+}13, LDSH16, LNA10a, Li12a, LHC19, LCCG14, MB16, NSH^{+}19b, PCSK16, Smi10, SEdS19, WSWT10, WJYJ13, WDP^{+}14, WAH19, WH17, XKB12, XLN18, YPQ^{+}16, ZCJ^{+}17, ZHM^{+}10]. Hierarchy [SWJ18]. High [ASFX10, AL14, AB17b, ASX19, BCK19, BTI19, BH13, BR12a, CLX13, Cer10, bCH10, CHAP16, CZZ10, CZW^{+}11, CCMW11, CSW17, CLLL18, CW19, CHZ16, DS12, DG18, DGHM19, FFS11, FLY12, FMD14, FFJT16, FK18, FTM12, FZW16, FCDA15, GS10, GKM19, GCBL15, GHJZ16, GBZ17, GWCL19, HL14a, HL18, HFZ18, HSZ15, HZ13, IKG^{+}10, JKL11, JR12, LLSS18, LZ10b, Li12a, LSY13, LSQ15, LL13, LFL15, Lin18, LLLW11, LC14, MB16, PGY14, PMZ19, QC15, RJ10, RGH13a, RSK17, RT17, SZ14, SL15, SMQ^{+}13, SJ14, SLC^{+}15, SZLI15, TWYZ11, VRS19, WLH12, WWL12, WM14, WPL15, YD16, YP17, YYY18, You19, Zha13b, ZC17b, ZH19, Zha11, ZC11, ZSPL17, ZB18, ZZS19, ZSCM18]. High-Breakdown [Cer10]. High-Dimensional [AL14, AB17b, BCK19, BR12a, CLX13, bCH10, CHAP16, CZZ10, CZW^{+}11, CCMW11, CSW17, CW19, DS12, DG18, FFJT16, FK18, FZW16, GS10, GKM19, GHJZ16, GBZ17, GWCL19, HL14a, HL18, HFZ18, HSZ15, HZ13, IKG^{+}10, JR12, LLSS18, LZ10b, Li12a, LSY13, LSQ15, LL13, LFL15, Lin18, LLLW11, MB16, PGY14, QC15, RGH13a, RSK17, RT17, SZ14, SL15, SZLI15, VRS19, WPL15, YD16, YP17, YYY18, Zha13b, ZC17b, ZH19, ZC11, ZSPL17, ZB18, ZZS19]. High-Frequency [ASFX10, ASX19, DGHM19, FLY12, FK18, JKL11, SJ14, TWYZ11, WM14]. High-Order [ZH19]. High-Resolution [FTM12]. High-Throughput [BTI19, FCDA15]. Higher [BML17, LRT12, SD17a]. Higher-Order [LRT12]. Highly [QLL10]. Hilbert [BCT18, SDWM18]. Hippocampal [AVWW14]. Histone [MML^{+}13]. History [Fuk15]. HIV [DGM12, Gil11, JPBM13, LHW^{+}13, LZNS17, MRB^{+}17, XAM^{+}14]. Homicide [SF13]. Homogeneity [HNW^{+}17, KFW15, LS19b, RK11]. Homologous [LQ11]. Hormone [LCCG14]. Hospital [GRR^{+}17]. Hotelling [CPPW11]. Hour [SJ14]. House [SQ10]. Households [KSGB10]. HPV [KWG15]. Huge [AB17b]. Human [EEdH19, MZW16]. Huntington [WGM12]. Hybrid [CGBY13, GK18, pKP12]. Hydrocyclone [OCAG19]. Hypercube [GS17a, Hun11, Qia12, XCQ15]. Hypergeometric [Wan15]. Hypergraphs [LMWA17]. Hypotheses [FSMS17, LN10, Lu16, MHZ15, SM12b, ZSR18]. Hypothesis [Cha17, HCKS18, LB17a, WY19, Zha11, ZB18]. Hyvärinen [SJDT19].

I/II [GY17, LGY18]. Ice [CHAP16]. Icing [LHH^{+}17b]. ICU [GCNC14, KHK19]. Ideas [Lit13]. Identifiability [DGYZ11, MDG17]. Identification [CP12, FMV16, HMM16, HHY19, JCL10, KYBB19, LG13, LLLW11, MVR12, SH13, SW11, SHM^{+}18, WBG^{+}18, WG11, XCC18, Sob12]. Identified [Sid13]. Identifiers [WCZ^{+}13]. Identify [BJT^{+}10, HDL^{+}16]. Identifying [FKH11, JLZ14, JL15, KPGJ12, SQ10, TPLG10, XS18, Zha13a, ZRY19, ZCJ^{+}17, ZCC^{+}17]. Idiosyncratic [KWX^{+}19]. if [LER^{+}12]. Ignorable [HFQ12]. Ignoring [LRZ17]. II [GY17, JLTY14, LGY18]. III [KSGB10]. Ill [CB18b]. Ill-Posed [CB18b]. Illness [LZ10a]. Illustrations [Pre10]. Image [CCL^{+}11a, CCL^{+}11b, Laz11, LNM11, She11, WZ17b, Wu11]. Images [COS^{+}10, FG13, LMH14]. Imaging [KOL^{+}12, LST^{+}11, PPLS18, SGVC13, YL13]. Imbalance [IKP11]. Imbalanced [CDQ13, CDQ14, JSPD19]. Immune [MMC^{+}12a]. Immunotherapy [LGY18]. Impact [PGW^{+}18, YPQ^{+}16]. Imperfect [KN18, Sid13]. Implications [HTFK10]. Importance [BG16, LTC13, TMP18, YYY18]. Imprinting [LCG^{+}15]. Improper [CH17]. Improve [MDR18]. Improved [Gil11, Tsi17, WLS17, WW15, YPOR18]. Improves [HSTP15]. Improving [AB17a, SS17]. Imputation [BR12b, HTFK10, KCK^{+}15, MVR17, MR17, RGH13a, WF12, Yan19, ZR15].

Imputed [ZK12]. INAR [PDF15]. Including [ZH18]. Inclusion [JZT19]. Income [JS11]. Incomplete [ACBK10, HSM^{+}15, HFQ12, JNR15, KN18, WDSL10, YHD^{+}10, Zub15]. Incorporating [PCC12, YL16, ZCJ^{+}17]. Increases [ZLP^{+}14]. Incremental [Ken19]. Independence [BFV16, FFS11, FMD14, PGY14, PWL16, PWXZ19, Zha19a]. Independent [HM15, LST^{+}11, MT17]. Index [Ano10g, Ano11i, Ano13a, DT15, FWYZ13, FMC11, JKL11, LWC10, LHH17a, LL19a, MS15, MMW^{+}17, NC10, RQJ15, WTM19]. Indexed [DDT17]. India [BIL^{+}11]. Indirect [AB13, SHS10, WTM19]. Individual [DGH^{+}10, ZLP^{+}14]. Individual-Level [DGH^{+}10]. Individualized [LS18a, ZZRK12, ZMHKK17]. Individuals [EUW17]. Induced [SMQ^{+}13]. Inducing [MN16]. Industrial [OCAG19]. Ineligibles [DIvdB13]. Inequality [DJP18]. Inexact [Plu17]. Infants [BSLR10, TNZM14]. Infection [CZW^{+}11, DDM^{+}10, MŞDN12]. Infection-Cardiovascular [MŞDN12]. Infectious [Ken15, NFG^{+}16, TYY^{+}19, YHD^{+}10]. Infer [KSGB10]. Inference [AIZ14, ACBK10, BBS17, BFH12, BDIK12, BC14, BL16, BCK19, BTI19, BKM17, BM10, BCS18, CZ13, CCF18, CCJ10, CJN18, CZ19, CLGS18, CL10, CGL10, CLF12, CGW17, CDS11, CZMHS17, DP11, DFRS14, DS12, DLR11, EG19, FOW11, FTZ15, FMGS16, FSMS17, Gil11, Hag17, HS12, HCCZ16, HH13, HV11, HILL16, HCKS18, Hos13, Ion12, JLY19, KC11, KM17, LHL15, LGR^{+}18, LNS10, LY16, LD17, LLQ17, LDL^{+}19, LJZ14, LH14, LM17, LvdL18, LSLR12, MHC13, ML13b, ML15, Mar15, MP13, MD19, MTC11, MM13, MDR18, MN16, MW17, NEB13, NH11, OR17, PRS10, PH18, PSV15, PSW13, RW10, RTT18, RKFL19, SB18, SLM11, SH15, SL14, TWYZ11, TZL17, TTLT16a, THY16, VRS19, WA18, Wan17a, WfTQ12, WBCD15, WLS17, WKKS12, WLNC14, XSWH17, XLN18, YJFL19, YKC^{+}15, Zaj12, ZC13, ZC17b, Zho11]. Inference [ZST12, ZL15b, ZZS19]. Inferences [CRL15, LRZ17]. Inferential [ML13b, ML15]. Inferring [FSS^{+}16, FSB^{+}19, HMM16, LSS^{+}19a]. Infinite [CLvdL19]. Infinite-Dimensional [CLvdL19]. Infinitesimal [LT10, LRT12]. Inflated [KKL15]. Influence [RPS19, TMP18]. Influenza [FBCA14, YHD^{+}10]. Information [CCMC16a, DDT17, DGH^{+}10, GS10, GDZ11, HQT16, JPCD19, JK16, JHZ16, LNP14, Ma15, MCW19, PCC12, SPU17, SH13, WYS19, YNLC19, ZLT10, ZCJ^{+}17, ZST12]. Information-Based [WYS19]. Informative [BTC10, DZ19, LDS10, MS18b, SSZL12, ZQ12]. Inhibition [LSLR12]. Inhomogeneous [RD12]. Initial [LER^{+}12]. Initiative [LZPH14, Pre10]. Inland [LDGX15]. Innovation [Pan11]. Instability [Sch11]. Institutions [ZNSR12]. Instrument [BSLR10]. Instrumental [Bla17, CW12, KZCS16, LHD19, LFL15, SW14, SHM^{+}18, Tan10, WFDS19]. Instruments [CMW^{+}19, ESR18, KZCS16, SHM^{+}18, WFDS19]. Insulin [WFZ18]. Insurance [FMC11, SY18]. Integer [Zub12]. Integral [OPG16]. Integrated [CDS11, HC15b, RWF^{+}13, ZA16]. Integration [CCZ16].

Integrative [HZZ^{+}17, SGVC13]. Intensities [HH12]. Intensity [FL10, Tad10]. Intensive [CMPS17]. Interaction [Bla17, CD18, EI19, HZ14, HZL^{+}17, LL19a, RJ10, ZSS19]. Interactions [Dav12, HRC12, LWC10, MM13, ST15, TAGT14, XCC18, ZMB^{+}16]. Interactive [LQS17, LLS17]. Interest [ML15]. Interference [AEI18, BF18, LH14, LSLR12]. Interim [IS12]. Intermediate [SLM11]. International [Dav13]. Interpretable [FMGS16, VRS19, ZLDT18]. Interpretation [yCtW13]. Interval [GZCL19, Gee14, LJGZH19, WW15, ZHS17]. Interval-Censored [GZCL19, LJGZH19, ZHS17]. Intervals [BBS17, Büh11, Düm11, GK11, HZ13, JZ15, LM11a, LM11b, PCB11, Pol11, Sam11, SM12a, Wan15, WN11, WFB13, YC11]. Intervention [FZ13, RSK17, VHJB13]. Interventions [Ken19, PHP^{+}12, YLRH19, ZSS19]. Intractability [Cha17]. Intractability-Countering [Cha17]. Intractable [LJSL16, PH18]. Intraday [ATT19, KLL17]. Intraparty [SQ10]. Intrinsic [SZI^{+}12, WLNC14]. Invalid [KZCS16, WFDS19]. Invariant [CLGS18, HL14a, PBP19]. Invasive [BSMR12]. Inverse [BF15, BDF16, IR15, LZL19, NZK11, ST18, Tad13a, TT12, YLRH19]. Investigate [EUW17]. Involvement [Sch15]. Involving [CDH12, LST^{+}11]. Ion [PJY16]. IPO [CGL10]. Irregularly [LT14]. ISBN [Pea19, Tri19]. IsoDOT [SLC^{+}15]. Isoform [SLC^{+}15]. Item [HS13, Ima11, PS10]. Iterated [GJL17]. Iterative [Sar12].

J [Pea19]. Jackknife [CCJ13a]. Jin [Dui16]. Jin-Ting [Dui16]. Job [GK18]. Joint [BAA^{+}15, CZ19, GDM17, HFSZ19, HH12, Hua14, HLG14, HMW17, LTJM15, LY16, LZS^{+}16, PGW^{+}18, RHCT14, SSZL12, WTM19, XCH^{+}17, YT17]. Jolly [WH17]. Jolly-Seber [WH17]. Journal [CDQ14]. Jump [BH18, DGHM19, JKL11, LTT17, TT12, ZFK14]. Jump-Diffusions [TT12].

Kalman [SSL^{+}10]. Kendall [ZLW10, JLZ14, ZJZ12]. Kernel [BCT18, CD11, CWZM18, FL14, Gee14, JT12, MMNS11, MYLC15, SDWM18, WRL10, WW15, WSW17]. Kernels [LMR14, PK12]. Key [Pan11]. Kidney [EUW17, Zub12]. Killed [LS19a]. Kinematic [KCZ^{+}13]. Kink [GJ18a]. Know [KBKS18, MSZ10]. Known [ZCJ^{+}17].

L [CDQ14]. Labor [FL10]. Laboratory [NFG^{+}16, TYY^{+}19, YHD^{+}10]. LADE [ZL15b]. LADE-Based [ZL15b]. Landmark [HSZ15, PCC12, PTC14, SKBM17]. Landmark-Based [HSZ15]. Landmark-Constrained [SKBM17]. Landmarks [SCK19]. Landscape [HH13]. Landscapes [FBM11]. Language [JML^{+}14]. Laplace [BPPD15, TT12]. Laplacians [DPS19]. Large [AS15, BFWE10, BCDS18, BGMP12, BKD^{+}17, BM10, Cai10, CL16a, CLL19, DBFG16, DVF11, Efr10b, Efr10c, FBM11, FSB^{+}19, GdMb13, Guh10, HLMH18, Hel10, HCKS18, HR11, KD13, LS19b, LCS^{+}13, LFU11, LH14, MSW^{+}18, PGY14, PPLS18, PKSR15, Sch10, TARS17, TWYZ11, TAGT14, Wan17a, WZM18, Wes10, XMZ12, ZBZ^{+}16, ZK17, LZW11]. Large-Margin [LZW11]. Large-Scale [BCDS18, BM10, Cai10, CL16a, Efr10b, Efr10c, Hel10, HCKS18, HR11, Sch10, TARS17, Wes10, ZK17]. Larvae [CCS14]. Lasso [CL11c, GHJZ16, HCCZ16, HWF15, LZL19, LC14, WFDS19, Zho14, CYZ14, JHZ16, RG18]. Lasso-Type [Zho14]. Late [LER^{+}12]. Latent [AB13, ACB^{+}14, CXCR13, CRVF10, EUW17, FJM19, KA18, MZ15, PHSS15, PS10, PCSK16, PSW13, Raf17, RDG10, RLP^{+}18, SC15, WK13, XS18, YOD11, ZEMD18, ZSY16]. Latin [GS17a, Qia12, XCQ15, Hun11]. Lattice [CY16, DLR11, QA10]. Layer [BJ11]. LD50 [PCB11]. Lead [Mor19]. Leadership [FSS^{+}16]. Learning [BGH^{+}18, BGC19, CCSN11, CZK17a, FTZ15, FO17, FZ13, LZZ12, LLS17, MMC^{+}12a, MJPW13b, MYT18, PBD13, TGP11, WFZ18, WK13, YYY18, ZZRK12, ZZLK15, ZMHKK17, ZZK15, ZZS19]. Least [AGT14, CGLY10, DGS19, KS11, PD14, SLW19]. Least-Square [AGT14]. Least-Squared [PD14]. Length [ACLZ14, yHQ12, QNLS11]. Length-Biased [yHQ12, QNLS11]. Lesion [ZLP^{+}14]. Letter [Ano11j, FHL15]. Letters [Ano10h, Ano10i]. Level [CCMC16a, CGW17, DHM11, DGH^{+}10, HCKS18, MW19, RSK17, SVTG17, You19]. Levels [SLW19, WZ17a]. Leverage [ASFL^{+}17, KX17, WM14]. Lévy [BIL^{+}11]. Lexis [CWH^{+}15, Fu16]. Li [SYS^{+}19]. Life [EEdH19, GAZ13, LDSH16]. Lifetime [CAJ14, HLSY16]. LIFT [JZ15]. Like [YZB^{+}16]. Likelihood [AO17, BC14, BL16, BGMP12, BS13, CW14, CCMC16a, CGSW10, CRL15, CLA10, CH17, DKK16, FHS13, GS10, HLSY16, HLSZ14, HZT11, yHQ12, HMW18b, KLN13, LNS10, LCG^{+}15, LRZ17, LJZ14, LCLZ18, NQS14, PRS10, PDF15, QNLS11, QP13, RW10, RMR19, Sar12, SPZ12, TLG16, WX15, WC18, XWG19, ZST12, ZZWJ12, ZHS17]. Likelihood-Based [HLSZ14, PRS10, SPZ12, ZZWJ12]. Likelihood-Free [BC14]. Likelihoods [HB14, SWW14, WSW17, ZS15b]. Limit [LD17]. Limited [GRH10]. Limiting [Han17b]. Limits [BBS17, EEdH19]. Linear [AY17, BY11, CL11b, CY12, CR10, CJN18, yCtW13, CS10, CRVF10, DNFZ10, DFRS14, FF13, FWYZ13, FS17, GHK10, GC11, GHJZ16, GWCL19, HCM13, HC15a, HD10, Hua14, JHZ16, KFE12, KMR17, KD14, Lei14, LNS10, LWC10, LC18, LSY13, LZ10c, LQ17, MHC13, PWL16, RN16, RMR19, RLY14, RWKS14, SZLI15, TPLG10, WYS19, WMBZ18, WLZ19, WQxYW19, YYY18, ZCL11, ZWMC17, ZYZL17, ZC17b, ZS15b, ZDP10, ZG18, ZB18, ZWZZ19, Tri19]. Linguistic [Mar19, PM19, TPAC19a, TPAC19b]. Link [God17, VT17]. Link-Tracing [VT17]. Linkage [BT11a, HRZ17, SF13, Sad17, SHF17]. Linked [EUW17]. Linking [GAZ13]. Links [KJ10, Zha13a]. Lists [HS12]. Little [Spi19, YZ19]. Liver [GSDR19]. Load [CGBY13]. Loblolly [FRG^{+}17, LHTB15]. Local [CTM10, CCS14, CWC^{+}14, yCtW13, CMPS17, LMZJ13, LSZD17, LJZ14, Ma15, MMM11a, MMM11b, MR17, PV13, Pap10, SL11, TGHS18].

Localized [CHP10, CL15b, ZA16]. Localizing [HCOW17, TP16a]. Locally [BB10, DPV11, DJW18a, HHY19, KRG11, XBZ^{+}14, ZD13]. Location [DYZ12, FZW16, KKL15, SWCB16]. Loci [JL15]. Log [RD12, VS15]. Log-Gaussian [RD12]. Log-Normal [VS15]. Logistic [JHH14, LL19a, PSW13, SH15, WZM18]. Logistic-Normal [SH15]. Logit [SWCB16]. Long [BBS17, CLM18, GLP19, PCC12, PAHJ11]. Long-Range [BBS17]. Long-Run [CLM18]. Long-Term [PCC12]. Longitudinal [BL16, CCZF15, CYC10, DBNZ16, DM18, DZ19, HBHC12, HZL^{+}17, HLG14, HMW18b, KWG15, KNK^{+}18, LVC14, LZP10, LZ11a, LWN^{+}18, LD15, LCCG14, LZS^{+}16, QJWG16, RHCT14, ŞM10, SMAC10, SWCB16, SSZL12, TLG16, WDSL10, WT13, XY12, ZPlW13, ZQ12, ZLL18, ZTS11]. Loop [MZN^{+}16]. Loss [DEV19, LB18, WJHZ13]. Low [DG16a, PT17, RD12, TWYZ11, ZKLI14]. Low-Frequency [TWYZ11]. Low-Rank [RD12]. Low-Resource [DG16a]. Lung [CWH^{+}15, MTY^{+}17]. Lyme [XXL^{+}19].

M [AIP19]. Machine [BGC19, CWC^{+}14, MYLC15, MYT18]. Machines [LZW11]. Macro [ZDP11]. Macro-Finance [ZDP11]. MAD [XMY^{+}15]. Magnetic [KOL^{+}12, LST^{+}11]. Magnetoencephalography [YZBE18]. mail [FSS^{+}16]. Main [MW19]. Maintenance [BRGS10]. Making [CZ13, Gne11, Mor15, ZMB^{+}16]. Malaria [BIL^{+}11, LS19a, SFM18]. Mallows [ZWZZ19]. Mallows-Type [ZWZZ19]. Malnutrition [FKH11]. Malware [BH18]. Mammal [LBS13]. Management [HMQA10, RGH13a]. Mandarin [HAME15]. Manifold [LSZD17, YZB^{+}16]. Manifold-Like [YZB^{+}16]. Manifold-Valued [LSZD17]. Manifolds [yCtW13]. Manolopoulou [BFH12]. MANOVA [VW11]. Manufacturing [HJCPV11]. Many [CJN18, HBHC12, MSZ10, ZSR18, ZSS19]. Mapping [DG16a, HSTP15]. Maps [FMP18, XBZ^{+}14]. Margin [LZW11]. Marginal [CWPC13, DKK16, Hos13, IR15, LJGZH19, ML15, Tan10, WTM19, XAM^{+}14, XQZ10]. Margins [SK12]. Marijuana [TEKM12]. Marine [LBS13]. Marked [HSM^{+}15]. Marker [CW17, SW17c, SHW17, WCZ^{+}13]. Marker-Assisted [WCZ^{+}13]. Markers [ZKLI14]. Market [FL10]. Markov [BL16, BCVD18, CSSK16, DGM12, FTM12, FFvSK17, HWZ^{+}13, IZR^{+}13, LBS13, LN10, MS18a, SD17a, SCM^{+}18, THY16]. Markov-Modulated [LBS13]. Marriage [Sob12]. Martingale [AI12, LS18b, SZ14]. Masking [Sna19]. Mass [NC10, VW15]. Massive [FLJL17, HZS18, Kat17, RKFL19, SVS16, SLS18]. Match [HJRS10]. Matched [FSMS17, FS17, Ros14, Zha19b]. Matches [KN18]. Matching [AI12, CMS17, IKP11, LM17, OR17, PKSR15, RK11, SYY15, Zub12, ZK17]. Matchings [Sad17]. Materials [CXCR13]. Matérn [AGS12, GKS10]. Mathematistry [Lit13]. Matrices [CZZ10, GBZ17, VH15, XMZ12, YP17, ZPS16]. Matrix [AGT14, BBX16, Bie19, CL11a, CLL11, CLX13, CCZ16, CY17, CLXY15, DJP18, FLY12, FK18, FCDA15, HFSZ19, HBI^{+}19, LS18b, LT12, MCW19, PD14, QC15, TWYZ11, Tri19, WQxYW19, XCC18, YKC^{+}15, YNLC19]. Matrix-Based [WQxYW19]. Matrix-Variate [HFSZ19, HBI^{+}19]. Matter [YZB^{+}16]. Matters [Sch15]. Maturity [DK18]. Max [PRS10]. Max-Stable [PRS10]. Maxillary [HGK12]. Maximization [QP13]. Maximum [ACF17, CCMC16a, CH17, LCLZ18, PLLG18, QNLS11, QP13, RMR19, ZHS17]. MaxT [Lu16]. MCMC [BH18, CMPS17, HB14, JSPD19, KM17, QKVT19]. Mean [ABvdW19, GSDR19, HFSZ19, Hua14, KY11a, LZP10, WPL15, WMBZ18, ZYH14, ZLL18]. Means [GCBL15, HDL^{+}16, MM13, ZSCM18]. Measurability [HH12]. Measure [DPV11, LSQ15, SFM18]. Measured [HBI^{+}19]. Measurement [BGK^{+}18, BR12b, CDH11, CMM^{+}13, GLS11, HBZ15, KBKS18, LS19a, MR11, MŞDN12, NSK16, SH13, SL14, SW14, SWW14, SW17c, WSW17, WMY15, YMSC15]. Measurements [AB13, FCDA15, PLGM11, YY16]. Measures [JZ15, LW19a, LCLZ18, RB18, ZSZ12]. Measuring [CZ19]. Mechanism [LRZ17]. Mechanisms [FMV16]. Mechanistic [BFH12, WDP^{+}14]. Media [PGW^{+}18]. Medial [SZI^{+}12]. Mediation [Lin12, VHJB13]. Medical [GAZ13, HZS18, LWN^{+}18, RS10]. Medicare [GRR^{+}17]. Medicine [GY17, LZNS17, MZC16]. MEG [ZS15a]. Melanogaster [ZC17a]. Melanoma [MZT^{+}17]. Membership [MHC11, MVR12]. Membrane [SFM18]. Membrane-Feeding [SFM18]. Memory [KPC^{+}17]. Men [Sob12]. Mendelian [KZCS16]. Mental [LZ10a]. Message [Wan17a]. Meta [BAN^{+}12, CXT14, FPSE15b, HL14a, HDL^{+}16, KJNW11, LHC19, LLgX14, LLX15, RDHL19, XSS11, YKC^{+}15]. Meta-Analysis [BAN^{+}12, CXT14, FPSE15b, LHC19, LLgX14, LLX15, RDHL19, XSS11]. Meta-Analytic [HDL^{+}16]. Meta-Elliptical [HL14a]. Meta-Regression [YKC^{+}15]. Metabolites [ADR^{+}12]. Method [AT11, BMW19, BCVD18, GLLL15, HZT11, JHZ16, KLN13, KOLL12, LQ11, LCS^{+}13, LX19, MW19, MTC11, OP13, ST17, TAGT14, WBCD15, Wol11, YY11, Vat19]. Methodology [JR14, JML^{+}14]. Methods [ADH10, Bla17, BR12b, CH19, CZ15, CH17, FL13, GDM17, GLS11, GDLMVF19, IKP11, KPBSK10, LFL15, OCAG19, PL19, Pea19, Pre10, RS10, RSK17, Spi19, SHM^{+}18, SFM18, YY16, YZ19, YRR19, YMSC15, ZZLK15, ZEL19a, ZEL19b, ZD14]. Metrics [CZ19, DPS19]. Mexican [MBCM13]. Mice [MJPW13b]. Micro [RSK17]. Micro-Level [RSK17]. Microarray [SW11]. Microbial [RBF^{+}17, WZ17a]. Microdata [DR10, KCK^{+}15]. Mid [BR12b, RDHL19]. Mid- [RDHL19]. Mid-Study [BR12b]. Migration [RWF^{+}13]. Million [FCSZ16]. Minicircles [PKM10]. Minimax [ACPLRC19, BT11a, CY12, CY16, DJW18a]. Minimization [CLL11, WAH19]. Minimum [ABvdW19, ACLZ14, QDC19]. Mining [FLJL17, JZ15]. Minkowski [DRC^{+}12]. Misclassification [GZJLM12, YMSC15]. Misconduct [HMM16]. Mises [SW19]. Mismeasured [LJGZH19]. MISR [WJYJ13]. Misreported [KPGJ12]. Missing [BR12b, CL10, CYC10, CWZ15, DYZ12, DW11, Efr11a, FOW11, FMPR12, GA12, Han14, KY11a, LDS10, LD15, LRZ17, MRB^{+}17, MRB12, MDG17, MR17, NTC13, QLL10, SS17, ST18, WRL10, WZL19, Yan19, ZS15b, Har19]. Missing-Data [LRZ17]. Missingness [HFQ12, LD15, RBB13]. Mission [Cre18a]. Misspecification [RPS19, RW13, ZST12]. Misspecified [AIZ14, FHS13]. Mitchell [Pea19]. Mixed [AY17, CW14, CR10, bCH10, Dai19, FOvS10, GA12, GMS14, HAME15, HMW17, HMW18b, IZR^{+}13, JRFN18, KOL^{+}12, LVC14, LMR^{+}19, LNS10, LLLW11, MVR12, MBCM13, MM15, MDCL13, SCM^{+}18, SW17a, SM12a, SWCB16, TPLG10, WfTQ12, YMSC15, ZYZL17, ZHM^{+}10, ZBM11, Zub12]. Mixed-Effects [CR10, HAME15, KOL^{+}12, LLLW11, WfTQ12, ZYZL17]. Mixing [HY12, LS15]. Mixture [CLF12, CMS^{+}18, FTM12, GJS16, Guh10, HNW^{+}17, HY12, HLW13, KS15, LC10, LTC13, LCG^{+}15, LLQ17, Ma13, MDG17, MH18, MR17, QP13, SH15, Tad10, WGM12]. Mixtures [ADR^{+}12, CD11, Dui16, LC18, LN14, VS15, Zha05]. MLE [XC14]. MNIR [Tad13b]. Mobile [BAWM18, CXCR13, YRR19]. Mobility [Zaj12]. Mobilization [Bla17]. Möbius [KJ10]. Mode [CP12]. Model [AL14, AVWW14, ADR^{+}12, ACLZ14, BBW14, BSK^{+}15, CW14, CCZF15, CDBG16, CLGS18, CHAP16, CCMC16a, CHC^{+}12, CYT12, CRL15, CRZZ16, CLQY16, CLLL18, CMS^{+}18, CXCR13, CMER11, CLZ15, DGM12, DBNZ16, DHM11, DT15, DV11, DG16a, DLP12, DVF11, Efr14a, FF13, FXZ17, FK18, FBCA14, FBM11, FOvS10, FJM19, Fu16, GHK10, GS10, GdMb13, GC11, GMS14, GH18, GDCL11, HH13, HFZ18, HRZ17, HH10, HH12, HHY19, HLR15, HQT16, HC17, IZR^{+}13, JWXJ13, JNR15, JRFN18, JR12, KC11, KOL^{+}12, KY11b, KPC^{+}17, KRG11, KW15, KLL17, KMR17, LMZJ13, LNP14, LZ11a, LCZ14, LCG^{+}15, LHH17a, LLQ17, LS18c, LHS18, LW19a, LHC19, LN10, LZWZ11, LYH13, LPH^{+}17, Ma15, MSW^{+}18, MG12, MM15, MYLC15, MML^{+}13, NTC13, NSH19a, NQS14, OPG16, PHSS15, PGW^{+}18, PPLS18, PD16, PDF15, Raf17, RBB13, RPS19, RHCT14, RSA16, SW17a, SW17b, SBG^{+}16]. Model [SJDT19, SH15, SGR10, SW14, SWCB16, SWW14, SS17, SZLI15, WBG^{+}18, WDSL10, WMA^{+}12, WS14, WC18, WGDG14b, WKKS12, WLNC14, WPS17b, WG11, WZL10, XBZ^{+}14, XKB12, XY12, XQZ10, YJFL19, YDZ16, Zha13a, ZWL^{+}15, ZWMC17, ZYZL17, ZST12, ZMW^{+}15, ZMB^{+}16, ZLLZ11, ZBM11, ZFK14, ZWZZ19, ZD14]. Model-Averaged [ZD14]. Model-Averaging [AL14]. Model-Based [DG16a, HH13, SBG^{+}16, WBG^{+}18, WG11, Zha13a]. Model-Free [CLZ15, WZL10, YDZ16, ZLLZ11]. Model-Robust [KW15]. Modeling [BGH^{+}18, BFH12, BAA^{+}15, BG11, BSZ12, Bro17, BB14, CKP19, CR10, CW17, CHP10, CM12, CIB^{+}18, CKY15, CGBY13, CT10, DM18, DK18, DDM^{+}10, Dup12, DDV17a, DDV17b, FPLM18, FLJL17, FO17, FH15, FF17b, FSS^{+}16, FK14a, GZJLM12, GCNC14, GPD14, HAME15, HCM13, Han17b, Han11, HZS18, Hos13, HLG14, HW19, JR14, JS11, JPBM13, JML^{+}14, JHH14, KNK^{+}18, KD13, KSKD12, LBS13, LNA10a, LLT15, LJGZH19, LSY13, LB17b, LCCG14, LLLW11, Ma13, MZN^{+}16, MMC^{+}12a, Mar19, MSB^{+}11, MJG18, PPD^{+}14, PM19, PCSK16, RSI14, RWF^{+}13, RRW17, RLM13, SD17a, SPU17, Smi10, SMAC10, SGVC13, SRR^{+}17, SJ14, TGHS18, TYY^{+}19, TPAC19a, TPAC19b, TEKM12, TMK^{+}12, Tsi17, WSWT10, WG14, WLSA17, WLXL14, XWK11, XLN18, YHD^{+}10, ZA16, ZTS11]. Models [AIZ14, AY17, AB17a, ASB11, BJT^{+}10, BSMR12, BFWE10, BL16, BCK19, BY11, BIL^{+}11, BD12, BA16, BM10, BGK^{+}18, BT11b, CLvdL19, CJN18, CL10, CGSW10, CGL10, CLXY15, CSSK16, CL16b, CFL18, CHZ16, CSS18, CS10, CRVF10, CMPS17, CZ15, CD18, CZMHS17, Cul11, Dai19, DRC^{+}12, DBFG16, DJP18, DT12, DG18, DTYG12, DLR11, EUW17, FFS11, FMD14, FRG^{+}17, FTM12, FWYZ13, FFvSK17, FHS13, FF17a, GNL19, GDM17, GBDL10, GKM19, GHP13, GDLMVF19, Guh10, GWCL19, HS13, Hag17, HC15a, HN18, HBHC12, HLMH18, HP10a, HNW^{+}17, HW10, HJ17, HD10, HY12, HLW13, Hua14, HMW17, HMW18b, HWZ^{+}13, IR15, Ion12, JL15, JHZ16, KS15, KCZ^{+}13, KLSY13, KA18, KKL15, KPBSK10, KHG18, KD14, LVC14, LSS11, LNP14, LDSH16, LMR^{+}19, Lei14, LNS10, LT12, LWC10, LCZ12, LCF14, LQS17, LCFW18, LC18, LL19a, LSY13]. Models [LSQ15, LLW14, LJZ14, LCCG14, LZ18, LZ10c, LSS^{+}19b, MR11, MS15, MS18c, MHC13, MVR12, MW10a, ML13b, ML15, MBCM13, MZ15, MWX14, MDG17, MH18, MŞDN12, MM13, MDCL13, MR17, NSB17, NSH^{+}19b, OKEK17, OP13, PSV15, PHP^{+}12, PJY16, Plu17, PSW13, PT17, QGJ19, QP13, RQJ15, RN16, RDLC10, RHCT14, RD12, RWKS14, RW13, SWSK19, SCM^{+}18, SW19, SFK12, SH13, ŞM10, SC15, SM12a, SZI^{+}12, SM16, SK12, Sob12, SLR^{+}15, SM12b, Tad10, Tan10, TLG16, TPLG10, TMP18, THY16, Tri19, TT10, Wan17a, WfTQ12, WX15, WZ17b, WKKS12, WZL18, WLZ19, WLSA17, WPS17b, WT13, WH17, XZX10, XTR^{+}14, XAM^{+}14, XCH^{+}17, XS18, XCM^{+}13, YY16, YOD11, YBT13, YTL15, You19, ZDP11, ZL14, ZCL11, ZPlW13, ZBZ^{+}16, ZYZL17, ZC17b, ZDD19, ZS15b, ZEMD18, ZDP10, ZHM^{+}10, ZH18, ZKLI14, ZL15b, ZSY16, ZB18]. Models [dCD14]. Moderate [HS12]. Moderate-Deviation-Based [HS12]. Moderation [BAWM18]. Modes [BBBH10]. Modification [HSR13]. Modifications [IS12, MML^{+}13]. Modified [Har19]. Modulated [LBS13]. Modules [ZC17a]. Molecular [TP16a]. Molecularly [GY17]. Molecule [CSSK16]. Moment [CSS18, ZEMD18]. Moments [OP13]. Monitoring [CIB^{+}18, ZZWJ12]. Monotone [BBS17, Cho17, GZJLM12, LD15]. Monotonic [CLQY16, IKP11]. Monotonicity [HRC12, Hua17]. Monte [BCVD18, Vat19, LCT16, LCM10, LX19, WBCD15, Zho14]. Monte-Carlo [LX19]. Mortality [FMGS16, GRR^{+}17]. Motility [BFH12, Ion12, MMC^{+}12a]. Motor [BGH^{+}18, LSLR12, YPQ^{+}16]. Movement [HJ17]. Moves [SY17]. Movie [PS10]. Moving [HP10a, WX15]. MPSS [DDM^{+}10]. MRI [FTM12, PT17, ZLP^{+}14]. MS [JNR15]. Much [JK16]. Multi [BJ11, BFT18, BMMS17, FTZ15, FWYZ13, Kat17, LCG^{+}15, NSB17, SO19, WH17, Zho11]. Multi-Agent [FTZ15]. Multi-Armed [BFT18]. Multi-Channel [SO19]. Multi-Dimensional [NSB17]. Multi-Domain [Zho11]. Multi-Index [FWYZ13]. Multi-Layer [BJ11]. Multi-Population [WH17]. Multi-Resolution [Kat17]. Multi-Sample [LCG^{+}15]. Multi-Scale [BMMS17]. Multicategory [ZWQ18]. Multiclass [PWL16, WZL10]. Multicomponent [CCS14]. Multidimensional [FDPS10, LLL^{+}10, MML12]. Multilab [FCDA15]. Multilevel [BL16, CWPC13, HBHC12, HBI^{+}19, ZK17]. Multimethod [OKEK17]. Multimodel [RGH13b]. Multinomial [ACBK10, Tad13a]. Multiple [AIP19, BBBH10, BFWE10, BCDS18, BT16, BR12b, CKP19, CL16a, CLQY16, CXCR13, FKSZ19, FS17, FH14, GZCL19, HLL10, HTFK10, HMW^{+}14, HDL^{+}16, IJ19, KLH11, KLM13, KRS^{+}17, LCT16, LQS17, LN10, LL19b, MZN^{+}16, MJ14, MR17, OST19, PSY19, PSV15, PPD15, RW10, RPS19, Ros14, SF13, SCG13, SW11, SM12b, TYY^{+}19, WF12, WZ17a, WB19a, WB19b, XCC18, XTR^{+}14, YTL15, ZLW10, Zha11, ZLP^{+}14, ZSS19, ZSP14, ZR15]. Multiple-Regime [YTL15]. Multiplicative [CMW^{+}19]. Multiply [Han14]. Multiproxy [JR14, LNA10b]. Multiresolution [CDBG16, KOLL12]. Multiresponses [TGZ15]. Multiscale [FF17a, GPD14, HPX12, HTFK10, TARS17, ZSPL17]. Multistate [WH17]. Multitask [FXZ17]. Multitrait [OKEK17]. Multitrait-Multimethod [OKEK17]. Multivariate [AGS12, BD12, Cer10, CDQ13, CDQ14, CF17, CMS^{+}18, CCS18, CHSV14, DYZ12, DP11, DLR11, FZW16, FFvSK17, GZJLM12, GKS10, HAME15, HG18, HLMH18, HMW18b, IKP11, Ima11, JP15, KMO15, KWG15, KMR17, LVC14, LDGX15, LS18b, LK19, LJGZH19, LLX15, LJW13, MS18c, MVR17, MT11, MJ14, MCIS19, MAS14, NLOP10, OP13, PCJ12, PK12, PPD15, SPC^{+}18, TZF^{+}15, WPL15, WLZ19, YKC^{+}15, dCD14]. Multivariate-Sign-Based [FZW16]. Music [RDLC10]. MWPCR [ZSPL17].

Naive [LZNS17, MYLC15]. Natural [GC17]. Nearest [DBFG16, SQC16]. Nearest-Neighbor [DBFG16]. Need [PD16]. Negative [ZPS16]. Neighbor [DBFG16, SQC16]. Neighborhood [RSK17]. Neighborhood-Specific [RSK17]. Neonatal [ZSL^{+}11]. Nested [BL16, CZ11, CZ13, HTGT13, PHP^{+}12, QA10, Tan10, XLN18, ZD13]. Net [Han11, PPD^{+}14]. Nets [FMV16]. Network [AB13, AEI18, CSW17, CL18, CIB^{+}18, CXCR13, Cho17, CWH18, FH15, GMS14, HCCZ16, HPX12, LHC19, LLLW11, MSZ10, PCSK16, WGE^{+}18, WLSA17, WLXL14, YJFL19, ZC13, ZC17a]. Networks [Bro17, CD18, CO10, DDV17a, DDV17b, FTZ15, FF17b, FSS^{+}16, FZ13, HP10a, HP10b, KLM13, LCZ12, LCZ14, LS18c, LLZ18, LB17b, MZ15, Raf17, RKFL19, RLM13, Sah10, SC15, Zho11]. Neural [LLZ18, SKS^{+}15]. Neuroimaging [JAW16, KJNW11, VRS19, ZS15a, ZLZ13, ZKLI14, ZFK14]. Neuronal [ZMB^{+}16]. Neutron [SMQ^{+}13]. Neutron-Induced [SMQ^{+}13]. Next [BA16]. Next-Generation [BA16]. NHANES [KSGB10]. Nigeria [MSK^{+}17]. Nilanjan [Pea19]. Nitrate [GBDL10]. NMR [ADR^{+}12]. Nocturnal [KRS^{+}17]. Nodal [FH15, VH15]. Node [SRR^{+}17]. Noise [BIL^{+}11]. Noises [ZL15b]. Noisy [ASFX10, BBBH10, CDH12, JKL11]. Non [HL18, MTY^{+}17, RMR19, RK11, WS14]. Non-Bipartite [RK11]. Non-Gaussian [HL18, RMR19, WS14]. Non-Small-Cell [MTY^{+}17]. Noncompliance [FMPR12, MP13]. Nonconvex [LB18, MFH11, SO11]. Nonemployment [FMPR12]. Nonhomogeneous [LBS13]. Nonignorable [GDM17, KY11a, KC10, LHD19, LD15, MRB^{+}17, MDG17, NC10, ZS15b]. Nonlinear [CLLL18, LW19a, MMC^{+}12a, RJ10, WX15, WGDG14b, YBT13, ZCL11]. Nonlinearity [ZSZ12]. Nonlocal [RT17]. Nonmonotone [CL10, ST18]. Nonparametric [AVWW14, BJQ17, BCR18, BFT18, Blu10, BGK^{+}18, BOSB16, Bro17, BR12b, CZ11, CLM14, CCF18, CDH11, CGSW10, DNFZ10, DQZ18, DM11, DH14, DZ15, DPT14, DDV17a, DDV17b, Efr11a, FOW11, FFS11, FMD14, FF17b, FK14a, GPPVW12, Har19, HLW13, HC15b, JYL15, KX17, Ken19, LMZJ13, LCAL12, LG13, LS18c, LB17b, LCLZ18, LT14, DPT19, MJ14, NLOP10, PARS19, PBD13, PK12, PAHJ11, PLLG18, QJWG16, RBB13, RDLC10, RBF^{+}17, RK11, SD17a, SH13, SWW14, SHW17, TT12, WRL10, WGM12, WMA^{+}12, WPL15, WT13, XMWT16b, YOD11, ZDD19, ZLL18, ZJZ12, ZD13]. Nonparametrics [FFJT16]. Nonrandom [RBB13]. Nonresponse [HBZ15, KC10, KBKS18, LHD19, NC10, Per10]. Nonreversible [BCVD18]. Nonstandard [MN16]. Nonstationary [BSZ12, MBCM13, MM15, RWS12]. Normal [AS15, CLF12, CRZZ16, FTM12, Han11, KS15, MDG17, SH15, VS15, WMBZ18]. Normality [RR18]. Normalization [Sha15]. Normalized [AVWW14, MS18c, ZL18]. Normalizing [LJSL16, PH18]. Normally [DH12]. Norman [Pea19]. North [CCS14, RLGL11]. Northwest [BIL^{+}11]. Novel [GC11, HNW^{+}17, LX19, WBG^{+}18, ZCJ^{+}17]. NPMLE [BKRF19]. ntrol [Cre18a]. Nuclear [CXCR13]. Nucleosome [XBZ^{+}14]. Nuisance [LQ11]. Nuisance-Parameter [LQ11]. Null [FSMS17, SM12b, ZG18]. Number [AB17b, AGS12, AS15, BBBH10, BJT^{+}10, CL18, GBP19, HLMH18, JCL10, KS15, LWC13, MH18, PGY14, TAGT14, Tra18, YZBE18]. Numbers [HLSZ14]. Numerical [OCAG19].

Object [CF17, CCS18, LMH14, Mar19, PM19, TPAC19a, TPAC19b, DPS19]. Object-Oriented [LMH14, DPS19]. Objective [BBS12, VS15, VW15]. Objects [FG13]. Observation [DZ19, FAIW^{+}11, MS18b, SSZL12]. Observational [BSLR10, ESR18, FMGS16, FSMS17, FS17, HRS14, HSR13, PKSR15, Ros10, Ros11, Ros14, Ros15a, Ros15b, WGM12, YLRH19, ZSL^{+}11, ZSR18, Zha19b, ZNSR12, Zub12, ZK17]. Observations [BBBH10, CM12, CWZ15, JKL11, RSI14, SSS18, XY12, ZDP10]. Observed [BIL^{+}11, CY16, HG18, LT14, MM15, WTM19, YKC^{+}15]. Ocean [HB14]. Odds [CHC^{+}12, CRL15, MW10a, SM16]. ODEs [LLLW11]. Off [ESR18, KX17, SL14]. Offshore [LDGX15]. Older [IZR^{+}13, MŞDN12]. On/Off [SL14]. Oncology [YY11]. One [Lu16, LvdL18, PSY19, SM12a, VW11]. One-Sided [Lu16, LvdL18, PSY19, SM12a]. One-Way [VW11]. Online [WG11]. Only [LLX15, WTM19]. Onset [SYS^{+}19]. Ontology [LN10]. Open [Dai19]. Operating [LHC19]. Operator [CY16]. Optimal [BPPD15, BG16, CCT15, CY17, CLQY16, CB18b, DNFZ10, DT12, DJW18a, FRG^{+}17, GJS16, GWCL19, HZ10, JCL10, JM14, KFE12, LZWZ11, LHW^{+}13, Pap12, PKSR15, RLY14, SWSK19, SB18, SDWM18, TPLG10, UL17, Wan15, WFZ18, WZM18, WZSS18, WYS19, WAH19, YBT13, ZYZL17, ZH19, ZZLK15, ZDP10, ZK17]. Optimality [MW10b]. Optimism [TR19]. Optimization [FMGS16, KM14, LCL14]. Optimize [ZB13]. Optimizing [MYT18, TNZM14]. Optional [MW11]. Options [ZWQ18]. Oracally [LYH13]. Oracle [CLZ10, LLSS18]. Order [CLF12, DP11, DHLL14, HLR15, LRT12, LC10, PKM10, PR14, RGH13b, RDHL19, SD17a, ZH19]. Order-Based [DHLL14]. Order-Restricted [DP11]. Ordered [FDPS10, Fuk15, IZR^{+}13, LB17a, MM13, YOD11]. Ordering [DO12]. Orders [HC15b]. Ordinal [DK18, HH10, LVC14, LS10, LZ18, PLGM11, SD17b]. Ordinary [BCdB14, CSW17, MWX14, WLXL14, WQxYW19]. Ordination [RBF^{+}17]. Oriented [LMH14, YYY18, DPS19]. Ørnulf [Pea19]. Orthant [Han11]. Orthogonal [MT11, QA10, SVTG17, SLW16, ST17, ZX17]. Orthogonality [BCdB14]. Osteoarthritis [LVC14]. Other [AB17a, CH17]. Outbreak [ASB11]. Outbreaks [NFG^{+}16]. Outcome [CZK17a, FH14, FMPR12, Fuk15, Hos13, KWG15, LCLZ18, ZZRK12, Zub15]. Outcomes [CDBG16, CW12, DGYZ11, FSMS17, HBHC12, JLTY14, KPGJ12, LVC14, LLL^{+}10, MP13, PKSR15, RS10, SHM^{+}18, WRL10, XTR^{+}14, ZSL^{+}11]. Outlier [ASB11, Cer10, JP15, SO11]. Outlier-Robust [ASB11]. Outliers [FR19, LB18]. Output [GH18, SWSK19]. Outputs [SLR^{+}15]. Over-Dispersed [HN18]. Overall [RLY14]. Overdispersed [KKL15].

Packing [He17]. Pair [BA16, PCJ12, SY18, SMAC10, Zha19b]. Pair-Copula [SMAC10]. Pair-Matched [Zha19b]. Paired [HS12, HLG14]. Pairs [LQ11]. Pairwise [BL16, DDT17, HMW18b, MH17, PWL16]. Paleoclimate [CT10, JR14, LNA10a, Smi10, WSWT10]. Paleomagnetic [SW19]. Pallid [WH17]. Pancreatic [LDSH16]. Panel [AB17b, GLL13, LQS17, MS18b, Sob12]. Paradata [NSK16]. Parallel [RLP^{+}18]. Parallelism [Zha13b]. Parameter [BBS12, CR10, DRC^{+}12, FL13, FHS13, HWF15, KOLL12, LQ11, LRZ17, SPZ12, VW15, WPS17b, WQxYW19, XCM^{+}13, ZLT10]. Parameterization [YZB^{+}16]. Parameters [CHC^{+}12, CXT14, DBNZ16, DFRS14, HV11, HZ13, LvdL18, ML15, PJY16, SLR^{+}15, Wan15]. Parametric [BCdB14, DZ15, FOW11, FF13, Hos13, LvdL18, SW17b]. Parametric-Rate [LvdL18]. Parametrically [DH14]. Parasites [LS19a]. Parsimonious [LZ17]. Partial [BG11, CL10, CGSW10, FWYZ13, HCM13, HMM16, HD10, yHQ12, KS11, LSQ15, MLT17, SHM^{+}18, WO19, XCM^{+}13]. Partially [BIL^{+}11, DGM12, FWYZ13, IZR^{+}13, LZ10c, Sid13, WLZ19, YKC^{+}15, ZCL11, ZWZZ19]. Participation [DIvdB13, KPGJ12]. Particle [BCVD18, BG16, GJL17, LPH^{+}17, Roc18, RLM13, YYC17]. Partition [DDT17, JL15, KM17, ZDP11]. Partitioning [LJW13, TPLG10]. Partitions [ZFW17]. Passing [Wan17a]. Past [LNA10b, WRCG13]. Path [MSK^{+}17, ZW14]. Path-Specific [MSK^{+}17]. Pathogen [ACBK10]. Pathogens [TYY^{+}19, YHD^{+}10]. Pathway [CPPW11, LCZ14, PCSK16, XCC18]. Pathways [TMK^{+}12]. Patient [HBHC12, KYBB19]. Patient-Reported [HBHC12]. Patients [LZNS17, MŞDN12, WFZ18, ZB13, ZLP^{+}14]. Pattern [HSM^{+}15, SD17b, SW11, XWG19]. Patterns [DHS11, Hah12, KA18]. PCA [HL14a]. PDE [ASS^{+}15]. Peer [LDL^{+}19]. Penalization [ASS^{+}15, HZZ^{+}17]. Penalize [FSLR19]. Penalized [BR12a, CW14, CWPC13, CH19, FHS13, KKC10, PL19, SO11, Spi19, SDWM18, XMZ12, YZ19, ZEL19a, ZEL19b]. Penalties [MFH11]. Penalty [HLL10]. PENCOMP [AD19]. Penetrance [SYS^{+}19]. People [MSZ10]. PEPFAR [MSK^{+}17]. Per-Protocol [GSH13]. Performance [SMQ^{+}13]. Perfringens [CMS^{+}18]. Perinatal [BSLR10]. Period [HN18, PAHJ11]. Periodicity [ATT19]. Periodograms [Li12b, Pap10]. Periodontal [RBB13]. Periods [FAIW^{+}11, FH14]. Permutation [ACCTW18, DR17, GJ18a, Hah12, KLH11, ST15]. Persistent [MJG18]. Personal [MSZ10]. Personalized [CZK17a, LZNS17, MZC16, WFZ18, ZSY16]. Perspective [Blu10, DGS19, HC15a, LL19b, MZW16, YHD^{+}10]. Perturbation [MTC11, PCSK16]. Perturbations [CB18b, HMQA10]. PET [CGO19]. Phase [GY17, HAME15, LGY18, TZL17, ZL14, JLTY14, ZB13]. Phenotype [LL19b]. Phenotypes [WBG^{+}18, XLN18, ZKLI14]. Phones [YRR19]. Phylogenetic [WBCD15, Wil19]. Physical [HBI^{+}19, Wu15]. Physics [MSW^{+}18]. Physiology [KRS^{+}17]. Pine [FRG^{+}17, LHTB15]. Pivotal [DV11]. Placebo [LZS^{+}16, RLP^{+}18, TPLG10]. Planar [SKBM17]. Plane [FSL17]. Planes [YT17]. Plant [XLN18]. Plausibility [Mar15]. PLEMT [HNW^{+}17]. Plot [LCAL12]. Plots [CCT15]. Point [CZMHS17, FSS^{+}16, GSDR19, Gne11, GPD14, Hah12, HLL10, HSM^{+}15, KJNW11, LLSS18, LG14, MJ14, MSB^{+}11, Sar12, SO19, XY12, XCQ15, XWG19, ZL18, ZMW^{+}15]. Points [SZ10]. Poisson [BGR13, DO12, FF17a, KCZ^{+}13, KA18, LBS13, Tad10, WLY^{+}14, ZZWJ12]. Policies [FL10]. Policing [MSM^{+}15]. Policy [Mor10]. Polls [SMRGG18]. Pollution [GDZ11]. Pólya [MW11, PSW13]. Pólya-Gamma [PSW13]. Polymerase [HV11]. Polynomial [PPLS18]. Polynomials [BJQ17]. Population [AT11, CAJ14, CCL^{+}11a, CCL^{+}11b, CWH18, DGH^{+}10, FTZ15, HGZ18, Laz11, LD17, LZ11b, LNM11, MHC11, MKG13, MDR18, RLY14, She11, TEKM12, VT17, WTM19, Wu11, WH17, YZBE18, ZTC^{+}13]. Population-Based [HGZ18]. Populations [BSK^{+}15, Bro17, bCH10, CCL^{+}11a, CCL^{+}11b, DDV17a, DDV17b, FMP18, FMGS16, FF17b, HBI^{+}19, Laz11, LB17b, LNM11, MHC11, MMW^{+}17, She11, WRCG13, Wu11]. Pore [BBBH10]. Portability [PD16]. Portfolio [FZY12, FLY12, GBZ17].

Portmanteau [FG12]. Posed [CB18b]. Position [Raf17]. Positioning [XBZ^{+}14]. Positive [IY14, PPD^{+}14, XMZ12]. Positive-Definite [XMZ12]. Possession [CDBG16]. Post [BCK19, CTM10, GRH10, HCKS18, TTLT16a, Wu15]. Post-Fisherian [Wu15]. Post-Selection [BCK19, HCKS18, TTLT16a]. Post-Stratification [CTM10]. Post-Vote [GRH10]. Posterior [GKM19, Guh10, HC15a, ML13b, Zha11]. Potential [DD16, Hos13]. Poverty [HBZ15]. Power [DPS19, GRH10, HSTP15, JT12, LDGX15, LMR14, LL19b, SdCG^{+}15, WL13, WFB13]. Powerful [Lit13, ZSCM18]. Powering [HG10]. pp [Pea19, Tri19]. PQL [HMW17]. Praise [Lit13]. Precipitation [KRG11]. Precision [CLL11, GY17, PPLS18, SB18, XCC18]. Predicting [CDBG16, CZW^{+}11, GR11]. Prediction [ANH15, BGK^{+}18, CY12, CG15, DTYG12, HH10, JRFN18, KLCT11, LRW13, LW19a, Lin18, MM15, PCC12, PBP19, Plu14, RD12, Wol11, ZSY16, ZSPL17]. Predictions [DNFZ10, RHCT14]. Predictive [BCR18, CDS11, HMW18a, JNR11, LGR^{+}18, MSM^{+}15, WZ17a, ZDP11]. Predictor [HCM13, IY14, PSSS17, YT17]. Predictor-Dependent [HCM13]. Predictors [AB17b, BGSB19, BF15, BDF16, Efr11a, GLS11, MQ15a, WSSQ16, YL16]. Preference [FDPS10, KYBB19]. Premature [BSLR10]. Presence [BF18, CDH11, DYZ12, DHM11, FR19, FJM19, FMPR12, GZJLM12, KHK19, LB18, LH14, MQ15a, PH18, PR14, PY12, SW17c, SHW17]. Presidential [Lin13]. Press [Pea19, Tri19]. Preterm [TNZM14]. Prevalence [DZ15, DG16a, Gil11, HMM16, LZ10a, MRB^{+}17]. Prevalent [yHQ12]. Prevention [DGM12, Pre10, WDSL10]. Price [KLL17, SY17]. Primary [BH13, SKS^{+}15]. Primate [BAA^{+}15]. Principal [ASX19, BGH^{+}18, BG11, BSB15, CL15b, DGM12, DGYZ11, EUW17, FCSZ16, FMV16, HPV14, HG18, HJCPV11, JWXJ13, LWC13, LG14, LL19b, NH11, PPY14, PY16, PSY19, SdCG^{+}15, SLW16, YZB^{+}16, ZSPL17]. Principle [ACLZ14]. Prior [Han11, JHZ16, ML13b, ML15, MH18, VW15]. Prior-Free [ML13b, ML15]. Priors [ACB^{+}14, BBS12, BPPD15, DRC^{+}12, FSLR19, GLP19, LC18, MN16, RT17, SFK12, TGHS18, ZPS16]. Privacy [WZ10]. Private [DJW18a, ZK17]. Probabilistic [BRGS10, BGB12, DT15, GBP19, HRZ17, KRG11, ML13b, ML15, MLC^{+}16, OCAG19, SGR10]. Probabilities [DRS18, WT13]. Probability [DTYG12, Hun11, IR15, LT14, SB18, SPU17, ST18, WZL10, YLRH19, Zha11]. Probability-Based [Hun11]. Probe [HH12]. Probing [KBKS18]. Probit [Gee14]. Problem [ACBK10, BGR13, DH14, FZW16, GC17, LQ11, MW11, WSG16]. Problems [GdMb13, GLS11, MN16, Zho14, Dai19]. Procedure [DW11, FDPS10, GAZ13, LCAL12, PWXZ19, TGZ15, Tra18]. Procedures [BT16, BB10, CZ13, DJW18a, TMP18, TTLT16a, Tsi17, WLNC14]. Process [BFWE10, CDBG16, CZMHS17, DBFG16, HTGT13, IZR^{+}13, MSB^{+}11, SHW17, WS14, ZA16, ZMW^{+}15]. Processes [AB13, BSZ12, BB14, CW17, CMS14, DPV11, FO17, FSS^{+}16, GZJLM12, GLS11, GSSVF13, HV11, HHY19, HW19, IY14, KJNW11, KWX^{+}19, KOLL12, LBS13, LG14, PRS10, PR14, RD12, RDG10, SD14, SPHL16, SW17c, Tad10, WG14, WDP^{+}14, ZPS16, ZD13]. Procurement [GSDR19]. Produced [PKSR15]. Profile [LPR12]. Profiles [LCCG14, MZT^{+}17]. Profiling [RS10]. Program [ADH10, GK18, MSK^{+}17, WH17]. Programming [FS17, RLY14, Zub12]. Progression [BMW19, LDS10]. Progression/Death [LDS10]. Progressive [KLN13]. Projected [WG14]. Projection [BY11, SZLI15, VRS19]. Projective [RWKS14]. Promoting [HZZ^{+}17]. Prompt [KHK19]. Prone [DDH15, LM17]. Propagated [Cul11]. Propagation [BC14, Ma15]. Propensity [CH19, GMS14, JLZ14, Ken19, LMZ18, MRB12, PL19, Per10, Spi19, YZ19, ZEL19a, ZEL19b, ZD14]. Proper [FG13, HSTP15, ZZS19]. Properties [LMR14, MW17, QZL^{+}10, SQC16, ZR15]. Property [CLZ10]. Proportion [FHG12a, GR11]. Proportional [CCZF15, CGSW10, CHC^{+}12, MW10a, SM16, SW14]. Proportions [HY12]. Proposals [LTC13, LCT16]. Prostate [WRL^{+}12a, ZCJ^{+}17]. Protein [CSSK16, MZN^{+}16, TMK^{+}12]. Proteins [CMS^{+}18]. Proteogenomics [NSH^{+}19b]. Proteomic [CPPW11]. Proteomics [LMZJ13]. Protocol [GSH13]. Prototypes [BT11a]. Provide [ZNSR12]. Provider [RS10]. Pseudo [RW10, ZS15b]. Pseudo-Likelihoods [ZS15b]. Pseudolikelihood [HNW^{+}17]. Pseudolikelihood-Based [HNW^{+}17]. Psychological [JPW^{+}17]. Public [DR10, GRR^{+}17]. Pull [SWCB16]. Pursuit [KFW15, LS19b, SH10, YDZ16, ZSP13, ZSP14, ZSY16]. Push [SWCB16].

Quadratic [HFZ18, LMR14]. Quadratically [FS17]. Qualitative [ZSS19]. Quality [KBKS18, LDSH16, OKEK17]. Quantification [ADR^{+}12, SWSK19, SCK19]. Quantifying [FBM11, GMS14, MSK^{+}17]. Quantile [ACLZ14, BCK19, CWZ15, DM18, DEV19, DT12, EG19, FXZ17, FZ16, FKH11, GLL13, Hag17, KY11b, KW15, LNP14, LLSS18, Li12b, LCF14, LLT15, MLT17, RFD11, SPHL16, WWL12, WZSS18, WSL^{+}16, WY13, WMY15, XSWH17, YT17]. Quantile-Optimal [WZSS18]. Quantile-Regression [EG19]. Quantiles [JR14, LLS17, LS17b, SYY15, WLH12, WL13]. Quantitation [HV11]. Quantitative [ESR18, JL15, KRG11]. Quasi [TLG16, WC18, XWG19, ZST12]. Quasi-Likelihood [TLG16, WC18, XWG19, ZST12]. Question [AIC14]. Questionnaire [CL10]. Queue [HLR15]. Queue-Reactive [HLR15].

R [Har19]. Race [KPC^{+}17]. Radial [RLM13]. Radiative [CMM^{+}13]. Radiological [PARS19, TARS17]. Radius [ABvdW19]. Random [AO17, AGS12, BJT^{+}10, CY16, CYC10, DDT17, DM15, DW11, Efr11a, FJM19, FSLR19, GKS10, HS12, Han17b, HV11, HZ10, KC11, KY11b, LHTB15, LMWA17, PGY14, PKM10, QLL10, RB18, SM12a, SSS18, ST18, TGHS18, WA18, WfTQ12, XG17, ZPS16, ZFW17]. Random-Effects [WfTQ12]. Randomization [BDIK12, BS19, BCS18, FMGS16, FSMS17, KZCS16, LDL^{+}19, LH14, NH11, Sna19]. Randomization-Based [NH11]. Randomized [BIZ15, DD16, GSH13, HTGT13, JTF^{+}19, MHZ15, MP13, MSM^{+}15, PTC14, RLY14, Tra18, VHJB13, WRL^{+}12a, WLS17]. Randomizing [QA10]. Range [BBS17]. Rank [ACCTW18, BAA^{+}15, CH12, CWPC13, DT15, DHLL14, HS12, LLL^{+}10, PLLG18, PD14, RD12, WT13, ZHM^{+}10, ZKLI14]. Rank-Based [ACCTW18, LLL^{+}10]. Rank-Tracking [WT13]. Ranked [WLS17]. Ranking [UL17]. Rao [GC11]. Rare [BAN^{+}12, LLgX14]. Rasch [GDCL11]. Rate [BCDS18, GRR^{+}17, HZZ10, LN14, LvdL18, SKS^{+}15, SLS18, TKPS18, WY13, Zha11, ZTS11]. Rates [KLH11, LS15]. Rating [SY18]. Ratio [CRL15, CLQY16, CLA10, HZT11, LCG^{+}15, LJZ14, TLG16, ZST12, dCD14]. Rational [Pap12]. Raton [Pea19, Tri19]. Ray [ABvdW19]. Reactions [HV11]. Reactive [HLR15]. Readability [KLCT11]. Readings [LSS^{+}19a]. Real [PGW^{+}18]. Realized [CHP10, TT12]. Reassessing [XAM^{+}14]. Reassessment [YY11]. Recapture [BSK^{+}15, YRR19]. Receiver [LHC19]. Reciprocal [SL15]. Recognition [KPC^{+}17]. Recommender [BQWS17]. Reconciliation [WAH19]. Reconciling [TMP18]. Reconstructing [WRCG13]. Reconstruction [CCS14, CSW17, COS^{+}10, CT10, LNA10a, LNA10b, SBG^{+}16, Smi10, WSWT10]. Reconstructions [JR14]. Record [HRZ17, RLGL11, SF13, Sad17, SHF17]. Recovery [CLX13]. Recurrence [ZCJ^{+}17]. Recurrent [CW17, SPHL16, SW17c, XCH^{+}17]. Recursive [CY17, HMW18a, Ma13]. Recursively [ZK12]. Reduced [BB14, CH12, CWPC13, VV15, ZHM^{+}10]. Reduced-Rank [CH12, CWPC13]. Reducing [SFM18]. Reduction [CH12, CZ14b, DW11, Efr10a, FWYZ13, FL14, GLLL15, HLMH18, HC17, LS18b, LLZ10, MZ12, QDC19, XZX10, YNLC19, ZZF10]. Reductions [BF15, BDF16]. Referenced [ZH18]. Refine [ZWQ18]. Refined [PKSR15, XCQ15]. Refinement [DW11]. Reflections [D’A19]. Refracted [COS^{+}10]. Regime [YTL15]. Regimen [LER^{+}12]. Regimes [BH18, LS18a, MYT18, WRL^{+}12a, WZSS18, XMWT16b, Zaj12, ZLDT18, ZZLK15, ZZS19]. Region [BTI19]. Regional [FL10]. Regions [BR12a, SSS18, ZDD19]. Regression [AL14, AVWW14, ACLZ14, ASS^{+}15, BJQ17, BCK19, BS13, CY12, CLM14, CCT15, CDH11, CJN18, CL10, CHC^{+}12, CH12, CCS14, CWZ15, yCtW13, CWZM18, DBNZ16, DEV19, DT15, DK18, DM11, DZ19, DT12, DVV14, DMVB18, DR17, DPT14, Efr11a, EG19, FOW11, FXZ17, FZ16, FKH11, Fu16, FL14, GLL13, GJ18a, GZCL19, GCNC14, GC11, GD15, GHJZ16, Hag17, HCM13, Han14, Han11, HFZ18, HH10, HY12, HLW13, Hua17, HYSM19, Ima11, KS15, Ken15, KY11b, KCZ^{+}13, KW15, KS11, LHL15, LSS11, LNP14, LLSS18, LGR^{+}18, LZP10, LCF14, LZ17, LL19a, LS19b, LFU11, LL13, LFL15, LSZD17, LZL19, Lin18, LZ18, LQ17, DPT19, MCIS19, MTC11, NZK11, NSB19, NSH^{+}19b, NEB13, PHSS15, Pap12, RFD11, RGL16, RWKS14, SW19, SFK12, SKS^{+}15, SC12, SO11, SZI^{+}12, Sob12, SDS19, SZLI15, SPHL16, SDWM18, Tad13a, TTLT16a]. Regression [TE11, Wan17a, WRL10, WF12, WMA^{+}12, WWL12, WS14, WZ17b, WZM18, WYS19, WSL^{+}16, WSW17, WY13, WMY15, XY12, XSWH17, YKC^{+}15, YYY18, YL16, ZC11, ZLZ13, ZHS17, ZG18, ZBM11, ZD13, ZKLI14, ZR15, ZSPL17, ZLWT17, AR15b]. Regressions [BF15, BDF16, LTT17, ZZF10]. Regressors [AIZ14, NL12]. Regularity [KA18]. Regularization [FL13, FHS13, FZ13, GNL19, HFZ18, LFL15, SH10, SL15, ZLT10]. Regularized [CPPW11, CCS14, COS^{+}10, HMW17, MTC11, SWJ18, ZW14]. Regulation [ZC17a]. Regulatory [LLLW11, WLXL14]. Reinforcement [ZZK15]. Reinforcements [ACB^{+}14]. Reject [ZWQ18]. Rejection [BCVD18]. Rejection-Free [BCVD18]. Rejective [FLL17]. Rejoinder [Air17, AR15a, CCJ13b, CCMC16b, CZK17b, CCL^{+}11b, Cre18b, DG16b, DJW18b, DDV17b, Efr10c, Efr14b, FHG12b, FPSE15a, FK14b, HP10b, LM11b, LNA10a, MMM11b, MMC^{+}12b, MQ15b, MJPW13a, MG17a, Tad13b, TPAC19a, TTLT16b, Wan17b, WRL^{+}12b, WB19b, WGDG14a, WPS17a, XMWT16a, ZEL19b]. Related [HSM^{+}15, KSKD12, ZCJ^{+}17]. Relatedness [GWCL19]. Relational [MZ15, VH15]. Relationships [DK18, KSGB10]. Relative [CGLY10, Ken15, RRW17]. Relative-Risk [Ken15]. Release [FSB^{+}19]. Releasing [DR10]. Relevance [LS17a]. Relevant [Tri19]. Reliability [Pre10, SMQ^{+}13]. Remote [Cre18a, NCB12]. Repeated [CM12, SSS18]. Replicate [BH13, PLLG18, YY16]. Replicated [FOvS10, KRS^{+}17, KHG18]. Reply [Fu15, YL15]. Reported [HBHC12]. Reporting [GRR^{+}17, NFG^{+}16, SEdS19]. Representation [AI12, SZI^{+}12]. Representations [LMWA17]. Representative [YLRH19]. Reproducibility [JPW^{+}17, PLLG18]. Reproducing [BCT18, SDWM18]. Rerandomization [MR15]. Resampling [CZ13, LCS^{+}13, MQ15a]. Resampling-Based [LCS^{+}13]. Research [JPCD19, Pre10, WGM12]. Residential [SWCB16]. Residual [MS18c, ZMHKK17]. Residuals [DV11, HS13, LZ18, WX15]. Resolution [BA16, FTM12, Kat17]. Resolving [GRH10]. Resonance [KOL^{+}12, LST^{+}11]. Resource [DG16a, RGH13a]. Respect [SLC^{+}15]. Respecting [Hua17]. Respiratory [TNZM14]. Respondent [CWH18, Gil11]. Respondent-Driven [CWH18, Gil11]. Responders [TPLG10]. Response [BIZ15, CYC10, DW11, HS13, HZH15, JPBM13, LHH17a, LZ17, MMC^{+}12a, PSSS17, Per10, PS10, RLP^{+}18, ST15, WMA^{+}12]. Response-Adaptive [HZH15]. Responses [KBKS18, MRB^{+}17, NTC13]. Restoration [Hua17]. Restricted [DP11, XS18]. Results [HILL16, LER^{+}12, RDHL19, YHD^{+}10]. Retesting [BTC10]. Retrieval [WJYJ13]. Return [You19]. Reusing [XCQ15]. Reversible [BH18]. Review [Ano13a, BKM17, Dai19, HILL16, Har19, Kim19, Pea19, Sha15, Sna19, Su19, SHM^{+}18, Tri19, Vat19, Yan19, Yu19, Zho19]. Reviews [Ano10b, Ano10c, Ano10d, Ano10e, Ano11b, Ano11c, Ano11d, Ano11e, Ano12a, Ano12b, Ano12c, Ano12d, Ano13b, Ano13c, Ano13d, Ano14a, Ano14b, Ano14c, Ano14d, Ano15d, Ano15a, Ano15b, Ano15c, Ano16c, Ano16a, Ano16b, Ano17a, Ano17b, Ano17c, Ano17d, Ano17e, Ano18a, Ano18b, Ano18c, Ano19a]. Revisited [BGR13]. Right [GZCL19, HLSY16, LW19a]. Right- [GZCL19]. Right-Censored [LW19a]. Rights [JML^{+}14]. Rilpivirine [LZNS17]. Ring [SBG^{+}16]. Risk [BKRF19, BGK^{+}18, CAJ14, DGH^{+}10, FKH11, GHP13, HCOW17, Ken15, LZPH14, MMW^{+}17, MVR12, MYLC15, MŞDN12, RS10, RRW17, SHW17, TGZ15, WFZ18, YLRH19, Zho19]. Risk-Adjusted [RS10, TGZ15]. Risks [LDSH16, LY16, MTY^{+}17, XAM^{+}14]. RNA [HSTP15, LNN19, SLC^{+}15]. RNA-Isoform [SLC^{+}15]. RNA-Seq [HSTP15, LNN19]. Road [BRGS10, GMS14]. Robust [AY17, ACF17, ASB11, BS13, CCR15, CCJ10, CH17, DYZ12, DR17, DVF11, FK18, FKSZ19, FPW10, Hag17, Han14, IR15, JR14, JTF^{+}19, KW15, LTT17, LL19a, MKG13, MD19, MTY^{+}17, QP13, SdCG^{+}15, SLW16, VW11, VV15, WJHZ13, WZL19, WZL10, YY11, Zho13, ZBM11]. Robustness [DH12, God17, KX17, LT10, LRT12, LLX15]. ROC [HC15b]. Rockfish [DK18]. Role [Cre18a, HBZ15]. Rolling [ACPLRC19, Pap10]. Root [LCL14]. Rotated [He17]. Rotations [DPT19, RG17b]. Rules [KM14, WFZ18, ZZRK12, ZMHKK17]. Run [CLM18, GLP19]. Runs [SVTG17].

Saddlepoint [MR11, PCB11, PDF15]. Safety [HZT11]. Salmonella [DDM^{+}10]. Sample [CLX13, CDQ13, CDQ14, CF17, CCS18, DV11, ESR18, FSL17, FZW16, GCBL15, JYL15, LCG^{+}15, LSS^{+}19a, LH14, MW11, MKG13, PCB11, SM12a, TE11, WZM18, YKC^{+}15, ZZWJ12]. Sampled [RQJ15]. Sampler [BCVD18, NSH19a, TY17]. Samples [LCM10, ZSL^{+}11]. Sampling [ACB^{+}14, AT11, BH18, BSK^{+}15, BG16, CKY15, CWH18, DR10, FLL17, GdMb13, GS17a, Gil11, yHQ12, JCRG17, KLSY13, LTC13, LJSL16, LCLZ18, MZN^{+}16, MDR18, QA10, TZF^{+}15, TZL17, TMP18, VT17, WLS17, XMY^{+}15, XSWH17, ZC13, Zho11]. Satellite [SSL^{+}10]. Scalable [CIB^{+}18, Ma15, NSH19a]. Scalar [WZ17b]. Scalar-on-Image [WZ17b]. Scale [BCDS18, BMMS17, BM10, Cai10, CL16a, Efr10b, Efr10c, HL14a, Hel10, HCKS18, HR11, KKL15, LZ10a, Sch10, TARS17, Wes10, XCH^{+}17, ZK17]. Scale-Change [XCH^{+}17]. Scale-Invariant [HL14a]. Scaled [SW19]. Scaling [FDPS10, PS10]. Scanning [MM19]. Scans [ACCTW18]. Scatter [DYZ12]. Scattering [ABvdW19]. Schedule [LCL14, ZB13]. Scheme [JP15, QA10]. Schemes [KLSY13]. Schizophrenia [LD15]. School [AR15b]. Schools [LZ10a, ZK17]. Science [JPW^{+}17, Yu19]. Sciences [RSA16]. Scientist [Cre18a, Lit13]. Sclerosis [ZLP^{+}14]. Score [GMS14, HFQ12, JLZ14, Ken19, LMZ18, MMW^{+}17, NQS14, SJDT19, VRS19, WZL19, WMY15, ZD14]. Scores [FMC11, MRB12]. Scoring [Cul11, MJPW13b]. Scott [Pea19]. Screen [CYT12]. Screen-Detected [CYT12]. Screening [BFV16, CLZ15, FFS11, FMD14, HZ14, HR11, KXZ19, LZ10a, LZZ12, MLT17, PWL16, PWXZ19, SVTG17, SZ14, XC14, ZSR18, ZX17, ZLLZ11].

Sea [RLM13, CCS14]. Searching [HH12]. Seasonal [DHS11]. Seber [WH17]. Second [PKM10, PR14, PSR16, RGH13b]. Second-Order [PKM10, PR14, RGH13b]. Secondary [GWZ13, MP13, WSL^{+}16]. Section [JML^{+}14]. Sectional [CAJ14]. SecYEG [BBBH10]. Sedative [TNZM14]. Seed [FRG^{+}17]. Segment [JCL10, THY16]. Segmentation [BJT^{+}10]. Segmented [ACLZ14]. SEIR [DLP12]. Seizure [SO19]. Select [bCH10]. Selected [HZ13, Li12a]. Selecting [LWC13, ZTC^{+}13]. Selection [ACLZ14, BCK19, BTI19, BR12a, BGSB19, CW14, CY17, CH12, CHZ16, CLZ10, CWZM18, DHM11, DG18, DVF11, Efr11b, Efr14a, FZY12, FLY12, FFJT16, FHS13, FJM19, GS10, GdMb13, GC11, HC15a, HFZ18, HLMH18, HCKS18, HWF15, HMW17, IY14, IKG^{+}10, JNR15, JHZ16, JR12, KHK19, KD14, LNP14, LZ10b, LL19a, LLZ18, Lin18, LLW14, LC14, MKG13, MW19, MG12, NC10, NSH19a, QDC19, QC15, RJ10, RSI14, RG14, Roc18, RR18, SFK12, SW17b, SPZ12, SL15, SWW14, TTLT16a, WBG^{+}18, WJHZ13, WYS19, WFB13, WSW17, WT10b, Wol11, WLNC14, WPS17b, WQxYW19, XXL^{+}19, XQZ10, YDZ16, YPOR18, ZWMC17, ZCJ^{+}17, ZQ12, ZSP13, ZD14]. Selection- [MG12]. Selection-Corrected [BTI19]. Selections [ZLT10]. Selective [DIvdB13]. Selectivity [CGL10]. Self [CGL10, CCSN11, FAIW^{+}11, FSS^{+}16, MSB^{+}11, Sha15, WLY^{+}14, ZL18]. Self-Controlled [FAIW^{+}11]. Self-Excited [WLY^{+}14]. Self-Exciting [FSS^{+}16, MSB^{+}11]. Self-Learning [CCSN11]. Self-Normalization [Sha15]. Self-Normalized [ZL18]. Self-Selectivity [CGL10]. Semi [MTY^{+}17]. Semi-Competing [MTY^{+}17]. Semicompeting [LDSH16]. Semigraphoid [LCZ14]. Semiparametric [CW17, CL10, CRL15, CLLL18, CKY15, CMER11, FLJL17, FPSE15b, GW15, GZCL19, Hos13, HFQ12, HY12, HC17, HYSM19, JS11, Ken15, KY11b, KY11a, KLSY13, KD14, LMR^{+}19, LZP10, LWC10, LZ11a, LLQ17, LWN^{+}18, LJGZH19, MZ12, MKG13, MMW^{+}17, MW10a, SPC^{+}18, SCM^{+}18, SH13, SLM11, SYS^{+}19, SS17, TLG16, TZL17, Wan17a, WGDG14b, WZL18, XZX10, XY12, ZL14, ZS15b, ZHS17, ZH18, ZTS11]. Semisupervised [Cul11]. Sense [PLGM11]. Sensing [Cre18a, NCB12]. Sensitive [LLZ18]. Sensitivity [FSMS17, FS17, GSH13, HRS14, HJRS10, HSR13, JPCD19, KYBB19, Ros10, Ros14, Ros15a, Sch11, SCHR13, SLC^{+}15, Zha19b]. Sensor [CXCR13, KLM13]. Sentiment [WSSQ16]. Sepsis [FMGS16]. Seq [MML^{+}13, LNN19, RSI14, DQZ18, HSTP15, KCP^{+}11]. Sequence [TZF^{+}15]. Sequences [Cra15]. Sequencing [BA16, DQZ18]. Sequential [LCT16, LJW13, LC14, PARS19, PBP19, RLP^{+}18, TTLT16a, Tra18, WBCD15, XMWT16b, ZR15]. Sequentially [WRL^{+}12a, XCQ15]. Serial [SMAC10]. Serially [LMR^{+}19]. Series [AB17b, ANH15, BS19, CKP19, CYZ14, CLLL18, CW19, DNFZ10, DHS11, FAIW^{+}11, FG12, FOvS10, KLN13, KMO15, KRS^{+}17, pKP12, LS18b, LK19, MBCM13, MT11, MM15, MJPW13b, MŞDN12, NL12, Pap10, PSY19, PPD15, RGL16, RWS12, SZ10, Sha15, SL14, TP16a, WAH19, Zha13b, ZWL^{+}15, ZL18]. Set [ACPLRC19, BML17, HZL^{+}17, SLW19, WLS17, WY19]. Set-Based [HZL^{+}17]. Set-Valued [SLW19]. Sets [AO17, BGMP12, CGW17, LRW13, SSS18, TYY^{+}19, Wil19]. Setting [GBW16, MB16]. Settings [CLX13, DG16a, JR12]. Several [GZZ19, MHC11]. Severe [FKH11, FMGS16]. Seychelles [XTR^{+}14]. Shahin [Mar19, PM19]. Shape [CDH11, DDH15, HSZ15, KKL15, KM14, SKBM17, VS15]. Shape-Constrained [CDH11]. Shapes [KSKD12]. Shared [BJT^{+}10]. Sharp [SPZ12]. Sharpen [MP13]. Sheet [CHAP16]. Shock [CMM^{+}13]. Short [BBS17, LZ10a, PCC12]. Short- [BBS17]. Short-Term [PCC12]. Shortcuts [BB10]. Shrinkage [BPPD15, CMW^{+}19, FFvSK17, GNL19, HCM13, HC15a, JLPZ17, TGHS18, WDSL10, YL13]. Side [SH13]. Sided [Lu16, LvdL18, PSY19, SM12a]. Sieve [BGC19, ZHS17]. Sigmoidal [ZA16]. Sign [FZW16, WFB13]. Signal [DKK16, HZT11, JZT19, LQ17, MM15, ZCC^{+}17]. Signaling [TMK^{+}12]. Signals [LSS^{+}19a, XBZ^{+}14]. Signatures [WZ17a]. Significance [LS17a, MML12, MG17a, MG17b, ZL11]. Significant [MQ15a]. Similarity [HZZ^{+}17]. Simple [LS15, Lit13, SW17b, TAGT14]. Simplex [BD12]. Simplicity [Lit13]. Simulated [LCL14, OP13]. Simulating [HLR15]. Simulation [GDLMVF19, Guh10, Zho14, Vat19]. Simulation-Based [GDLMVF19]. Simulations [MSW^{+}18]. Simultaneous [BBW14, CH12, GBW16, KCK^{+}15, KKC10, MVR17, MRB^{+}17, QDC19, WKKS12, ZC17b, ZCC^{+}17, ZYH14, ZSP13]. Simultaneously [JHH14, ZB13]. Single [BBBH10, CSSK16, LWC10, LHH17a, LPH^{+}17, MMW^{+}17]. Single-Index [LWC10, LHH17a, MMW^{+}17]. Single-Molecule [CSSK16]. Singular [DJP18, ZH19]. Sink [CCS14]. Size [CY17, CWH18, DDH15, ESR18, EG19, FSL17, MSZ10, VT17, ZFW17]. Size-Dependent [ZFW17]. Sizes [ZZWJ12]. Skagerrak [CCS14]. Skellam [KLL17]. Skewed [WBG^{+}18]. Skinny [NSH19a]. Slab [RG18, SFK12]. Sleep [MJPW13b, WBG^{+}18]. Sliced [LZL19, Qia12]. Slicing [JYL15, ZZF10]. Slope [COS^{+}10, BGSB19]. Small [ABvdW19, DHM11, DM15, JNR11, LZ10a, MTY^{+}17, NC10, PCB11, SM12a, TGHS18, TYY^{+}19]. Small-Angle [ABvdW19]. Small-Area [DHM11, LZ10a]. Small-Sample [SM12a]. Smoke [WLSA17]. Smoking [GDM17]. Smooth [Sar12, WPS17b, ZYH14]. Smoothed [CWH^{+}15, WMY15]. Smoother [SSL^{+}10]. Smoothing [CWPC13, DHS11, SLR^{+}15, TARS17, TKPS18, WPS17b]. Smoothly [GSDR19]. Smoothly-Changing [GSDR19]. SNAP [KPGJ12]. Snippet [DM18]. SNP [BML17]. SNP-Set [BML17]. SNPs [JHH14]. SNR [PT17]. Social [FTZ15, RSA16, SL14]. Soft [LZW11]. Solution [SH10]. Solutions [GC17, PPLS18]. Some [DVV14, GC17, KZCS16, RSI14, Ros11, Ros15b, RDHL19, WFDS19]. Sorbent [KKLL17]. Sound [Mar19, PM19, TPAC19a, TPAC19b]. Source [CCS14, CXCR13, FG13, PPD^{+}14]. Sources [CCMC16a, FPSE15b, HMW^{+}14, LST^{+}11, YZBE18]. Southeastern [FRG^{+}17]. Space [BGMP12, DLP12, FL13, HG10, KPBSK10, LCCG14, LC14, NTC13, SC15, ST17, SDWM18, WSG16, WG14, ZX14]. Space-Filling [ST17, ZX14]. Space-Time [BGMP12, HG10, WG14]. Spaces [BBS12, BCT18, GdMb13, LZ10b, VW15, YT17]. Span [EEdH19]. Sparse [BCK19, CL11a, CLL11, CL11b, CLX13, CCZF15, CH12, CHZ16, CZMHS17, FFS11, FMD14, FZ13, GBW16, HL14a, HMW18b, HDL^{+}16, JCL10, KD13, LT12, LCZ12, LL13, LZL19, NSB17, Per12, PKSR15, Plu14, QDC19, RLY14, SZLI15, WGDG14b, WLXL14, XC14, YP17, YL16, ZH19, ZCC^{+}17, ZYH14, ZBHD15, ZLL18]. Sparsely [RQJ15]. SparseNet [MFH11]. Sparsity [HZZ^{+}17, NSH^{+}19b, RG17b, YYY18, ZSY16]. Spatial [ASS^{+}15, BFWE10, BDIK12, BWH16, CHAP16, CXCR13, DGH^{+}10, FRG^{+}17, FMP18, Hah12, Han17b, HP10a, HW19, KC11, KJNW11, Kat17, KHG18, LS19b, LT14, Mar19, NCB12, PM19, RFD11, RBB13, SWSK19, SSS18, Tad10, TARS17, TPAC19a, TPAC19b, XXL^{+}19, YL13]. Spatially [DGH^{+}10, LHTB15, RPS19, ZBZ^{+}16, ZHM^{+}10, ZH18, ZFK14]. Spatio [BB14, CZMHS17, GBDL10, GDZ11, Han17b, HW10, KOL^{+}12, MMC^{+}12a, PBG16, TYY^{+}19, ZMW^{+}15]. Spatio-Dynamic [MMC^{+}12a]. Spatio-Spectral [KOL^{+}12]. Spatio-Temporal [BB14, CZMHS17, GBDL10, GDZ11, Han17b, HW10, PBG16, TYY^{+}19, ZMW^{+}15]. Spatiotemporal [BMW19, FF17a, LG14, RD12, SRR^{+}17]. Species [ACB^{+}14, BSMR12, BFT18]. Specific [BQWS17, CXT14, HLSZ14, HSTP15, MSK^{+}17, RSK17, SO19, SYS^{+}19]. Specificity [SLC^{+}15]. Spectra [ADR^{+}12, FOvS10]. Spectral [CKP19, KOL^{+}12, KRS^{+}17, LW10a, RWS12, STFP12, dCD14]. Speed [SGR10]. Speeding [QKVT19]. Sphere [FPLM18, He17, DPT19, PBG16]. Spherical [DPT14, RWKS14]. Spike [RG18, SFK12]. Spike-and-Slab [RG18, SFK12]. Spiked [YZBE18]. Spillover [VHJB13]. Splicing [HH12]. Spline [CMM^{+}13, CWPC13, CH19, GHJZ16, KKC10, PL19, Spi19, SLR^{+}15, YZ19, ZEL19a, ZEL19b]. Spline-Based [CMM^{+}13]. Spline-Lasso [GHJZ16]. Splines [FSB^{+}19]. Split [ZSL^{+}11]. SPReM [SZLI15]. Square [AGT14, LCL14]. Square-Root [LCL14]. Squared [ABvdW19, Dui16, PD14, WJHZ13, Zha05]. Squares [DGS19, KS11]. Stabilized [SQC16]. Stable [PRS10, Zub15]. Stage [BF18, BS13, EUW17, HCCZ16, LTC13, LCLZ18, SCG13]. Stamps [KPGJ12]. Standard [LHW^{+}13, MDR18, SFM18]. Standardization [GRR^{+}17, WTM19]. Stars [PAHJ11]. Start [MS18b]. State [DLP12, HMQA10, KPBSK10, LCCG14, NTC13, OCAG19]. State-Space [DLP12, KPBSK10, NTC13]. Statement [LS17a]. States [Lin13]. Static [CLGS18]. Stationarity [DPV11, Pap10]. Stationary [ANH15, DPV11, GLS11, HHY19, LS18b]. Statistic [YP17]. Statistical [BTI19, BKD^{+}17, Cai10, CDQ14, CLXY15, CMS17, CT10, CO10, Cre18a, Efr10b, Efr10c, GBDL10, HMQA10, Hel10, HP10a, HP10b, HW10, HSTP15, JLY19, JML^{+}14, KLCT11, KCP^{+}11, KSKD12, LS17a, LNA10a, LCZ14, Lit13, MMM11a, MMM11b, MHC13, MJPW13b, MG17a, MG17b, NCB12, Pea19, RSI14, Sah10, Sch10, Smi10, SL11, SQC16, SFM18, THY16, WSWT10, WZ10, Wes10, YJFL19, ZZLK15, ZW14, Kim19]. Statisticians [BKM17, LS17a]. Statistics [FG12, Gel11, LLX15, Mor15, Mor10, Nus18, Pan11, Rod13, Utt17, Ros14, Dav13]. Status [HH10]. Steady [HMQA10]. Step [LYH13]. Stepwise [DKK16]. Stick [DBNZ16, RDG10, SD14]. Stick-Breaking [RDG10, SD14]. Stochastic [BIL^{+}11, CDBG16, CYT12, GLS11, GSSVF13, HC15b, KLL17, LCS^{+}13, LCL14, STFP12, Tsi17, XWG19, ZTS11]. Stochastically [YOD11]. Stock [MM15]. Stocks [CLM18]. Straight [FG13]. Strains [CMS^{+}18]. Strata [BOSB16, NH11]. Strategies [GdMb13, RSI14]. Stratification [BG11, BOSB16, CTM10, DGM12, DGYZ11, EUW17, FMV16, LZ11b]. Stratified [CKY15]. Stream [CO10, HP10a, HP10b, Sah10]. Strengthened [ESR18]. Stringing [CCMW11]. Stroke [YPQ^{+}16]. Strong [CLZ10]. Stronger [BSLR10]. Structural [CYZ14, IR15, LQS17, LSS^{+}19b, PHP^{+}12, PPD15, SD17b, SWJ18, Tan10, WZL18, XAM^{+}14, YTL15, YMSC15, Zho11, Zho13, ZSP14]. Structure [CMER11, GBZ17, Tad13b, YL16, ZCL11, ZQ12]. Structured [ACCTW18, BKD^{+}17, CCZ16, FCDA15, FOvS10, LZ10b, Per12, SFK12, SH15, WMA^{+}12]. Structures [AB17b, HZZ^{+}17, LG13, RJ10, SZI^{+}12, XWK11, XS18, YZB^{+}16]. Student [Zaj12]. Studentized [Hah12]. Studies [ADH10, BML17, CZ11, CZ13, CG15, CPPW11, ESR18, FSMS17, FS17, FK14a, GDM17, GWZ13, HRS14, HZL^{+}17, HSR13, HGZ18, HDL^{+}16, IS12, JHH14, KWG15, LZ11b, LD15, LLX15, LL19b, MMC^{+}12a, Ros10, Ros11, Ros14, Ros15a, Ros15b, RLP^{+}18, TZL17, WZL19, WT13, YHD^{+}10, YLRH19, ZL14, ZL11, ZSR18, Zha19b, ZTS11, ZK17, Pea19]. Study [AVWW14, BSLR10, BSMR12, BH13, BR12b, CXT14, FMGS16, HZ10, yHQ12, JTF^{+}19, KHK19, KCZ^{+}13, LER^{+}12, LZNS17, MBCM13, PKSR15, SL14, TP16a, VHJB13, XAM^{+}14, YKC^{+}15, ZSL^{+}11, ZTC^{+}13, ZNSR12, Zub12, ZK17, DRC^{+}12, XTR^{+}14]. Study-Specific [CXT14]. Sturgeon [WH17]. Subcortical [SZI^{+}12]. Subdata [WYS19]. Subdistribution [BKRF19]. Subgroup [FSL17, HQT16, MH17, SH15]. Subgroups [EUW17]. Subject [HZ10]. Submatrices [XCC18]. Subpopulations [RLY14]. Subsampled [SVS16]. Subsampling [CDQ13, CDQ14, QKVT19, WZM18]. Subset [Li12a, LSY13]. Subsets [LRZ17]. Subspace [CZ14b]. Subspaces [PBD13]. Substitute [Bri17]. Subtypes [HDL^{+}16, LW19b, MZT^{+}17]. Sufficient [BF15, BDF16, CZ14b, FWYZ13, HC17, QDC19, YNLC19]. Sum [TGZ15]. Summarizing [FMC11, TMP18]. Summary [CCMC16a, HC15a, LHC19, LLX15]. Summary-Level [CCMC16a]. Sunter [SF13]. Supercomputers [SRR^{+}17]. Superior [Ros14]. Supersaturated [JM14]. Supervised [WK13]. Support [BWH16, CLX13, CWC^{+}14]. Sure [BFV16, PWL16, PWXZ19, JLPZ17, TR19, XKB12]. SURE-Type [JLPZ17]. Surface [MZ15, RLM13, SH10]. Surfaces [YZB^{+}16]. Surgeons [PKSR15]. Surgery [Zub12]. Surrogate [LZ18, MSW^{+}18]. Surveillance [CXCR13, LG14, RD12, TYY^{+}19]. Survey [BWH16, CMS17, JCRG17, NSK16, OKEK17, Per10, RGH13a, TARS17]. Surveys [BSK^{+}15, KBKS18, LBS13, LHD19, RW10]. Survival [CAJ14, CYT12, DTYG12, EUW17, HBHC12, HRZ17, HQT16, IKG^{+}10, LHTB15, LY16, LN14, LCLZ18, PCC12, PTC14, VS15, WY13, XZX10, XSWH17, YY16, ZH18, ZK12]. Survivor [KHK19]. Sweeping [PPD^{+}14]. Swiss [XAM^{+}14]. Switching [BR12b]. Syllables [MBCM13]. Symmetric [MKG13]. Symmetry [GA12]. Synchronicity [LSS^{+}19a]. Synchrony [SKS^{+}15]. Syndrome [TNZM14, SYS^{+}19]. Syntactic [GDCL11]. Syntax [SCM^{+}18]. Synthesis [DR10, JPCD19]. Synthetic [ADH10, RSK17]. System [BMMS17, BQWS17, CTM10, HMQA10, ZK17]. Systematic [SL14]. Systems [CB18b, HW10, Ion12, SF13, WQxYW19, Yu19].

Tables [CLA10, DD16, HS13, KD13, LLgX14]. Tail [MW17, Zho19]. Tailed [MBCM13, WLH12, ZL15b]. Tailored [CMS^{+}18]. Tails [Dup12, HL14b]. Taiwan [CWH^{+}15]. Tale [DD16, LHD19]. Tangential [FPLM18]. Target [ZTC^{+}13]. Targeted [CWC^{+}14, GY17]. Targets [KLM13]. Tau [JLZ14, ZLW10, ZJZ12]. Tavakoli [Mar19, PM19]. Taxic [MMC^{+}12a]. Taxonomic [WZ17a]. Technique [BIZ15, Ima11]. Techniques [LLLW11]. Temperature [Dup12, HCOW17, JR14, RLM13]. Temperatures [SRR^{+}17]. Template [BA16]. Template-Based [BA16]. Temporal [BB14, CZMHS17, GBDL10, GDZ11, Han17b, HW10, KD13, PBG16, TYY^{+}19, ZS15a, ZMW^{+}15]. Ten [Lit13]. Tensor [LZ17, PMZ19, SL19, YD16, ZLZ13]. Tensors [SDT10, YL13, ZBHD15]. Tent [Rod13]. Term [CMER11, PCC12]. Terminal [CW17, KNK^{+}18, SSZL12, SW17c, SHW17]. Test [BB10, Büh11, Cha17, CPPW11, CF17, CCS18, Düm11, GJ18a, GK11, GCBL15, Hah12, HZ10, HJRS10, HNW^{+}17, HZT11, LM11a, LM11b, LS10, LCG^{+}15, LLL^{+}10, MR11, MQ15a, Pol11, RGL16, Sam11, SW17b, WPL15, WN11, YC11, ZLW10, ZSR18, ZST12, ZSCM18]. Test-Control [HZ10]. Testing [BBBH10, BBW14, BML17, BCDS18, BGC19, CLX13, CL16a, CDH11, CW19, DHM11, DJP18, DM11, DZ15, DPV11, FKSZ19, FZ16, FG12, FH15, HRC12, HCKS18, KS15, LSS11, Lei14, LC10, LB17a, LN10, LHW^{+}13, Ma13, MHZ15, MB16, MAS14, PGY14, PR14, PSSS17, PY12, PD14, RPS19, RW13, SCG13, SD17b, SZ10, ST15, SW11, SM12b, VH15, WO19, WY19, XCC18, ZL18, Zha11, ZSS19, ZB18]. Tests [BKD^{+}17, BS19, CZZ10, CDQ13, CDQ14, CLQY16, CLA10, CS10, DT15, DR17, DGHM19, FZW16, GHK10, GZZ19, HZL^{+}17, HD10, JYL15, KLH11, LB17a, LHC19, LMR14, LL19b, Lu16, PARS19, PAHJ11, RLY14, RK11, SCG13, TLG16, VW11, WZL19, ZL11, ZC11, ZJZ12]. Text [AB17a, RSA16, Tad13a]. Their [Pre10]. Themselves [Ros15b]. Theorems [LD17]. Theoretically [UL17]. Theory [Dai19, EEdH19, HS13, HH13, PPLS18, Zho19]. Therapies [WFZ18]. Therapy [JPBM13, MSK^{+}17]. Thermal [HMQA10]. Thermodynamic [OPG16]. Threat [SL14]. Three [CLA10, XWK11]. Three-Dimensional [XWK11]. Three-Way [CLA10]. Threshold [LSS11, LZ11a, WLY^{+}14, YTL15]. Thresholded [FL13]. Thresholding [CL11a, CLL19, Sar12]. Thresholds [You19]. Throughput [BTI19, FCDA15, Zha11]. Tiers [MR15]. Tilt [HNW^{+}17]. Time [ATT19, AB17b, ANH15, BGMP12, BS19, BAWM18, BGK^{+}18, CKP19, CYZ14, CY17, CP12, CLLL18, CW19, CKY15, DNFZ10, DGM12, DHS11, FG12, FOvS10, GSH13, HG10, IS12, KLN13, KA18, KMO15, KRS^{+}17, pKP12, LS18b, LK19, LW19a, LZPH14, LCLZ18, MS18b, MBCM13, MT11, MM15, MJPW13b, NQS14, NL12, OCAG19, Pap10, PCC12, PGW^{+}18, PR14, PSY19, PHP^{+}12, PPD15, RHCT14, RGL16, RWS12, SZ10, Sha15, SY17, SL14, SS17, SW11, TP16a, WG14, WAH19, WT13, XAM^{+}14, XCH^{+}17, Zha13b, ZPlW13, ZWL^{+}15, ZC17a, ZL18, ZZWJ12, ZHS17]. Time-Average [CY17]. Time-Course [SW11]. Time-Dependent [LCLZ18, OCAG19, XAM^{+}14]. Time-Frequency [LK19]. Time-to-Event [BGK^{+}18, GSH13, IS12, LW19a, LZPH14, RHCT14]. Time-Varying [ATT19, BAWM18, CP12, DGM12, PR14, WT13, ZPlW13, ZC17a, ZZWJ12]. Times [CKY15, DZ19, HBHC12, MS18b, SSZL12, XMWT16b]. Ting [Dui16]. Tissue [FTM12]. Tobacco [ADH10]. Tolerance [SM12a]. Tools [MDR18]. Topic [AB17a, FLJL17]. Topological [HGK12]. Topology [Yu19]. Total [GBDL10, HLL10, WZ17b]. Toxicity [FK14a, LTJM15]. Toxicology [DP11]. Trace [WAH19, YDZ16]. Tracing [VT17]. Tracking [CG15, DLP12, KLM13, LPH^{+}17, PGW^{+}18, Tad10, WT13, Zaj12]. Tractable [RR18]. Trade [ESR18, KX17]. Trade-Off [ESR18, KX17]. Trading [BS19]. Traffic [GMS14]. Training [FMPR12, GK18]. Trait [JL15, TZF^{+}15]. Trait-Dependent [TZF^{+}15]. Traits [BFWE10, GWZ13, ZLW10]. Trajectories [CZW^{+}11, DM18, HLG14]. Trajectory [LWN^{+}18]. Transcriptomic [HDL^{+}16]. Transform [DV11]. Transformation [DTYG12, FF13, Gee14, KJ10, KLSY13, LZ10c, SC12, WL13, WW15, WT13, ZL14]. Transformation-Based [WW15]. Transformations [CDH12, FF13, RWKS14]. Transformed [LN10, SdCG^{+}15]. Transforms [TT12]. Transition [XMWT16b]. Transmission [Ken15, KWG15, SFM18]. Transnormal [FXZ17]. Transportation [CSSK16]. Treating [WFZ18]. Treatment [CH19, Cho17, DIvdB13, DFM19, Fra15, FL10, FH14, GW15, GSH13, GLS11, HRS14, HLG14, JTF^{+}19, KHK19, LS18a, LTJM15, LHD19, LHW^{+}13, LZS^{+}16, LZNS17, MTY^{+}17, MYT18, OR17, PL19, PTC14, RLY14, Sid13, Sob12, Spi19, SHM^{+}18, TNZM14, TAGT14, WA18, WRL^{+}12a, WLS17, WFZ18, WZSS18, XAM^{+}14, XMWT16b, YZ19, Zaj12, ZLDT18, ZZRK12, ZZLK15, ZMHKK17, ZEL19a, ZEL19b, ZZS19, ZNSR12]. Treatment-Naive [LZNS17]. Treatments [LDS10, SHM^{+}18]. Tree [BKD^{+}17, FOvS10, LN10, SBG^{+}16, WMA^{+}12]. Tree-Ring [SBG^{+}16]. Tree-Structured [BKD^{+}17, FOvS10, WMA^{+}12]. Tree-Transformed [LN10]. Treed [KKLL17]. Trees [LHTB15, Lin18, MW11, SDS19, TGP11, TE11, Wil19, ZK12, ZZK15]. Trend [GSDR19, PAHJ11, PY12]. Trends [CW19, CWH^{+}15, DLP12, MJG18]. Trial [HTGT13, LD15, LGY18, PTC14, VCBT19, WDSL10, WRL^{+}12a, WGM12]. Trials [BFWE10, DGM12, GSH13, JLTY14, KYBB19, LER^{+}12, LZS^{+}16, MHZ15, MZC16, MSM^{+}15, RLY14, VHJB13]. Trimming [DJP18]. Tropical [RLGL11]. Truncated [DGYZ11]. Tsunami [PPD^{+}14]. Tukey [XG17]. Tumor [XMY^{+}15]. Tuned [TR19]. Tuning [CH17, HWF15, RKFL19]. Tuning-Free [RKFL19]. Tunnel [LHH^{+}17b]. Tweedie [Efr11b]. Twin [LSS^{+}19b]. Two [BF18, BS13, CLX13, CDQ13, CDQ14, CF17, CCS18, DD16, FZW16, GCBL15, HCCZ16, HSZ15, LTJM15, LS10, LTC13, LHS18, LYH13, LCLZ18, MW11, RLY14, SCG13, SVTG17, SM12a, TZL17, ZL14]. Two-by-Two [DD16]. Two-Dimensional [HSZ15]. Two-Level [SVTG17]. Two-Phase [TZL17, ZL14]. Two-Sample [CLX13, CDQ13, CDQ14, CF17, CCS18, FZW16, GCBL15]. Two-Sided [SM12a]. Two-Stage [BF18, BS13, HCCZ16, LTC13, LCLZ18, SCG13]. Two-Step [LYH13]. Two-Way [LHS18]. Type [Dui16, JLPZ17, WFZ18, Zha05, Zho14, ZWZZ19].

U.K. [SQ10, WLSA17]. U.S. [CTM10, You19]. Ultra [CLLL18, CHZ16, FFS11, FMD14, LC14, WWL12]. Ultra-High [CLLL18, CHZ16, LC14, WWL12]. Ultra-High-Dimensional [FFS11, FMD14]. Ultrahigh [CFL18, CLZ15, HZ14, KXZ19, LHL15, LLW14, PWL16, QDC19, XC14, ZLLZ11]. Ultrahigh-Dimensional [CFL18, HZ14, LHL15, LLW14, PWL16, XC14, ZLLZ11]. Unbiased [QLL10, SDS19]. Uncertain [CB18b, DM15, FBM11, MHC11]. Uncertainty [BMMS17, SWSK19, SCK19, VCBT19, WRCG13, WLNC14, ZD14]. Uncovering [PS10]. Under-Reporting [SEdS19]. Undergoing [TNZM14]. Undermine [Ros15b]. Undernutrition [FPSE15b, WKKS12]. Understanding [LHS18, YPQ^{+}16]. Undirected [ZSP13, ZSP14]. Unequal [GNL19]. Unified [GWZ13, HZH15, KLSY13, LZW11, NSB17, ZH18]. Uniformly [GW15, KLH11, ZSS19]. Unifying [BSK^{+}15, HAME15, XSS11]. Unimodal [TEKM12]. Unit [Gee14, KC10, LHD19, WW15]. Units [FMP18, LSLR12]. Universal [BBS17]. Unknown [CXT14, GBP19, HW19, LW19b, MKG13, MS18b]. Unmeasured [GK18, HRS14, PSR16]. Unordered [BD12, LQ11]. Unreplicated [HFSZ19]. Unspecified [WfTQ12, ZL15b]. Unstructured [WSSQ16]. Unsupervised [ZL18]. Unusual [ZLP^{+}14]. Upscaling [BMMS17]. Upstream [SHM19]. Usage [SLC^{+}15]. Use [BCT18, DR10, HSTP15, SZ14, TEKM12, WFDS19]. Using [ACLZ14, BGH^{+}18, BJQ17, BGC19, BMW19, BH18, BG16, CAJ14, bCH10, CHAP16, CCMC16a, CLQY16, CZK17a, DHS11, DH13, DGH^{+}10, FLY12, FRG^{+}17, FTM12, FMV16, FSS^{+}16, FSB^{+}19, FMC11, GBDL10, GDZ11, Han17b, HZS18, HWF15, HYSM19, JT12, JLTY14, JKL11, KLM13, KSGB10, KC10, KSKD12, LNS10, LZ10a, LLX15, MS18c, MVR12, MRB12, MZ15, MLC^{+}16, MP13, MDR18, NSK16, PCB11, PPD^{+}14, Per10, Plu14, PSW13, RS10, RJ10, RHCT14, RD12, RLY14, RWKS14, SWSK19, SBG^{+}16, SM12a, SO11, SGR10, SMAC10, SHW17, SHM^{+}18, Tan10, TGHS18, THY16, VS15, WRL10, WJYJ13, WG14, WBCD15, WLS17, WSW17, XWK11, YZBE18, YZB^{+}16, ZSL^{+}11, ZWL^{+}15, ZZRK12, Zub12, HMW17, RLM13, WA18]. Utilities [MTY^{+}17]. Utility [LTJM15, SW17c]. Utilization [HZS18].

Vaccine [BGC19]. Valid [AGS12, BCK19]. Validated [CG15]. Validating [Pap10]. Validation [CL18, MHC13, MMNS11, MRB12, SHS10, TYY^{+}19]. Value [Ber17, CCL^{+}11a, CCL^{+}11b, EEdH19, GC17, God17, JPCD19, Laz11, LNA10b, LNM11, She11, Wu11, ZH19, Zha19b, ZZS19, LX19]. Valued [BWH16, LSZD17, SLW19, YPOR18]. Values [AEI18, Bri17, Cai10, Efr10b, Efr10c, Hel10, LS17a, MR17, NTC13, RDHL19, Sch10, Wes10, ZSS19]. Variability [HC15b, XKBS17]. Variable [Bla17, BR12a, CH12, CHZ16, CLZ10, CW12, CWZM18, GdMb13, GCNC14, GC11, HLMH18, IKG^{+}10, JHZ16, KA18, KXZ19, KD14, LZ10b, LL19a, Lin18, MLT17, MW19, PAHJ11, Per10, QDC19, RJ10, RSI14, RG14, Roc18, RR18, SZ14, SL15, SWW14, WBG^{+}18, WJHZ13, WSW17, Wol11, WQxYW19, XXL^{+}19, YDZ16, YPOR18, ZD14]. Variable-Domain [GCNC14]. Variables [AS15, CY16, bCH10, CRVF10, FBM11, FJM19, KZCS16, LS10, LFL15, MB16, PHSS15, PSW13, SLM11, SW14, SHM^{+}18, Tan10, WC18, WFDS19]. Variance [BOSB16, CFL18, DJP18, Dav12, FLL17, GSDR19, GSSVF13, HGK12, HD10, LPR12, PR14, RN16, SMRGG18]. Variants [JLZ14, MCIS19]. Variate [HFSZ19, HBI^{+}19]. Variates [LCT16]. Variation [DFM19, HLL10, JCL10, WZ17b]. Variational [BKM17, BM10, FOW11, HYSM19, WB19c]. Variations [Mar19, PM19, TPAC19a, TPAC19b]. Varying [ATT19, BAWM18, CP12, CHZ16, DGM12, FMD14, FZ16, HLMH18, HHY19, HY12, JWXJ13, LHH17a, LLW14, LJZ14, MS15, NSH^{+}19b, PR14, ŞM10, WT13, ZPlW13, ZC17a, ZZWJ12, ZFK14, ZWZZ19]. Varying-Coefficient [HLMH18, HHY19, LHH17a, LJZ14, ZWZZ19]. Varying-Sparsity [NSH^{+}19b]. Vast [FZY12, FLY12]. Vector [CWC^{+}14, DFRS14, FPLM18, FDPS10, GKM19, WPL15]. Vectors [GBW16, PGY14]. Velocity [ASS^{+}15]. Verbal [MLC^{+}16]. Versus [GK18, LER^{+}12, LZNS17, ZA16]. Very [bCH10]. Via [ASS^{+}15, CLL19, LX19, RMR19, SWW14, VRS19, BIL^{+}11, BT11a, BR12a, BS13, CW14, CSSK16, CMW^{+}19, CMPS17, DKK16, FFJT16, GMS14, GLLL15, HV11, HFZ18, HB14, JLZ14, JYL15, JHZ16, JTF^{+}19, JHH14, KJNW11, LNP14, LZZ12, LMZ18, LZL19, MLT17, MT17, MD19, MR17, PBD13, PD14, QP13, RG17b, SPU17, SD17b, SK12, SLR^{+}15, TWYZ11, Wan17a, WZ17b, ZLT10, ZD13]. Viable [WRL^{+}12a]. VIF [LFU11]. Violence [Sid13]. Violent [Tad10]. Viral [CZW^{+}11]. Virginia [XXL^{+}19].

Virtual [Wu15]. Virulent [CMS^{+}18]. Visual [BMW19, MHC13, SKS^{+}15]. Visualization [CGW17, XKBS17]. Vocalization [SCM^{+}18]. Volatility [ATT19, CLM18, CP12, CHP10, CDS11, FLY12, FK18, JS11, KLL17, SJ14, TWYZ11, TT12]. Volume [Ano10g, Ano11i, ACPLRC19, GMS14]. Vote [GRH10]. Vote-Choice [GRH10]. Voter [Bla17, PGW^{+}18]. Voting [SQ10, JML^{+}14]. Voucher [ZK17]. vs [Sch12].

W [CDQ14]. Wages [FMPR12, Sob12]. Wait [LHS18]. Waiting [LHS18]. Walks [Han17b]. Wang [AIP19]. Wanna [AR15b]. War [Fuk15]. Warming [MMM11a, MMM11b, SL11]. Water [COS^{+}10]. Wave [COS^{+}10]. Wavelet [FOvS10, GSSVF13, JK16, MS18a]. Wavelet-Domain [MS18a]. Wavelet-Variance-Based [GSSVF13]. Waves [Dup12]. Way [CLA10, LHS18, VW11]. Weather [BRGS10]. Weight [LZWZ11, WSG16]. Weighted [BCDS18, BKRF19, BGMP12, CCJ10, CCJ13a, CYC10, CZK17a, CCS18, FG12, MW10b, QZL^{+}10, Ros14, RLM13, WRL10, YP17, YLRH19, YL13, ZZRK12, ZMHKK17, ZSPL17]. Weighting [KC10, LMZ18, LM17, ST18]. Weights [AVWW14, IR15, SB18, Zub15]. Weyl [DRC^{+}12]. Where [Tsi17]. Whether [CCS14]. While [LS10]. White [YZB^{+}16]. Whittle [WX15]. Wide [BA16, ZL11]. Wild [pKP12, Pea19, Sha10]. Wildlife [BSK^{+}15]. Wind [HG10, JT12, LDGX15, LHH^{+}17b, SGR10, You19]. Window [PPD^{+}14]. Winsorizing [DJP18]. Winter [BRGS10]. Wireless [KLM13]. Withdraw [CGL10]. Within [KSGB10, NH11, YKC^{+}15, ZB13, ZNSR12]. Within-Study [YKC^{+}15]. Without [SH13]. Women [LZPH14, Pre10]. Won’t [GC17]. Word [KLCT11]. Work [GC17]. World [Pan11].

X [ABvdW19]. X-Ray [ABvdW19]. xiii [Tri19]. xvii [Pea19].

Year [Dav13]. Yields [YPOR18]. Yixin [AIP19].

Z [CDQ14]. Zero [KKL15]. Zero-Inflated [KKL15]. Zeros [HBHC12]. Zhang [Dui16]. Zhou [YZ19, Spi19]. Zipf [DO12].

References

Airoldi:2013:ELP

[AB13] Edoardo M. Airoldi and Alexander W. Blocker. Estimating latent processes on a network from indirect measurements. Journal of the American Statistical Association, 108(501):149–164, 2013. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Airoldi:2017:IET

[AB17a] Edoardo M. Airoldi and Jonathan M. Bischof. Improving and evaluating topic models and other models of text. Journal of the American Statistical Association, 111(516):1381–1403, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See comments [Tad17, GK17, Ble17] and rejoinder [Air17].

Ando:2017:CHN

[AB17b] Tomohiro Ando and Jushan Bai. Clustering huge number of financial time series: A panel data approach with highdimensional predictors and factor structures. Journal of the American Statistical Association, 112(519):1182–1198, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Alsaker:2019:MMS

[ABvdW19] Cody Alsaker, F. Jay Breidt, and Mark J. van der Woerd. Minimum mean squared error estimation of the radius of gyration in small-angle X-ray scattering experiments. Journal of the American Statistical Association, 114(525):39–47, 2019. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Airoldi:2014:GSS

[ACB^{+}14] Edoardo M. Airoldi, Thiago Costa, Federico Bassetti, Fabrizio Leisen, and Michele Guindani. Generalized species sampling priors with latent beta reinforcements. Journal of the American Statistical Association, 109(508):1466–1480, 2014. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ahn:2010:BII

[ACBK10] Kwang Woo Ahn, Kung-Sik Chan, Ying Bai, and Michael Kosoy. Bayesian inference with incomplete multinomial data: a problem in pathogen diversity. Journal of the American Statistical Association, 105(490):600–611, June 2010. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Arias-Castro:2018:DFD

[ACCTW18] Ery Arias-Castro, Rui M. Castro, Ervin Tánczos, and Meng Wang. Distribution-free detection of structured anomalies: Permutation and rank-based scans. Journal of the American Statistical Association, 113(522):789–801, 2018. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Alfons:2017:RMA

[ACF17] Andreas Alfons, Christophe Croux, and Peter Filzmoser. Robust maximum association estimators. Journal of the American Statistical Association, 112(517):436–445, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Aue:2014:SMS

[ACLZ14] Alexander Aue, Rex C. Y. Cheung, Thomas C. M. Lee, and Ming Zhong. Segmented model selection in quantile regression using the minimum description length principle. Journal of the American Statistical Association, 109(507):1241–1256, 2014. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Arias-Castro:2019:MEV

[ACPLRC19] Ery Arias-Castro, Beatriz Pateiro-López, and Alberto Rodríguez-Casal. Minimax estimation of the volume of a set under the rolling ball condition. Journal of the American Statistical Association, 114(527):1162–1173, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Antonelli:2019:DP

[AD19] Joseph Antonelli and Michael J. Daniels. Discussion of PEN-COMP. Journal of the American Statistical Association, 114 (525):24–27, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Abadie:2010:SCM

[ADH10] Alberto Abadie, Alexis Diamond, and Jens Hainmueller. Synthetic control methods for comparative case studies: Estimating the effect of california’s tobacco control program. Journal of the American Statistical Association, 105(490):493–505, June 2010. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Astle:2012:BMN

[ADR^{+}12] William Astle, Maria De Iorio, Sylvia Richardson, David Stephens, and Timothy Ebbels. A Bayesian model of NMR spectra for the deconvolution and quantification of metabolites in complex biological mixtures. Journal of the American Statistical Association, 107(500):1259–1271, 2012. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Athey:2018:EVN

[AEI18] Susan Athey, Dean Eckles, and Guido W. Imbens. Exact pvalues for network interference. Journal of the American Statistical Association, 113(521):230–240, 2018. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Apanasovich:2012:VMR

[AGS12] Tatiyana V. Apanasovich, Marc G. Genton, and Ying Sun. A valid Matérn class of cross-covariance functions for multivariate random fields with any number of components. Journal of the American Statistical Association, 107(497):180–193, 2012. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Allen:2014:GLS

[AGT14] Genevera I. Allen, Logan Grosenick, and Jonathan Taylor. A generalized least-square matrix decomposition. Journal of the American Statistical Association, 109(505):145–159, 2014. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Abadie:2012:MRM

[AI12] Alberto Abadie and Guido W. Imbens. A martingale representation for matching estimators. Journal of the American Statistical Association, 107(498):833–843, 2012. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ayra:2014:FFQ

[AIC14] Eduardo S. Ayra, David Ríos Insua, and Javier Cano. To fuel or not to fuel? Is that the question? Journal of the American Statistical Association, 109(506):465–476, 2014. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Athey:2019:CBM

[AIP19] Susan Athey, Guido W. Imbens, and Michael Pollmann. Comment on: “The Blessings of Multiple Causes” by Yixin Wang and David M. Blei. Journal of the American Statistical Association, 114(528):1602–1604, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [WB19a].

Airoldi:2017:R

[Air17] Edoardo M. Airoldi. Rejoinder. Journal of the American Statistical Association, 111(516):1410–1412, 2017. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [AB17a, Tad17].

Abadie:2014:IMM

[AIZ14] Alberto Abadie, Guido W. Imbens, and Fanyin Zheng. Inference for misspecified models with fixed regressors. Journal of the American Statistical Association, 109(508):1601–1614, 2014. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ando:2014:MAA

[AL14] Tomohiro Ando and Ker-Chau Li. A model-averaging approach for high-dimensional regression. Journal of the American Statistical Association, 109(505):254–265, 2014. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Almirall:2012:C

[ALM12] Daniel Almirall, Daniel J. Lizotte, and Susan A. Murphy. Comment. Journal of the American Statistical Association, 107(498): 509–512, 2012. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Aue:2015:PSF

[ANH15] Alexander Aue, Diogo Dubart Norinho, and Siegfried Hörmann. On the prediction of stationary functional time series. Journal of the American Statistical Association, 110(509):378–392, 2015. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:2010:EC

[Ano10a] Anonymous. 2010 editorial collaborators. Journal of the American Statistical Association, 105(492):1638–1643, December 2010. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:2010:BRa

[Ano10b] Anonymous. Book reviews. Journal of the American Statistical Association, 105(489):437–445, March 2010. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:2010:BRb

[Ano10c] Anonymous. Book reviews. Journal of the American Statistical Association, 105(490):873–879, June 2010. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:2010:BRc

[Ano10d] Anonymous. Book reviews. Journal of the American Statistical Association, 105(491):1276–1283, September 2010. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:2010:BRd

[Ano10e] Anonymous. Book reviews. Journal of the American Statistical Association, 105(492):1626–1631, December 2010. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:2010:C

[Ano10f] Anonymous. Correction. Journal of the American Statistical Association, 105(490):882, June 2010. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:2010:IV

[Ano10g] Anonymous. Index to volume 105 (2010). Journal of the American Statistical Association, 105(492):1644–1651, December 2010. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:2010:LEa

[Ano10h] Anonymous. Letters to the Editor. Journal of the American Statistical Association, 105(490):880–881, June 2010. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:2010:LEb

[Ano10i] Anonymous. Letters to the Editor. Journal of the American Statistical Association, 105(492):1632–1635, December 2010. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:2011:EC

[Ano11a] Anonymous. 2011 editorial collaborators. Journal of the American Statistical Association, 106(496):1649–1654, December 2011. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:2011:BRa

[Ano11b] Anonymous. Book reviews. Journal of the American Statistical Association, 106(493):375–382, March 2011. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:2011:BRb

[Ano11c] Anonymous. Book reviews. Journal of the American Statistical Association, 106(494):763–772, June 2011. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:2011:BRc

[Ano11d] Anonymous. Book reviews. Journal of the American Statistical Association, 106(495):1219–1224, September 2011. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:2011:BRd

[Ano11e] Anonymous. Book reviews. Journal of the American Statistical Association, 106(496):1637–1648, December 2011. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:2011:Ca

[Ano11f] Anonymous. Correction. Journal of the American Statistical Association, 106(493):384, March 2011. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:2011:Cb

[Ano11g] Anonymous. Correction. Journal of the American Statistical Association, 106(493):385, March 2011. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:2011:Cc

[Ano11h] Anonymous. Correction. Journal of the American Statistical Association, 106(494):773, June 2011. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:2011:IV

[Ano11i] Anonymous. Index to volume 106 (2011). Journal of the American Statistical Association, 106(496):1655–1661, December 2011. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:2011:LE

[Ano11j] Anonymous. Letter to the Editor. Journal of the American Statistical Association, 106(493):383, March 2011. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:2012:BRa

[Ano12a] Anonymous. Book reviews. Journal of the American Statistical Association, 107(497):429–438, 2012. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:2012:BRb

[Ano12b] Anonymous. Book reviews. Journal of the American Statistical Association, 107(498):845–853, 2012. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:2012:BRc

[Ano12c] Anonymous. Book reviews. Journal of the American Statistical Association, 107(499):1253–1256, 2012. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:2012:BRd

[Ano12d] Anonymous. Book reviews. Journal of the American Statistical Association, 107(500):1653–1655, 2012. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:2012:C

[Ano12e] Anonymous. Correction. Journal of the American Statistical Association, 107(499):1257, 2012. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:2012:EBE

[Ano12f] Anonymous. Editorial Board EOV. Journal of the American Statistical Association, 107(500):ebi, 2012. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:2012:EC

[Ano12g] Anonymous. Editorial collaborators. Journal of the American Statistical Association, 107(500):1657–1662, 2012. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:2012:Ea

[Ano12h] Anonymous. Erratum. Journal of the American Statistical Association, 107(499):1258, 2012. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:2012:Eb

[Ano12i] Anonymous. Erratum. Journal of the American Statistical Association, 107(500):1656, 2012. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:2013:BRI

[Ano13a] Anonymous. Book review index. Journal of the American Statistical Association, 108(501):353–358, 2013. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:2013:BRa

[Ano13b] Anonymous. Book reviews. Journal of the American Statistical Association, 108(502):750–754, 2013. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:2013:BRb

[Ano13c] Anonymous. Book reviews. Journal of the American Statistical Association, 108(503):1132–1137, 2013. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:2013:BRc

[Ano13d] Anonymous. Book reviews. Journal of the American Statistical Association, 108(504):1545–1554, 2013. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:2013:EBE

[Ano13e] Anonymous. Editorial Board EOV. Journal of the American Statistical Association, 108(504):ebi, 2013. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:2013:EC

[Ano13f] Anonymous. Editorial collaborators. Journal of the American Statistical Association, 108(504):1555–1561, 2013. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:2014:BRa

[Ano14a] Anonymous. Book reviews. Journal of the American Statistical Association, 109(505):448–454, 2014. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:2014:BRb

[Ano14b] Anonymous. Book reviews. Journal of the American Statistical Association, 109(506):864–870, 2014. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:2014:BRc

[Ano14c] Anonymous. Book reviews. Journal of the American Statistical Association, 109(507):1325–1337, 2014. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:2014:BRd

[Ano14d] Anonymous. Book reviews. Journal of the American Statistical Association, 109(508):1712–1720, 2014. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:2014:C

[Ano14e] Anonymous. Correction. Journal of the American Statistical Association, 109(508):1721, 2014. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:2014:EBE

[Ano14f] Anonymous. Editorial Board EOV. Journal of the American Statistical Association, 109(508):ebi, 2014. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:2014:EC

[Ano14g] Anonymous. Editorial collaborators. Journal of the American Statistical Association, 109(508):1722–1729, 2014. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:2015:BRa

[Ano15a] Anonymous. Book reviews. Journal of the American Statistical Association, 110(509):449–456, 2015. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:2015:BRb

[Ano15b] Anonymous. Book reviews. Journal of the American Statistical Association, 110(510):878–888, 2015. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:2015:BRc

[Ano15c] Anonymous. Book reviews. Journal of the American Statistical Association, 110(511):1320–1323, 2015. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:2015:BR

[Ano15d] Anonymous. Book reviews. Journal of the American Statistical Association, 110(512):1818–1824, 2015. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/01621459.2015.
1121043.

Anonymous:2015:EBE

[Ano15e] Anonymous. Editorial Board EOV. Journal of the American Statistical Association, 110(512):ebi, 2015. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/01621459.2015.
1121640.

Anonymous:2015:EC

[Ano15f] Anonymous. Editorial collaborators. Journal of the American Statistical Association, 110(512):1825–1831, 2015. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/01621459.2015.1121639.

Anonymous:2016:BRa

[Ano16a] Anonymous. Book reviews. Journal of the American Statistical Association, 111(513):438–445, 2016. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:2016:BRb

[Ano16b] Anonymous. Book reviews. Journal of the American Statistical Association, 111(514):912–919, 2016. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:2016:BR

[Ano16c] Anonymous. Book reviews. Journal of the American Statistical Association, 111(515):1362–1369, 2016. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:2017:BRa

[Ano17a] Anonymous. Book reviews. Journal of the American Statistical Association, 111(516):1840–1851, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:2017:BRb

[Ano17b] Anonymous. Book reviews. Journal of the American Statistical Association, 112(517):457–464, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:2017:BRc

[Ano17c] Anonymous. Book reviews. Journal of the American Statistical Association, 112(518):878–882, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:2017:BRd

[Ano17d] Anonymous. Book reviews. Journal of the American Statistical Association, 112(519):1370–1379, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:2017:BRe

[Ano17e] Anonymous. Book reviews. Journal of the American Statistical Association, 112(520):1771–1783, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:2017:Ca

[Ano17f] Anonymous. Correction. Journal of the American Statistical Association, 111(516):1852, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:2017:Cb

[Ano17g] Anonymous. Correction. Journal of the American Statistical Association, 112(517):465, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:2017:Cc

[Ano17h] Anonymous. Correction. Journal of the American Statistical Association, 112(517):466–469, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:2017:EBEa

[Ano17i] Anonymous. Editorial Board EOV. Journal of the American Statistical Association, 111(516):ebi, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:2017:EBEb

[Ano17j] Anonymous. Editorial Board EOV. Journal of the American Statistical Association, 112(520):ebi, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:2017:ECa

[Ano17k] Anonymous. Editorial collaborators. Journal of the American Statistical Association, 111(516):1853–1861, 2017. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:2017:ECb

[Ano17l] Anonymous. Editorial collaborators. Journal of the American Statistical Association, 112(520):1784–1791, 2017. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:2018:BRb

[Ano18a] Anonymous. Book reviews. Journal of the American Statistical Association, 113(521):480–485, 2018. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:2018:BRc

[Ano18b] Anonymous. Book reviews. Journal of the American Statistical Association, 113(522):948–953, 2018. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:2018:BRd

[Ano18c] Anonymous. Book reviews. Journal of the American Statistical Association, 113(523):1391–1394, 2018. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:2018:C

[Ano18d] Anonymous. Corrigendum. Journal of the American Statistical Association, 113(521):486, 2018. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:2018:E

[Ano18e] Anonymous. Erratum. Journal of the American Statistical Association, 113(521):487, 2018. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:2019:BR

[Ano19a] Anonymous. Book reviews. Journal of the American Statistical Association, 114(525):477–483, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:2019:Cb

[Ano19b] Anonymous. Correction. Journal of the American Statistical Association, 114(525):485, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:2019:Cc

[Ano19c] Anonymous. Correction. Journal of the American Statistical Association, 114(525):486, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:2019:Ca

[Ano19d] Anonymous. Corrigendum. Journal of the American Statistical Association, 114(525):484, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:2019:ECa

[Ano19e] Anonymous. Editorial collaborators. Journal of the American Statistical Association, 114(525):487–494, 2019. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anonymous:2019:ECb

[Ano19f] Anonymous. Editorial collaborators. Journal of the American Statistical Association, 114(528):W1930–W1938, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Adusumilli:2017:ELR

[AO17] Karun Adusumilli and Taisuke Otsu. Empirical likelihood for random sets. Journal of the American Statistical Association, 112(519):1064–1075, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Angrist:2015:R

[AR15a] Joshua D. Angrist and Miikka Rokkanen. Rejoinder. Journal of the American Statistical Association, 110(512):1348–1349, 2015. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/01621459.2015.1106189.

Angrist:2015:WGR

[AR15b] Joshua D. Angrist and Miikka Rokkanen. Wanna get away? Regression discontinuity estimation of exam school effects away from the cutoff. Journal of the American Statistical Association, 110(512):1331–1344, 2015. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/01621459.2015.
1012259.

Azriel:2015:EDL

[AS15] David Azriel and Armin Schwartzman. The empirical distribution of a large number of correlated normal variables. Journal of the American Statistical Association, 110(511):1217–1228, 2015. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Alimadad:2011:ORF

[ASB11] Azadeh Alimadad and Matias Salibian-Barrera. An outlierrobust fit for generalized additive models with applications to disease outbreak detection. Journal of the American Statistical Association, 106(494):719–731, June 2011. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ait-Sahalia:2017:ECD

[ASFL^{+}17] Yacine Aït-Sahalia, Jianqing Fan, Roger J. A. Laeven, Christina Dan Wang, and Xiye Yang. Estimation of the continuous and discontinuous leverage effects. Journal of the American Statistical Association, 112(520):1744–1758, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ait-Sahalia:2010:HFC

[ASFX10] Yacine Aït-Sahalia, Jianqing Fan, and Dacheng Xiu. Highfrequency covariance estimates with noisy and asynchronous financial data. Journal of the American Statistical Association, 105(492):1504–1517, December 2010. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Azzimonti:2015:BFV

[ASS^{+}15] Laura Azzimonti, Laura M. Sangalli, Piercesare Secchi, Maurizio Domanin, and Fabio Nobile. Blood flow velocity field estimation via spatial regression with PDE penalization. Journal of the American Statistical Association, 110(511):1057–1071, 2015. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Axeft-Sahalia:2019:PCA

[ASX19] Yacine Aït-Sahalia and Dacheng Xiu. Principal component analysis of high-frequency data. Journal of the American Statistical Association, 114(525):287–303, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Antal:2011:DBM

[AT11] Erika Antal and Yves Tillé. A direct bootstrap method for complex sampling designs from a finite population. Journal of the American Statistical Association, 106(494):534–543, June 2011. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Andersen:2019:TVP

[ATT19] Torben G. Andersen, Martin Thyrsgaard, and Viktor Todorov. Time-varying periodicity in intraday volatility. Journal of the American Statistical Association, 114(528):1695–1707, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Antoniano-Villalobos:2014:BNR

[AVWW14] Isadora Antoniano-Villalobos, Sara Wade, and Stephen G. Walker. A Bayesian nonparametric regression model with normalized weights: A study of hippocampal atrophy in Alzheimer’s disease. Journal of the American Statistical Association, 109 (506):477–490, 2014. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Agostinelli:2017:CRE

[AY17] Claudio Agostinelli and Víctor J. Yohai. Composite robust estimators for linear mixed models. Journal of the American Statistical Association, 111(516):1764–1774, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Blocker:2016:TBM

[BA16] Alexander W. Blocker and Edoardo M. Airoldi. Template-based models for genome-wide analysis of next-generation sequencing data at base-pair resolution. Journal of the American Statistical Association, 111(515):967–987, 2016. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Barney:2015:JBM

[BAA^{+}15] Bradley J. Barney, Federica Amici, Filippo Aureli, Josep Call, and Valen E. Johnson. Joint Bayesian modeling of binomial and rank data for primate cognition. Journal of the American Statistical Association, 110(510):573–582, 2015. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Bhaumik:2012:MAR

[BAN^{+}12] Dulal K. Bhaumik, Anup Amatya, Sharon-Lise T. Normand, Joel Greenhouse, Eloise Kaizar, Brian Neelon, and Robert D. Gibbons. Meta-analysis of rare binary adverse event data. Journal of the American Statistical Association, 107(498):555–567, 2012. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Boruvka:2018:ATV

[BAWM18] Audrey Boruvka, Daniel Almirall, Katie Witkiewitz, and Susan A. Murphy. Assessing time-varying causal effect moderation in mobile health. Journal of the American Statistical Association, 113(523):1112–1121, 2018. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Brannath:2010:SLC

[BB10] Werner Brannath and Frank Bretz. Shortcuts for locally consonant closed test procedures. Journal of the American Statistical Association, 105(490):660–669, June 2010. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Brynjarsdottir:2014:DRM

[BB14] Jenný Brynjarsdóttir and L. Mark Berliner. Dimension-reduced modeling of spatio-temporal processes. Journal of the American Statistical Association, 109(508):1647–1659, 2014. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Balabdaoui:2010:DSM

[BBBH10] Fadoua Balabdaoui, Kathrin Bissantz, Nicolai Bissantz, and Hajo Holzmann. Demonstrating single and multiple currents through the E. coli-SecYEG-Pore: Testing for the number of modes of noisy observations. Journal of the American Statistical Association, 105(489):136–146, March 2010. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Berger:2012:OPD

[BBS12] James O. Berger, Jose M. Bernardo, and Dongchu Sun. Objective priors for discrete parameter spaces. Journal of the American Statistical Association, 107(498):636–648, 2012. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Bagchi:2017:IMF

[BBS17] Pramita Bagchi, Moulinath Banerjee, and Stilian A. Stoev. Inference for monotone functions under short- and long-range dependence: Confidence intervals and new universal limits. Journal of the American Statistical Association, 111(516):1634–1647, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Bar:2014:BMS

[BBW14] Haim Y. Bar, James G. Booth, and Martin T. Wells. A bivariate model for simultaneous testing in bioinformatics data. Journal of the American Statistical Association, 109(506):537–547, 2014. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Bien:2016:CBC

[BBX16] Jacob Bien, Florentina Bunea, and Luo Xiao. Convex banding of the covariance matrix. Journal of the American Statistical Association, 111(514):834–??, 2016. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Barthelme:2014:EPL

[BC14] Simon Barthelmé and Nicolas Chopin. Expectation propagation for likelihood-free inference. Journal of the American Statistical Association, 109(505):315–333, 2014. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Belloni:2015:C

[BC15] Alexandre Belloni and Victor Chernozhukov. Comment. Journal of the American Statistical Association, 110(512):1449–1451, 2015. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/01621459.2015.1098545.

Brunel:2014:PEO

[BCdB14] Nicolas J-B. Brunel, Quentin Clairon, and Florence d’Alché Buc. Parametric estimation of ordinary differential equations with orthogonality conditions. Journal of the American Statistical Association, 109(505):173–185, 2014. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Basu:2018:WFD

[BCDS18] Pallavi Basu, T. Tony Cai, Kiranmoy Das, and Wenguang Sun. Weighted false discovery rate control in large-scale multiple testing. Journal of the American Statistical Association, 113(523): 1172–1183, 2018. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chan:2010:UEM

[bCH10] Yao ban Chan and Peter Hall. Using evidence of mixed populations to select variables for clustering very high-dimensional data. Journal of the American Statistical Association, 105(490): 798–809, June 2010. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Belloni:2019:VPS

[BCK19] Alexandre Belloni, Victor Chernozhukov, and Kengo Kato. Valid post-selection inference in high-dimensional approximately sparse quantile regression models. Journal of the American Statistical Association, 114(526):749–758, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Bassetti:2018:BNC

[BCR18] Federico Bassetti, Roberto Casarin, and Francesco Ravazzolo. Bayesian nonparametric calibration and combination of predictive distributions. Journal of the American Statistical Association, 113(522):675–685, 2018. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Bugni:2018:IUC

[BCS18] Federico A. Bugni, Ivan A. Canay, and Azeem M. Shaikh. Inference under covariate-adaptive randomization. Journal of the American Statistical Association, 113(524):1784–1796, 2018. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Berrendero:2018:URK

[BCT18] José R. Berrendero, Antonio Cuevas, and José L. Torrecilla. On the use of reproducing kernel Hilbert spaces in functional classification. Journal of the American Statistical Association, 113(523):1210–1218, 2018. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Bouchard-Cxftxe:2018:BPS

[BCVD18] Alexandre Bouchard-Côté, Sebastian J. Vollmer, and Arnaud Doucet. The bouncy particle sampler: A nonreversible rejectionfree Markov Chain Monte Carlo method. Journal of the American Statistical Association, 113(522):855–867, 2018. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Bhattacharya:2012:SFM

[BD12] Anirban Bhattacharya and David B. Dunson. Simplex factor models for multivariate unordered categorical data. Journal of the American Statistical Association, 107(497):362–377, 2012. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Bura:2016:SRR

[BDF16] Efstathia Bura, Sabrina Duarte, and Liliana Forzani. Sufficient reductions in regressions with exponential family inverse predictors. Journal of the American Statistical Association, 111(515): 1313–1329, 2016. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Barrios:2012:CSC

[BDIK12] Thomas Barrios, Rebecca Diamond, Guido W. Imbens, and Michal Kolesár. Clustering, spatial correlations, and randomization inference. Journal of the American Statistical Association, 107(498):578–591, 2012. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Berry:2017:V

[Ber17] Donald Berry. A p-value to die for. Journal of the American Statistical Association, 112(519):895–897, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Bura:2015:SRR

[BF15] Efstathia Bura and Liliana Forzani. Sufficient reductions in regressions with elliptically contoured inverse predictors. Journal of the American Statistical Association, 110(509):420–434, 2015. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Basse:2018:ATS

[BF18] Guillaume Basse and Avi Feller. Analyzing two-stage experiments in the presence of interference. Journal of the American Statistical Association, 113(521):41–55, 2018. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Bao:2012:CMM

[BFH12] Le Bao, John Fricks, and Murali Haran. Comment on the mechanistic modeling and inference for cell motility by Manolopoulou et al. Journal of the American Statistical Association, 107(499): 869–871, 2012. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Battiston:2018:MAB

[BFT18] Marco Battiston, Stefano Favaro, and Yee Whye Teh. Multiarmed bandit for species discovery: A Bayesian nonparametric approach. Journal of the American Statistical Association, 113 (521):455–466, 2018. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Barut:2016:CSI

[BFV16] Emre Barut, Jianqing Fan, and Anneleen Verhasselt. Conditional sure independence screening. Journal of the American Statistical Association, 111(515):1266–1277, 2016. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Banerjee:2010:HSP

[BFWE10] Sudipto Banerjee, Andrew O. Finley, Patrik Waldmann, and Tore Ericsson. Hierarchical spatial process models for multiple traits in large genetic trials. Journal of the American Statistical Association, 105(490):506–521, June 2010. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Bartolucci:2011:MPC

[BG11] Francesco Bartolucci and Leonardo Grilli. Modeling partial compliance through copulas in a principal stratification framework. Journal of the American Statistical Association, 106(494):469– 479, June 2011. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Bunch:2016:AOI

[BG16] Pete Bunch and Simon Godsill. Approximations of the optimal importance density using Gaussian particle flow importance sampling. Journal of the American Statistical Association, 111 (514):748–??, 2016. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Brannath:2012:PFC

[BGB12] W. Brannath, G. Gutjahr, and P. Bauer. Probabilistic foundation of confirmatory adaptive designs. Journal of the American Statistical Association, 107(498):824–832, 2012. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Benkeser:2019:ETV

[BGC19] David Benkeser, Peter B. Gilbert, and Marco Carone. Estimating and testing vaccine sieve effects using machine learning. Journal of the American Statistical Association, 114(527):1038–1049, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Backenroth:2018:MML

[BGH^{+}18] Daniel Backenroth, Jeff Goldsmith, Michelle D. Harran, Juan C. Cortes, John W. Krakauer, and Tomoko Kitago. Modeling motor learning using heteroscedastic functional principal components analysis. Journal of the American Statistical Association, 113(523):1003–1015, 2018. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Braun:2018:NAM

[BGK^{+}18] Danielle Braun, Malka Gorfine, Hormuzd A. Katki, Argyrios Ziogas, and Giovanni Parmigiani. Nonparametric adjustment for measurement error in time-to-event data: Application to risk prediction models. Journal of the American Statistical Association, 113(521):14–25, 2018. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Bevilacqua:2012:ESS

[BGMP12] Moreno Bevilacqua, Carlo Gaetan, Jorge Mateu, and Emilio Porcu. Estimating space and space-time covariance functions for large data sets: a weighted composite likelihood approach. Journal of the American Statistical Association, 107(497):268–280, 2012. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Brown:2013:PCD

[BGR13] Lawrence D. Brown, Eitan Greenshtein, and Ya’cov Ritov. The Poisson compound decision problem revisited. Journal of the American Statistical Association, 108(502):741–749, 2013. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Brzyski:2019:GSA

[BGSB19] Damian Brzyski, Alexej Gossmann, Weijie Su, and Małgorzata Bogdan. Group SLOPE — adaptive selection of groups of predictors. Journal of the American Statistical Association, 114(525): 419–433, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Bogomolov:2013:DFR

[BH13] Marina Bogomolov and Ruth Heller. Discovering findings that replicate from a primary study of high dimension to a follow-up study. Journal of the American Statistical Association, 108(504): 1480–1492, 2013. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Bolton:2018:MFD

[BH18] Alexander D. Bolton and Nicholas A. Heard. Malware family discovery using reversible jump MCMC sampling of regimes. Journal of the American Statistical Association, 113(524):1490–1502, 2018. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Bien:2019:GGB

[Bie19] Jacob Bien. Graph-guided banding of the covariance matrix. Journal of the American Statistical Association, 114(526):782– 792, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Bhadra:2011:MNI

[BIL^{+}11] Anindya Bhadra, Edward L. Ionides, Karina Laneri, Mercedes Pascual, Menno Bouma, and Ramesh C. Dhiman. Malaria in Northwest India: Data analysis via partially observed stochastic differential equation models driven by Lévy noise. Journal of the American Statistical Association, 106(494):440–451, June 2011. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Blair:2015:DAR

[BIZ15] Graeme Blair, Kosuke Imai, and Yang-Yang Zhou. Design and analysis of the randomized response technique. Journal of the American Statistical Association, 110(511):1304–1319, 2015. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ba:2011:MLD

[BJ11] Shan Ba and V. Roshan Joseph. Multi-layer designs for computer experiments. Journal of the American Statistical Association, 106(495):1139–1149, September 2011. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Brown:2016:C

[BJ16] Lawrence D. Brown and Kory D. Johnson. Comment. Journal of the American Statistical Association, 111(514):614–??, 2016. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Barrientos:2017:FNR

[BJQ17] Andrés F. Barrientos, Alejandro Jara, and Fernando A. Quintana. Fully nonparametric regression for bounded data using dependent Bernstein polynomials. Journal of the American Statistical Association, 112(518):806–825, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Baladandayuthapani:2010:BRS

[BJT^{+}10] Veerabhadran Baladandayuthapani, Yuan Ji, Rajesh Talluri, Luis E. Nieto-Barajas, and Jeffrey S. Morris. Bayesian random segmentation models to identify shared copy number aberrations for array CGH data. Journal of the American Statistical Association, 105(492):1358–1375, December 2010. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Bharath:2017:STL

[BKD^{+}17] Karthik Bharath, Prabhanjan Kambadur, Dipak. K. Dey, Arvind Rao, and Veerabhadran Baladandayuthapani. Statistical tests for large tree-structured data. Journal of the American Statistical Association, 112(520):1733–1743, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Blei:2017:VIR

[BKM17] David M. Blei, Alp Kucukelbir, and Jon D. McAuliffe. Variational inference: A review for statisticians. Journal of the American Statistical Association, 112(518):859–877, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Bellach:2019:WNS

[BKRF19] Anna Bellach, Michael R. Kosorok, Ludger Rüschendorf, and Jason P. Fine. Weighted NPMLE for the subdistribution of a competing risk. Journal of the American Statistical Association, 114(525):259–270, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Bartolucci:2016:PLI

[BL16] Francesco Bartolucci and Monia Lupparelli. Pairwise likelihood inference for nested hidden Markov chain models for multilevel longitudinal data. Journal of the American Statistical Association, 111(513):216–228, 2016. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Blackwell:2017:IVM

[Bla17] Matthew Blackwell. Instrumental variable methods for conditional effects and causal interaction in voter mobilization experiments. Journal of the American Statistical Association, 112(518): 590–599, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Blei:2013:C

[Ble13] David M. Blei. Comment. Journal of the American Statistical Association, 108(503):771–772, 2013. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [Tad13a].

Blei:2017:C

[Ble17] David M. Blei. Comment. Journal of the American Statistical Association, 111(516):1408–1410, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [AB17a].

Blum:2010:ABC

[Blu10] Michael G. B. Blum. Approximate Bayesian computation: a nonparametric perspective. Journal of the American Statistical Association, 105(491):1178–1187, September 2010. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Braun:2010:VIL

[BM10] Michael Braun and Jon McAuliffe. Variational inference for large-scale models of discrete choice. Journal of the American Statistical Association, 105(489):324–335, March 2010. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Brown:2015:C

[BM15] Lawrence D. Brown and Daniel McCarthy. Comment. Journal of the American Statistical Association, 110(512):1446–1449, 2015. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/01621459.2015.1099536.

Barnett:2017:GHC

[BML17] Ian Barnett, Rajarshi Mukherjee, and Xihong Lin. The generalized higher criticism for testing SNP-set effects in genetic association studies. Journal of the American Statistical Association, 112(517):64–76, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Bhat:2017:UUD

[BMMS17] K. Sham Bhat, David S. Mebane, Priyadarshi Mahapatra, and Curtis B. Storlie. Upscaling uncertainty with dynamic discrepancy for a multi-scale carbon capture system. Journal of the American Statistical Association, 112(520):1453–1467, 2017.

CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Berchuck:2019:DGP

[BMW19] Samuel I. Berchuck, Jean-Claude Mwanza, and Joshua L. Warren. Diagnosing glaucoma progression with visual field data using a spatiotemporal boundary detection method. Journal of the American Statistical Association, 114(527):1063–1074, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Breidt:2016:NVE

[BOSB16] F. Jay Breidt, Jean D. Opsomer, and Ismael Sanchez-Borrego. Nonparametric variance estimation under fine stratification: An alternative to collapsed strata. Journal of the American Statistical Association, 111(514):822–??, 2016. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Bhattacharya:2015:DLP

[BPPD15] Anirban Bhattacharya, Debdeep Pati, Natesh S. Pillai, and David B. Dunson. Dirichlet–Laplace priors for optimal shrinkage. Journal of the American Statistical Association, 110(512): 1479–1490, 2015. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/01621459.2014.960967.

Bi:2017:GSR

[BQWS17] Xuan Bi, Annie Qu, Junhui Wang, and Xiaotong Shen. A groupspecific recommender system. Journal of the American Statistical Association, 112(519):1344–1353, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Bondell:2012:CHD

[BR12a] Howard D. Bondell and Brian J. Reich. Consistent highdimensional Bayesian variable selection via penalized credible regions. Journal of the American Statistical Association, 107(500): 1610–1624, 2012. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Burgette:2012:NBM

[BR12b] Lane F. Burgette and Jerome P. Reiter. Nonparametric Bayesian multiple imputation for missing data due to mid-study switching of measurement methods. Journal of the American Statistical Association, 107(498):439–449, 2012. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Berrocal:2010:PWF

[BRGS10] Veronica J. Berrocal, Adrian E. Raftery, Tilmann Gneiting, and Richard C. Steed. Probabilistic weather forecasting for winter road maintenance. Journal of the American Statistical Association, 105(490):522–537, June 2010. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Briggs:2017:SV

[Bri17] William M. Briggs. The substitute for p-values. Journal of the American Statistical Association, 112(519):897–898, 2017. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Broderick:2017:CNB

[Bro17] Tamara Broderick. Comment: Nonparametric Bayes modeling of populations of networks. Journal of the American Statistical Association, 112(520):1534–1537, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Bondell:2013:ERR

[BS13] Howard D. Bondell and Leonard A. Stefanski. Efficient robust regression via two-stage generalized empirical likelihood. Journal of the American Statistical Association, 108(502):644–655, 2013. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Bojinov:2019:TSE

[BS19] Iavor Bojinov and Neil Shephard. Time series experiments and causal estimands: Exact randomization tests and trading. Journal of the American Statistical Association, 114(528):1665–1682, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Boente:2015:EFP

[BSB15] Graciela Boente and Matías Salibian-Barrera. S-estimators for functional principal component analysis. Journal of the American Statistical Association, 110(511):1100–1111, 2015. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Borchers:2015:UMC

[BSK^{+}15] D. L. Borchers, B. C. Stevenson, D. Kidney, L. Thomas, and T. A. Marques. A unifying model for capture-recapture and distance sampling surveys of wildlife populations. Journal of the American Statistical Association, 110(509):195–204, 2015. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Baiocchi:2010:BSI

[BSLR10] Mike Baiocchi, Dylan S. Small, Scott Lorch, and Paul R. Rosenbaum. Building a stronger instrument in an observational study of perinatal care for premature infants. Journal of the American Statistical Association, 105(492):1285–1296, December 2010. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Balderama:2012:ABM

[BSMR12] Earvin Balderama, Frederic Paik Schoenberg, Erin Murray, and Philip W. Rundel. Application of branching models in the study of invasive species. Journal of the American Statistical Association, 107(498):467–476, 2012. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Bornn:2012:MNP

[BSZ12] Luke Bornn, Gavin Shaddick, and James V. Zidek. Modeling nonstationary processes through dimension expansion. Journal of the American Statistical Association, 107(497):281–289, 2012. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Bien:2011:HCP

[BT11a] Jacob Bien and Robert Tibshirani. Hierarchical clustering with prototypes via minimax linkage. Journal of the American Statistical Association, 106(495):1075–1084, September 2011. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Breitung:2011:GED

[BT11b] Jörg Breitung and Jörn Tenhofen. GLS estimation of dynamic factor models. Journal of the American Statistical Association, 106(495):1150–1166, September 2011. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Bennett:2016:GPM

[BT16] Christopher J. Bennett and Brennan S. Thompson. Graphical procedures for multiple comparisons under general dependence. Journal of the American Statistical Association, 111(515):1278– 1288, 2016. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Bilder:2010:IR

[BTC10] Christopher R. Bilder, Joshua M. Tebbs, and Peng Chen. Informative retesting. Journal of the American Statistical Association, 105(491):942–955, September 2010. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Benjamini:2019:SCS

[BTI19] Yuval Benjamini, Jonathan Taylor, and Rafael A. Irizarry. Selection-corrected statistical inference for region detection with high-throughput assays. Journal of the American Statistical Association, 114(527):1351–1365, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Buhlmann:2011:CAC

[Büh11] Peter Bühlmann. Comment: “Adaptive Confidence Intervals for the Test Error in Classification”. Journal of the American Statistical Association, 106(495):916–918, September 2011. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Barut:2015:C

[BW15] Emre Barut and Huixia Judy Wang. Comment. Journal of the American Statistical Association, 110(512):1442–1445, 2015. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/01621459.2015.1100619.

Bradley:2016:BSC

[BWH16] Jonathan R. Bradley, Christopher K. Wikle, and Scott H. Holan. Bayesian spatial change of support for count-valued survey data with application to the American community survey. Journal of the American Statistical Association, 111(514):472–??, 2016. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Bergesio:2011:PEG

[BY11] Andrea Bergesio and Victor J. Yohai. Projection estimators for generalized linear models. Journal of the American Statistical Association, 106(494):661–671, June 2011. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Cai:2010:CCV

[Cai10] T. Tony Cai. Comment: “Correlated z-Values and the Accuracy of Large-Scale Statistical Estimates”. Journal of the American Statistical Association, 105(491):1055–1056, September 2010. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Carone:2014:ELR

[CAJ14] Marco Carone, Masoud Asgharian, and Nicholas P. Jewell. Estimating the lifetime risk of dementia in the Canadian elderly population using cross-sectional cohort survival data. Journal of the American Statistical Association, 109(505):24–35, 2014. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chevallier:2018:C

[CB18a] Frédéric Chevallier and François-Marie Bréon. Comment. Journal of the American Statistical Association, 113(521):173–175, 2018. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Clairon:2018:OCA

[CB18b] Quentin Clairon and Nicolas J.-B. Brunel. Optimal control and additive perturbations help in estimating ill-posed and uncertain dynamical systems. Journal of the American Statistical Association, 113(523):1195–1209, 2018. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Calonico:2018:EBE

[CCF18] Sebastian Calonico, Matias D. Cattaneo, and Max H. Farrell. On the effect of bias estimation on coverage accuracy in nonparametric inference. Journal of the American Statistical Association, 113(522):767–779, 2018. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Cattaneo:2010:RDD

[CCJ10] Matias D. Cattaneo, Richard K. Crump, and Michael Jansson. Robust data-driven inference for density-weighted average derivatives. Journal of the American Statistical Association, 105(491):1070–1083, September 2010. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Cattaneo:2013:GJE

[CCJ13a] Matias D. Cattaneo, Richard K. Crump, and Michael Jansson. Generalized jackknife estimators of weighted average derivatives. Journal of the American Statistical Association, 108(504):1243– 1256, 2013. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See comments [Zen13, Det13, Mam13, Che13] and reply [CCJ13b].

Cattaneo:2013:R

[CCJ13b] Matias D. Cattaneo, Richard K. Crump, and Michael Jansson. Rejoinder. Journal of the American Statistical Association, 108 (504):1265–1268, 2013. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [CCJ13a].

Crainiceanu:2011:PVD

[CCL^{+}11a] Ciprian M. Crainiceanu, Brian S. Caffo, Sheng Luo, Vadim M. Zipunnikov, and Naresh M. Punjabi. Population value decomposition, a framework for the analysis of image populations. Journal of the American Statistical Association, 106(495):775–790, September 2011. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Crainiceanu:2011:RPV

[CCL^{+}11b] Ciprian M. Crainiceanu, Brian S. Caffo, Sheng Luo, Vadim M. Zipunnikov, and Naresh M. Punjabi. Rejoinder: “Population Value Decomposition, a Framework for the Analysis of Image Populations”. Journal of the American Statistical Association, 106(495):803–806, September 2011. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chatterjee:2016:CML

[CCMC16a] Nilanjan Chatterjee, Yi-Hau Chen, Paige Maas, and Raymond J. Carroll. Constrained maximum likelihood estimation for model calibration using summary-level information from external big data sources. Journal of the American Statistical Association, 111(513):107–117, 2016. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chatterjee:2016:R

[CCMC16b] Nilanjan Chatterjee, Yi-Hau Chen, Paige Maas, and Raymond J. Carroll. Rejoinder. Journal of the American Statistical Association, 111(513):130–131, 2016. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chen:2011:SHD

[CCMW11] Kun Chen, Kehui Chen, Hans-Georg Müller, and Jane-Ling Wang. Stringing high-dimensional data for functional analysis. Journal of the American Statistical Association, 106(493):275– 284, March 2011. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Calvet:2015:RF

[CCR15] Laurent E. Calvet, Veronika Czellar, and Elvezio Ronchetti. Robust filtering. Journal of the American Statistical Association, 110(512):1591–1606, 2015. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/01621459.2014.983520.

Chen:2014:SSR

[CCS14] Kun Chen, Kung-Sik Chan, and Nils Chr. Stenseth. Source– sink reconstruction through regularized multicomponent regression analysis-with application to assessing whether North Sea cod larvae contributed to local fjord cod in Skagerrak. Journal of the American Statistical Association, 109(506):560–573, 2014. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chen:2018:WEC

[CCS18] Hao Chen, Xu Chen, and Yi Su. A weighted edge-count two-sample test for multivariate and object data. Journal of the American Statistical Association, 113(523):1146–1155, 2018. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chen:2011:DCS

[CCSN11] Aiyou Chen, Jin Cao, Larry Shepp, and Tuan Nguyen. Distinct counting with a self-learning bitmap. Journal of the American Statistical Association, 106(495):879–890, September 2011. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Calonico:2015:ODD

[CCT15] Sebastian Calonico, Matias D. Cattaneo, and Rocío Titiunik. Optimal data-driven regression discontinuity plots. Journal of the American Statistical Association, 110(512):1753–1769, 2015. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/01621459.2015.1017578.

Cai:2016:SMC

[CCZ16] Tianxi Cai, T. Tony Cai, and Anru Zhang. Structured matrix completion with applications to genomic data integration. Journal of the American Statistical Association, 111(514):621– ??, 2016. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Cao:2015:APH

[CCZF15] Hongyuan Cao, Mathew M. Churpek, Donglin Zeng, and Jason P. Fine. Analysis of the proportional hazards model with sparse longitudinal covariates. Journal of the American Statistical Association, 110(511):1187–1196, 2015. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Canale:2011:BKM

[CD11] Antonio Canale and David B. Dunson. Bayesian kernel mixtures for counts. Journal of the American Statistical Association, 106 (496):1528–1539, December 2011. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Crane:2018:EEM

[CD18] Harry Crane and Walter Dempsey. Edge exchangeable models for interaction networks. Journal of the American Statistical Association, 113(523):1311–1326, 2018. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Cervone:2016:MSP

[CDBG16] Daniel Cervone, Alex D’Amour, Luke Bornn, and Kirk Goldsberry. A multiresolution stochastic process model for predicting basketball possession outcomes. Journal of the American Statistical Association, 111(514):585–??, 2016. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Carroll:2011:TES

[CDH11] Raymond J. Carroll, Aurore Delaigle, and Peter Hall. Testing and estimating shape-constrained nonparametric density and regression in the presence of measurement error. Journal of the American Statistical Association, 106(493):191–202, March 2011. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Carroll:2012:DWC

[CDH12] Raymond Carroll, Aurore Delaigle, and Peter Hall. Deconvolution when classifying noisy data involving transformations. Journal of the American Statistical Association, 107(499):1166–1177, 2012. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chen:2013:ESI

[CDQ13] Lisha Chen, Winston Wei Dou, and Zhihua Qiao. Ensemble subsampling for imbalanced multivariate two-sample tests. Journal of the American Statistical Association, 108(504):1308–1323, 2013. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See erratum [CDQ14].

Chen:2014:ESI

[CDQ14] L. Chen, W. W. Dou, and Z. Qiao. “Ensemble Subsampling for Imbalanced Multivariate Two-Sample Tests,” Chen, L., Dou, W. W., and Qiao, Z. (2013), Journal of the American Statistical Association, 108, 1308–1323. Journal of the American Statistical Association, 109(506):871, 2014. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [CDQ13].

Corradi:2011:PII

[CDS11] Valentina Corradi, Walter Distaso, and Norman R. Swanson. Predictive inference for integrated volatility. Journal of the American Statistical Association, 106(496):1496–1512, December 2011. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Cerioli:2010:MOD

[Cer10] Andrea Cerioli. Multivariate outlier detection with highbreakdown estimators. Journal of the American Statistical Association, 105(489):147–156, March 2010. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chen:2017:NGB

[CF17] Hao Chen and Jerome H. Friedman. A new graph-based twosample test for multivariate and object data. Journal of the American Statistical Association, 112(517):397–409, 2017. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chen:2018:EVE

[CFL18] Zhao Chen, Jianqing Fan, and Runze Li. Error variance estimation in ultrahigh-dimensional additive models. Journal of the American Statistical Association, 113(521):315–327, 2018. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chang:2015:TCV

[CG15] Lo-Bin Chang and Donald Geman. Tracking cross-validated estimates of prediction error as studies accumulate. Journal of the American Statistical Association, 110(511):1239–1247, 2015. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Cho:2013:MFD

[CGBY13] Haeran Cho, Yannig Goude, Xavier Brossat, and Qiwei Yao. Modeling and forecasting daily electricity load curves: a hybrid approach. Journal of the American Statistical Association, 108 (501):7–21, 2013. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chen:2010:SSF

[CGL10] Rong Chen, Re-Jin Guo, and Ming Lin. Self-selectivity in firm’s decision to withdraw IPO: Bayesian inference for hazard models of bankruptcy with feedback. Journal of the American Statistical Association, 105(492):1297–1309, December 2010. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chen:2010:LAR

[CGLY10] Kani Chen, Shaojun Guo, Yuanyuan Lin, and Zhiliang Ying. Least absolute relative error estimation. Journal of the American Statistical Association, 105(491):1104–1112, September 2010. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chen:2019:FDA

[CGO19] Yakuan Chen, Jeff Goldsmith, and R. Todd Ogden. Functional data analysis of dynamic PET data. Journal of the American Statistical Association, 114(526):595–609, 2019. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chen:2010:GPL

[CGSW10] Kani Chen, Shaojun Guo, Liuquan Sun, and Jane-Ling Wang. Global partial likelihood for nonparametric proportional hazards models. Journal of the American Statistical Association, 105 (490):750–760, June 2010. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chen:2017:DLS

[CGW17] Yen-Chi Chen, Christopher R. Genovese, and Larry Wasserman. Density level sets: Asymptotics, inference, and visualization. Journal of the American Statistical Association, 112(520):1684– 1696, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chen:2012:SRR

[CH12] Lisha Chen and Jianhua Z. Huang. Sparse reduced-rank regression for simultaneous dimension reduction and variable selection. Journal of the American Statistical Association, 107(500):1533– 1545, 2012. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Coretto:2017:RIM

[CH17] Pietro Coretto and Christian Hennig. Robust improper maximum likelihood: Tuning, computation, and a comparison with other methods for robust Gaussian clustering. Journal of the American Statistical Association, 111(516):1648–1659, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chen:2019:CPS

[CH19] Qingxia Chen and Frank E. Harrell Jr. Comment: Penalized spline of propensity methods for treatment comparison. Journal of the American Statistical Association, 114(525):28–30, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chakrabarty:2017:NBT

[Cha17] Dalia Chakrabarty. A new Bayesian test to test for the intractability-countering hypothesis. Journal of the American Statistical Association, 112(518):561–577, 2017. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chang:2016:CIS

[CHAP16] Won Chang, Murali Haran, Patrick Applegate, and David Pollard. Calibrating an ice sheet model using high-dimensional binary spatial data. Journal of the American Statistical Association, 111(513):57–72, 2016. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chen:2012:ERP

[CHC^{+}12] Ying Qing Chen, Nan Hu, Su-Chun Cheng, Philippa Musoke, and Lue Ping Zhao. Estimating regression parameters in an extended proportional odds model. Journal of the American Statistical Association, 107(497):318–330, 2012. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chen:2013:C

[Che13] Xiaohong Chen. Comment. Journal of the American Statistical Association, 108(504):1262–1264, 2013. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [CCJ13a].

Choi:2017:EMT

[Cho17] David Choi. Estimation of monotone treatment effects in network experiments. Journal of the American Statistical Association, 112(519):1147–1155, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chen:2010:LRV

[CHP10] Ying Chen, Wolfgang Karl Härdle, and Uta Pigorsch. Localized realized volatility modeling. Journal of the American Statistical Association, 105(492):1376–1393, December 2010. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Claeskens:2014:MFH

[CHSV14] Gerda Claeskens, Mia Hubert, Leen Slaets, and Kaveh Vakili. Multivariate functional halfspace depth. Journal of the American Statistical Association, 109(505):411–423, 2014. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Cheng:2016:FVS

[CHZ16] Ming-Yen Cheng, Toshio Honda, and Jin-Ting Zhang. Forward variable selection for sparse ultra-high dimensional varying coefficient models. Journal of the American Statistical Association, 111(515):1209–1221, 2016. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chen:2018:SBM

[CIB^{+}18] Xi Chen, Kaoru Irie, David Banks, Robert Haslinger, Jewell Thomas, and Mike West. Scalable Bayesian modeling, monitoring, and analysis of dynamic network flow data. Journal of the American Statistical Association, 113(522):519–533, 2018. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Cattaneo:2018:ILR

[CJN18] Matias D. Cattaneo, Michael Jansson, and Whitney K. Newey. Inference in linear regression models with many covariates and heteroscedasticity. Journal of the American Statistical Association, 113(523):1350–1361, 2018. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Cadonna:2019:BSM

[CKP19] Annalisa Cadonna, Athanasios Kottas, and Raquel Prado. Bayesian spectral modeling for multiple time series. Journal of the American Statistical Association, 114(528):1838–1853, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chiou:2015:SAF

[CKY15] Sy Han Chiou, Sangwook Kang, and Jun Yan. Semiparametric accelerated failure time modeling for clustered failure times from stratified sampling. Journal of the American Statistical Association, 110(510):621–629, 2015. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chatterjee:2010:ISR

[CL10] Nilanjan Chatterjee and Yan Li. Inference in semiparametric regression models under partial questionnaire design and nonmonotone missing data. Journal of the American Statistical Association, 105(490):787–797, June 2010. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Cai:2011:ATS

[CL11a] Tony Cai and Weidong Liu. Adaptive thresholding for sparse covariance matrix estimation. Journal of the American Statistical Association, 106(494):672–684, June 2011. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Cai:2011:DEA

[CL11b] Tony Cai and Weidong Liu. A direct estimation approach to sparse linear discriminant analysis. Journal of the American Statistical Association, 106(496):1566–1577, December 2011. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chatterjee:2011:BLE

[CL11c] A. Chatterjee and S. N. Lahiri. Bootstrapping lasso estimators. Journal of the American Statistical Association, 106(494):608– 625, June 2011. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chatterjee:2015:C

[CL15a] A. Chatterjee and S. N. Lahiri. Comment. Journal of the American Statistical Association, 110(512):1434–1438, 2015. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/01621459.2015.1102143.

Chen:2015:LFP

[CL15b] Kehui Chen and Jing Lei. Localized functional principal component analysis. Journal of the American Statistical Association, 110(511):1266–1275, 2015. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Cai:2016:LSM

[CL16a] T. Tony Cai and Weidong Liu. Large-scale multiple testing of correlations. Journal of the American Statistical Association, 111 (513):229–240, 2016. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chen:2016:DCM

[CL16b] Ziqi Chen and Chenlei Leng. Dynamic covariance models. Journal of the American Statistical Association, 111(515):1196–1207, 2016. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chen:2018:NCV

[CL18] Kehui Chen and Jing Lei. Network cross-validation for determining the number of communities in network data. Journal of the American Statistical Association, 113(521):241–251, 2018. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Cheng:2010:LRT

[CLA10] Philip E. Cheng, Michelle Liou, and John A. D. Aston. Likelihood ratio tests with three-way tables. Journal of the American Statistical Association, 105(490):740–749, June 2010. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chen:2012:ION

[CLF12] Jiahua Chen, Pengfei Li, and Yuejiao Fu. Inference on the order of a normal mixture. Journal of the American Statistical Association, 107(499):1096–1105, 2012. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chan:2018:IIE

[CLGS18] Joshua Chan, Roberto Leon-Gonzalez, and Rodney W. Strachan. Invariant inference and efficient computation in the static factor model. Journal of the American Statistical Association, 113 (522):819–828, 2018. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Cai:2011:CMA

[CLL11] Tony Cai, Weidong Liu, and Xi Luo. A constrained ell_{1} minimization approach to sparse precision matrix estimation. Journal of the American Statistical Association, 106(494):594–607, June 2011. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Cao:2019:LCE

[CLL19] Yuanpei Cao, Wei Lin, and Hongzhe Li. Large covariance estimation for compositional data via composition-adjusted thresholding. Journal of the American Statistical Association, 114(526): 759–772, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chen:2018:SUH

[CLLL18] Jia Chen, Degui Li, Oliver Linton, and Zudi Lu. Semiparametric ultra-high dimensional model averaging of nonlinear dynamic time series. Journal of the American Statistical Association, 113 (522):919–932, 2018. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Cai:2014:ACB

[CLM14] T. Tony Cai, Mark Low, and Zongming Ma. Adaptive confidence bands for nonparametric regression functions. Journal of the American Statistical Association, 109(507):1054–1070, 2014. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Carvalho:2018:LRV

[CLM18] Carlos M. Carvalho, Hedibert F. Lopes, and Robert E. McCulloch. On the long-run volatility of stocks. Journal of the American Statistical Association, 113(523):1050–1069, 2018. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chen:2016:UMD

[CLQY16] Baojiang Chen, Pengfei Li, Jing Qin, and Tao Yu. Using a monotonic density ratio model to find the asymptotically optimal combination of multiple diagnostic tests. Journal of the American Statistical Association, 111(514):861–??, 2016. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Carone:2019:TCE

[CLvdL19] Marco Carone, Alexander R. Luedtke, and Mark J. van der Laan. Toward computerized efficient estimation in infinite-dimensional models. Journal of the American Statistical Association, 114 (527):1174–1190, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Cai:2013:TSC

[CLX13] Tony Cai, Weidong Liu, and Yin Xia. Two-sample covariance matrix testing and support recovery in high-dimensional and sparse settings. Journal of the American Statistical Association, 108(501):265–277, 2013. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chen:2015:SAM

[CLXY15] Yunxiao Chen, Jingchen Liu, Gongjun Xu, and Zhiliang Ying. Statistical analysis of Q-matrix based diagnostic classification models. Journal of the American Statistical Association, 110 (510):850–866, 2015. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Choi:2010:VSS

[CLZ10] Nam Hee Choi, William Li, and Ji Zhu. Variable selection with the strong heredity constraint and its oracle property. Journal of the American Statistical Association, 105(489):354–364, March 2010. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Cui:2015:MFF

[CLZ15] Hengjian Cui, Runze Li, and Wei Zhong. Model-free feature screening for ultrahigh dimensional discriminant analysis. Journal of the American Statistical Association, 110(510):630–641, 2015. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chen:2016:C

[CLZ^{+}16] Jingxiang Chen, Yufeng Liu, Donglin Zeng, Rui Song, Yingqi Zhao, and Michael R. Kosorok. Comment. Journal of the American Statistical Association, 111(515):942–947, 2016. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chen:2012:MRF

[CM12] Kehui Chen and Hans-Georg Müller. Modeling repeated functional observations. Journal of the American Statistical Association, 107(500):1599–1609, 2012. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Cruz-Marcelo:2011:ETS

[CMER11] Alejandro Cruz-Marcelo, Katherine B. Ensor, and Gary L. Rosner. Estimating the term structure with a semiparametric Bayesian hierarchical model: An application to corporate bonds. Journal of the American Statistical Association, 106(494):387– 395, June 2011. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chakraborty:2013:SBE

[CMM^{+}13] Avishek Chakraborty, Bani K. Mallick, Ryan G. Mcclarren, Carolyn C. Kuranz, Derek Bingham, Michael J. Grosskopf, Erica M. Rutter, Hayes F. Stripling, and R. Paul Drake. Spline-based emulators for radiative shock experiments with measurement error. Journal of the American Statistical Association, 108(502):

411–428, 2013. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Conrad:2017:AAE

[CMPS17] Patrick R. Conrad, Youssef M. Marzouk, Natesh S. Pillai, and Aaron Smith. Accelerating asymptotically exact MCMC for computationally intensive models via local approximations. Journal of the American Statistical Association, 111(516):1591–1607, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Crawford:2014:EGB

[CMS14] Forrest W. Crawford, Vladimir N. Minin, and Marc A. Suchard. Estimation for general birth–death processes. Journal of the American Statistical Association, 109(506):730–747, 2014. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Conti:2017:SMA

[CMS17] Pier Luigi Conti, Daniela Marella, and Mauro Scanu. Statistical matching analysis for complex survey data with applications. Journal of the American Statistical Association, 111(516): 1715–1725, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chen:2018:TMM

[CMS^{+}18] Kun Chen, Neha Mishra, Joan Smyth, Haim Bar, Elizabeth Schifano, Lynn Kuo, and Ming-Hui Chen. A tailored multivariate mixture model for detecting proteins of concordant change among virulent strains of Clostridium Perfringens. Journal of the American Statistical Association, 113(522):546–559, 2018. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chen:2019:CCA

[CMW^{+}19] Yang Chen, Xiao-Li Meng, Xufei Wang, David A. van Dyk, Herman L. Marshall, and Vinay L. Kashyap. Calibration concordance for astronomical instruments via multiplicative shrinkage. Journal of the American Statistical Association, 114(527):1018– 1037, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Cressie:2010:CSD

[CO10] Noel Cressie and David O’Donnell. Comment: Statistical dependence in stream networks. Journal of the American Statistical Association, 105(489):18–21, March 2010. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Choudhury:2010:RRW

[COS^{+}10] Kingshuk Roy Choudhury, Finbarr O’Sullivan, Mayukh Samanta, Guillemette Caulliez, and Victor Shrira. Regularized reconstruction of wave height and slope fields from refracted images of water. Journal of the American Statistical Association, 105(489): 36–47, March 2010. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chandler:2012:MIV

[CP12] Gabriel Chandler and Wolfgang Polonik. Mode identification of volatility in time-varying autoregression. Journal of the American Statistical Association, 107(499):1217–1229, 2012. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chen:2011:RHT

[CPPW11] Lin S. Chen, Debashis Paul, Ross L. Prentice, and Pei Wang. A regularized Hotelling’s T^{2} test for pathway analysis in proteomic studies. Journal of the American Statistical Association, 106 (496):1345–1360, December 2011. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Cao:2010:LME

[CR10] J. Cao and J. O. Ramsay. Linear mixed-effects modeling by parameter cascading. Journal of the American Statistical Association, 105(489):365–374, March 2010. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Christensen:2018:C

[CR18] William F. Christensen and C. Shane Reese. Comment. Journal of the American Statistical Association, 113(521):171–173, 2018. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Crane:2015:CCD

[Cra15] Harry Crane. Clustering from categorical data sequences. Journal of the American Statistical Association, 110(510):810–823, 2015. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Cressie:2018:MCN

[Cre18a] Noel Cressie. Mission CO_{2} ntrol: A statistical scientist’s role in remote sensing of atmospheric carbon dioxide. Journal of the American Statistical Association, 113(521):152–168, 2018. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Cressie:2018:R

[Cre18b] Noel Cressie. Rejoinder. Journal of the American Statistical Association, 113(521):178–181, 2018. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chen:2015:LIS

[CRL15] Hua Yun Chen, Daniel E. Rader, and Mingyao Li. Likelihood inferences on semiparametric odds ratio model. Journal of the American Statistical Association, 110(511):1125–1135, 2015. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Conne:2010:GFG

[CRVF10] David Conne, Elvezio Ronchetti, and Maria-Pia Victoria-Feser. Goodness of fit for generalized linear latent variables models. Journal of the American Statistical Association, 105(491):1126– 1134, September 2010. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chen:2016:ANE

[CRZZ16] Mengjie Chen, Zhao Ren, Hongyu Zhao, and Harrison Zhou. Asymptotically normal and efficient estimation of covariateadjusted Gaussian graphical model. Journal of the American Statistical Association, 111(513):394–406, 2016. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Christensen:2010:AGF

[CS10] Ronald Christensen and Siu Kei Sun. Alternative goodness-of-fit tests for linear models. Journal of the American Statistical Association, 105(489):291–301, March 2010. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chib:2018:BEC

[CSS18] Siddhartha Chib, Minchul Shin, and Anna Simoni. Bayesian estimation and comparison of moment condition models. Journal of the American Statistical Association, 113(524):1656–1668, 2018. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chen:2016:ASM

[CSSK16] Yang Chen, Kuang Shen, Shu-Ou Shan, and S. C. Kou. Analyzing single-molecule protein transportation experiments via hierarchical hidden Markov models. Journal of the American Statistical Association, 111(515):951–966, 2016. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chen:2017:NRH

[CSW17] Shizhe Chen, Ali Shojaie, and Daniela M. Witten. Network reconstruction from high-dimensional ordinary differential equations. Journal of the American Statistical Association, 112(520): 1697–1707, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Cressie:2010:CHS

[CT10] Noel Cressie and Martin P. Tingley. Comment: “Hierarchical Statistical Modeling for Paleoclimate Reconstruction”. Journal of the American Statistical Association, 105(491):895–900, September 2010. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Cai:2017:C

[CT17] Tianxi Cai and Lu Tian. Comment. Journal of the American Statistical Association, 111(516):1521–1524, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [CZK17a, CZK17b].

Chen:2010:LPS

[CTM10] Song Xi Chen, Cheng Yong Tang, and Vincent T. Mule, Jr. Local post-stratification in dual system accuracy and coverage evaluation for the U.S. Census. Journal of the American Statistical Association, 105(489):105–119, March 2010. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Culp:2011:PSS

[Cul11] Mark Culp. On propagated scoring for semisupervised additive models. Journal of the American Statistical Association, 106 (493):248–259, March 2011. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chaffee:2012:C

[CvdL12] Paul Chaffee and Mark van der Laan. Comment. Journal of the American Statistical Association, 107(498):513–517, 2012. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Clarke:2012:IVE

[CW12] Paul S. Clarke and Frank Windmeijer. Instrumental variable estimators for binary outcomes. Journal of the American Statistical Association, 107(500):1638–1652, 2012. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Cai:2014:SMC

[CW14] Zongwu Cai and Xian Wang. Selection of mixed copula model via penalized likelihood. Journal of the American Statistical Association, 109(506):788–801, 2014. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chan:2017:SME

[CW17] Kwun Chuen Gary Chan and Mei-Cheng Wang. Semiparametric modeling and estimation of the terminal behavior of recurrent marker processes before failure events. Journal of the American Statistical Association, 112(517):351–362, 2017. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chen:2019:TTH

[CW19] Likai Chen and Wei Biao Wu. Testing for trends in highdimensional time series. Journal of the American Statistical Association, 114(526):869–881, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chen:2014:TLS

[CWC^{+}14] Tianle Chen, Yuanjia Wang, Huaihou Chen, Karen Marder, and Donglin Zeng. Targeted local support vector machine for agedependent classification. Journal of the American Statistical Association, 109(507):1174–1187, 2014. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chien:2015:SLD

[CWH^{+}15] Li-Chu Chien, Yuh-Jenn Wu, Chao A. Hsiung, Lu-Hai Wang, and I-Shou Chang. Smoothed Lexis diagrams with applications to lung and breast cancer trends in Taiwan. Journal of the American Statistical Association, 110(511):1000–1012, 2015. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Crawford:2018:HPS

[CWH18] Forrest W. Crawford, Jiacheng Wu, and Robert Heimer. Hidden population size estimation from respondent-driven sampling: A network approach. Journal of the American Statistical Association, 113(522):755–766, 2018. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chen:2013:MAR

[CWPC13] Huaihou Chen, Yuanjia Wang, Myunghee Cho Paik, and H. Alex Choi. A marginal approach to reduced-rank penalized spline smoothing with application to multilevel functional data. Journal of the American Statistical Association, 108(504):1216–1229, 2013. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chen:2015:EQR

[CWZ15] Xuerong Chen, Alan T. K. Wan, and Yong Zhou. Efficient quantile regression analysis with missing observations. Journal of the American Statistical Association, 110(510):723–741, 2015. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Crawford:2018:BAK

[CWZM18] Lorin Crawford, Kris C. Wood, Xiang Zhou, and Sayan Mukherjee. Bayesian approximate kernel regression with variable selection. Journal of the American Statistical Association, 113(524): 1710–1721, 2018. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Cheng:2013:LSM

[CXCR13] Jerry Q. Cheng, Minge Xie, Rong Chen, and Fred Roberts. A latent source model to detect multiple spatial clusters with application in a mobile sensor network for surveillance of nuclear materials. Journal of the American Statistical Association, 108 (503):902–913, 2013. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Claggett:2014:MAF

[CXT14] Brian Claggett, Minge Xie, and Lu Tian. Meta-analysis with fixed, unknown, study-specific parameters. Journal of the American Statistical Association, 109(508):1660–1671, 2014. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Cai:2012:MAP

[CY12] T. Tony Cai and Ming Yuan. Minimax and adaptive prediction for functional linear regression. Journal of the American Statistical Association, 107(499):1201–1216, 2012. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Cai:2016:MAE

[CY16] T. Tony Cai and Ming Yuan. Minimax and adaptive estimation of covariance operator for random variables observed on a lattice graph. Journal of the American Statistical Association, 111(513): 253–265, 2016. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chan:2017:AOB

[CY17] Kin Wai Chan and Chun Yip Yau. Automatic optimal batch size selection for recursive estimators of time-average covariance matrix. Journal of the American Statistical Association, 112(519): 1076–1089, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chen:2010:WGE

[CYC10] Baojiang Chen, Grace Y. Yi, and Richard J. Cook. Weighted generalized estimating functions for longitudinal response and covariate data that are missing at random. Journal of the American Statistical Association, 105(489):336–353, March 2010. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chen:2012:SMC

[CYT12] Hsiu-Hsi Chen, Amy Ming-Fang Yen, and László Tabár. A stochastic model for calibrating the survival benefit of screendetected cancers. Journal of the American Statistical Association, 107(500):1339–1359, 2012. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chan:2014:GLS

[CYZ14] Ngai Hang Chan, Chun Yip Yau, and Rong-Mao Zhang. Group LASSO for structural break time series. Journal of the American Statistical Association, 109(506):590–599, 2014. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Cai:2011:NEB

[CZ11] Tianxi Cai and Yingye Zheng. Nonparametric evaluation of biomarker accuracy under nested case-control studies. Journal of the American Statistical Association, 106(494):569–580, June 2011. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Cai:2013:RPM

[CZ13] Tianxi Cai and Yingye Zheng. Resampling procedures for making inference under nested case-control studies. Journal of the American Statistical Association, 108(504):1532–1544, 2013. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chen:2014:C

[CZ14a] Huaihou Chen and Donglin Zeng. Comment. Journal of the American Statistical Association, 109(508):1350–1353, 2014. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Cook:2014:FEC

[CZ14b] R. Dennis Cook and Xin Zhang. Fused estimators of the central subspace in sufficient dimension reduction. Journal of the American Statistical Association, 109(506):815–827, 2014. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Cook:2015:FEM

[CZ15] R. Dennis Cook and Xin Zhang. Foundations for envelope models and methods. Journal of the American Statistical Association, 110(510):599–611, 2015. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chakraborty:2019:DMM

[CZ19] Shubhadeep Chakraborty and Xianyang Zhang. Distance metrics for measuring joint dependence with application to causal inference. Journal of the American Statistical Association, 114(528): 1638–1650, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chen:2017:PDF

[CZK17a] Guanhua Chen, Donglin Zeng, and Michael R. Kosorok. Personalized dose finding using outcome weighted learning. Journal of the American Statistical Association, 111(516):1509–1521, 2017.

CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See comments [CT17, FY17, LvdL17, WMS17, Ogb17, Qia17, Ros17] and rejoinder [CZK17b].

Chen:2017:R

[CZK17b] Guanhua Chen, Donglin Zeng, and Michael R. Kosorok. Rejoinder. Journal of the American Statistical Association, 111(516): 1543–1547, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [CT17, CZK17a, FY17, LvdL17, Ogb17, Qia17, Ros17, WMS17].

Cseke:2017:SAI

[CZMHS17] Botond Cseke, Andrew Zammit-Mangion, Tom Heskes, and Guido Sanguinetti. Sparse approximate inference for spatiotemporal point process models. Journal of the American Statistical Association, 111(516):1746–1763, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chen:2011:PVI

[CZW^{+}11] Minhua Chen, Aimee Zaas, Christopher Woods, Geoffrey S. Ginsburg, Joseph Lucas, David Dunson, and Lawrence Carin. Predicting viral infection from high-dimensional biomarker trajectories. Journal of the American Statistical Association, 106 (496):1259–1279, December 2011. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chen:2010:THD

[CZZ10] Song Xi Chen, Li-Xin Zhang, and Ping-Shou Zhong. Tests for high-dimensional covariance matrices. Journal of the American Statistical Association, 105(490):810–819, June 2010. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

DAmour:2019:CRD

[D’A19] Alexander D’Amour. Comment: Reflections on the Deconfounder. Journal of the American Statistical Association, 114 (528):1597–1601, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [WB19a].

Dai:2019:BRA

[Dai19] Hongying Dai. Book review: Asymptotic Analysis of Mixed Effects Models: Theory, Applications, and Open Problems. Journal of the American Statistical Association, 114(527):1418–1420, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Davies:2012:IAV

[Dav12] P. L. Davies. Interactions in the analysis of variance. Journal of the American Statistical Association, 107(500):1502–1509, 2012. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Davidian:2013:IYS

[Dav13] Marie Davidian. The International Year of Statistics: A celebration and a call to action. Journal of the American Statistical Association, 108(504):1141–1146, 2013. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Datta:2016:HNN

[DBFG16] Abhirup Datta, Sudipto Banerjee, Andrew O. Finley, and Alan E. Gelfand. Hierarchical nearest-neighbor Gaussian process models for large geostatistical datasets. Journal of the American Statistical Association, 111(514):800–??, 2016. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Das:2016:FER

[DBNZ16] Ritabrata Das, Moulinath Banerjee, Bin Nan, and Huiyong Zheng. Fast estimation of regression parameters in a broken-stick model for longitudinal data. Journal of the American Statistical Association, 111(515):1132–1143, 2016. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

deCarvalho:2014:SDR

[dCD14] Miguel de Carvalho and Anthony C. Davison. Spectral density ratio models for multivariate extremes. Journal of the American Statistical Association, 109(506):764–776, 2014. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ding:2016:PTT

[DD16] Peng Ding and Tirthankar Dasgupta. A potential tale of twoby-two tables from completely randomized experiments. Journal of the American Statistical Association, 111(513):157–168, 2016. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Du:2015:SSA

[DDH15] Jiejun Du, Ian L. Dryden, and Xianzheng Huang. Size and shape analysis of error-prone shape data. Journal of the American Statistical Association, 110(509):368–377, 2015. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Dhavala:2010:BMM

[DDM^{+}10] Soma S. Dhavala, Sujay Datta, Bani K. Mallick, Raymond J. Carroll, Sangeeta Khare, Sara D. Lawhon, and L. Garry Adams. Bayesian modeling of MPSS data: Gene expression analysis of bovine salmonella infection. Journal of the American Statistical Association, 105(491):956–967, September 2010. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Dahl:2017:RPD

[DDT17] David B. Dahl, Ryan Day, and Jerry W. Tsai. Random partition distribution indexed by pairwise information. Journal of the American Statistical Association, 112(518):721–732, 2017. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Durante:2017:NBM

[DDV17a] Daniele Durante, David B. Dunson, and Joshua T. Vogelstein. Nonparametric Bayes modeling of populations of networks. Journal of the American Statistical Association, 112(520):1516–1530, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Durante:2017:RNB

[DDV17b] Daniele Durante, David B. Dunson, and Joshua T. Vogelstein. Rejoinder: Nonparametric Bayes modeling of populations of networks. Journal of the American Statistical Association, 112(520): 1547–1552, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Dette:2013:C

[Det13] Holger Dette. Comment. Journal of the American Statistical Association, 108(504):1258–1260, 2013. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [CCJ13a].

DeBacker:2019:ALF

[DEV19] Mickaël De Backer, Anouar El Ghouch, and Ingrid Van Keilegom. An adapted loss function for censored quantile regression.

Journal of the American Statistical Association, 114(527):1126– 1137, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ding:2019:DTE

[DFM19] Peng Ding, Avi Feller, and Luke Miratrix. Decomposing treatment effect variation. Journal of the American Statistical Association, 114(525):304–317, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Davison:2014:ADI

[DFRS14] A. C. Davison, D. A. S. Fraser, N. Reid, and N. Sartori. Accurate directional inference for vector parameters in linear exponential families. Journal of the American Statistical Association, 109 (505):302–314, 2014. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Diggle:2016:MBG

[DG16a] Peter J. Diggle and Emanuele Giorgi. Model-based geostatistics for prevalence mapping in low-resource settings. Journal of the American Statistical Association, 111(515):1096–1120, 2016. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Diggle:2016:R

[DG16b] Peter J. Diggle and Emanuele Giorgi. Rejoinder. Journal of the American Statistical Association, 111(515):1119–1120, 2016. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Devijver:2018:BDC

[DG18] Emilie Devijver and Mélina Gallopin. Block-diagonal covariance selection for high-dimensional Gaussian graphical models. Journal of the American Statistical Association, 113(521):306–314, 2018. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Diggle:2010:EIL

[DGH^{+}10] Peter J. Diggle, Yongtao Guan, Anthony C. Hart, Fauzia Paize, and Michelle Stanton. Estimating individual-level risk in spatial epidemiology using spatially aggregated information on the population at risk. Journal of the American Statistical Association, 105(492):1394–1402, December 2010. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Dovonon:2019:BHF

[DGHM19] Prosper Dovonon, Sílvia Gonçalves, Ulrich Hounyo, and Nour Meddahi. Bootstrapping high-frequency jump tests. Journal of the American Statistical Association, 114(526):793–803, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Dai:2012:PHM

[DGM12] James Y. Dai, Peter B. Gilbert, and Benoît R. Mâsse. Partially hidden Markov model for time-varying principal stratification in HIV prevention trials. Journal of the American Statistical Association, 107(497):52–65, 2012. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Daouia:2019:ENP

[DGS19] Abdelaati Daouia, Irène Gijbels, and Gilles Stupfler. Extremiles: A new perspective on asymmetric least squares. Journal of the American Statistical Association, 114(527):1366–1381, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ding:2011:IEC

[DGYZ11] Peng Ding, Zhi Geng, Wei Yan, and Xiao-Hua Zhou. Identifiability and estimation of causal effects by principal stratification with outcomes truncated by death. Journal of the American Statistical Association, 106(496):1578–1591, December 2011. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Delaigle:2012:CRA

[DH12] Aurore Delaigle and Peter Hall. Comment: Robustness to assumption of normally distributed errors. Journal of the American Statistical Association, 107(499):1036–1039, 2012. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Delaigle:2013:CUC

[DH13] Aurore Delaigle and Peter Hall. Classification using censored functional data. Journal of the American Statistical Association, 108(504):1269–1283, 2013. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Delaigle:2014:PAN

[DH14] Aurore Delaigle and Peter Hall. Parametrically assisted nonparametric estimation of a density in the deconvolution problem. Journal of the American Statistical Association, 109(506): 717–729, 2014. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Deng:2014:BAO

[DHLL14] Ke Deng, Simeng Han, Kate J. Li, and Jun S. Liu. Bayesian aggregation of order-based rank data. Journal of the American Statistical Association, 109(507):1023–1039, 2014. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Datta:2011:MST

[DHM11] Gauri S. Datta, Peter Hall, and Abhyuday Mandal. Model selection by testing for the presence of small-area effects, and application to area-level data. Journal of the American Statistical Association, 106(493):362–374, March 2011. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

DeLivera:2011:FTS

[DHS11] Alysha M. De Livera, Rob J. Hyndman, and Ralph D. Snyder. Forecasting time series with complex seasonal patterns using exponential smoothing. Journal of the American Statistical Association, 106(496):1513–1527, December 2011. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Dias:2013:TES

[DIvdB13] Monica Costa Dias, Hidehiko Ichimura, and Gerard J. van den Berg. Treatment evaluation with selective participation and ineligibles. Journal of the American Statistical Association, 108 (502):441–455, 2013. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Davidov:2018:TIC

[DJP18] Ori Davidov, Casey M. Jelsema, and Shyamal Peddada. Testing for inequality constraints in singular models by trimming or winsorizing the variance matrix. Journal of the American Statistical Association, 113(522):906–918, 2018. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Duchi:2018:MOP

[DJW18a] John C. Duchi, Michael I. Jordan, and Martin J. Wainwright. Minimax optimal procedures for locally private estimation. Journal of the American Statistical Association, 113(521):182–201, 2018. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Duchi:2018:R

[DJW18b] John C. Duchi, Michael I. Jordan, and Martin J. Wainwright. Rejoinder. Journal of the American Statistical Association, 113 (521):212–215, 2018. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

DeYoreo:2018:MDO

[DK18] Maria DeYoreo and Athanasios Kottas. Modeling for dynamic ordinal regression relationships: An application to estimating maturity of rockfish in California. Journal of the American Statistical Association, 113(521):68–80, 2018. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Du:2016:SSE

[DKK16] Chao Du, Chu-Lan Michael Kao, and S. C. Kou. Stepwise signal extraction via marginal likelihood. Journal of the American Statistical Association, 111(513):314–330, 2016. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Dukic:2012:TEG

[DLP12] Vanja Dukic, Hedibert F. Lopes, and Nicholas G. Polson. Tracking epidemics with Google flu trends data and a state-space SEIR model. Journal of the American Statistical Association, 107(500):1410–1426, 2012. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Dobra:2011:BIG

[DLR11] Adrian Dobra, Alex Lenkoski, and Abel Rodriguez. Bayesian inference for general Gaussian graphical models with application to multivariate lattice data. Journal of the American Statistical Association, 106(496):1418–1433, December 2011. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Delaigle:2011:NRA

[DM11] Aurore Delaigle and Alexander Meister. Nonparametric regression analysis for group testing data. Journal of the American Statistical Association, 106(494):640–650, June 2011. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Datta:2015:SAE

[DM15] Gauri Sankar Datta and Abhyuday Mandal. Small area estimation with uncertain random effects. Journal of the American Statistical Association, 110(512):1735–1744, 2015. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/01621459.2015.1016526.

Dawson:2018:DMC

[DM18] Matthew Dawson and Hans-Georg Müller. Dynamic modeling of conditional quantile trajectories, with application to longitudinal snippet data. Journal of the American Statistical Association, 113(524):1612–1624, 2018. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Dette:2018:ERC

[DMVB18] Holger Dette, Kathrin Möllenhoff, Stanislav Volgushev, and Frank Bretz. Equivalence of regression curves. Journal of the American Statistical Association, 113(522):711–729, 2018. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Dabo-Niang:2010:CNO

[DNFZ10] Sophie Dabo-Niang, Christian Francq, and Jean-Michel Zakoïan. Combining nonparametric and optimal linear time series predictions. Journal of the American Statistical Association, 105(492): 1554–1565, December 2010. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Dyer:2012:COZ

[DO12] Justin S. Dyer and Art B. Owen. Correct ordering in the Zipf– Poisson ensemble. Journal of the American Statistical Association, 107(500):1510–1517, 2012. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Davidov:2011:ORI

[DP11] Ori Davidov and Shyamal Peddada. Order-restricted inference for multivariate binary data with application to toxicology. Journal of the American Statistical Association, 106(496):1394–1404, December 2011. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Dryden:2019:DOO

[DPS19] Ian L. Dryden, Simon P. Preston, and Katie E. Severn. Discussion: Object-Oriented Data Analysis, Power Metrics, and Graph Laplacians. Journal of the American Statistical Association, 114(527):1097–1098, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

DiMarzio:2014:NRS

[DPT14] Marco Di Marzio, Agnese Panzera, and Charles C. Taylor. Nonparametric regression for spherical data. Journal of the American Statistical Association, 109(506):748–763, 2014. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Marzio:2019:NRS

[DPT19] Marco Di Marzio, Agnese Panzera, and Charles C. Taylor. Nonparametric rotations for sphere–sphere regression. Journal of the American Statistical Association, 114(525):466–476, 2019. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Dette:2011:MSL

[DPV11] Holger Dette, Philip Preuß, and Mathias Vetter. A measure of stationarity in locally stationary processes with applications to testing. Journal of the American Statistical Association, 106 (495):1113–1124, September 2011. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Dadaneh:2018:BSB

[DQZ18] Siamak Zamani Dadaneh, Xiaoning Qian, and Mingyuan Zhou. BNP-Seq: Bayesian nonparametric differential expression analysis of sequencing count data. Journal of the American Statistical Association, 113(521):81–94, 2018. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Drechsler:2010:SSN

[DR10] Jörg Drechsler and Jerome P. Reiter. Sampling with synthesis: a new approach for releasing public use census microdata. Journal of the American Statistical Association, 105(492):1347–1357, December 2010. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

DiCiccio:2017:RPT

[DR17] Cyrus J. DiCiccio and Joseph P. Romano. Robust permutation tests for correlation and regression coefficients. Journal of the American Statistical Association, 112(519):1211–1220, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Danaher:2012:MWP

[DRC^{+}12] Michelle R. Danaher, Anindya Roy, Zhen Chen, Sunni L. Mumford, and Enrique F. Schisterman. Minkowski–Weyl priors for models with parameter constraints: An analysis of the BioCycle Study. Journal of the American Statistical Association, 107(500): 1395–1409, 2012. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Dombry:2018:PCE

[DRS18] Clément Dombry, Mathieu Ribatet, and Stilian Stoev. Probabilities of concurrent extremes. Journal of the American Statistical Association, 113(524):1565–1582, 2018. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Desai:2012:CDI

[DS12] Keyur H. Desai and John D. Storey. Cross-dimensional inference of dependent high-dimensional data. Journal of the American Statistical Association, 107(497):135–151, 2012. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Dette:2012:ODQ

[DT12] Holger Dette and Matthias Trampisch. Optimal designs for quantile regression models. Journal of the American Statistical Association, 107(499):1140–1151, 2012. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

DeNeve:2015:RFR

[DT15] Jan De Neve and Olivier Thas. A regression framework for rank tests based on the probabilistic index model. Journal of the American Statistical Association, 110(511):1276–1283, 2015. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ding:2012:CDT

[DTYG12] A. Adam Ding, Shaonan Tian, Yan Yu, and Hui Guo. A class of discrete transformation survival models with application to default probability prediction. Journal of the American Statistical Association, 107(499):990–1003, 2012. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Du:2014:C

[Du14] Pang Du. Comment. Journal of the American Statistical Association, 109(508):1349–1350, 2014. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Duivesteijn:2016:CJT

[Dui16] Wouter Duivesteijn. Correction to Jin-Ting Zhang’s “Approximate and Asymptotic Distributions of Chi-Squared-Type Mixtures With Applications”. Journal of the American Statistical Association, 111(515):1370–1371, 2016. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [Zha05].

Dumbgen:2011:CAC

[Düm11] Lutz Dümbgen. Comment: “Adaptive Confidence Intervals for the Test Error in Classification”. Journal of the American Statistical Association, 106(495):919, September 2011. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Dunson:2014:C

[Dun14] David B. Dunson. Comment. Journal of the American Statistical Association, 109(507):890–891, 2014. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Dupuis:2012:MWE

[Dup12] Debbie J. Dupuis. Modeling waves of extreme temperature: The changing tails of four cities. Journal of the American Statistical Association, 107(497):24–39, 2012. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Delgado:2011:APT

[DV11] Miguel A. Delgado and Carlos Velasco. An asymptotically pivotal transform of the residuals sample autocorrelations with application to model checking. Journal of the American Statistical Association, 106(495):946–958, September 2011. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Dupuis:2011:FRM

[DVF11] Debbie J. Dupuis and Maria-Pia Victoria-Feser. Fast robust model selection in large datasets. Journal of the American Statistical Association, 106(493):203–212, March 2011. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Dette:2014:SCC

[DVV14] Holger Dette, Ria Van Hecke, and Stanislav Volgushev. Some comments on copula-based regression. Journal of the American Statistical Association, 109(507):1319–1324, 2014. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ding:2011:FRP

[DW11] Xiaobo Ding and Qihua Wang. Fusion-refinement procedure for dimension reduction with missing response at random. Journal of the American Statistical Association, 106(495):1193–1207, September 2011. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Danilov:2012:REM

[DYZ12] Mike Danilov, Víctor J. Yohai, and Ruben H. Zamar. Robust estimation of multivariate location and scatter in the presence of missing data. Journal of the American Statistical Association, 107(499):1178–1186, 2012. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Delaigle:2015:NPE

[DZ15] Aurore Delaigle and Wen-Xin Zhou. Nonparametric and parametric estimators of prevalence from group testing data with aggregated covariates. Journal of the American Statistical Association, 110(512):1785–1796, 2015. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/01621459.2015.
1054491.

Deng:2019:CAR

[DZ19] Shirong Deng and Xingqiu Zhao. Covariate-adjusted regression for distorted longitudinal data with informative observation times. Journal of the American Statistical Association, 114(527): 1241–1250, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Einmahl:2019:LHL

[EEdH19] Jesson J. Einmahl, John H. J. Einmahl, and Laurens de Haan. Limits to human life span through extreme value theory. Journal of the American Statistical Association, 114(527):1075–1080, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Efromovich:2010:DRA

[Efr10a] Sam Efromovich. Dimension reduction and adaptation in conditional density estimation. Journal of the American Statistical Association, 105(490):761–774, June 2010. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Efron:2010:CVA

[Efr10b] Bradley Efron. Correlated z-values and the accuracy of largescale statistical estimates. Journal of the American Statistical Association, 105(491):1042–1055, September 2010. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Efron:2010:RCV

[Efr10c] Bradley Efron. Rejoinder: “Correlated z-Values and the Accuracy of Large-Scale Statistical Estimates”. Journal of the American Statistical Association, 105(491):1067–1069, September 2010. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Efromovich:2011:NRP

[Efr11a] Sam Efromovich. Nonparametric regression with predictors missing at random. Journal of the American Statistical Association, 106(493):306–319, March 2011. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Efron:2011:TFS

[Efr11b] Bradley Efron. Tweedie’s formula and selection bias. Journal of the American Statistical Association, 106(496):1602–1614, December 2011. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Efron:2014:EAA

[Efr14a] Bradley Efron. Estimation and accuracy after model selection. Journal of the American Statistical Association, 109(507):991– 1007, 2014. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Efron:2014:R

[Efr14b] Bradley Efron. Rejoinder. Journal of the American Statistical Association, 109(507):1021–1022, 2014. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Escanciano:2019:QRI

[EG19] J. C. Escanciano and S. C. Goh. Quantile-regression inference with adaptive control of size. Journal of the American Statistical Association, 114(527):1382–1393, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Egami:2019:CIF

[EI19] Naoki Egami and Kosuke Imai. Causal interaction in factorial experiments: Application to conjoint analysis. Journal of the American Statistical Association, 114(526):529–540, 2019. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ertefaie:2018:QET

[ESR18] Ashkan Ertefaie, Dylan S. Small, and Paul R. Rosenbaum. Quantitative evaluation of the trade-off of strengthened instruments and sample size in observational studies. Journal of the American Statistical Association, 113(523):1122–1134, 2018. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Egleston:2017:LCS

[EUW17] Brian L. Egleston, Robert G. Uzzo, and Yu-Ning Wong. Latent class survival models linked by principal stratification to investigate heterogeneous survival subgroups among individuals with early-stage kidney cancer. Journal of the American Statistical Association, 112(518):534–546, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Farrington:2011:SCC

[FAIW^{+}11] C. Paddy Farrington, Karim Anaya-Izquierdo, Heather J. Whitaker, Mounia N. Hocine, Ian Douglas, and Liam Smeeth. Self-controlled case series analysis with event-dependent observation periods. Journal of the American Statistical Association, 106(494):417–426, June 2011. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Farah:2014:BEC

[FBCA14] Marian Farah, Paul Birrell, Stefano Conti, and Daniela De Angelis. Bayesian emulation and calibration of a dynamic epidemic model for A/H1N1 influenza. Journal of the American Statistical Association, 109(508):1398–1411, 2014. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Finley:2011:HMQ

[FBM11] Andrew O. Finley, Sudipto Banerjee, and David W. MacFarlane. A hierarchical model for quantifying forest variables over large heterogeneous landscapes with uncertain forest areas. Journal of the American Statistical Association, 106(493):31–48, March 2011. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Franks:2015:ESC

[FCDA15] Alexander M. Franks, Gábor Csárdi, D. Allan Drummond, and Edoardo M. Airoldi. Estimating a structured covariance matrix from multilab measurements in high-throughput biology. Journal of the American Statistical Association, 110(509):27–44, 2015. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Fisher:2016:FEB

[FCSZ16] Aaron Fisher, Brian Caffo, Brian Schwartz, and Vadim Zipunnikov. Fast, exact bootstrap principal component analysis for p > 1 million. Journal of the American Statistical Association, 111(514):846–??, 2016. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Fong:2010:BVM

[FDPS10] Duncan K. H. Fong, Wayne S. DeSarbo, Joonwook Park, and Crystal J. Scott. A Bayesian vector multidimensional scaling procedure for the analysis of ordered preference data. Journal of the American Statistical Association, 105(490):482–492, June 2010. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Fan:2013:LTM

[FF13] Chunpeng Fan and Jason P. Fine. Linear transformation model with parametric covariate transformations. Journal of the American Statistical Association, 108(502):701–712, 2013. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Fonseca:2017:DMS

[FF17a] Thaís C. O. Fonseca and Marco A. R. Ferreira. Dynamic multiscale spatiotemporal models for Poisson data. Journal of the American Statistical Association, 112(517):215–234, 2017. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Foti:2017:CNB

[FF17b] Nicholas J. Foti and Emily B. Fox. Comment: Nonparametric Bayes modeling of populations of networks. Journal of the American Statistical Association, 112(520):1539–1543, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Fan:2016:FAN

[FFJT16] Jianqing Fan, Yang Feng, Jiancheng Jiang, and Xin Tong. Feature augmentation via nonparametrics and selection (FANS) in high-dimensional classification. Journal of the American Statistical Association, 111(513):275–287, 2016. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Fan:2011:NIS

[FFS11] Jianqing Fan, Yang Feng, and Rui Song. Nonparametric independence screening in sparse ultra-high-dimensional additive models. Journal of the American Statistical Association, 106(494): 544–557, June 2011. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Fiecas:2017:SEM

[FFvSK17] Mark Fiecas, Jürgen Franke, Rainer von Sachs, and Joseph Tadjuidje Kamgaing. Shrinkage estimation for multivariate hidden Markov models. Journal of the American Statistical Association, 112(517):424–435, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Fisher:2012:NWP

[FG12] Thomas J. Fisher and Colin M. Gallagher. New weighted portmanteau statistics for time series goodness of fit testing. Journal of the American Statistical Association, 107(498):777–787, 2012. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Friedenberg:2013:SSD

[FG13] David A. Friedenberg and Christopher R. Genovese. Straight to the source: Detecting aggregate objects in astronomical images with proper error control. Journal of the American Statistical Association, 108(502):456–468, 2013. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Frolich:2014:TEM

[FH14] Markus Frölich and Martin Huber. Treatment evaluation with multiple outcome periods under endogeneity and attrition. Journal of the American Statistical Association, 109(508):1697–1711, 2014. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Fosdick:2015:TMD

[FH15] Bailey K. Fosdick and Peter D. Hoff. Testing and modeling dependencies between a network and nodal attributes. Journal of the American Statistical Association, 110(511):1047–1056, 2015. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Fan:2012:EFD

[FHG12a] Jianqing Fan, Xu Han, and Weijie Gu. Estimating false discovery proportion under arbitrary covariance dependence. Journal of the American Statistical Association, 107(499):1019–1035, 2012. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Fan:2012:R

[FHG12b] Jianqing Fan, Xu Han, and Weijie Gu. Rejoinder. Journal of the American Statistical Association, 107(499):1046–1048, 2012. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Fienberg:2015:LE

[FHL15] Stephen E. Fienberg, James S. Hodges, and Liying Luo. Letter to the Editor. Journal of the American Statistical Association, 110(509):457, 2015. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Flynn:2013:ERP

[FHS13] Cheryl J. Flynn, Clifford M. Hurvich, and Jeffrey S. Simonoff. Efficiency for regularization parameter selection in penalized likelihood estimation of misspecified models. Journal of the American Statistical Association, 108(503):1031–1043, 2013. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Frot:2019:GMS

[FJM19] Benjamin Frot, Luke Jostins, and Gilean McVean. Graphical model selection for Gaussian conditional random fields in the presence of latent variables. Journal of the American Statistical Association, 114(526):723–734, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Fronczyk:2014:BNM

[FK14a] Kassandra Fronczyk and Athanasios Kottas. A Bayesian nonparametric modeling framework for developmental toxicity studies. Journal of the American Statistical Association, 109(507): 873–888, 2014. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Fronczyk:2014:R

[FK14b] Kassandra Fronczyk and Athanasios Kottas. Rejoinder. Journal of the American Statistical Association, 109(507):891–893, 2014. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Fan:2018:RHD

[FK18] Jianqing Fan and Donggyu Kim. Robust high-dimensional volatility matrix estimation for high-frequency factor model. Journal of the American Statistical Association, 113(523):1268– 1283, 2018. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Fenske:2011:IRF

[FKH11] Nora Fenske, Thomas Kneib, and Torsten Hothorn. Identifying risk factors for severe childhood malnutrition by boosting additive quantile regression. Journal of the American Statistical Association, 106(494):494–510, June 2011. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Fan:2019:FFA

[FKSZ19] Jianqing Fan, Yuan Ke, Qiang Sun, and Wen-Xin Zhou. FarmTest: Factor-adjusted robust multiple testing with approximate false discovery control. Journal of the American Statistical Association, 114(528):1880–1893, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Frolich:2010:ERT

[FL10] Markus Frölich and Michael Lechner. Exploiting regional treatment intensity for the evaluation of labor market policies. Journal of the American Statistical Association, 105(491):1014–1029, September 2010. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Fan:2013:AER

[FL13] Yingying Fan and Jinchi Lv. Asymptotic equivalence of regularization methods in thresholded parameter space. Journal of the American Statistical Association, 108(503):1044–1061, 2013. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Fukumizu:2014:GBK

[FL14] Kenji Fukumizu and Chenlei Leng. Gradient-based kernel dimension reduction for regression. Journal of the American Statistical Association, 109(505):359–370, 2014. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Fang:2017:MMA

[FLJL17] Ethan X. Fang, Min-Dian Li, Michael I. Jordan, and Han Liu. Mining massive amounts of genomic data: A semiparametric topic modeling approach. Journal of the American Statistical Association, 112(519):921–932, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Fuller:2017:BVE

[FLL17] Wayne A. Fuller, Jason C. Legg, and Yang Li. Bootstrap variance estimation for rejective sampling. Journal of the American Statistical Association, 112(520):1562–1570, 2017. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Fan:2012:VVM

[FLY12] Jianqing Fan, Yingying Li, and Ke Yu. Vast volatility matrix estimation using high-frequency data for portfolio selection. Journal of the American Statistical Association, 107(497):412–428, 2012. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Frees:2011:SIS

[FMC11] Edward W. Frees, Glenn Meyers, and A. David Cummings. Summarizing insurance scores using a Gini index. Journal of the American Statistical Association, 106(495):1085–1098, September 2011. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Fan:2014:NIS

[FMD14] Jianqing Fan, Yunbei Ma, and Wei Dai. Nonparametric independence screening in sparse ultra-high-dimensional varying coefficient models. Journal of the American Statistical Association, 109(507):1270–1284, 2014. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Fogarty:2016:DOI

[FMGS16] Colin B. Fogarty, Mark E. Mikkelsen, David F. Gaieski, and Dylan S. Small. Discrete optimization for interpretable study populations and randomization inference in an observational study of severe sepsis mortality. Journal of the American Statistical Association, 111(514):447–??, 2016. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Fattorini:2018:DBM

[FMP18] L. Fattorini, M. Marcheselli, and L. Pratelli. Design-based maps for finite populations of spatial units. Journal of the American Statistical Association, 113(522):686–697, 2018. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Frumento:2012:EET

[FMPR12] Paolo Frumento, Fabrizia Mealli, Barbara Pacini, and Donald B. Rubin. Evaluating the effect of training on wages in the presence of noncompliance, nonemployment, and missing outcome data. Journal of the American Statistical Association, 107(498):450– 466, 2012. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Forastiere:2016:IEC

[FMV16] Laura Forastiere, Fabrizia Mealli, and Tyler J. VanderWeele. Identification and estimation of causal mechanisms in clustered encouragement designs: Disentangling bed nets using Bayesian principal stratification. Journal of the American Statistical Association, 111(514):510–??, 2016. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Fiecas:2017:MED

[FO17] Mark Fiecas and Hernando Ombao. Modeling the evolution of dynamic brain processes during an associative learning experiment. Journal of the American Statistical Association, 111(516): 1440–1453, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Freyermuth:2010:TSW

[FOvS10] Jean-Marc Freyermuth, Hernando Ombao, and Rainer von Sachs. Tree-structured wavelet estimation in a mixed effects model for spectra of replicated time series. Journal of the American Statistical Association, 105(490):634–646, June 2010. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Faes:2011:VBI

[FOW11] C. Faes, J. T. Ormerod, and M. P. Wand. Variational Bayesian inference for parametric and nonparametric regression with missing data. Journal of the American Statistical Association, 106 (495):959–971, September 2011. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Fan:2018:MTV

[FPLM18] Minjie Fan, Debashis Paul, Thomas C. M. Lee, and Tomoko Matsuo. Modeling tangential vector fields on a sphere. Journal of the American Statistical Association, 113(524):1625–1636, 2018. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Finucane:2015:R

[FPSE15a] Mariel M. Finucane, Christopher J. Paciorek, Gretchen A. Stevens, and Majid Ezzati. Rejoinder. Journal of the American Statistical Association, 110(511):906–909, 2015. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Finucane:2015:SBD

[FPSE15b] Mariel M. Finucane, Christopher J. Paciorek, Gretchen A. Stevens, and Majid Ezzati. Semiparametric Bayesian density estimation with disparate data sources: A meta-analysis of global childhood undernutrition. Journal of the American Statistical Association, 110(511):889–901, 2015. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Field:2010:BRE

[FPW10] C. A. Field, Zhen Pang, and A. H. Welsh. Bootstrapping robust estimates for clustered data. Journal of the American Statistical Association, 105(492):1606–1616, December 2010. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Fearnhead:2019:CDP

[FR19] Paul Fearnhead and Guillem Rigaill. Changepoint detection in the presence of outliers. Journal of the American Statistical Association, 114(525):169–183, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Frandsen:2015:TEC

[Fra15] Brigham R. Frandsen. Treatment effects with censoring and endogeneity. Journal of the American Statistical Association, 110(512):1745–1752, 2015. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/01621459.2015.
1017577.

Farjat:2017:OSD

[FRG^{+}17] Alfredo Farjat, Brian J. Reich, Joseph Guinness, Ross Whetten, Steven McKeand, and Fikret Isik. Optimal seed deployment under climate change using spatial models: Application to loblolly pine in the Southeastern US. Journal of the American Statistical Association, 112(519):909–920, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Fogarty:2017:SAM

[FS17] Colin B. Fogarty and Dylan S. Small. Sensitivity analysis for multiple comparisons in matched observational studies through quadratically constrained linear programming. Journal of the American Statistical Association, 111(516):1820–1830, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Francom:2019:IAR

[FSB^{+}19] Devin Francom, Bruno Sansó, Vera Bulaevskaya, Donald Lucas, and Matthew Simpson. Inferring atmospheric release characteristics in a large computer experiment using Bayesian adaptive splines. Journal of the American Statistical Association, 114(528):1450–1465, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Fan:2017:CPA

[FSL17] Ailin Fan, Rui Song, and Wenbin Lu. Change-plane analysis for subgroup detection and sample size calculation. Journal of the American Statistical Association, 112(518):769–778, 2017. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Fuglstad:2019:CPP

[FSLR19] Geir-Arne Fuglstad, Daniel Simpson, Finn Lindgren, and Håvard Rue. Constructing priors that penalize the complexity of Gaussian random fields. Journal of the American Statistical Association, 114(525):445–452, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Fogarty:2017:RIS

[FSMS17] Colin B. Fogarty, Pixu Shi, Mark E. Mikkelsen, and Dylan S. Small. Randomization inference and sensitivity analysis for composite null hypotheses with binary outcomes in matched observational studies. Journal of the American Statistical Association, 112(517):321–331, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Fox:2016:MMN

[FSS^{+}16] Eric W. Fox, Martin B. Short, Frederic P. Schoenberg, Kathryn D. Coronges, and Andrea L. Bertozzi. Modeling email networks and inferring leadership using self-exciting point processes. Journal of the American Statistical Association, 111 (514):564–??, 2016. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Feng:2012:MTC

[FTM12] Dai Feng, Luke Tierney, and Vincent Magnotta. MRI tissue classification using high-resolution Bayesian hidden Markov normal mixture models. Journal of the American Statistical Association, 107(497):102–119, 2012. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Fan:2015:MAI

[FTZ15] Jianqing Fan, Xin Tong, and Yao Zeng. Multi-agent inference in social networks: A finite population learning approach. Journal of the American Statistical Association, 110(509):149–158, 2015. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Fu:2015:R

[Fu15] Wenjiang J. Fu. Reply. Journal of the American Statistical Association, 110(509):458, 2015. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Fu:2016:CEC

[Fu16] Wenjiang Fu. Constrained estimators and consistency of a regression model on a Lexis diagram. Journal of the American Statistical Association, 111(513):180–199, 2016. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Fukumoto:2015:WHD

[Fuk15] Kentaro Fukumoto. What happens depends on when it happens: Copula-based ordered event history analysis of civil war duration and outcome. Journal of the American Statistical Association, 110(509):83–92, 2015. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Feng:2013:PSD

[FWYZ13] Zhenghui Feng, Xuerong Meggie Wen, Zhou Yu, and Lixing Zhu. On partial sufficient dimension reduction with applications to partially linear multi-index models. Journal of the American Statistical Association, 108(501):237–246, 2013. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Fan:2017:MQR

[FXZ17] Jianqing Fan, Lingzhou Xue, and Hui Zou. Multitask quantile regression under the transnormal model. Journal of the American Statistical Association, 111(516):1726–1735, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Fan:2017:C

[FY17] Jun Fan and Ming Yuan. Comment. Journal of the American Statistical Association, 111(516):1524–1525, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [CZK17a, CZK17b].

Fu:2013:LSC

[FZ13] Fei Fu and Qing Zhou. Learning sparse causal Gaussian networks with experimental intervention: Regularization and coordinate descent. Journal of the American Statistical Association, 108 (501):288–300, 2013. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Feng:2016:ETV

[FZ16] Xingdong Feng and Liping Zhu. Estimation and testing of varying coefficients in quantile regression. Journal of the American Statistical Association, 111(513):266–274, 2016. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Feng:2016:MSB

[FZW16] Long Feng, Changliang Zou, and Zhaojun Wang. Multivariatesign-based high-dimensional tests for the two-sample location problem. Journal of the American Statistical Association, 111 (514):721–??, 2016. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Fan:2012:VPS

[FZY12] Jianqing Fan, Jingjin Zhang, and Ke Yu. Vast portfolio selection with gross-exposure constraints. Journal of the American Statistical Association, 107(498):592–606, 2012. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Gaugler:2012:MED

[GA12] Trent Gaugler and Michael G. Akritas. Mixed effects designs: The symmetry assumption and missing data. Journal of the American Statistical Association, 107(499):1230–1238, 2012. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Gutman:2013:BPF

[GAZ13] Roee Gutman, Christopher C. Afendulis, and Alan M. Zaslavsky. A Bayesian procedure for file linking to analyze end-of-life medical costs. Journal of the American Statistical Association, 108 (501):34–47, 2013. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ghosh:2010:STA

[GBDL10] Sujit K. Ghosh, Prakash V. Bhave, Jerry M. Davis, and Hyeyoung Lee. Spatio-temporal analysis of total nitrate concentrations using dynamic statistical models. Journal of the American Statistical Association, 105(490):538–551, June 2010. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Geng:2019:PCD

[GBP19] Junxian Geng, Anirban Bhattacharya, and Debdeep Pati. Probabilistic community detection with unknown number of communities. Journal of the American Statistical Association, 114(526): 893–905, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Gaynanova:2016:SSE

[GBW16] Irina Gaynanova, James G. Booth, and Martin T. Wells. Simultaneous sparse estimation of canonical vectors in the p ≥ N setting. Journal of the American Statistical Association, 111 (514):696–??, 2016. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Guo:2017:DSH

[GBZ17] Shaojun Guo, John Leigh Box, and Wenyang Zhang. A dynamic structure for high-dimensional covariance matrices and its application in portfolio allocation. Journal of the American Statistical Association, 112(517):235–253, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ghosh:2011:RBB

[GC11] Joyee Ghosh and Merlise A. Clyde. Rao–Blackwellization for Bayesian variable selection and model averaging in linear and binary regression: a novel data augmentation approach. Journal of the American Statistical Association, 106(495):1041–1052, September 2011. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Gelman:2017:SNS

[GC17] Andrew Gelman and John Carlin. Some natural solutions to the p-value communication problem — and why they won’t work.

Journal of the American Statistical Association, 112(519):899– 901, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Gregory:2015:TST

[GCBL15] Karl Bruce Gregory, Raymond J. Carroll, Veerabhadran Baladandayuthapani, and Soumendra N. Lahiri. A two-sample test for equality of means in high dimension. Journal of the American Statistical Association, 110(510):837–849, 2015. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Gellar:2014:VDF

[GCNC14] Jonathan E. Gellar, Elizabeth Colantuoni, Dale M. Needham, and Ciprian M. Crainiceanu. Variable-domain functional regression for modeling ICU data. Journal of the American Statistical Association, 109(508):1425–1439, 2014. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Guhaniyogi:2015:BCR

[GD15] Rajarshi Guhaniyogi and David B. Dunson. Bayesian compressed regression. Journal of the American Statistical Association, 110(512):1500–1514, 2015. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/01621459.2014.969425.

Gutman:2011:RME

[GDCL11] Roee Gutman, Gayle DeDe, David Caplan, and Jun S. Liu. Rasch model and its extensions for analysis of aphasic deficits in syntactic comprehension. Journal of the American Statistical Association, 106(496):1304–1316, December 2011. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Guerrier:2019:SBB

[GDLMVF19] Stéphane Guerrier, Elise Dupuis-Lozeron, Yanyuan Ma, and Maria-Pia Victoria-Feser. Simulation-based bias correction methods for complex models. Journal of the American Statistical Association, 114(525):146–157, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Gaskins:2017:BMN

[GDM17] J. T. Gaskins, M. J. Daniels, and B. H. Marcus. Bayesian methods for nonignorable dropout in joint models in smoking cessation studies. Journal of the American Statistical Association, 111(516):1454–1465, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Garcia-donato:2013:SSB

[GdMb13] G. García-donato and M. A. Martínez-beneito. On sampling strategies in Bayesian variable selection problems with large model spaces. Journal of the American Statistical Association, 108(501):340–352, 2013. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Greven:2011:AEC

[GDZ11] Sonja Greven, Francesca Dominici, and Scott Zeger. An approach to the estimation of chronic air pollution effects using spatio-temporal information. Journal of the American Statistical Association, 106(494):396–406, June 2011. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Geenens:2014:PTK

[Gee14] Gery Geenens. Probit transformation for kernel density estimation on the unit interval. Journal of the American Statistical Association, 109(505):346–358, 2014. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Geller:2011:SAE

[Gel11] Nancy L. Geller. Statistics: An all-encompassing discipline. Journal of the American Statistical Association, 106(496):1225– 1229, December 2011. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Guinness:2018:CCE

[GH18] Joseph Guinness and Dorit Hammerling. Compression and conditional emulation of climate model output. Journal of the American Statistical Association, 113(521):56–67, 2018. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Guo:2016:SLH

[GHJZ16] Jianhua Guo, Jianchang Hu, Bing-Yi Jing, and Zhen Zhang. Spline-lasso in high-dimensional linear regression. Journal of the American Statistical Association, 111(513):288–297, 2016. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Gabrys:2010:TEC

[GHK10] Robertas Gabrys, Lajos Horváth, and Piotr Kokoszka. Tests for error correlation in the functional linear model. Journal of the American Statistical Association, 105(491):1113–1125, September 2010. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Gorfine:2013:FMF

[GHP13] Malka Gorfine, Li Hsu, and Giovanni Parmigiani. Frailty models for familial risk with application to breast cancer. Journal of the American Statistical Association, 108(504):1205–1215, 2013. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Gile:2011:IIR

[Gil11] Krista J. Gile. Improved inference for respondent-driven sampling data with application to HIV prevalence estimation. Journal of the American Statistical Association, 106(493):135–146, March 2011. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ganong:2018:PTR

[GJ18a] Peter Ganong and Simon Jäger. A permutation test for the regression kink design. Journal of the American Statistical Association, 113(522):494–504, 2018. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Genton:2018:C

[GJ18b] Marc G. Genton and Jaehong Jeong. Comment. Journal of the American Statistical Association, 113(521):176–178, 2018. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Guarniero:2017:IAP

[GJL17] Pieralberto Guarniero, Adam M. Johansen, and Anthony Lee. The iterated auxiliary particle filter. Journal of the American Statistical Association, 112(520):1636–1647, 2017. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Goos:2016:ODM

[GJS16] Peter Goos, Bradley Jones, and Utami Syafitri. I-optimal design of mixture experiments. Journal of the American Statistical Association, 111(514):899–911, 2016. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Goldberg:2011:CAC

[GK11] Yair Goldberg and Michael R. Kosorok. Comment: “Adaptive Confidence Intervals for the Test Error in Classification”. Journal of the American Statistical Association, 106(495):920–924, September 2011. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Goeva:2017:C

[GK17] Aleksandrina Goeva and Eric D. Kolaczyk. Comment. Journal of the American Statistical Association, 111(516):1405–1408, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [AB17a].

Glynn:2018:FDV

[GK18] Adam N. Glynn and Konstantin Kashin. Front-door versus backdoor adjustment with unmeasured confounding: Bias formulas for front-door and hybrid adjustments with application to a job training program. Journal of the American Statistical Association, 113(523):1040–1049, 2018. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ghosh:2019:HDP

[GKM19] Satyajit Ghosh, Kshitij Khare, and George Michailidis. Highdimensional posterior consistency in Bayesian vector autoregressive models. Journal of the American Statistical Association, 114 (526):735–748, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Gneiting:2010:MCC

[GKS10] Tilmann Gneiting, William Kleiber, and Martin Schlather. Matérn cross-covariance functions for multivariate random fields. Journal of the American Statistical Association, 105(491):1167– 1177, September 2010. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Gupta:2014:C

[GL14] Shuva Gupta and S. N. Lahiri. Comment. Journal of the American Statistical Association, 109(507):1013–1015, 2014. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Galvao:2013:ECQ

[GLL13] Antonio F. Galvao, Carlos Lamarche, and Luiz Renato Lima. Estimation of censored quantile regression for panel data with fixed effects. Journal of the American Statistical Association, 108(503):1075–1089, 2013. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Guo:2015:GDR

[GLLL15] Zifang Guo, Lexin Li, Wenbin Lu, and Bing Li. Groupwise dimension reduction via envelope method. Journal of the American Statistical Association, 110(512):1515–1527, 2015. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/01621459.2014.970687.

Giannone:2019:PLR

[GLP19] Domenico Giannone, Michele Lenza, and Giorgio E. Primiceri. Priors for the long run. Journal of the American Statistical Association, 114(526):565–580, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Guan:2016:C

[GLR16] Qian Guan, Eric B. Laber, and Brian J. Reich. Comment. Journal of the American Statistical Association, 111(515):936–942, 2016. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Guan:2011:CDT

[GLS11] Yongtao Guan, Yehua Li, and Rajita Sinha. Cocaine dependence treatment data: Methods for measurement error problems with predictors derived from stationary stochastic processes. Journal of the American Statistical Association, 106(494):480–493, June 2011. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Graham:2014:QCE

[GMS14] Daniel J. Graham, Emma J. McCoy, and David A. Stephens. Quantifying causal effects of road network capacity expansions on traffic volume and density via a mixed model propensity score estimator. Journal of the American Statistical Association, 109(508):1440–1449, 2014. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Gneiting:2011:MEP

[Gne11] Tilmann Gneiting. Making and evaluating point forecasts. Journal of the American Statistical Association, 106(494):746–762, June 2011. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Gan:2019:BRG

[GNL19] Lingrui Gan, Naveen N. Narisetty, and Feng Liang. Bayesian regularization for graphical models with unequal shrinkage. Journal of the American Statistical Association, 114(527):1218–1231, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Godolphin:2017:LBV

[God17] J. D. Godolphin. A link between the E-value and the robustness of block designs. Journal of the American Statistical Association, 111(516):1736–1745, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Gu:2014:BMM

[GPD14] Kelvin Gu, Debdeep Pati, and David B. Dunson. Bayesian multiscale modeling of closed curves in point clouds. Journal of the American Statistical Association, 109(508):1481–1494, 2014. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Genovese:2012:GNF

[GPPVW12] Christopher R. Genovese, Marco Perone-Pacifico, Isabella Verdinelli, and Larry Wasserman. The geometry of nonparametric filament estimation. Journal of the American Statistical Association, 107(498):788–799, 2012. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ghosal:2011:PFD

[GR11] Subhashis Ghosal and Anindya Roy. Predicting false discovery proportion under dependence. Journal of the American Statistical Association, 106(495):1208–1218, September 2011. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Glynn:2010:RCE

[GRH10] Adam N. Glynn, Thomas S. Richardson, and Mark S. Handcock. Resolving contested elections: The limited power of post-vote vote-choice data. Journal of the American Statistical Association, 105(489):84–91, March 2010. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Grimmer:2013:C

[Gri13] Justin Grimmer. Comment. Journal of the American Statistical Association, 108(503):770–771, 2013. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [Tad13a].

George:2017:MRE

[GRR^{+}17] E. I. George, V. Rocková, P. R. Rosenbaum, V. A. Satopää, and J. H. Silber. Mortality rate estimation and standardization for public reporting: Medicare’s hospital compare. Journal of the American Statistical Association, 112(519):933–947, 2017. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Gao:2010:CLB

[GS10] Xin Gao and Peter X.-K. Song. Composite likelihood Bayesian information criteria for model selection in high-dimensional data. Journal of the American Statistical Association, 105(492):1531– 1540, December 2010. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Garg:2017:HLH

[GS17a] Vikram V. Garg and Roy H. Stogner. Hierarchical latin hypercube sampling. Journal of the American Statistical Association, 112(518):673–682, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Greven:2017:C

[GS17b] Sonja Greven and Fabian Scheipl. Comment. Journal of the American Statistical Association, 111(516):1568–1573, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [WPS17b, WPS17a].

Gao:2019:VCP

[GSDR19] Zhenguo Gao, Zuofeng Shang, Pang Du, and John L. Robertson. Variance change point detection under a smoothly-changing mean trend with application to liver procurement. Journal of the American Statistical Association, 114(526):773–781, 2019. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Gilbert:2013:SAP

[GSH13] Peter B. Gilbert, Bryan E. Shepherd, and Michael G. Hudgens. Sensitivity analysis of per-protocol time-to-event treatment efficacy in randomized clinical trials. Journal of the American Statistical Association, 108(503):789–800, 2013. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Guerrier:2013:WVB

[GSSVF13] Stéphane Guerrier, Jan Skaloud, Yannick Stebler, and Maria-Pia Victoria-Feser. Wavelet-variance-based estimation for composite stochastic processes. Journal of the American Statistical Association, 108(503):1021–1030, 2013. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Guha:2010:PSC

[Guh10] Subharup Guha. Posterior simulation in countable mixture models for large datasets. Journal of the American Statistical Association, 105(490):775–786, June 2010. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Gelman:2014:C

[GV14] Andrew Gelman and Aki Vehtari. Comment. Journal of the American Statistical Association, 109(507):1015–1016, 2014. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Galvao:2015:USE

[GW15] Antonio F. Galvao and Liang Wang. Uniformly semiparametric efficient estimation of treatment effects with a continuous treatment. Journal of the American Statistical Association, 110(512): 1528–1542, 2015. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/01621459.2014.978005.

Guo:2019:OEG

[GWCL19] Zijian Guo, Wanjie Wang, T. Tony Cai, and Hongzhe Li. Optimal estimation of genetic relatedness in high-dimensional linear models. Journal of the American Statistical Association, 114 (525):358–369, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ghosh:2013:UAS

[GWZ13] Arpita Ghosh, Fred A. Wright, and Fei Zou. Unified analysis of secondary traits in case-control association studies. Journal of the American Statistical Association, 108(502):566–576, 2013. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Guo:2017:BPI

[GY17] Beibei Guo and Ying Yuan. Bayesian phase I/II biomarkerbased dose finding for precision medicine with molecularly targeted agents. Journal of the American Statistical Association, 112(518):508–520, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Gao:2019:SRA

[GZCL19] Fei Gao, Donglin Zeng, David Couper, and D. Y. Lin. Semiparametric regression analysis of multiple right- and intervalcensored events. Journal of the American Statistical Association, 114(527):1232–1240, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Garcia-Zattera:2012:MMM

[GZJLM12] María José García-Zattera, Alejandro Jara, Emmanuel Lesaffre, and Guillermo Marshall. Modeling of multivariate monotone disease processes in the presence of misclassification. Journal of the American Statistical Association, 107(499):976–989, 2012. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Guo:2019:NTE

[GZZ19] Jia Guo, Bu Zhou, and Jin-Ting Zhang. New tests for equality of several covariance functions for functional data. Journal of the American Statistical Association, 114(527):1251–1263, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hagemann:2017:CRB

[Hag17] Andreas Hagemann. Cluster-robust bootstrap inference in quantile regression models. Journal of the American Statistical Association, 112(517):446–456, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hahn:2012:SPT

[Hah12] Ute Hahn. A Studentized permutation test for the comparison of spatial point patterns. Journal of the American Statistical Association, 107(498):754–764, 2012. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hadjipantelis:2015:UAP

[HAME15] P. Z. Hadjipantelis, J. A. D. Aston, H. G. Müller, and J. P. Evans. Unifying amplitude and phase analysis: A compositional data approach to functional multivariate mixed-effects modeling of Mandarin Chinese. Journal of the American Statistical Association, 110(510):545–559, 2015. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hans:2011:ENR

[Han11] Chris Hans. Elastic net regression modeling with the orthant normal prior. Journal of the American Statistical Association, 106(496):1383–1393, December 2011. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Han:2014:MRE

[Han14] Peisong Han. Multiply robust estimation in regression analysis with missing data. Journal of the American Statistical Association, 109(507):1159–1173, 2014. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Handcock:2017:C

[Han17a] Mark S. Handcock. Comment. Journal of the American Statistical Association, 112(520):1537–1539, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hanks:2017:MSC

[Han17b] Ephraim M. Hanks. Modeling spatial covariance using the limiting distribution of spatio-temporal random walks. Journal of the American Statistical Association, 112(518):497–507, 2017. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hardt:2018:C

[Har18] Moritz Hardt. Comment. Journal of the American Statistical Association, 113(521):207–208, 2018. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Harel:2019:BRM

[Har19] Ofer Harel. Book review: Missing and Modified Data in Nonparametric Estimation: With R Examples. Journal of the American Statistical Association, 114(527):1421–1423, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Herbei:2014:EOC

[HB14] Radu Herbei and L. Mark Berliner. Estimating ocean circulation: An MCMC approach with approximated likelihoods via the Bernoulli factory. Journal of the American Statistical Association, 109(507):944–954, 2014. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hatfield:2012:MBM

[HBHC12] Laura A. Hatfield, Mark E. Boye, Michelle D. Hackshaw, and Bradley P. Carlin. Multilevel Bayesian models for survival times and longitudinal patient-reported outcomes with many zeros. Journal of the American Statistical Association, 107(499):875– 885, 2012. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Huang:2019:MMV

[HBI^{+}19] Lei Huang, Jiawei Bai, Andrada Ivanescu, Tamara Harris, Mathew Maurer, Philip Green, and Vadim Zipunnikov. Multilevel matrix-variate analysis and its application to accelerometrymeasured physical activity in clinical populations. Journal of the American Statistical Association, 114(526):553–564, 2019. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hokayem:2015:RCN

[HBZ15] Charles Hokayem, Christopher Bollinger, and James P. Ziliak. The role of CPS nonresponse in the measurement of poverty. Journal of the American Statistical Association, 110(511):935– 945, 2015. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hahn:2015:DSS

[HC15a] P. Richard Hahn and Carlos M. Carvalho. Decoupling shrinkage and selection in Bayesian linear models: A posterior summary perspective. Journal of the American Statistical Association, 110 (509):435–448, 2015. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hwang:2015:IBN

[HC15b] Beom Seuk Hwang and Zhen Chen. An integrated Bayesian nonparametric approach for stochastic and variability orders in ROC curve estimation: An application to endometriosis diagnosis. Journal of the American Statistical Association, 110(511): 923–934, 2015. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Huang:2017:ESE

[HC17] Ming-Yueh Huang and Chin-Tsang Chiang. An effective semiparametric estimation approach for the sufficient dimension reduction model. Journal of the American Statistical Association, 112(519):1296–1310, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Han:2016:EDA

[HCCZ16] Sung Won Han, Gong Chen, Myun-Seok Cheon, and Hua Zhong. Estimation of directed acyclic graphs through two-stage adaptive lasso for gene network inference. Journal of the American Statistical Association, 111(515):1004–1019, 2016. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Heller:2018:PSI

[HCKS18] Ruth Heller, Nilanjan Chatterjee, Abba Krieger, and Jianxin Shi. Post-selection inference following aggregate level hypothesis testing in large-scale genomic data. Journal of the American Statistical Association, 113(524):1770–1783, 2018. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hahn:2013:PFM

[HCM13] P. Richard Hahn, Carlos M. Carvalho, and Sayan Mukherjee. Partial factor modeling: Predictor-dependent shrinkage for linear regression. Journal of the American Statistical Association, 108(503):999–1008, 2013. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hardle:2017:LTR

[HCOW17] Wolfgang Karl Härdle, Brenda López Cabrera, Ostap Okhrin, and Weining Wang. Localizing temperature risk. Journal of the American Statistical Association, 111(516):1491–1508, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Huang:2010:AVT

[HD10] Li-Shan Huang and Philip W. Davidson. Analysis of variance and F -tests for partial linear models with applications to environmental health data. Journal of the American Statistical Association, 105(491):991–1004, September 2010. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Huo:2016:MAF

[HDL^{+}16] Zhiguang Huo, Ying Ding, Silvia Liu, Steffi Oesterreich, and George Tseng. Meta-analytic framework for sparse K-means to identify disease subtypes in multiple transcriptomic studies. Journal of the American Statistical Association, 111(513):27–42, 2016. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

He:2017:RSP

[He17] Xu He. Rotated sphere packing designs. Journal of the American Statistical Association, 112(520):1612–1622, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Heller:2010:CCV

[Hel10] Ruth Heller. Comment: “Correlated z-Values and the Accuracy of Large-Scale Statistical Estimates”. Journal of the American Statistical Association, 105(491):1057–1059, September 2010. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hero:2018:C

[Her18] Alfred Hero. Comment. Journal of the American Statistical Association, 113(521):203–204, 2018. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hu:2012:SDB

[HFQ12] Zonghui Hu, Dean A. Follmann, and Jing Qin. Semiparametric double balancing score estimation for incomplete data with ignorable missingness. Journal of the American Statistical Association, 107(497):247–257, 2012. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hornstein:2019:JMC

[HFSZ19] Michael Hornstein, Roger Fan, Kerby Shedden, and Shuheng Zhou. Joint mean and covariance estimation with unreplicated matrix-variate data. Journal of the American Statistical Association, 114(526):682–696, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hao:2018:MSH

[HFZ18] Ning Hao, Yang Feng, and Hao Helen Zhang. Model selection for high-dimensional quadratic regression via regularization. Journal of the American Statistical Association, 113(522):615–625, 2018. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hering:2010:PST

[HG10] Amanda S. Hering and Marc G. Genton. Powering up with spacetime wind forecasting. Journal of the American Statistical Association, 105(489):92–104, March 2010. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Happ:2018:MFP

[HG18] Clara Happ and Sonja Greven. Multivariate functional principal component analysis for data observed on different (dimensional) domains. Journal of the American Statistical Association, 113 (522):649–659, 2018. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Heo:2012:TAV

[HGK12] Giseon Heo, Jennifer Gamble, and Peter T. Kim. Topological analysis of variance and the maxillary complex. Journal of the American Statistical Association, 107(498):477–492, 2012. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hsu:2018:EHF

[HGZ18] Li Hsu, Malka Gorfine, and David Zucker. On estimation of the hazard function from population-based case-control studies. Journal of the American Statistical Association, 113(522):560– 570, 2018. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hong:2010:PFS

[HH10] Hyokyoung Grace Hong and Xuming He. Prediction of functional status for the elderly based on a new ordinal regression model. Journal of the American Statistical Association, 105 (491):930–941, September 2010. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hu:2012:SAS

[HH12] Jianhua Hu and Xuming He. Searching for alternative splicing with a joint model on probe measurability and expression intensities. Journal of the American Statistical Association, 107(499): 935–945, 2012. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hanks:2013:CTM

[HH13] Ephraim M. Hanks and Mevin B. Hooten. Circuit theory and model-based inference for landscape connectivity. Journal of the American Statistical Association, 108(501):22–33, 2013. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hu:2019:EIV

[HHY19] Lixia Hu, Tao Huang, and Jinhong You. Estimation and identification of a varying-coefficient additive model for locally stationary processes. Journal of the American Statistical Association, 114(527):1191–1204, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hannig:2016:GFI

[HILL16] Jan Hannig, Hari Iyer, Randy C. S. Lai, and Thomas C. M. Lee. Generalized fiducial inference: A review and new results. Journal of the American Statistical Association, 111(515):1346– 1361, 2016. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hooten:2017:BFM

[HJ17] Mevin B. Hooten and Devin S. Johnson. Basis function models for animal movement. Journal of the American Statistical Association, 112(518):578–589, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Huynh:2011:FPC

[HJCPV11] Kim P. Huynh, David T. Jacho-Chávez, Robert J. Petrunia, and Marcel Voia. Functional principal component analysis of density families with categorical and continuous data on Canadian entrant manufacturing firms. Journal of the American Statistical Association, 106(495):858–878, September 2011. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hjort:2014:C

[Hjo14] Nils Lid Hjort. Comment. Journal of the American Statistical Association, 109(507):1017–1020, 2014. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Heller:2010:SAC

[HJRS10] Ruth Heller, Shane T. Jensen, Paul R. Rosenbaum, and Dylan S. Small. Sensitivity analysis for the cross-match test, with applications in genomics. Journal of the American Statistical Association, 105(491):1005–1013, September 2010. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Han:2014:SIS

[HL14a] Fang Han and Han Liu. Scale-invariant sparse PCA on highdimensional meta-elliptical data. Journal of the American Statistical Association, 109(505):275–287, 2014. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Harvey:2014:FHT

[HL14b] Andrew Harvey and Alessandra Luati. Filtering with heavy tails. Journal of the American Statistical Association, 109(507):1112– 1122, 2014. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Han:2016:C

[HL16] Peisong Han and Jerald F. Lawless. Comment. Journal of the American Statistical Association, 111(513):118–121, 2016. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Han:2018:EHD

[HL18] Fang Han and Han Liu. ECA: High-dimensional elliptical component analysis in non-Gaussian distributions. Journal of the American Statistical Association, 113(521):252–268, 2018. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Huang:2014:JMC

[HLG14] Hui Huang, Yehua Li, and Yongtao Guan. Joint modeling and clustering paired generalized longitudinal trajectories with application to cocaine abuse treatment data. Journal of the American Statistical Association, 109(508):1412–1424, 2014. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Harchaoui:2010:MCP

[HLL10] Z. Harchaoui and C. Lévy-Leduc. Multiple change-point estimation with a total variation penalty. Journal of the American Statistical Association, 105(492):1480–1493, December 2010. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

He:2018:DRV

[HLMH18] Kejun He, Heng Lian, Shujie Ma, and Jianhua Z. Huang. Dimensionality reduction and variable selection in multivariate varyingcoefficient models with a large number of covariates. Journal of the American Statistical Association, 113(522):746–754, 2018. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Huang:2015:SAO

[HLR15] Weibing Huang, Charles-Albert Lehalle, and Mathieu Rosenbaum. Simulating and analyzing order book data: The queuereactive model. Journal of the American Statistical Association, 110(509):107–122, 2015. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

He:2016:ELR

[HLSY16] Shuyuan He, Wei Liang, Junshan Shen, and Grace Yang. Empirical likelihood for right censored lifetime data. Journal of the American Statistical Association, 111(514):646–??, 2016. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hu:2014:LBF

[HLSZ14] Y. J. Hu, D. Y. Lin, W. Sun, and D. Zeng. A likelihood-based framework for association analysis of allele-specific copy numbers. Journal of the American Statistical Association, 109(508): 1533–1545, 2014. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Huang:2013:NMR

[HLW13] Mian Huang, Runze Li, and Shaoli Wang. Nonparametric mixture of regression models. Journal of the American Statistical Association, 108(503):929–941, 2013. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hallin:2015:EAI

[HM15] Marc Hallin and Chintan Mehta. R-estimation for asymmetric independent component analysis. Journal of the American Statistical Association, 110(509):218–232, 2015. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Held:2016:C

[HM16] Leonhard Held and Stefanie Muff. Comment. Journal of the American Statistical Association, 111(515):1108–1110, 2016. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hahn:2016:BPI

[HMM16] P. Richard Hahn, Jared S. Murray, and Ioanna Manolopoulou. A Bayesian partial identification approach to inferring the prevalence of accounting misconduct. Journal of the American Statistical Association, 111(513):14–26, 2016. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Haaland:2010:SAT

[HMQA10] Ben Haaland, Wanli Min, Peter Z. G. Qian, and Yasuo Amemiya. A statistical approach to thermal management of data centers under steady state and system perturbations. Journal of the American Statistical Association, 105(491):1030–1041, September 2010. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Huang:2014:NEA

[HMW^{+}14] Hui Huang, Xiaomei Ma, Rasmus Waagepetersen, Theodore R. Holford, Rong Wang, Harvey Risch, Lloyd Mueller, and Yongtao Guan. A new estimation approach for combining epidemiological data from multiple sources. Journal of the American Statistical Association, 109(505):11–23, 2014. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hui:2017:JSM

[HMW17] Francis K. C. Hui, Samuel Müller, and A. H. Welsh. Joint selection in mixed models using regularized PQL. Journal of the American Statistical Association, 112(519):1323–1333, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hahn:2018:RBP

[HMW18a] P. Richard Hahn, Ryan Martin, and Stephen G. Walker. On recursive Bayesian predictive distributions. Journal of the American Statistical Association, 113(523):1085–1093, 2018. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hui:2018:SPL

[HMW18b] Francis K. C. Hui, Samuel Müller, and A. H. Welsh. Sparse pairwise likelihood estimation for multivariate longitudinal mixed models. Journal of the American Statistical Association, 113 (524):1759–1769, 2018. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Harnau:2018:DAP

[HN18] Jonas Harnau and Bent Nielsen. Over-dispersed age-period-cohort models. Journal of the American Statistical Association, 113(524):1722–1732, 2018. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hong:2017:PNP

[HNW^{+}17] Chuan Hong, Yang Ning, Shuang Wang, Hao Wu, Raymond J. Carroll, and Yong Chen. PLEMT: A novel pseudolikelihoodbased EM test for homogeneity in generalized exponential tilt mixture models. Journal of the American Statistical Association, 112(520):1393–1404, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hodges:2015:C

[Hod15] Jim Hodges. Comment. Journal of the American Statistical Association, 110(511):903–905, 2015. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hoshino:2013:SBE

[Hos13] Takahiro Hoshino. Semiparametric Bayesian estimation for marginal parametric potential outcome modeling: Application to causal inference. Journal of the American Statistical Association, 108(504):1189–1204, 2013. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hoef:2010:MAA

[HP10a] Jay M. Ver Hoef and Erin E. Peterson. A moving average approach for spatial statistical models of stream networks. Journal of the American Statistical Association, 105(489):6–18, March 2010. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hoef:2010:RCS

[HP10b] Jay M. Ver Hoef and Erin E. Peterson. Rejoinder: Comment: Statistical dependence in stream networks. Journal of the American Statistical Association, 105(489):22–24, March 2010. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hallin:2014:EEP

[HPV14] Marc Hallin, Davy Paindaveine, and Thomas Verdebout. Efficient R-estimation of principal and common principal components. Journal of the American Statistical Association, 109(507): 1071–1083, 2014. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ho:2012:MCB

[HPX12] Qirong Ho, Ankur P. Parikh, and Eric P. Xing. A multiscale community blockmodel for network exploration. Journal of the American Statistical Association, 107(499):916–934, 2012. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Huang:2016:EEC

[HQT16] Chiung-Yu Huang, Jing Qin, and Huei-Ting Tsai. Efficient estimation of the Cox model with auxiliary subgroup survival information. Journal of the American Statistical Association, 111 (514):787–??, 2016. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hero:2011:LSC

[HR11] Alfred Hero and Bala Rajaratnam. Large-scale correlation screening. Journal of the American Statistical Association, 106 (496):1540–1552, December 2011. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Haneuse:2016:C

[HR16] Sebastien Haneuse and Claudia Rivera. Comment. Journal of the American Statistical Association, 111(513):121–122, 2016. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Han:2012:TGE

[HRC12] Summer S. Han, Philip S. Rosenberg, and Nilanjan Chatterjee. Testing for gene-environment and gene-gene interactions under monotonicity constraints. Journal of the American Statistical Association, 107(500):1441–1452, 2012. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hansen:2014:CTA

[HRS14] Ben B. Hansen, Paul R. Rosenbaum, and Dylan S. Small. Clustered treatment assignments and sensitivity to unmeasured biases in observational studies. Journal of the American Statistical Association, 109(505):133–144, 2014. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hof:2017:PRL

[HRZ17] Michel H. Hof, Anita C. Ravelli, and Aeilko H. Zwinderman. A probabilistic record linkage model for survival data. Journal of the American Statistical Association, 112(520):1504–1515, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hall:2012:MDB

[HS12] Peter Hall and Michael G. Schimek. Moderate-deviation-based inference for random degeneration in paired rank lists. Journal of the American Statistical Association, 107(498):661–672, 2012. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Haberman:2013:GRG

[HS13] Shelby J. Haberman and Sandip Sinharay. Generalized residuals for general models for contingency tables with application to item response theory. Journal of the American Statistical Association, 108(504):1435–1444, 2013. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Heaton:2015:AIM

[HSM^{+}15] Matthew J. Heaton, Stephan R. Sain, Andrew J. Monaghan, Olga V. Wilhelmi, and Mary H. Hayden. An analysis of an incomplete marked point pattern of heat-related 911 calls. Journal of the American Statistical Association, 110(509):123–135, 2015. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hsu:2013:EMD

[HSR13] Jesse Y. Hsu, Dylan S. Small, and Paul R. Rosenbaum. Effect modification and design sensitivity in observational studies. Journal of the American Statistical Association, 108(501):135– 148, 2013. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hu:2015:PUA

[HSTP15] Yi-Juan Hu, Wei Sun, Jung-Ying Tzeng, and Charles M. Perou. Proper use of allele-specific expression improves statistical power for cis-eQTL mapping with RNA-Seq data. Journal of the American Statistical Association, 110(511):962–974, 2015. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Huang:2015:CHD

[HSZ15] Chao Huang, Martin Styner, and Hongtu Zhu. Clustering highdimensional landmark-based two-dimensional shape data. Journal of the American Statistical Association, 110(511):946–961, 2015. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Holan:2010:BMM

[HTFK10] Scott H. Holan, Daniell Toth, Marco A. R. Ferreira, and Alan F. Karr. Bayesian multiscale multiple imputation with implications for data confidentiality. Journal of the American Statistical Association, 105(490):564–577, June 2010. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ho:2013:NDP

[HTGT13] Man-Wai Ho, Wanzhu Tu, Pulak Ghosh, and Ram C. Tiwari. A nested Dirichlet process analysis of cluster randomized trial data with application in geriatric care assessment. Journal of the American Statistical Association, 108(501):48–68, 2013. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Huang:2014:JEM

[Hua14] Alan Huang. Joint estimation of the mean and error distribution in generalized linear models. Journal of the American Statistical Association, 109(505):186–196, 2014. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Huang:2017:RMR

[Hua17] Yijian Huang. Restoration of monotonicity respecting in dynamic regression. Journal of the American Statistical Association, 112(518):613–622, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hudgens:2015:C

[Hud15] Michael G. Hudgens. Comment. Journal of the American Statistical Association, 110(512):1345–1347, 2015. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.tandfonline.com/doi/abs/10.1080/01621459.
2015.1033058.

Hung:2011:APB

[Hun11] Ying Hung. Adaptive probability-based Latin hypercube designs. Journal of the American Statistical Association, 106(493):213– 219, March 2011. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hanlon:2011:IQP

[HV11] Bret Hanlon and Anand N. Vidyashankar. Inference for quantitation parameters in polymerase chain reactions via branching processes with random effects. Journal of the American Statistical Association, 106(494):525–533, June 2011. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hooten:2010:SAB

[HW10] Mevin B. Hooten and Christopher K. Wikle. Statistical agentbased models for discrete spatio-temporal systems. Journal of the American Statistical Association, 105(489):236–248, March 2010. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Huser:2019:MSP

[HW19] Raphaël Huser and Jennifer L. Wadsworth. Modeling spatial processes with unknown extremal dependence class. Journal of the American Statistical Association, 114(525):434–444, 2019. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hui:2015:TPS

[HWF15] Francis K. C. Hui, David I. Warton, and Scott D. Foster. Tuning parameter selection for the adaptive lasso using ERIC. Journal of the American Statistical Association, 110(509):262–269, 2015. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hung:2013:HMM

[HWZ^{+}13] Ying Hung, Yijie Wang, Veronika Zarnitsyna, Cheng Zhu, and C. F. Jeff Wu. Hidden Markov models with applications in Cell adhesion experiments. Journal of the American Statistical Association, 108(504):1469–1479, 2013. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Huang:2012:MRM

[HY12] Mian Huang and Weixin Yao. Mixture of regression models with varying mixing proportions: a semiparametric approach. Journal of the American Statistical Association, 107(498):711–724, 2012. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hui:2019:SRU

[HYSM19] Francis K. C. Hui, C. You, H. L. Shang, and Samuel Müller. Semiparametric regression using variational approximations. Journal of the American Statistical Association, 114(528):1765– 1777, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hedayat:2010:OEC

[HZ10] A. S. Hedayat and Wei Zheng. Optimal and efficient crossover designs for test-control study when subject effects are random. Journal of the American Statistical Association, 105(492):1581– 1592, December 2010. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hwang:2013:EBC

[HZ13] J. T. Gene Hwang and Zhigen Zhao. Empirical Bayes confidence intervals for selected parameters in high-dimensional data. Journal of the American Statistical Association, 108(502):607–618, 2013. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hao:2014:ISU

[HZ14] Ning Hao and Hao Helen Zhang. Interaction screening for ultrahigh-dimensional data. Journal of the American Statistical Association, 109(507):1285–1301, 2014. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hu:2015:UFC

[HZH15] Jianhua Hu, Hongjian Zhu, and Feifang Hu. A unified family of covariate-adjusted response-adaptive designs based on efficiency and ethics. Journal of the American Statistical Association, 110 (509):357–367, 2015. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

He:2017:SBT

[HZL^{+}17] Zihuai He, Min Zhang, Seunggeun Lee, Jennifer A. Smith, Sharon L. R. Kardia, V. Diez Roux, and Bhramar Mukherjee. Set-based tests for the gene-environment interaction in longitudinal studies. Journal of the American Statistical Association, 112 (519):966–978, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hilton:2018:MHH

[HZS18] Ross P. Hilton, Yuchen Zheng, and Nicoleta Serban. Modeling heterogeneity in healthcare utilization using massive medical claims data. Journal of the American Statistical Association, 113 (521):111–121, 2018. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Huang:2011:LRT

[HZT11] Lan Huang, Jyoti Zalkikar, and Ram C. Tiwari. A likelihood ratio test based method for signal detection with application to FDA’s drug safety data. Journal of the American Statistical Association, 106(496):1230–1241, December 2011. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hu:2010:FDR

[HZZ10] James X. Hu, Hongyu Zhao, and Harrison H. Zhou. False discovery rate control with groups. Journal of the American Statistical Association, 105(491):1215–1227, September 2010. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Huang:2017:PSS

[HZZ^{+}17] Yuan Huang, Qingzhao Zhang, Sanguo Zhang, Jian Huang, and Shuangge Ma. Promoting similarity of sparsity structures in integrative analysis with penalization. Journal of the American Statistical Association, 112(517):342–350, 2017. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Imai:2019:CCM

[IJ19] Kosuke Imai and Zhichao Jiang. Comment: The Challenges of Multiple Causes. Journal of the American Statistical Association, 114(528):1605–1610, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [WB19a, WB19b].

Ishwaran:2010:HDV

[IKG^{+}10] Hemant Ishwaran, Udaya B. Kogalur, Eiran Z. Gorodeski, Andy J. Minn, and Michael S. Lauer. High-dimensional variable selection for survival data. Journal of the American Statistical Association, 105(489):205–217, March 2010. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Iacus:2011:MMM

[IKP11] Stefano M. Iacus, Gary King, and Giuseppe Porro. Multivariate matching methods that are monotonic imbalance bounding. Journal of the American Statistical Association, 106(493):345– 361, March 2011. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Imai:2011:MRA

[Ima11] Kosuke Imai. Multivariate regression analysis for the item count technique. Journal of the American Statistical Association, 106 (494):407–416, June 2011. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ionides:2012:CCM

[Ion12] Edward L. Ionides. Comment: Cell motility models and inference for dynamic systems. Journal of the American Statistical Association, 107(499):865–868, 2012. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Imai:2015:REI

[IR15] Kosuke Imai and Marc Ratkovic. Robust estimation of inverse probability weights for marginal structural models. Journal of the American Statistical Association, 110(511):1013–1023, 2015. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Irle:2012:IDM

[IS12] Sebastian Irle and Helmut Schäfer. Interim design modifications in time-to-event studies. Journal of the American Statistical Association, 107(497):341–348, 2012. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ing:2014:PSP

[IY14] Ching-Kang Ing and Chiao-Yi Yang. Predictor selection for positive autoregressive processes. Journal of the American Statistical Association, 109(505):243–253, 2014. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ip:2013:POM

[IZR^{+}13] Edward Ip, Qiang Zhang, Jack Rejeski, Tammy Harris, and Stephen Kritchevsky. Partially ordered mixed hidden Markov model for the disablement process of older adults. Journal of the American Statistical Association, 108(502):370–384, 2013. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Jiang:2016:FAD

[JAW16] Ci-Ren Jiang, John A. D. Aston, and Jane-Ling Wang. A functional approach to deconvolve dynamic neuroimaging data. Journal of the American Statistical Association, 111(513):1–13, 2016. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Jeng:2010:OSS

[JCL10] X. Jessie Jeng, T. Tony Cai, and Hongzhe Li. Optimal sparse segment identification with application in copy number variation analysis. Journal of the American Statistical Association, 105(491):1156–1166, September 2010. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Juillard:2017:EUC

[JCRG17] Hélène Juillard, Guillaume Chauvet, and Anne Ruiz-Gazen. Estimation under cross-classified sampling with application to a childhood survey. Journal of the American Statistical Association, 112(518):850–858, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Jung:2014:BDA

[JHH14] Yoonsuh Jung, Jianhua Z. Huang, and Jianhua Hu. Biomarker detection in association studies: Modeling SNPs simultaneously via logistic ANOVA. Journal of the American Statistical Association, 109(508):1355–1367, 2014. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Jiang:2016:VSP

[JHZ16] Yuan Jiang, Yunxiao He, and Heping Zhang. Variable selection with prior information for generalized linear models via the prior LASSO method. Journal of the American Statistical Association, 111(513):355–376, 2016. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Jin:2012:C

[Jin12] Jiashun Jin. Comment. Journal of the American Statistical Association, 107(499):1042–1045, 2012. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Jentsch:2016:HMI

[JK16] Carsten Jentsch and Claudia Kirch. How much information does dependence between wavelet coefficients contain? Journal of the American Statistical Association, 111(515):1330–1345, 2016. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Jing:2011:EJA

[JKL11] Bing-Yi Jing, Xin-Bing Kong, and Zhi Liu. Estimating the jump activity index under noisy observations using high-frequency data. Journal of the American Statistical Association, 106(494): 558–568, June 2011. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Jiang:2015:BPM

[JL15] Bo Jiang and Jun S. Liu. Bayesian partition models for identifying expression quantitative trait loci. Journal of the American Statistical Association, 110(512):1350–1361, 2015. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/01621459.2015.1049746.

Jing:2017:STD

[JLPZ17] Bing-Yi Jing, Zhouping Li, Guangming Pan, and Wang Zhou. On SURE-type double shrinkage estimation. Journal of the American Statistical Association, 111(516):1696–1704, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Jin:2014:UDA

[JLTY14] Ick Hoon Jin, Suyu Liu, Peter F. Thall, and Ying Yuan. Using data augmentation to facilitate conduct of Phase I–II clinical trials with delayed outcomes. Journal of the American Statistical Association, 109(506):525–536, 2014. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Jordan:2019:CED

[JLY19] Michael I. Jordan, Jason D. Lee, and Yun Yang. Communicationefficient distributed statistical inference. Journal of the American Statistical Association, 114(526):668–681, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Jiang:2014:IGV

[JLZ14] Yuan Jiang, Ni Li, and Heping Zhang. Identifying genetic variants for addiction via propensity score adjusted generalized Kendall’s tau. Journal of the American Statistical Association, 109(507):905–930, 2014. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Jones:2014:OSD

[JM14] Bradley Jones and Dibyen Majumdar. Optimal supersaturated designs. Journal of the American Statistical Association, 109 (508):1592–1600, 2014. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Joyce:2014:SMM

[JML^{+}14] Patrick M. Joyce, Donald Malec, Roderick J. A. Little, Aaron Gilary, Alfredo Navarro, and Mark E. Asiala. Statistical modeling methodology for the Voting Rights Act Section 203 Language Assistance Determinations. Journal of the American Statistical Association, 109(505):36–47, 2014. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Jiang:2011:BPS

[JNR11] Jiming Jiang, Thuan Nguyen, and J. Sunil Rao. Best predictive small area estimation. Journal of the American Statistical Association, 106(494):732–745, June 2011. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Jiang:2015:MAM

[JNR15] Jiming Jiang, Thuan Nguyen, and J. Sunil Rao. The E-MS algorithm: Model selection with incomplete data. Journal of the American Statistical Association, 110(511):1136–1147, 2015. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Jobe:2015:CBO

[JP15] J. Marcus Jobe and Michael Pokojovy. A cluster-based outlier detection scheme for multivariate data. Journal of the American Statistical Association, 110(512):1543–1551, 2015. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/01621459.2014.983231.

Jensen:2013:BHM

[JPBM13] Shane T. Jensen, Jared Park, Alexander F. Braunstein, and Jon Mcauliffe. Bayesian hierarchical modeling of the HIV evolutionary response to therapy. Journal of the American Statistical Association, 108(504):1230–1242, 2013. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Jackson:2019:VIS

[JPCD19] Christopher Jackson, Anne Presanis, Stefano Conti, and Daniela De Angelis. Value of information: Sensitivity analysis and research design in Bayesian evidence synthesis. Journal of the American Statistical Association, 114(528):1436–1449, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Johnson:2017:RPS

[JPW^{+}17] Valen E. Johnson, Richard D. Payne, Tianying Wang, Alex Asher, and Soutrik Mandal. On the reproducibility of psychological science. Journal of the American Statistical Association, 112(517):1–10, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Johnson:2012:BMS

[JR12] Valen E. Johnson and David Rossell. Bayesian model selection in high-dimensional settings. Journal of the American Statistical Association, 107(498):649–660, 2012. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Janson:2014:MRM

[JR14] Lucas Janson and Bala Rajaratnam. A methodology for robust multiproxy paleoclimate reconstructions and modeling of temperature conditional quantiles. Journal of the American Statistical Association, 109(505):63–77, 2014. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Jiang:2018:CMM

[JRFN18] Jiming Jiang, J. Sunil Rao, Jie Fan, and Thuan Nguyen. Classified mixed model prediction. Journal of the American Statistical Association, 113(521):269–279, 2018. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Jensen:2011:SBM

[JS11] Shane T. Jensen and Stephen H. Shore. Semiparametric Bayesian modeling of income volatility heterogeneity. Journal of the American Statistical Association, 106(496):1280–1290, December 2011. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Johndrow:2019:MIC

[JSPD19] James E. Johndrow, Aaron Smith, Natesh Pillai, and David B. Dunson. MCMC for imbalanced categorical data. Journal of the American Statistical Association, 114(527):1394–1403, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Jeon:2012:UCK

[JT12] Jooyoung Jeon and James W. Taylor. Using conditional kernel density estimation for wind power density forecasting. Journal of the American Statistical Association, 107(497):66–79, 2012. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Jiang:2019:RAA

[JTF^{+}19] Fei Jiang, Lu Tian, Haoda Fu, Takahiro Hasegawa, and L. J. Wei. Robust alternatives to ANCOVA for estimating the treatment effect via a randomized comparative study. Journal of the American Statistical Association, 114(528):1854–1864, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Jiang:2013:PVC

[JWXJ13] Qian Jiang, Hansheng Wang, Yingcun Xia, and Guohua Jiang. On a principal varying coefficient model. Journal of the American Statistical Association, 108(501):228–236, 2013. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Jiang:2015:NST

[JYL15] Bo Jiang, Chao Ye, and Jun S. Liu. Nonparametric K-sample tests via dynamic slicing. Journal of the American Statistical Association, 110(510):642–653, 2015. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Jiang:2015:ADC

[JZ15] Wenxin Jiang and Yu Zhao. On asymptotic distributions and confidence intervals for LIFT measures in data mining. Journal of the American Statistical Association, 110(512):1717–1725, 2015. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/01621459.2014.993080.

Jeng:2019:ESI

[JZT19] X. Jessie Jeng, Teng Zhang, and Jung-Ying Tzeng. Efficient signal inclusion with genomic applications. Journal of the American Statistical Association, 114(528):1787–1799, 2019. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kim:2018:LVP

[KA18] Sungduk Kim and Paul S. Albert. Latent variable Poisson models for assessing the regularity of circadian patterns over time. Journal of the American Statistical Association, 113(523):992–1002, 2018. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Karwa:2018:C

[Kar18] Vishesh Karwa. Comment. Journal of the American Statistical Association, 113(521):204–207, 2018. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Katzfuss:2017:MRA

[Kat17] Matthias Katzfuss. A multi-resolution approximation for massive spatial datasets. Journal of the American Statistical Association, 112(517):201–214, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kuha:2018:EPD

[KBKS18] Jouni Kuha, Sarah Butt, Myrsini Katsikatsou, and Chris J. Skinner. The effect of probing “don’t know” responses on measurement quality and nonresponse in surveys. Journal of the American Statistical Association, 113(521):26–40, 2018. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kott:2010:UCW

[KC10] Phillip S. Kott and Ted Chang. Using calibration weighting to adjust for nonignorable unit nonresponse. Journal of the American Statistical Association, 105(491):1265–1275, September 2010. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kang:2011:BIS

[KC11] Emily L. Kang and Noel Cressie. Bayesian inference for the spatial random effects model. Journal of the American Statistical Association, 106(495):972–983, September 2011. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kim:2015:SEI

[KCK^{+}15] Hang J. Kim, Lawrence H. Cox, Alan F. Karr, Jerome P. Reiter, and Quanli Wang. Simultaneous edit-imputation for continuous microdata. Journal of the American Statistical Association, 110 (511):987–999, 2015. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kuan:2011:SFA

[KCP^{+}11] Pei Fen Kuan, Dongjun Chung, Guangjin Pan, James A. Thomson, Ron Stewart, and Sündüz Keleş. A statistical framework for the analysis of ChIP-Seq data. Journal of the American Statistical Association, 106(495):891–903, September 2011. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kim:2013:BHP

[KCZ^{+}13] Sungduk Kim, Zhen Chen, Zhiwei Zhang, Bruce G. Simons-Morton, and Paul S. Albert. Bayesian hierarchical Poisson regression models: An application to a driving study with kinematic events. Journal of the American Statistical Association, 108(502):494–503, 2013. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kunihama:2013:BMT

[KD13] Tsuyoshi Kunihama and David B. Dunson. Bayesian modeling of temporal dependence in large sparse contingency tables. Journal of the American Statistical Association, 108(504):1324–1338, 2013. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kundu:2014:BVS

[KD14] Suprateek Kundu and David B. Dunson. Bayes variable selection in semiparametric linear models. Journal of the American Statistical Association, 109(505):437–447, 2014. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kenah:2015:SRR

[Ken15] Eben Kenah. Semiparametric relative-risk regression for infectious disease transmission data. Journal of the American Statistical Association, 110(509):313–325, 2015. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kennedy:2019:NCE

[Ken19] Edward H. Kennedy. Nonparametric causal effects based on incremental propensity score interventions. Journal of the American Statistical Association, 114(526):645–656, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Killick:2012:ODC

[KFE12] R. Killick, P. Fearnhead, and I. A. Eckley. Optimal detection of changepoints with a linear computational cost. Journal of the American Statistical Association, 107(500):1590–1598, 2012. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ke:2015:HP

[KFW15] Zheng Tracy Ke, Jianqing Fan, and Yichao Wu. Homogeneity pursuit. Journal of the American Statistical Association, 110 (509):175–194, 2015. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Krupskii:2018:FCM

[KHG18] Pavel Krupskii, Raphaël Huser, and Marc G. Genton. Factor copula models for replicated spatial data. Journal of the American Statistical Association, 113(521):467–479, 2018. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kennedy:2019:SCE

[KHK19] Edward H. Kennedy, Steve Harris, and Luke J. Keele. Survivor– complier effects in the presence of selection on treatment, with application to a study of prompt ICU admission. Journal of the American Statistical Association, 114(525):93–104, 2019. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kim:2019:BRS

[Kim19] Jae-Kwang Kim. Book review: Statistical Data Fusion. Journal of the American Statistical Association, 114(527):1425–1426, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kato:2010:FDC

[KJ10] Shogo Kato and M. C. Jones. A family of distributions on the circle with links to, and applications arising from, Möbius transformation. Journal of the American Statistical Association, 105 (489):249–262, March 2010. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kang:2011:MAF

[KJNW11] Jian Kang, Timothy D. Johnson, Thomas E. Nichols, and Tor D. Wager. Meta analysis of functional neuroimaging data via Bayesian spatial point processes. Journal of the American Statistical Association, 106(493):124–134, March 2011. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Krivobokova:2010:SCB

[KKC10] Tatyana Krivobokova, Thomas Kneib, and Gerda Claeskens. Simultaneous confidence bands for penalized spline estimators.

Journal of the American Statistical Association, 105(490):852– 863, June 2010. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Klein:2015:BGA

[KKL15] Nadja Klein, Thomas Kneib, and Stefan Lang. Bayesian generalized additive models for location, scale, and shape for zeroinflated and overdispersed count data. Journal of the American Statistical Association, 110(509):405–419, 2015. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Konomi:2017:BTC

[KKLL17] Bledar A. Konomi, Georgios Karagiannis, Kevin Lai, and Guang Lin. Bayesian treed calibration: An application to carbon capture with AX sorbent. Journal of the American Statistical Association, 112(517):37–53, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kidwell:2011:SEW

[KLCT11] Paul Kidwell, Guy Lebanon, and Kevyn Collins-Thompson. Statistical estimation of word acquisition with application to readability prediction. Journal of the American Statistical Association, 106(493):21–30, March 2011. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kaizar:2011:PMT

[KLH11] Eloise E. Kaizar, Yan Li, and Jason C. Hsu. Permutation multiple tests of binary features do not uniformly control error rates. Journal of the American Statistical Association, 106(495):1067– 1074, September 2011. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Koopman:2017:ISV

[KLL17] Siem Jan Koopman, Rutger Lit, and André Lucas. Intraday stochastic volatility in discrete price changes: The dynamic Skellam model. Journal of the American Statistical Association, 112(520):1490–1503, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Katenka:2013:TMT

[KLM13] Natallia Katenka, Elizaveta Levina, and George Michailidis. Tracking multiple targets using binary decisions from wireless sensor networks. Journal of the American Statistical Association, 108(502):398–410, 2013. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kim:2013:PBE

[KLN13] Young Min Kim, Soumendra N. Lahiri, and Daniel J. Nordman. A progressive block empirical likelihood method for time series. Journal of the American Statistical Association, 108(504):1506– 1516, 2013. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kim:2013:UAS

[KLSY13] Jane Paik Kim, Wenbin Lu, Tony Sit, and Zhiliang Ying. A unified approach to semiparametric transformation models under general biased sampling schemes. Journal of the American Statistical Association, 108(501):217–227, 2013. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Koenker:2014:COS

[KM14] Roger Koenker and Ivan Mizera. Convex optimization, shape constraints, compound decisions, and empirical Bayes rules. Journal of the American Statistical Association, 109(506):674– 685, 2014. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kuipers:2017:PMI

[KM17] Jack Kuipers and Giusi Moffa. Partition MCMC for inference on acyclic digraphs. Journal of the American Statistical Association, 112(517):282–299, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kirch:2015:DCM

[KMO15] Claudia Kirch, Birte Muhsal, and Hernando Ombao. Detection of changes in multivariate time series with application to EEG data. Journal of the American Statistical Association, 110(511): 1197–1216, 2015. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kowal:2017:BMF

[KMR17] Daniel R. Kowal, David S. Matteson, and David Ruppert. A Bayesian multivariate functional dynamic linear model. Journal of the American Statistical Association, 112(518):733–744, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Keich:2018:CFI

[KN18] Uri Keich and William Stafford Noble. Controlling the FDR in imperfect matches to an incomplete database. Journal of the American Statistical Association, 113(523):973–982, 2018. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kneib:2017:C

[Kne17] Thomas Kneib. Comment. Journal of the American Statistical Association, 111(516):1563–1565, 2017. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [WPS17b, WPS17a].

Kong:2018:CML

[KNK^{+}18] Shengchun Kong, Bin Nan, John D. Kalbfleisch, Rajiv Saran, and Richard Hirth. Conditional modeling of longitudinal data with terminal event. Journal of the American Statistical Association, 113(521):357–368, 2018. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kang:2012:SSM

[KOL^{+}12] Hakmook Kang, Hernando Ombao, Crystal Linkletter, Nicole Long, and David Badre. Spatio-spectral mixed-effects model for functional magnetic resonance imaging data. Journal of the American Statistical Association, 107(498):568–577, 2012. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kou:2012:MMP

[KOLL12] S. C. Kou, Benjamin P. Olding, Martin Lysy, and Jun S. Liu. A multiresolution method for parameter estimation of diffusion processes. Journal of the American Statistical Association, 107(500): 1558–1574, 2012. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kou:2012:C

[Kou12] Samuel Kou. Comment. Journal of the American Statistical Association, 107(499):868–869, 2012. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Koyama:2010:AMS

[KPBSK10] Shinsuke Koyama, Lucia Castellanos Pérez-Bolde, Cosma Rohilla Shalizi, and Robert E. Kass. Approximate methods for state-space models. Journal of the American Statistical Association, 105(489):170–180, March 2010. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kim:2017:BRM

[KPC^{+}17] Sungmin Kim, Kevin Potter, Peter F. Craigmile, Mario Peruggia, and Trisha Van Zandt. A Bayesian race model for recognition memory. Journal of the American Statistical Association, 112 (517):77–91, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kreider:2012:IES

[KPGJ12] Brent Kreider, John V. Pepper, Craig Gundersen, and Dean Jolliffe. Identifying the effects of SNAP (food stamps) on child health outcomes when participation is endogenous and misreported. Journal of the American Statistical Association, 107 (499):958–975, 2012. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kleiber:2011:GMA

[KRG11] William Kleiber, Adrian E. Raftery, and Tilmann Gneiting. Geostatistical model averaging for locally calibrated probabilistic quantitative precipitation forecasting. Journal of the American Statistical Association, 106(496):1291–1303, December 2011. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Krafty:2017:CSA

[KRS^{+}17] Robert T. Krafty, Ori Rosen, David S. Stoffer, Daniel J. Buysse, and Martica H. Hall. Conditional spectral analysis of replicated multiple time series with application to nocturnal physiology. Journal of the American Statistical Association, 112(520):1405– 1416, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kramer:2011:DFP

[KS11] Nicole Krämer and Masashi Sugiyama. The degrees of freedom of partial least squares regression. Journal of the American Statistical Association, 106(494):697–705, June 2011. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kasahara:2015:TNC

[KS15] Hiroyuki Kasahara and Katsumi Shimotsu. Testing the number of components in normal mixture regression models. Journal of the American Statistical Association, 110(512):1632–1645, 2015. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/01621459.2014.986272.

Katki:2010:UDF

[KSGB10] Hormuzd A. Katki, Christopher L. Sanders, Barry I. Graubard, and Andrew W. Bergen. Using DNA fingerprints to infer familial relationships within NHANES III households. Journal of the American Statistical Association, 105(490):552–563, June 2010. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kurtek:2012:SMC

[KSKD12] Sebastian Kurtek, Anuj Srivastava, Eric Klassen, and Zhaohua Ding. Statistical modeling of curves using shapes and related features. Journal of the American Statistical Association, 107(499): 1152–1165, 2012. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kuhnert:2018:C

[Kuh18] Petra M. Kuhnert. Comment. Journal of the American Statistical Association, 113(521):168–170, 2018. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kong:2015:MRD

[KW15] Linglong Kong and Douglas P. Wiens. Model-robust designs for quantile regression. Journal of the American Statistical Association, 110(509):233–245, 2015. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kong:2015:ALM

[KWG15] Xiangrong Kong, Mei-Cheng Wang, and Ronald Gray. Analysis of longitudinal multivariate outcome data from couples cohort studies: Application to HPV transmission dynamics. Journal of the American Statistical Association, 110(510):472–485, 2015.

CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kong:2019:FIE

[KWX^{+}19] Xinbing Kong, Jiangyan Wang, Jinbao Xing, Chao Xu, and Chao Ying. Factor and idiosyncratic empirical processes. Journal of the American Statistical Association, 114(527):1138–1146, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kalnina:2017:NEL

[KX17] Ilze Kalnina and Dacheng Xiu. Nonparametric estimation of the leverage effect: A trade-off between robustness and efficiency. Journal of the American Statistical Association, 112(517):384– 396, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kong:2019:CCD

[KXZ19] Efang Kong, Yingcun Xia, and Wei Zhong. Composite coefficient of determination and its application in ultrahigh dimensional variable screening. Journal of the American Statistical Association, 114(528):1740–1751, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kim:2011:SEM

[KY11a] Jae Kwang Kim and Cindy Long Yu. A semiparametric estimation of mean functionals with nonignorable missing data. Journal of the American Statistical Association, 106(493):157–165, March 2011. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kim:2011:SAR

[KY11b] Mi-Ok Kim and Yunwen Yang. Semiparametric approach to a random effects quantile regression model. Journal of the American Statistical Association, 106(496):1405–1417, December 2011. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Knox:2019:DIS

[KYBB19] Dean Knox, Teppei Yamamoto, Matthew A. Baum, and Adam J. Berinsky. Design, identification, and sensitivity analysis for patient preference trials. Journal of the American Statistical Association, 114(528):1532–1546, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kang:2016:IVE

[KZCS16] Hyunseung Kang, Anru Zhang, T. Tony Cai, and Dylan S. Small. Instrumental variables estimation with some invalid instruments and its application to Mendelian randomization. Journal of the American Statistical Association, 111(513):132–144, 2016. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lazar:2011:CPV

[Laz11] Nicole A. Lazar. Comment: “Population Value Decomposition, a Framework for the Analysis of Image Populations”. Journal of the American Statistical Association, 106(495):791–796, September 2011. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Li:2017:ATF

[LB17a] Ang Li and Rina Foygel Barber. Accumulation tests for FDR control in ordered hypothesis testing. Journal of the American Statistical Association, 112(518):837–849, 2017. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Linderman:2017:CDN

[LB17b] Scott W. Linderman and David M. Blei. Comment: A discussion of “nonparametric Bayes modeling of populations of networks”. Journal of the American Statistical Association, 112(520):1543– 1547, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Li:2018:BPO

[LB18] Alexander Hanbo Li and Jelena Bradic. Boosting in the presence of outliers: Adaptive classification with nonconvex loss functions. Journal of the American Statistical Association, 113(522):660– 674, 2018. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Langrock:2013:MMN

[LBS13] Roland Langrock, David L. Borchers, and Hans J. Skaug. Markov-modulated nonhomogeneous Poisson processes for modeling detections in surveys of marine mammal abundance. Journal of the American Statistical Association, 108(503):840–851, 2013. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Li:2010:TOF

[LC10] Pengfei Li and Jiahua Chen. Testing the order of a finite mixture. Journal of the American Statistical Association, 105(491): 1084–1092, September 2010. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Luo:2014:SLE

[LC14] Shan Luo and Zehua Chen. Sequential lasso cum EBIC for feature selection with ultra-high dimensional feature space. Journal of the American Statistical Association, 109(507):1229–1240, 2014. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Li:2018:MPG

[LC18] Yingbo Li and Merlise A. Clyde. Mixtures of g-priors in generalized linear models. Journal of the American Statistical Association, 113(524):1828–1845, 2018. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Li:2012:DCN

[LCAL12] Jun Li, Juan A. Cuesta-Albertos, and Regina Y. Liu. DDclassifier: Nonparametric classification procedure based on DDplot. Journal of the American Statistical Association, 107(498): 737–753, 2012. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Liu:2014:MBL

[LCCG14] Ziyue Liu, Anne R. Cappola, Leslie J. Crofford, and Wensheng Guo. Modeling bivariate longitudinal hormone profiles by hierarchical state space models. Journal of the American Statistical Association, 109(505):108–118, 2014. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Li:2014:QAR

[LCF14] Ruosha Li, Yu Cheng, and Jason P. Fine. Quantile association regression models. Journal of the American Statistical Association, 109(505):230–242, 2014. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Li:2018:EBD

[LCFW18] Quefeng Li, Guang Cheng, Jianqing Fan, and Yuyan Wang. Embracing the blessing of dimensionality in factor models. Journal of the American Statistical Association, 113(521):380–389, 2018. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Li:2015:LRT

[LCG^{+}15] Shaoting Li, Jiahua Chen, Jianhua Guo, Bing-Yi Jing, Shui-Ying Tsang, and Hong Xue. Likelihood ratio test for multi-sample mixture model and its application to genetic imprinting. Journal of the American Statistical Association, 110(510):867–877, 2015. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Liang:2014:SSA

[LCL14] Faming Liang, Yichen Cheng, and Guang Lin. Simulated stochastic approximation annealing for global optimization with a square-root cooling schedule. Journal of the American Statistical Association, 109(506):847–863, 2014. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Liu:2018:NML

[LCLZ18] Dandan Liu, Tianxi Cai, Anna Lok, and Yingye Zheng. Nonparametric maximum likelihood estimators of time-dependent accuracy measures for survival outcome under two-stage sampling designs. Journal of the American Statistical Association, 113(522):882–892, 2018. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lin:2010:GMC

[LCM10] Ming Lin, Rong Chen, and Per Mykland. On generating Monte Carlo samples of continuous diffusion bridges. Journal of the American Statistical Association, 105(490):820–838, June 2010. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Liang:2013:RBS

[LCS^{+}13] Faming Liang, Yichen Cheng, Qifan Song, Jincheol Park, and Ping Yang. A resampling-based stochastic approximation method for analysis of large geostatistical data. Journal of the American Statistical Association, 108(501):325–339, 2013. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Li:2016:ESM

[LCT16] Wentao Li, Rong Chen, and Zhiqiang Tan. Efficient sequential Monte Carlo with multiple proposals and control variates. Journal of the American Statistical Association, 111(513):298–313, 2016. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Li:2012:SEC

[LCZ12] Bing Li, Hyonho Chun, and Hongyu Zhao. Sparse estimation of conditional graphical models with application to gene networks. Journal of the American Statistical Association, 107(497):152– 167, 2012. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Li:2014:ASM

[LCZ14] Bing Li, Hyonho Chun, and Hongyu Zhao. On an additive semigraphoid model for statistical networks with application to pathway analysis. Journal of the American Statistical Association, 109(507):1188–1204, 2014. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Linero:2015:FBA

[LD15] Antonio R. Linero and Michael J. Daniels. A flexible Bayesian approach to monotone missing data in longitudinal studies with nonignorable missingness with application to an acute schizophrenia clinical trial. Journal of the American Statistical Association, 110(509):45–55, 2015. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Li:2017:GFF

[LD17] Xinran Li and Peng Ding. General forms of finite population central limit theorems with applications to causal inference. Journal of the American Statistical Association, 112(520):1759–1769, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lee:2015:PCE

[LDGX15] Giwhyun Lee, Yu Ding, Marc G. Genton, and Le Xie. Power curve estimation with multivariate environmental factors for inland and offshore wind farms. Journal of the American Statistical Association, 110(509):56–67, 2015. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Li:2019:RIP

[LDL^{+}19] Xinran Li, Peng Ding, Qian Lin, Dawei Yang, and Jun S. Liu. Randomization inference for peer effects. Journal of the American Statistical Association, 114(528):1651–1664, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lee:2010:CET

[LDS10] Keunbaik Lee, Michael J. Daniels, and Daniel J. Sargent. Causal effects of treatments for informative missing data due to progression/death. Journal of the American Statistical Association, 105 (491):912–929, September 2010. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lee:2016:HMS

[LDSH16] Kyu Ha Lee, Francesca Dominici, Deborah Schrag, and Sebastien Haneuse. Hierarchical models for semicompeting risks data with application to quality of end-of-life care for pancreatic cancer. Journal of the American Statistical Association, 111(515):1075– 1095, 2016. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Leeb:2015:C

[Lee15] Hannes Leeb. Comment. Journal of the American Statistical Association, 110(512):1457–1459, 2015. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/01621459.2015.
1109516.

Lei:2014:AGT

[Lei14] Jing Lei. Adaptive global testing for functional linear models. Journal of the American Statistical Association, 109(506):624– 634, 2014. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lemieux:2015:C

[Lem15] Thomas Lemieux. Comment. Journal of the American Statistical Association, 110(512):1347–1348, 2015. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/01621459.2015.
1054490.

Li:2012:EEE

[LER^{+}12] Li Li, Joseph J. Eron, Heather Ribaudo, Roy M. Gulick, and Brent A. Johnson. Evaluating the effect of early versus late ARV regimen change if failure on an initial regimen: Results from the AIDS clinical trials group study A5095. Journal of the American Statistical Association, 107(498):542–554, 2012. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lin:2015:RMH

[LFL15] Wei Lin, Rui Feng, and Hongzhe Li. Regularization methods for high-dimensional instrumental variables regression with an application to genetical genomics. Journal of the American Statistical Association, 110(509):270–288, 2015. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lin:2011:VRF

[LFU11] Dongyu Lin, Dean P. Foster, and Lyle H. Ungar. VIF regression: a fast regression algorithm for large data. Journal of the American Statistical Association, 106(493):232–247, March 2011. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Li:2013:NIC

[LG13] Bo Li and Marc G. Genton. Nonparametric identification of copula structures. Journal of the American Statistical Association, 108(502):666–675, 2013. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Li:2014:FPC

[LG14] Yehua Li and Yongtao Guan. Functional principal component analysis of spatiotemporal point processes with applications in disease surveillance. Journal of the American Statistical Association, 109(507):1205–1215, 2014. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lei:2018:DFP

[LGR^{+}18] Jing Lei, Max G’Sell, Alessandro Rinaldo, Ryan J. Tibshirani, and Larry Wasserman. Distribution-free predictive inference for regression. Journal of the American Statistical Association, 113(523):1094–1111, 2018. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Liu:2018:BPI

[LGY18] Suyu Liu, Beibei Guo, and Ying Yuan. A Bayesian phase I/II trial design for immunotherapy. Journal of the American Statistical Association, 113(523):1016–1027, 2018. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Liu:2014:LSR

[LH14] Lan Liu and Michael G. Hudgens. Large sample randomization inference of causal effects in the presence of interference. Journal of the American Statistical Association, 109(505):288–301, 2014. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lian:2019:BHS

[LHC19] Qinshu Lian, James S. Hodges, and Haitao Chu. A Bayesian hierarchical summary receiver operating characteristic model for network meta-analysis of diagnostic tests. Journal of the American Statistical Association, 114(527):949–961, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lesage:2019:CTI

[LHD19] Éric Lesage, David Haziza, and Xavier D’Haultfoeuille. A cautionary tale on instrumental calibration for the treatment of nonignorable unit nonresponse in surveys. Journal of the American Statistical Association, 114(526):906–915, 2019. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Li:2017:FVC

[LHH17a] Jialiang Li, Chao Huang, and Zhub Hongtu. A functional varying-coefficient single-index model for functional response data. Journal of the American Statistical Association, 112(519): 1169–1181, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). For the the Alzheimer’s Disease Neuroimaging Initiative.

Li:2017:DAI

[LHH^{+}17b] Xiaodong Li, Xu He, Yuanzhen He, Hui Zhang, Zhong Zhang, and Dennis K. J. Lin. The design and analysis for the icing wind tunnel experiment of a new deicing coating. Journal of the American Statistical Association, 112(520):1417–1429, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lai:2015:GFI

[LHL15] Randy C. S. Lai, Jan Hannig, and Thomas C. M. Lee. Generalized fiducial inference for ultrahigh-dimensional regression. Journal of the American Statistical Association, 110(510):760– 772, 2015. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Li:2018:WWT

[LHS18] Gen Li, Jianhua Z. Huang, and Haipeng Shen. To wait or not to wait: Two-way functional hazards model for understanding waiting in call centers. Journal of the American Statistical Association, 113(524):1503–1514, 2018. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Li:2015:SAL

[LHTB15] Jie Li, Yili Hong, Ram Thapa, and Harold E. Burkhart. Survival analysis of loblolly pine trees with spatially correlated random effects. Journal of the American Statistical Association, 110 (510):486–502, 2015. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Liu:2013:OAG

[LHW^{+}13] Tao Liu, Joseph W. Hogan, Lisa Wang, Shangxuan Zhang, and Rami Kantor. Optimal allocation of gold standard testing under constrained availability: Application to assessment of HIV treatment failure. Journal of the American Statistical Association, 108(504):1173–1188, 2013. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Li:2012:BCH

[Li12a] Longhai Li. Bias-corrected hierarchical Bayesian classification with a selected subset of high-dimensional features. Journal of the American Statistical Association, 107(497):120–134, 2012. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Li:2012:QP

[Li12b] Ta-Hsin Li. Quantile periodograms. Journal of the American Statistical Association, 107(498):765–776, 2012. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lindquist:2012:FCM

[Lin12] Martin A. Lindquist. Functional causal mediation analysis with an application to brain connectivity. Journal of the American Statistical Association, 107(500):1297–1309, 2012. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Linzer:2013:DBF

[Lin13] Drew A. Linzer. Dynamic Bayesian forecasting of Presidential elections in the States. Journal of the American Statistical Association, 108(501):124–134, 2013. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Linero:2018:BRT

[Lin18] Antonio R. Linero. Bayesian regression trees for highdimensional prediction and variable selection. Journal of the American Statistical Association, 113(522):626–636, 2018. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Little:2013:PSM

[Lit13] Roderick J. Little. In praise of simplicity not mathematistry! Ten simple powerful ideas for the statistical scientist. Journal of the American Statistical Association, 108(502):359–369, 2013. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Li:2019:MBS

[LJGZH19] Li Li, Alejandro Jara, María José García-Zattera, and Timothy E. Hanson. Marginal Bayesian semiparametric modeling of mismeasured multivariate interval-censored data. Journal of the American Statistical Association, 114(525):129–145, 2019. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Liang:2016:AEA

[LJSL16] Faming Liang, Ick Hoon Jin, Qifan Song, and Jun S. Liu. An adaptive exchange algorithm for sampling from distributions with intractable normalizing constants. Journal of the American Statistical Association, 111(513):377–393, 2016. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lu:2013:MDE

[LJW13] Luo Lu, Hui Jiang, and Wing H. Wong. Multivariate density estimation by Bayesian sequential partitioning. Journal of the American Statistical Association, 108(504):1402–1410, 2013. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Liu:2014:LEL

[LJZ14] Xu Liu, Hongmei Jiang, and Yong Zhou. Local empirical likelihood inference for varying-coefficient density-ratio models based on case-control data. Journal of the American Statistical Association, 109(506):635–646, 2014. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Louis:2016:C

[LK16] Thomas A. Louis and Niels Keiding. Comment. Journal of the American Statistical Association, 111(513):123–124, 2016. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Li:2019:ABT

[LK19] Zeda Li and Robert T. Krafty. Adaptive Bayesian time-frequency analysis of multivariate time series. Journal of the American Statistical Association, 114(525):453–465, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lin:2013:HDS

[LL13] Wei Lin and Jinchi Lv. High-dimensional sparse additive hazards regression. Journal of the American Statistical Association, 108 (501):247–264, 2013. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Li:2019:RVI

[LL19a] Yang Li and Jun S. Liu. Robust variable and interaction selection for logistic regression and general index models. Journal of the American Statistical Association, 114(525):271–286, 2019. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Liu:2019:GPP

[LL19b] Zhonghua Liu and Xihong Lin. A geometric perspective on the power of principal component association tests in multiple phenotype studies. Journal of the American Statistical Association, 114(527):975–990, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Liu:2014:EMA

[LLgX14] Dungang Liu, Regina Y. Liu, and Min ge Xie. Exact metaanalysis approach for discrete data and its application to 2 × 2 tables with rare events. Journal of the American Statistical Association, 109(508):1450–1465, 2014. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Liu:2010:RBT

[LLL^{+}10] Aiyi Liu, Qizhai Li, Chunling Liu, Kai Yu, and Kai F. Yu. A rank-based test for comparison of multidimensional outcomes. Journal of the American Statistical Association, 105(490):578– 587, June 2010. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lu:2011:HDO

[LLLW11] Tao Lu, Hua Liang, Hongzhe Li, and Hulin Wu. Highdimensional ODEs coupled with mixed-effects modeling techniques for dynamic gene regulatory network identification. Journal of the American Statistical Association, 106(496):1242–1258, December 2011. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Li:2017:SIG

[LLQ17] Pengfei Li, Yukun Liu, and Jing Qin. Semiparametric inference in a genetic mixture model. Journal of the American Statistical Association, 112(519):1250–1260, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Linn:2017:ILQ

[LLS17] Kristin A. Linn, Eric B. Laber, and Leonard A. Stefanski. Interactive Q-learning for quantiles. Journal of the American Statistical Association, 112(518):638–649, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lee:2018:OEC

[LLSS18] Sokbae Lee, Yuan Liao, Myung Hwan Seo, and Youngki Shin. Oracle estimation of a change point in high-dimensional quantile regression. Journal of the American Statistical Association, 113(523):1184–1194, 2018. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Li:2015:QCQ

[LLT15] Guodong Li, Yang Li, and Chih-Ling Tsai. Quantile correlations and quantile autoregressive modeling. Journal of the American Statistical Association, 110(509):246–261, 2015. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Liu:2014:FSV

[LLW14] Jingyuan Liu, Runze Li, and Rongling Wu. Feature selection for varying coefficient models with ultrahigh-dimensional covariates. Journal of the American Statistical Association, 109(505):266– 274, 2014. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Liu:2015:MMA

[LLX15] Dungang Liu, Regina Y. Liu, and Minge Xie. Multivariate metaanalysis of heterogeneous studies using only summary statistics: Efficiency and robustness. Journal of the American Statistical Association, 110(509):326–340, 2015. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Li:2010:GDR

[LLZ10] Lexin Li, Bing Li, and Li-Xing Zhu. Groupwise dimension reduction. Journal of the American Statistical Association, 105(491): 1188–1201, September 2010. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Liang:2018:BNN

[LLZ18] Faming Liang, Qizhai Li, and Lei Zhou. Bayesian neural networks for selection of drug sensitive genes. Journal of the American Statistical Association, 113(523):955–972, 2018. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Laber:2011:ACI

[LM11a] Eric B. Laber and Susan A. Murphy. Adaptive confidence intervals for the test error in classification. Journal of the American Statistical Association, 106(495):904–913, September 2011. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Laber:2011:RAC

[LM11b] Eric B. Laber and Susan A. Murphy. Rejoinder: Adaptive confidence intervals for the test error in classification. Journal of the American Statistical Association, 106(495):940–945, September 2011. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lockwood:2017:MWF

[LM17] J. R. Lockwood and Daniel F. McCaffrey. Matching and weighting with functions of error-prone covariates for causal inference. Journal of the American Statistical Association, 111(516):1831– 1839, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lu:2014:OOD

[LMH14] Xiaosun Lu, J. S. Marron, and Perry Haaland. Object-oriented data analysis of cell images. Journal of the American Statistical Association, 109(506):548–559, 2014. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lindsay:2014:KDF

[LMR14] Bruce G. Lindsay, Marianthi Markatou, and Surajit Ray. Kernels, degrees of freedom, and power properties of quadratic distance goodness-of-fit tests. Journal of the American Statistical Association, 109(505):395–410, 2014. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lee:2019:BSF

[LMR^{+}19] Wonyul Lee, Michelle F. Miranda, Philip Rausch, Veerabhadran Baladandayuthapani, Massimo Fazio, J. Crawford Downs, and Jeffrey S. Morris. Bayesian semiparametric functional mixed models for serially correlated functional data, with application to glaucoma data. Journal of the American Statistical Association, 114(526):495–513, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lunagomez:2017:GRR

[LMWA17] Simón Lunagómez, Sayan Mukherjee, Robert L. Wolpert, and Edoardo M. Airoldi. Geometric representations of random hypergraphs. Journal of the American Statistical Association, 112 (517):363–383, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Li:2015:C

[LMZ15] Sai Li, Ritwik Mitra, and Cun-Hui Zhang. Comment. Journal of the American Statistical Association, 110(512):1455–1456, 2015. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/01621459.2015.1106404.

Li:2018:BCP

[LMZ18] Fan Li, Kari Lock Morgan, and Alan M. Zaslavsky. Balancing covariates via propensity score weighting. Journal of the American Statistical Association, 113(521):390–400, 2018. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lee:2013:NBM

[LMZJ13] Juhee Lee, Peter Müller, Yitan Zhu, and Yuan Ji. A nonparametric Bayesian model for local clustering with application to proteomics. Journal of the American Statistical Association, 108 (503):775–788, 2013. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Liang:2010:HMM

[LN10] Kun Liang and Dan Nettleton. A hidden Markov model approach to testing multiple hypotheses on a tree-transformed gene ontology graph. Journal of the American Statistical Association, 105(492):1444–1454, December 2010. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lijoi:2014:CHR

[LN14] Antonio Lijoi and Bernardo Nipoti. A class of hazard rate mixtures for combining survival data from different experiments. Journal of the American Statistical Association, 109(506):802– 814, 2014. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Li:2010:RCH

[LNA10a] Bo Li, Douglas W. Nychka, and Caspar M. Ammann. Rejoinder: “Comment: Hierarchical Statistical Modeling for Paleoclimate Reconstruction”. Journal of the American Statistical Association, 105(491):910–911, September 2010. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Li:2010:VMR

[LNA10b] Bo Li, Douglas W. Nychka, and Caspar M. Ammann. The value of multiproxy reconstruction of past climate. Journal of the American Statistical Association, 105(491):883–895, September 2010. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lock:2011:CPV

[LNM11] E. F. Lock, A. B. Nobel, and J. S. Marron. Comment: “Population Value Decomposition, a Framework for the Analysis of Image Populations”. Journal of the American Statistical Association, 106(495):798–802, September 2011. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Landau:2019:FBA

[LNN19] Will Landau, Jarad Niemi, and Dan Nettleton. Fully Bayesian analysis of RNA-seq counts for the detection of gene expression heterosis. Journal of the American Statistical Association, 114 (526):610–621, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lee:2014:MSB

[LNP14] Eun Ryung Lee, Hohsuk Noh, and Byeong U. Park. Model selection via Bayesian information criterion for quantile regression models. Journal of the American Statistical Association, 109 (505):216–229, 2014. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lele:2010:ELI

[LNS10] Subhash R. Lele, Khurram Nadeem, and Byron Schmuland. Estimability and likelihood inference for generalized linear mixed models using data cloning. Journal of the American Statistical Association, 105(492):1617–1625, December 2010. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lysy:2017:MCA

[LPH^{+}17] Martin Lysy, Natesh S. Pillai, David B. Hill, M. Gregory Forest, John W. R. Mellnik, Paula A. Vasquez, and Scott A. McKinley. Model comparison and assessment for single particle tracking in biological fluids. Journal of the American Statistical Association, 111(516):1413–1426, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Luati:2012:VP

[LPR12] Alessandra Luati, Tommaso Proietti, and Marco Reale. The variance profile. Journal of the American Statistical Association, 107(498):607–621, 2012. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Li:2011:NNP

[LQ11] Pengfei Li and Jing Qin. A new nuisance-parameter elimination method with application to the unordered homologous chromosome pairs problem. Journal of the American Statistical Association, 106(496):1476–1484, December 2011. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Luo:2017:FFL

[LQ17] Ruiyan Luo and Xin Qi. Function-on-function linear regression by signal compression. Journal of the American Statistical Association, 112(518):690–705, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Li:2017:PDM

[LQS17] Degui Li, Junhui Qian, and Liangjun Su. Panel data models with interactive fixed effects and multiple structural breaks. Journal of the American Statistical Association, 111(516):1804–1819, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

LaVecchia:2012:HOI

[LRT12] Davide La Vecchia, Elvezio Ronchetti, and Fabio Trojani. Higher-order infinitesimal robustness. Journal of the American Statistical Association, 107(500):1546–1557, 2012. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lei:2013:DFP

[LRW13] Jing Lei, James Robins, and Larry Wasserman. Distribution-free prediction sets. Journal of the American Statistical Association, 108(501):278–287, 2013. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Little:2017:CIM

[LRZ17] Roderick J. Little, Donald B. Rubin, and Sahar Z. Zangeneh. Conditions for ignoring the missing-data mechanism in likelihood inferences for parameter subsets. Journal of the American Statistical Association, 112(517):314–320, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Li:2010:TAB

[LS10] Chun Li and Bryan E. Shepherd. Test of association between two ordinal variables while adjusting for covariates. Journal of the American Statistical Association, 105(490):612–620, June 2010. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lee:2015:SFM

[LS15] Stephen S. M. Lee and Mehdi Soleymani. A simple formula for mixing estimators with different convergence rates. Journal of the American Statistical Association, 110(512):1463–1478, 2015. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/01621459.2014.960966.

Laber:2017:SSD

[LS17a] Eric B. Laber and Kerby Shedden. Statistical significance and the dichotomization of evidence: The relevance of the ASA statement on statistical significance and p-values for statisticians. Journal of the American Statistical Association, 112(519):902– 904, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

LopezCabrera:2017:FGQ

[LS17b] Brenda López Cabrera and Franziska Schulz. Forecasting generalized quantiles of electricity demand: A functional data approach. Journal of the American Statistical Association, 112 (517):127–136, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Laber:2018:FFC

[LS18a] Eric B. Laber and Ana-Maria Staicu. Functional feature construction for individualized treatment regimes. Journal of the American Statistical Association, 113(523):1219–1227, 2018. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lee:2018:MDD

[LS18b] Chung Eun Lee and Xiaofeng Shao. Martingale difference divergence matrix and its application to dimension reduction for stationary multivariate time series. Journal of the American Statistical Association, 113(521):216–229, 2018. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Li:2018:NGM

[LS18c] Bing Li and Eftychia Solea. A nonparametric graphical model for functional data with application to brain networks based on fMRI. Journal of the American Statistical Association, 113(524): 1637–1655, 2018. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lee:2019:EMA

[LS19a] Kwonsang Lee and Dylan S. Small. Estimating the malaria attributable fever fraction accounting for parasites being killed by fever and measurement error. Journal of the American Statistical Association, 114(525):79–92, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Li:2019:SHP

[LS19b] Furong Li and Huiyan Sang. Spatial homogeneity pursuit of regression coefficients for large datasets. Journal of the American Statistical Association, 114(527):1050–1062, 2019. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Luo:2012:IIB

[LSLR12] Xi Luo, Dylan S. Small, Chiang-Shan R. Li, and Paul R. Rosenbaum. Inference with interference between units in an fMRI experiment of motor inhibition. Journal of the American Statistical Association, 107(498):530–541, 2012. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Liang:2015:EMP

[LSQ15] Faming Liang, Qifan Song, and Peihua Qiu. An equivalent measure of partial correlation coefficients for high-dimensional Gaussian graphical models. Journal of the American Statistical Association, 110(511):1248–1265, 2015. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lee:2011:TTE

[LSS11] Sokbae Lee, Myung Hwan Seo, and Youngki Shin. Testing for threshold effects in regression models. Journal of the American Statistical Association, 106(493):220–231, March 2011. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lee:2017:C

[LSS17] Sokbae Lee, Myung Hwan Seo, and Youngki Shin. Correction. Journal of the American Statistical Association, 112(518):883, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Li:2019:IBS

[LSS^{+}19a] Qian Li, Damla Sentürk, Catherine A. Sugar, Shafali Jeste, Charlotte DiStefano, Joel Frohlich, and Donatello Telesca. Inferring brain signals synchronicity from a sample of EEG readings. Journal of the American Statistical Association, 114(527): 991–1001, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Luo:2019:FFS

[LSS^{+}19b] S. Luo, R. Song, M. Styner, J. H. Gilmore, and H. Zhu. FSEM: Functional structural equation models for twin functional data. Journal of the American Statistical Association, 114(525):344– 357, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lee:2011:ICA

[LST^{+}11] Seonjoo Lee, Haipeng Shen, Young Truong, Mechelle Lewis, and Xuemei Huang. Independent component analysis involving autocorrelated sources with an application to functional magnetic resonance imaging. Journal of the American Statistical Association, 106(495):1009–1024, September 2011. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Liang:2013:BSM

[LSY13] Faming Liang, Qifan Song, and Kai Yu. Bayesian subset modeling for high-dimensional generalized linear models. Journal of the American Statistical Association, 108(502):589–606, 2013. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lin:2017:ELR

[LSZD17] Lizhen Lin, Brian St. Thomas, Hongtu Zhu, and David B. Dunson. Extrinsic local regression on manifold-valued data. Journal of the American Statistical Association, 112(519):1261–1273, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

LaVecchia:2010:IRD

[LT10] Davide La Vecchia and Fabio Trojani. Infinitesimal robustness for diffusions. Journal of the American Statistical Association, 105(490):703–712, June 2010. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Leng:2012:SMG

[LT12] Chenlei Leng and Cheng Yong Tang. Sparse matrix graphical models. Journal of the American Statistical Association, 107(499):1187–1200, 2012. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lu:2014:NEP

[LT14] Zudi Lu and Dag Tjøstheim. Nonparametric estimation of probability density functions for irregularly observed spatial data. Journal of the American Statistical Association, 109(508):1546– 1564, 2014. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Li:2013:TSI

[LTC13] Wentao Li, Zhiqiang Tan, and Rong Chen. Two-stage importance sampling with mixture proposals. Journal of the American Statistical Association, 108(504):1350–1365, 2013. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lee:2015:BDF

[LTJM15] Juhee Lee, Peter F. Thall, Yuan Ji, and Peter Müller. Bayesian dose-finding in two treatment cycles based on the joint utility of efficacy and toxicity. Journal of the American Statistical Association, 110(510):711–722, 2015. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Li:2017:RJR

[LTT17] Jia Li, Viktor Todorov, and George Tauchen. Robust jump regressions. Journal of the American Statistical Association, 112 (517):332–341, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lu:2016:EMT

[Lu16] Zeng-Hua Lu. Extended MaxT tests of one-sided hypotheses. Journal of the American Statistical Association, 111(513):423– 437, 2016. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Laffont:2014:MAL

[LVC14] Celine Marielle Laffont, Marc Vandemeulebroecke, and Didier Concordet. Multivariate analysis of longitudinal ordinal data with mixed effects models, with application to clinical outcomes in osteoarthritis. Journal of the American Statistical Association, 109(507):955–966, 2014. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Luedtke:2017:C

[LvdL17] Alexander R. Luedtke and Mark J. van der Laan. Comment. Journal of the American Statistical Association, 111(516):1526– 1530, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [CZK17a, CZK17b].

Luedtke:2018:PRI

[LvdL18] Alexander R. Luedtke and Mark J. van der Laan. Parametricrate inference for one-sided differentiable parameters. Journal of the American Statistical Association, 113(522):780–788, 2018. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lee:2010:SCA

[LW10a] Ann B. Lee and Larry Wasserman. Spectral connectivity analysis. Journal of the American Statistical Association, 105(491): 1241–1255, September 2010. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Liang:2010:C

[LW10b] Hua Liang and Hulin Wu. Correction. Journal of the American Statistical Association, 105(492):1636, December 2010. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Li:2019:PAM

[LW19a] Gang Li and Xiaoyan Wang. Prediction accuracy measures for a nonlinear model and for right-censored time-to-event data. Journal of the American Statistical Association, 114(528):1815–1825, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Luo:2019:BEC

[LW19b] Xiangyu Luo and Yingying Wei. Batch effects correction with unknown subtypes. Journal of the American Statistical Association, 114(526):581–594, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Li:2010:GFL

[LWC10] Yehua Li, Naisyin Wang, and Raymond J. Carroll. Generalized functional linear models with semiparametric single-index interactions. Journal of the American Statistical Association, 105 (490):621–633, June 2010. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Li:2013:SNP

[LWC13] Yehua Li, Naisyin Wang, and Raymond J. Carroll. Selecting the number of principal components in functional data. Journal of the American Statistical Association, 108(504):1284–1294, 2013. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Li:2018:SEL

[LWN^{+}18] Liang Li, Chih-Hsien Wu, Jing Ning, Xuelin Huang, Ya-Chen Tina Shih, and Yu Shen. Semiparametric estimation of longitudinal medical cost trajectory. Journal of the American Statistical Association, 113(522):582–592, 2018. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Liu:2019:AEV

[LX19] Yaowu Liu and Jun Xie. Accurate and efficient P -value calculation via Gaussian approximation: A novel Monte-Carlo method. Journal of the American Statistical Association, 114(525):384– 392, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Li:2016:JIC

[LY16] Gang Li and Qing Yang. Joint inference for competing risks survival data. Journal of the American Statistical Association, 111(515):1289–1300, 2016. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Liu:2013:OET

[LYH13] Rong Liu, Lijian Yang, and Wolfgang K. Härdle. Oracally efficient two-step estimation of generalized additive model. Journal of the American Statistical Association, 108(502):619–631, 2013. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Li:2010:USS

[LZ10a] Fan Li and Alan M. Zaslavsky. Using a short screening scale for small-area estimation of mental illness prevalence for schools. Journal of the American Statistical Association, 105(492):1323– 1332, December 2010. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Li:2010:BVS

[LZ10b] Fan Li and Nancy R. Zhang. Bayesian variable selection in structured high-dimensional covariate spaces with applications in genomics. Journal of the American Statistical Association, 105(491):1202–1214, September 2010. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lu:2010:EPL

[LZ10c] Wenbin Lu and Hao Helen Zhang. On estimation of partially linear transformation models. Journal of the American Statistical Association, 105(490):683–691, June 2010. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Li:2011:STM

[LZ11a] Jialiang Li and Wenyang Zhang. A semiparametric threshold model for censored longitudinal data analysis. Journal of the American Statistical Association, 106(494):685–696, June 2011. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lin:2011:CPS

[LZ11b] D. Y. Lin and D. Zeng. Correcting for population stratification in genomewide association studies. Journal of the American Statistical Association, 106(495):997–1008, September 2011. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Li:2017:PTR

[LZ17] Lexin Li and Xin Zhang. Parsimonious tensor response regression. Journal of the American Statistical Association, 112(519): 1131–1146, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Liu:2018:RDO

[LZ18] Dungang Liu and Heping Zhang. Residuals and diagnostics for ordinal regression models: A surrogate approach. Journal of the American Statistical Association, 113(522):845–854, 2018. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lin:2019:SSI

[LZL19] Qian Lin, Zhigen Zhao, and Jun S. Liu. Sparse sliced inverse regression via lasso. Journal of the American Statistical Association, 114(528):1726–1739, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Liu:2017:CSP

[LZNS17] Wei Liu, Zhiwei Zhang, Lei Nie, and Guoxing Soon. A case study in personalized medicine: Rilpivirine versus Efavirenz for treatment-naive HIV patients. Journal of the American Statistical Association, 112(520):1381–1392, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Leng:2010:SMC

[LZP10] Chenlei Leng, Weiping Zhang, and Jianxin Pan. Semiparametric mean–covariance regression analysis for longitudinal data. Journal of the American Statistical Association, 105(489):181–193, March 2010. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Liu:2014:ERT

[LZPH14] Dandan Liu, Yingye Zheng, Ross L. Prentice, and Li Hsu. Estimating risk with time-to-event data: An application to the Women’s Health Initiative. Journal of the American Statistical Association, 109(506):514–524, 2014. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Liu:2016:JET

[LZS^{+}16] Wei Liu, Zhiwei Zhang, R. Jason Schroeder, Martin Ho, Bo Zhang, Cynthia Long, Hui Zhang, and Telba Z. Irony. Joint estimation of treatment and placebo effects in clinical trials with longitudinal blinding assessments. Journal of the American Statistical Association, 111(514):538–??, 2016. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Liu:2011:HSC

[LZW11] Yufeng Liu, Hao Helen Zhang, and Yichao Wu. Hard or soft classification? Large-margin unified machines. Journal of the American Statistical Association, 106(493):166–177, March 2011. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Liang:2011:OWC

[LZWZ11] Hua Liang, Guohua Zou, Alan T. K. Wan, and Xinyu Zhang. Optimal weight choice for frequentist model average estimators. Journal of the American Statistical Association, 106(495):1053– 1066, September 2011. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Li:2012:FSD

[LZZ12] Runze Li, Wei Zhong, and Liping Zhu. Feature screening via distance correlation learning. Journal of the American Statistical Association, 107(499):1129–1139, 2012. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ma:2013:ATC

[Ma13] Li Ma. Adaptive testing of conditional association through recursive mixture modeling. Journal of the American Statistical Association, 108(504):1493–1505, 2013. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ma:2015:SBM

[Ma15] Li Ma. Scalable Bayesian model averaging through local information propagation. Journal of the American Statistical Association, 110(510):795–809, 2015. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Mammen:2013:C

[Mam13] Enno Mammen. Comment. Journal of the American Statistical Association, 108(504):1260–1262, 2013. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [CCJ13a].

Martin:2015:PFE

[Mar15] Ryan Martin. Plausibility functions and exact frequentist inference. Journal of the American Statistical Association, 110(512): 1552–1561, 2015. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/01621459.2014.983232.

Marron:2019:DSM

[Mar19] J. S. Marron. Discussion: A Spatial Modeling Approach for Linguistic Object Data: Analysing Dialect Sound Variations Across Great Britain, by Shahin Tavakoli et al. Journal of the American Statistical Association, 114(527):1102, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Minas:2014:AMG

[MAS14] Giorgos Minas, John A. D. Aston, and Nigel Stallard. Adaptive multivariate global testing. Journal of the American Statistical Association, 109(506):613–623, 2014. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Mandozzi:2016:HTH

[MB16] Jacopo Mandozzi and Peter Bühlmann. Hierarchical testing in the high-dimensional setting with correlated variables. Journal of the American Statistical Association, 111(513):331–343, 2016. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Martinez:2013:SMF

[MBCM13] Josue G. Martinez, Kirsten M. Bohn, Raymond J. Carroll, and Jeffrey S. Morris. A study of Mexican free-tailed bat chirp syllables: Bayesian functional mixed models for nonstationary acoustic time series. Journal of the American Statistical Association, 108(502):514–526, 2013. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Mazumder:2019:CFM

[MCIS19] Rahul Mazumder, Arkopal Choudhury, Garud Iyengar, and Bodhisattva Sen. A computational framework for multivariate convex regression and its variants. Journal of the American Statistical Association, 114(525):318–331, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Mao:2019:MCC

[MCW19] Xiaojun Mao, Song Xi Chen, and Raymond K. W. Wong. Matrix completion with covariate information. Journal of the American Statistical Association, 114(525):198–210, 2019. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Miller:2019:RBI

[MD19] Jeffrey W. Miller and David B. Dunson. Robust Bayesian inference via coarsening. Journal of the American Statistical Association, 114(527):1113–1125, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Murray:2013:BGC

[MDCL13] Jared S. Murray, David B. Dunson, Lawrence Carin, and Joseph E. Lucas. Bayesian Gaussian copula factor models for mixed data. Journal of the American Statistical Association, 108 (502):656–665, 2013. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Miao:2017:INN

[MDG17] Wang Miao, Peng Ding, and Zhi Geng. Identifiability of normal and normal mixture models with nonignorable missing data. Journal of the American Statistical Association, 111(516):1673– 1683, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Mukerjee:2018:UST

[MDR18] Rahul Mukerjee, Tirthankar Dasgupta, and Donald B. Rubin. Using standard tools from finite population sampling to improve causal inference for complex experiments. Journal of the American Statistical Association, 113(522):868–881, 2018. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Mazumder:2011:SCD

[MFH11] Rahul Mazumder, Jerome H. Friedman, and Trevor Hastie. SparseNet: Coordinate descent with nonconvex penalties. Journal of the American Statistical Association, 106(495):1125–1138, September 2011. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Marchenko:2012:HSM

[MG12] Yulia V. Marchenko and Marc G. Genton. A Heckman selectiont model. Journal of the American Statistical Association, 107 (497):304–317, 2012. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

McShane:2017:RSS

[MG17a] Blakeley B. McShane and David Gal. Rejoinder: Statistical significance and the dichotomization of evidence. Journal of the American Statistical Association, 112(519):904–908, 2017. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

McShane:2017:SSD

[MG17b] Blakeley B. McShane and David Gal. Statistical significance and the dichotomization of evidence. Journal of the American Statistical Association, 112(519):885–895, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ma:2017:CPF

[MH17] Shujie Ma and Jian Huang. A concave pairwise fusion approach to subgroup analysis. Journal of the American Statistical Association, 112(517):410–423, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Miller:2018:MMP

[MH18] Jeffrey W. Miller and Matthew T. Harrison. Mixture models with a prior on the number of components. Journal of the American Statistical Association, 113(521):340–356, 2018. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ma:2011:DES

[MHC11] Yanyuan Ma, Jeffrey D. Hart, and Raymond J. Carroll. Density estimation in several populations with uncertain population membership. Journal of the American Statistical Association, 106(495):1180–1192, September 2011. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Majumder:2013:VVS

[MHC13] Mahbubul Majumder, Heike Hofmann, and Dianne Cook. Validation of visual statistical inference, applied to linear models.

Journal of the American Statistical Association, 108(503):942– 956, 2013. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ma:2015:THC

[MHZ15] Wei Ma, Feifang Hu, and Lixin Zhang. Testing hypotheses of covariate-adaptive randomized clinical trials. Journal of the American Statistical Association, 110(510):669–680, 2015. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Matteson:2014:NAM

[MJ14] David S. Matteson and Nicholas A. James. A nonparametric approach for multiple change point analysis of multivariate data. Journal of the American Statistical Association, 109(505):334– 345, 2014. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Mueller:2018:MPT

[MJG18] Jonas Mueller, Tommi Jaakkola, and David Gifford. Modeling persistent trends in distributions. Journal of the American Statistical Association, 113(523):1296–1310, 2018. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

McShane:2013:R

[MJPW13a] Blakeley B. McShane, Shane T. Jensen, Allan I. Pack, and Abraham J. Wyner. Rejoinder. Journal of the American Statistical Association, 108(504):1165–1172, 2013. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [MJPW13b].

McShane:2013:SLT

[MJPW13b] Blakeley B. McShane, Shane T. Jensen, Allan I. Pack, and Abraham J. Wyner. Statistical learning with time series dependence: An application to scoring sleep in mice. Journal of the American Statistical Association, 108(504):1147–1162, 2013. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [She13, ZW13, MJPW13a].

Ma:2013:SER

[MKG13] Yanyuan Ma, Mijeong Kim, and Marc G. Genton. Semiparametric efficient and robust estimation of an unknown symmetric population under arbitrary sample selection bias. Journal of the American Statistical Association, 108(503):1090–1104, 2013.

CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Martin:2013:C

[ML13a] Ryan Martin and Chuanhai Liu. Correction. Journal of the American Statistical Association, 108(503):1138–1139, 2013. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Martin:2013:IMF

[ML13b] Ryan Martin and Chuanhai Liu. Inferential models: a framework for prior-free posterior probabilistic inference. Journal of the American Statistical Association, 108(501):301–313, 2013. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Martin:2015:MIM

[ML15] Ryan Martin and Chuanhai Liu. Marginal inferential models: Prior-free probabilistic inference on interest parameters. Journal of the American Statistical Association, 110(512):1621–1631, 2015. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/01621459.2014.985827.

McCormick:2016:PCD

[MLC^{+}16] Tyler H. McCormick, Zehang Richard Li, Clara Calvert, Amelia C. Crampin, Kathleen Kahn, and Samuel J. Clark. Probabilistic cause-of-death assignment using verbal autopsies. Journal of the American Statistical Association, 111(515):1036–1049, 2016. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ma:2017:VSQ

[MLT17] Shujie Ma, Runze Li, and Chih-Ling Tsai. Variable screening via quantile partial correlation. Journal of the American Statistical Association, 112(518):650–663, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Morrissette:2013:EIC

[MM13] Jason L. Morrissette and Michael P. Mcdermott. Estimation and inference concerning ordered means in analysis of covariance models with interactions. Journal of the American Statistical Association, 108(503):832–839, 2013. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

McElroy:2015:MEP

[MM15] Tucker McElroy and Brian Monsell. Model estimation, prediction, and signal extraction for nonstationary stock and flow time series observed at mixed frequencies. Journal of the American Statistical Association, 110(511):1284–1303, 2015. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ma:2019:FES

[MM19] Li Ma and Jialiang Mao. Fisher exact scanning for dependency. Journal of the American Statistical Association, 114(525):245– 258, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Manolopoulou:2012:BSD

[MMC^{+}12a] Ioanna Manolopoulou, Melanie P. Matheu, Michael D. Cahalan, Mike West, and Thomas B. Kepler. Bayesian spatio-dynamic modeling in cell motility studies: Learning nonlinear taxic fields guiding the immune response. Journal of the American Statistical Association, 107(499):855–865, 2012. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Manolopoulou:2012:R

[MMC^{+}12b] Ioanna Manolopoulou, Melanie P. Matheu, Michael D. Cahalan, Mike West, and Thomas B. Kepler. Rejoinder. Journal of the American Statistical Association, 107(499):871–874, 2012. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Maitra:2012:BSC

[MML12] Ranjan Maitra, Volodymyr Melnykov, and Soumendra N. Lahiri. Bootstrapping for significance of compact clusters in multidimensional datasets. Journal of the American Statistical Association, 107(497):378–392, 2012. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Mitra:2013:BGM

[MML^{+}13] Riten Mitra, Peter Müller, Shoudan Liang, Lu Yue, and Yuan Ji. A Bayesian graphical model for ChIP-seq data on histone modifications. Journal of the American Statistical Association, 108 (501):69–80, 2013. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Magnus:2011:GWL

[MMM11a] Jan R. Magnus, Bertrand Melenberg, and Chris Muris. Global warming and local dimming: The statistical evidence. Journal of the American Statistical Association, 106(494):452–464, June 2011. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Magnus:2011:RGW

[MMM11b] Jan R. Magnus, Bertrand Melenberg, and Chris Muris. Rejoinder: Global warming and local dimming: The statistical evidence. Journal of the American Statistical Association, 106(494): 457–468, June 2011. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Mammen:2011:DVK

[MMNS11] Enno Mammen, María Dolores Martínez Miranda, Jens Perch Nielsen, and Stefan Sperlich. Do-validation for kernel density estimation. Journal of the American Statistical Association, 106 (494):651–660, June 2011. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ma:2017:SSI

[MMW^{+}17] Shujie Ma, Yanyuan Ma, Yanqing Wang, Eli S. Kravitz, and Raymond J. Carroll. A semiparametric single-index risk score across populations. Journal of the American Statistical Association, 112(520):1648–1662, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Muller:2016:CIP

[MN16] Ulrich K. Müller and Andriy Norets. Coverage inducing priors in nonstandard inference problems. Journal of the American Statistical Association, 111(515):1233–1241, 2016. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Morton:2010:SEP

[Mor10] Sally C. Morton. Statistics: From evidence to policy. Journal of the American Statistical Association, 105(489):1–5, March 2010. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Morganstein:2015:SMB

[Mor15] David Morganstein. Statistics: Making better decisions. Journal of the American Statistical Association, 110(512):1325–1330, 2015. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/01621459.2015.1106790.

Moraga:2016:C

[Mor16] Paula Moraga. Comment. Journal of the American Statistical Association, 111(515):1110–1111, 2016. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

MorrisseyLaVange:2019:CL

[Mor19] Lisa Morrissey LaVange. Choose to lead. Journal of the American Statistical Association, 114(528):1427–1435, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Mealli:2013:USO

[MP13] Fabrizia Mealli and Barbara Pacini. Using secondary outcomes to sharpen inference in randomized experiments with noncompliance. Journal of the American Statistical Association, 108(503): 1120–1131, 2013. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Muller:2014:C

[MQ14] Peter Müller and Fernando Quintana. Comment. Journal of the American Statistical Association, 109(507):889, 2014. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

McKeague:2015:ART

[MQ15a] Ian W. McKeague and Min Qian. An adaptive resampling test for detecting the presence of significant predictors. Journal of the American Statistical Association, 110(512):1422–1433, 2015. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/01621459.2015.1095099.

McKeague:2015:R

[MQ15b] Ian W. McKeague and Min Qian. Rejoinder. Journal of the American Statistical Association, 110(512):1459–1462, 2015. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/01621459.2015.1107431.

Ma:2011:STM

[MR11] Yanyuan Ma and Elvezio Ronchetti. Saddlepoint test in measurement error models. Journal of the American Statistical Association, 106(493):147–156, March 2011. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Morgan:2015:RBT

[MR15] Kari Lock Morgan and Donald B. Rubin. Rerandomization to balance tiers of covariates. Journal of the American Statistical Association, 110(512):1412–1421, 2015. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/01621459.2015.
1079528.

Murray:2017:MIM

[MR17] Jared S. Murray and Jerome P. Reiter. Multiple imputation of missing categorical and continuous values via Bayesian mixture models with local dependence. Journal of the American Statistical Association, 111(516):1466–1479, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

McCandless:2012:AMC

[MRB12] Lawrence C. McCandless, Sylvia Richardson, and Nicky Best. Adjustment for missing confounders using external validation data and propensity scores. Journal of the American Statistical Association, 107(497):40–51, 2012. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Marra:2017:SEA

[MRB^{+}17] Giampiero Marra, Rosalba Radice, Till Bärnighausen, Simon N. Wood, and Mark E. McGovern. A simultaneous equation approach to estimating HIV prevalence with nonignorable missing responses. Journal of the American Statistical Association, 112 (518):484–496, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ma:2015:VIC

[MS15] Shujie Ma and Peter X.-K. Song. Varying index coefficient models. Journal of the American Statistical Association, 110(509): 341–356, 2015. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ma:2018:EFA

[MS18a] Li Ma and Jacopo Soriano. Efficient functional ANOVA through wavelet-domain Markov groves. Journal of the American Statistical Association, 113(522):802–818, 2018. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ma:2018:AGT

[MS18b] Ling Ma and Rajeshwari Sundaram. Analysis of gap times based on panel count data with informative observation times and unknown start time. Journal of the American Statistical Association, 113(521):294–305, 2018. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Mainassara:2018:DCM

[MS18c] Yacouba Boubacar Maïnassara and Bruno Saussereau. Diagnostic checking in multivariate ARMA models with dependent errors using normalized residual autocorrelations. Journal of the American Statistical Association, 113(524):1813–1827, 2018. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Mohler:2011:SEP

[MSB^{+}11] G. O. Mohler, M. B. Short, P. J. Brantingham, F. P. Schoenberg, and G. E. Tita. Self-exciting point process modeling of crime. Journal of the American Statistical Association, 106(493):100– 108, March 2011. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Mohammed:2012:MEC

[MŞDN12] Sandra M. Mohammed, Damla Şentürk, Lorien S. Dalrymple, and Danh V. Nguyen. Measurement error case series models with application to infection-cardiovascular risk in older patients on dialysis. Journal of the American Statistical Association, 107(500):1310–1323, 2012. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Miles:2017:QAP

[MSK^{+}17] Caleb H. Miles, Ilya Shpitser, Phyllis Kanki, Seema Meloni, and Eric J. Tchetgen Tchetgen. Quantifying an adherence pathspecific effect of antiretroviral therapy in the Nigeria PEPFAR program. Journal of the American Statistical Association, 112 (520):1443–1452, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Mohler:2015:RCF

[MSM^{+}15] G. O. Mohler, M. B. Short, Sean Malinowski, Mark Johnson, G. E. Tita, Andrea L. Bertozzi, and P. J. Brantingham. Randomized controlled field trials of predictive policing. Journal of the American Statistical Association, 110(512):1399–1411, 2015. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/01621459.2015.1077710.

Mak:2018:ESM

[MSW^{+}18] Simon Mak, Chih-Li Sung, Xingjian Wang, Shiang-Ting Yeh, Yu-Hung Chang, V. Roshan Joseph, Vigor Yang, and C. F. Jeff Wu. An efficient surrogate model for emulation and physics extraction of large eddy simulations. Journal of the American Statistical Association, 113(524):1443–1456, 2018. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

McCormick:2010:HMP

[MSZ10] Tyler H. McCormick, Matthew J. Salganik, and Tian Zheng. How many people do you know?: Efficiently estimating personal network size. Journal of the American Statistical Association, 105(489):59–70, March 2010. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Matteson:2011:DOC

[MT11] David S. Matteson and Ruey S. Tsay. Dynamic orthogonal components for multivariate time series. Journal of the American Statistical Association, 106(496):1450–1463, December 2011. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Matteson:2017:ICA

[MT17] David S. Matteson and Ruey S. Tsay. Independent component analysis via distance covariance. Journal of the American Statistical Association, 112(518):623–637, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Minnier:2011:PMI

[MTC11] Jessica Minnier, Lu Tian, and Tianxi Cai. A perturbation method for inference on regularized regression estimates. Journal of the American Statistical Association, 106(496):1371–1382, December 2011. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Murray:2017:RTC

[MTY^{+}17] Thomas A. Murray, Peter F. Thall, Ying Yuan, Sarah McAvoy, and Daniel R. Gomez. Robust treatment comparison based on utilities of semi-competing risks in non-small-cell lung cancer. Journal of the American Statistical Association, 112(517):11–23, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Manrique-Vallier:2012:EID

[MVR12] Daniel Manrique-Vallier and Jerome P. Reiter. Estimating identification disclosure risk using mixed membership models. Journal of the American Statistical Association, 107(500):1385–1394, 2012. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Manrique-Vallier:2017:BSE

[MVR17] Daniel Manrique-Vallier and Jerome P. Reiter. Bayesian simultaneous edit and imputation for multivariate categorical data. Journal of the American Statistical Association, 112(520):1708– 1719, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Mao:2010:SEE

[MW10a] Meng Mao and Jane-Ling Wang. Semiparametric efficient estimation for a class of generalized proportional odds cure models. Journal of the American Statistical Association, 105(489):302– 311, March 2010. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Morgan:2010:WOD

[MW10b] John P. Morgan and Xiaowei Wang. Weighted optimality in designed experimentation. Journal of the American Statistical Association, 105(492):1566–1580, December 2010. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ma:2011:COP

[MW11] Li Ma and Wing Hung Wong. Coupling optional Pólya trees and the two sample problem. Journal of the American Statistical Association, 106(496):1553–1565, December 2011. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Muller:2017:FAI

[MW17] Ulrich K. Müller and Yulong Wang. Fixed-k asymptotic inference about tail properties. Journal of the American Statistical Association, 112(519):1334–1343, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Mak:2019:CNM

[MW19] Simon Mak and C. F. Jeff Wu. cmenet: A new method for bilevel variable selection of conditional main effects. Journal of the American Statistical Association, 114(526):844–856, 2019. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Miao:2014:GOD

[MWX14] Hongyu Miao, Hulin Wu, and Hongqi Xue. Generalized ordinary differential equation models. Journal of the American Statistical Association, 109(508):1672–1682, 2014. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Minnier:2015:RCA

[MYLC15] Jessica Minnier, Ming Yuan, Jun S. Liu, and Tianxi Cai. Risk classification with an adaptive naive Bayes kernel machine model. Journal of the American Statistical Association, 110(509):393– 404, 2015. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Murray:2018:BML

[MYT18] Thomas A. Murray, Ying Yuan, and Peter F. Thall. A Bayesian machine learning approach for optimizing dynamic treatment regimes. Journal of the American Statistical Association, 113 (523):1255–1267, 2018. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ma:2012:SAD

[MZ12] Yanyuan Ma and Liping Zhu. A semiparametric approach to dimension reduction. Journal of the American Statistical Association, 107(497):168–179, 2012. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

McCormick:2015:LSM

[MZ15] Tyler H. McCormick and Tian Zheng. Latent surface models for networks using aggregated relational data. Journal of the American Statistical Association, 110(512):1684–1695, 2015. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/01621459.2014.991395.

Minsker:2016:ACT

[MZC16] Stanislav Minsker, Ying-Qi Zhao, and Guang Cheng. Active clinical trials for personalized medicine. Journal of the American Statistical Association, 111(514):875–??, 2016. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Maadooliat:2016:CEM

[MZN^{+}16] Mehdi Maadooliat, Lan Zhou, Seyed Morteza Najibi, Xin Gao, and Jianhua Z. Huang. Collective estimation of multiple bivariate density functions with application to angular-sampling-based protein loop modeling. Journal of the American Statistical Association, 111(513):43–56, 2016. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Mauguen:2017:DCS

[MZT^{+}17] Audrey Mauguen, Emily C. Zabor, Nancy E. Thomas, Marianne Berwick, Venkatraman E. Seshan, and Colin B. Begg. Defining cancer subtypes with distinctive etiologic profiles: An application to the epidemiology of melanoma. Journal of the American Statistical Association, 112(517):54–63, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Mefford:2016:CHG

[MZW16] Joel A. Mefford, Noah A. Zaitlen, and John S. Witte. Comment: A human genetics perspective. Journal of the American Statistical Association, 111(513):124–127, 2016. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Nandram:2010:BAB

[NC10] Balgobin Nandram and Jai Won Choi. A Bayesian analysis of body mass index data from small domains under nonignorable nonresponse and selection. Journal of the American Statistical Association, 105(489):120–135, March 2010. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Nguyen:2012:SSD

[NCB12] Hai Nguyen, Noel Cressie, and Amy Braverman. Spatial statistical data fusion for remote sensing applications. Journal of the American Statistical Association, 107(499):1004–1018, 2012.

CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Noh:2013:CBR

[NEB13] Hohsuk Noh, Anouar El Ghouch, and Taoufik Bouezmarni. Copula-based regression estimation and inference. Journal of the American Statistical Association, 108(502):676–688, 2013. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Noufaily:2016:DID

[NFG^{+}16] Angela Noufaily, Paddy Farrington, Paul Garthwaite, Doyo Gragn Enki, Nick Andrews, and Andre Charlett. Detection of infectious disease outbreaks from laboratory data with reporting delays. Journal of the American Statistical Association, 111(514):488– ??, 2016. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Nolen:2011:RBI

[NH11] Tracy L. Nolen and Michael G. Hudgens. Randomization-based inference within principal strata. Journal of the American Statistical Association, 106(494):581–593, June 2011. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Nordman:2012:BBT

[NL12] Daniel J. Nordman and Soumendra N. Lahiri. Block bootstraps for time series with fixed regressors. Journal of the American Statistical Association, 107(497):233–246, 2012. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Nevalainen:2010:NAC

[NLOP10] Jaakko Nevalainen, Denis Larocque, Hannu Oja, and Ilkka Pörsti. Nonparametric analysis of clustered multivariate data. Journal of the American Statistical Association, 105(490):864– 872, June 2010. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Narisetty:2017:EDF

[NN17] Naveen N. Narisetty and Vijayan N. Nair. Extremal depth for functional data and applications. Journal of the American Statistical Association, 111(516):1705–1714, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ning:2014:SEE

[NQS14] Jing Ning, Jing Qin, and Yu Shen. Score estimating equations from embedded likelihood functions under accelerated failure time model. Journal of the American Statistical Association, 109(508):1625–1635, 2014. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ni:2017:SMD

[NSB17] Yang Ni, Francesco C. Stingo, and Veerabhadran Baladandayuthapani. Sparse multi-dimensional graphical models: A unified Bayesian framework. Journal of the American Statistical Association, 112(518):779–793, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ni:2019:BGR

[NSB19] Yang Ni, Francesco C. Stingo, and Veerabhadran Baladandayuthapani. Bayesian graphical regression. Journal of the American Statistical Association, 114(525):184–197, 2019. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Narisetty:2019:SGC

[NSH19a] Naveen N. Narisetty, Juan Shen, and Xuming He. Skinny Gibbs: A consistent and scalable Gibbs sampler for model selection. Journal of the American Statistical Association, 114(527):1205– 1217, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ni:2019:BHV

[NSH^{+}19b] Yang Ni, Francesco C. Stingo, Min Jin Ha, Rehan Akbani, and Veerabhadran Baladandayuthapani. Bayesian hierarchical varying-sparsity regression models with application to cancer proteogenomics. Journal of the American Statistical Association, 114(525):48–60, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

NobregaDaSilva:2016:UBP

[NSK16] Damião Nóbrega Da Silva, Chris Skinner, and Jae Kwang Kim. Using binary paradata to correct for measurement error in survey data analysis. Journal of the American Statistical Association, 111(514):526–??, 2016. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Naranjo:2013:ESS

[NTC13] Arlene Naranjo, A. Alexandre Trindade, and George Casella. Extending the state-space model to accommodate missing values in responses and covariates. Journal of the American Statistical Association, 108(501):202–216, 2013. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Nussbaum:2018:SEN

[Nus18] Barry D. Nussbaum. Statistics: Essential now more than ever. Journal of the American Statistical Association, 113(522):489– 493, 2018. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Nadkarni:2011:IRE

[NZK11] Nivedita V. Nadkarni, Yingqi Zhao, and Michael R. Kosorok. Inverse regression estimation for censored data. Journal of the American Statistical Association, 106(493):178–190, March 2011. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Oates:2019:BPN

[OCAG19] Chris J. Oates, Jon Cockayne, Robert G. Aykroyd, and Mark Girolami. Bayesian probabilistic numerical methods in timedependent state estimation for industrial hydrocyclone equipment. Journal of the American Statistical Association, 114(528): 1518–1531, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ogburn:2017:C

[Ogb17] Elizabeth L. Ogburn. Comment. Journal of the American Statistical Association, 111(516):1534–1537, 2017. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [CZK17a, CZK17b].

Oberski:2017:EQS

[OKEK17] D. L. Oberski, A. Kirchner, S. Eckman, and F. Kreuter. Evaluating the quality of survey and administrative data with generalized multitrait-multimethod models. Journal of the American Statistical Association, 112(520):1477–1489, 2017. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Oh:2013:SMM

[OP13] Dong Hwan Oh and Andrew J. Patton. Simulated method of moments estimation for copula-based multivariate models. Journal of the American Statistical Association, 108(502):689–700, 2013. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Oates:2016:CTI

[OPG16] Chris J. Oates, Theodore Papamarkou, and Mark Girolami. The controlled thermodynamic integral for Bayesian model evidence evaluation. Journal of the American Statistical Association, 111 (514):634–??, 2016. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Otsu:2017:BIM

[OR17] Taisuke Otsu and Yoshiyasu Rai. Bootstrap inference of matching estimators for average treatment effects. Journal of the American Statistical Association, 112(520):1720–1732, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ogburn:2019:CBM

[OST19] Elizabeth L. Ogburn, Ilya Shpitser, and Eric J. Tchetgen Tchetgen. Comment on “Blessings of Multiple Causes”. Journal of the American Statistical Association, 114(528):1611–1615, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [WB19a, WB19b].

Park:2011:ALP

[PAHJ11] Cheolwoo Park, Jeongyoun Ahn, Martin Hendry, and Woncheol Jang. Analysis of long period variable stars with nonparametric tests for trend detection. Journal of the American Statistical Association, 106(495):832–845, September 2011. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Pantula:2011:SKI

[Pan11] Sastry G. Pantula. Statistics: a key to innovation in a datacentric world! Journal of the American Statistical Association, 106(493):1–5, March 2011. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Paparoditis:2010:VSA

[Pap10] Efstathios Paparoditis. Validating stationarity assumptions in time series analysis by rolling local periodograms. Journal of the American Statistical Association, 105(490):839–851, June 2010. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Papp:2012:ODR

[Pap12] Dávid Papp. Optimal designs for rational function regression. Journal of the American Statistical Association, 107(497):400– 411, 2012. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Padilla:2019:SNT

[PARS19] Oscar Hernan Madrid Padilla, Alex Athey, Alex Reinhart, and James G. Scott. Sequential nonparametric tests for a change in distribution: An application to detecting radiological anomalies. Journal of the American Statistical Association, 114(526):514– 528, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Page:2013:CBN

[PBD13] Garritt Page, Abhishek Bhattacharya, and David Dunson. Classification via Bayesian nonparametric learning of affine subspaces. Journal of the American Statistical Association, 108 (501):187–201, 2013. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Porcu:2016:STC

[PBG16] Emilio Porcu, Moreno Bevilacqua, and Marc G. Genton. Spatiotemporal covariance and cross-covariance functions of the great circle distance on a sphere. Journal of the American Statistical Association, 111(514):888–??, 2016. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Pfister:2019:ICP

[PBP19] Niklas Pfister, Peter Bühlmann, and Jonas Peters. Invariant causal prediction for sequential data. Journal of the American Statistical Association, 114(527):1264–1276, 2019. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Paige:2011:SSL

[PCB11] Robert L. Paige, Phillip L. Chapman, and Ronald W. Butler. Small sample LD50 confidence intervals using saddlepoint approximations. Journal of the American Statistical Association, 106(493):334–344, March 2011. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Parast:2012:LPL

[PCC12] Layla Parast, Su-Chun Cheng, and Tianxi Cai. Landmark prediction of long-term survival incorporating short-term event time information. Journal of the American Statistical Association, 107(500):1492–1501, 2012. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Panagiotelis:2012:PCC

[PCJ12] Anastasios Panagiotelis, Claudia Czado, and Harry Joe. Pair copula constructions for multivariate discrete data. Journal of the American Statistical Association, 107(499):1063–1072, 2012. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Pham:2016:PDT

[PCSK16] Lisa M. Pham, Luis Carvalho, Scott Schaus, and Eric D. Kolaczyk. Perturbation detection through modeling of gene expression on a latent biological pathway network: A Bayesian hierarchical approach. Journal of the American Statistical Association, 111(513):73–92, 2016. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Portier:2014:BTR

[PD14] François Portier and Bernard Delyon. Bootstrap testing of the rank of a matrix via least-squared constrained estimation. Journal of the American Statistical Association, 109(505):160–172, 2014. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Patel:2016:CAN

[PD16] Chirag J. Patel and Francesca Dominici. Comment: Addressing the need for portability in big data model building and calibration. Journal of the American Statistical Association, 111(513): 127–129, 2016. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Pedeli:2015:LEI

[PDF15] Xanthi Pedeli, Anthony C. Davison, and Konstantinos Fokianos. Likelihood estimation for the INAR(p) model by saddlepoint approximation. Journal of the American Statistical Association, 110(511):1229–1238, 2015. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Pearce:2019:BRH

[Pea19] Neil Pearce. Book review: Handbook of Statistical Methods for Case-Control Studies, by Ørnulf Borgan, Norman E. Breslow, Nilanjan Chatterjee, Mitchell H. Gail, Alastair Scott, and Christopher J. Wild, eds. Boca Raton, FL: Chapman & Hall/CRC Press, 2018, xvii + 536 pp., $119.95 (H), ISBN: 978-1-49-876858-0. Journal of the American Statistical Association, 114(528):1926– 1928, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Peress:2010:CSN

[Per10] Michael Peress. Correcting for survey nonresponse using variable response propensity. Journal of the American Statistical Association, 105(492):1418–1430, December 2010. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Percival:2012:SSA

[Per12] Daniel Percival. Structured, sparse aggregation. Journal of the American Statistical Association, 107(498):814–823, 2012. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Pareek:2018:TIM

[PGW^{+}18] Bhuvanesh Pareek, Pulak Ghosh, Hugh N. Wilson, Emma K. Macdonald, and Paul Baines. Tracking the impact of media on voter choice in real time: A Bayesian dynamic joint model. Journal of the American Statistical Association, 113(524):1457–1475, 2018. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Pan:2014:TIA

[PGY14] Guangming Pan, Jiti Gao, and Yanrong Yang. Testing independence among a large number of high-dimensional random vectors. Journal of the American Statistical Association, 109(506): 600–612, 2014. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Park:2018:BIP

[PH18] Jaewoo Park and Murali Haran. Bayesian inference in the presence of intractable normalizing functions. Journal of the American Statistical Association, 113(523):1372–1390, 2018. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Picciotto:2012:SNC

[PHP^{+}12] Sally Picciotto, Miguel A. Hernán, John H. Page, Jessica G. Young, and James M. Robins. Structural nested cumulative failure time models to estimate the effects of interventions. Journal of the American Statistical Association, 107(499):886–900, 2012. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Pan:2015:RAA

[PHSS15] Deng Pan, Haijin He, Xinyuan Song, and Liuquan Sun. Regression analysis of additive hazards model with latent variables. Journal of the American Statistical Association, 110(511):1148– 1159, 2015. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Plumlee:2016:CFP

[PJY16] Matthew Plumlee, V. Roshan Joseph, and Hui Yang. Calibrating functional parameters in the ion channel models of cardiac cells. Journal of the American Statistical Association, 111(514):500– ??, 2016. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Panaretos:2012:NCM

[PK12] Victor M. Panaretos and Kjell Konis. Nonparametric construction of multivariate kernels. Journal of the American Statistical Association, 107(499):1085–1095, 2012. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Panaretos:2010:SOC

[PKM10] Victor M. Panaretos, David Kraus, and John H. Maddocks. Second-order comparison of Gaussian random functions and the geometry of DNA minicircles. Journal of the American Statistical Association, 105(490):670–682, June 2010. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kreiss:2012:HWB

[pKP12] Jens peter Kreiss and Efstathios Paparoditis. The hybrid wild bootstrap for time series. Journal of the American Statistical Association, 107(499):1073–1084, 2012. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Pimentel:2015:LSO

[PKSR15] Samuel D. Pimentel, Rachel R. Kelz, Jeffrey H. Silber, and Paul R. Rosenbaum. Large, sparse optimal matching with refined covariate balance in an observational study of the health outcomes produced by new surgeons. Journal of the American Statistical Association, 110(510):515–527, 2015. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Pimentel:2017:C

[PKSR17] Samuel D. Pimentel, Rachel R. Kelz, Jeffrey H. Silber, and Paul R. Rosenbaum. Correction. Journal of the American Statistical Association, 112(520):1770, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Papadogeorgou:2019:DPS

[PL19] Georgia Papadogeorgou and Fan Li. Discussion of “Penalized Spline of Propensity Methods for Treatment Comparison”. Journal of the American Statistical Association, 114(525):32–35, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [ZEL19a, ZEL19b].

Peng:2011:FAB

[PLGM11] Limin Peng, Ruosha Li, Ying Guo, and Amita Manatunga. A framework for assessing broad sense agreement between ordinal and continuous measurements. Journal of the American Statistical Association, 106(496):1592–1601, December 2011. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Philtron:2018:MRR

[PLLG18] Daisy Philtron, Yafei Lyu, Qunhua Li, and Debashis Ghosh. Maximum rank reproducibility: A nonparametric approach to assessing reproducibility in replicate experiments. Journal of the American Statistical Association, 113(523):1028–1039, 2018. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Plumlee:2014:FPD

[Plu14] Matthew Plumlee. Fast prediction of deterministic functions using sparse grid experimental designs. Journal of the American Statistical Association, 109(508):1581–1591, 2014. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Plumlee:2017:BCI

[Plu17] Matthew Plumlee. Bayesian calibration of inexact computer models. Journal of the American Statistical Association, 112 (519):1274–1285, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Petersen:2019:DSM

[PM19] Alexander Petersen and Hans-Georg Müller. Discussion: A Spatial Modeling Approach for Linguistic Object Data: Analyzing Dialect Sound Variations Across Great Britain, by Shahin Tavakoli et al. Journal of the American Statistical Association, 114(527):1099–1101, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Pan:2019:CAT

[PMZ19] Yuqing Pan, Qing Mai, and Xin Zhang. Covariate-adjusted tensor classification in high dimensions. Journal of the American Statistical Association, 114(527):1305–1319, 2019. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Polonik:2011:CAC

[Pol11] Wolfgang Polonik. Comment: “Adaptive Confidence Intervals for the Test Error in Classification”. Journal of the American Statistical Association, 106(495):936–940, September 2011. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Politis:2014:C

[Pol14] Dimitris N. Politis. Comment. Journal of the American Statistical Association, 109(507):1010–1013, 2014. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Percival:2014:ATS

[PPD^{+}14] Daniel M. Percival, Donald B. Percival, Donald W. Denbo, Edison Gica, Paul Y. Huang, Harold O. Mofjeld, and Michael C.

Spillane. Automated tsunami source modeling using the sweeping window positive elastic net. Journal of the American Statistical Association, 109(506):491–499, 2014. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Preuss:2015:DMS

[PPD15] Philip Preuss, Ruprecht Puchstein, and Holger Dette. Detection of multiple structural breaks in multivariate time series. Journal of the American Statistical Association, 110(510):654–668, 2015. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Parker:2018:PAS

[PPLS18] Albert E. Parker, Betsey Pitts, Lindsey Lorenz, and Philip S. Stewart. Polynomial accelerated solutions to a large Gaussian model for imaging biofilms: In theory and finite precision. Journal of the American Statistical Association, 113(524):1431–1442, 2018. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Panaretos:2014:PF

[PPY14] Victor M. Panaretos, Tung Pham, and Zhigang Yao. Principal flows. Journal of the American Statistical Association, 109(505): 424–436, 2014. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Patilea:2014:TSO

[PR14] Valentin Patilea and Hamdi Raïssi. Testing second-order dynamics for autoregressive processes in presence of time-varying variance. Journal of the American Statistical Association, 109(507): 1099–1111, 2014. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Prentice:2010:CDP

[Pre10] Ross L. Prentice. Chronic disease prevention research methods and their reliability, with illustrations from the Women’s Health Initiative. Journal of the American Statistical Association, 105 (492):1431–1443, December 2010. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Padoan:2010:LBI

[PRS10] S. A. Padoan, M. Ribatet, and S. A. Sisson. Likelihood-based inference for max-stable processes. Journal of the American Statistical Association, 105(489):263–277, March 2010. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Peress:2010:SCU

[PS10] Michael Peress and Arthur Spirling. Scaling the critics: Uncovering the latent dimensions of movie criticism with an item response approach. Journal of the American Statistical Association, 105(489):71–83, March 2010. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Pimentel:2016:CSC

[PSR16] Samuel D. Pimentel, Dylan S. Small, and Paul R. Rosenbaum. Constructed second control groups and attenuation of unmeasured biases. Journal of the American Statistical Association, 111(515):1157–1167, 2016. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Patilea:2017:TPE

[PSSS17] Valentin Patilea, César Sánchez-Sellero, and Matthieu Saumard. Testing the predictor effect on a functional response. Journal of the American Statistical Association, 111(516):1684–1695, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Peterson:2015:BIM

[PSV15] Christine Peterson, Francesco C. Stingo, and Marina Vannucci. Bayesian inference of multiple Gaussian graphical models. Journal of the American Statistical Association, 110(509):159–174, 2015. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Polson:2013:BIL

[PSW13] Nicholas G. Polson, James G. Scott, and Jesse Windle. Bayesian inference for logistic models using Pólya-gamma latent variables. Journal of the American Statistical Association, 108(504):1339– 1349, 2013. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Pena:2019:FMT

[PSY19] Daniel Peña, Ezequiel Smucler, and Victor J. Yohai. Forecasting multiple time series with one-sided dynamic principal components. Journal of the American Statistical Association, 114(528):

1683–1694, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Polzehl:2017:LSD

[PT17] Jörg Polzehl and Karsten Tabelow. Low SNR in diffusion MRI models. Journal of the American Statistical Association, 111 (516):1480–1490, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Parast:2014:LES

[PTC14] Layla Parast, Lu Tian, and Tianxi Cai. Landmark estimation of survival and treatment effect in a randomized clinical trial. Journal of the American Statistical Association, 109(505):384– 394, 2014. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Paindaveine:2013:DLD

[PV13] Davy Paindaveine and Germain Van Bever. From depth to local depth: A focus on centrality. Journal of the American Statistical Association, 108(503):1105–1119, 2013. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Pan:2016:UDM

[PWL16] Rui Pan, Hansheng Wang, and Runze Li. Ultrahigh-dimensional multiclass linear discriminant analysis by pairwise sure independence screening. Journal of the American Statistical Association, 111(513):169–179, 2016. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Pan:2019:GSI

[PWXZ19] Wenliang Pan, Xueqin Wang, Weinan Xiao, and Hongtu Zhu. A generic sure independence screening procedure. Journal of the American Statistical Association, 114(526):928–937, 2019. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Perron:2012:TTP

[PY12] Pierre Perron and Tomoyoshi Yabu. Testing for trend in the presence of autoregressive error: A comment. Journal of the American Statistical Association, 107(498):844, 2012. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Pena:2016:GDP

[PY16] Daniel Peña and Victor J. Yohai. Generalized dynamic principal components. Journal of the American Statistical Association, 111(515):1121–1131, 2016. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Qian:2010:NLS

[QA10] Peter Z. G. Qian and Mingyao Ai. Nested lattice sampling: a new sampling scheme derived by randomizing nested orthogonal arrays. Journal of the American Statistical Association, 105(491): 1147–1155, September 2010. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Qiu:2015:BSH

[QC15] Yumou Qiu and Song Xi Chen. Bandwidth selection for highdimensional covariance matrix estimation. Journal of the American Statistical Association, 110(511):1160–1174, 2015. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Qian:2019:SMD

[QDC19] Wei Qian, Shanshan Ding, and R. Dennis Cook. Sparse minimum discrepancy approach to sufficient dimension reduction with simultaneous variable selection in ultrahigh dimension. Journal of the American Statistical Association, 114(527):1277–1290, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Qiao:2019:FGM

[QGJ19] Xinghao Qiao, Shaojun Guo, and Gareth M. James. Functional graphical models. Journal of the American Statistical Association, 114(525):211–222, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Qian:2012:SLH

[Qia12] Peter Z. G. Qian. Sliced latin hypercube designs. Journal of the American Statistical Association, 107(497):393–399, 2012. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Qian:2017:C

[Qia17] Min Qian. Comment. Journal of the American Statistical Association, 111(516):1538–1541, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [CZK17a, CZK17b].

Quintana:2016:BNL

[QJWG16] Fernando A. Quintana, Wesley O. Johnson, L. Elaine Waetjen, and Ellen B. Gold. Bayesian nonparametric longitudinal data analysis. Journal of the American Statistical Association, 111(515):1168–1181, 2016. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Quiroz:2019:SME

[QKVT19] Matias Quiroz, Robert Kohn, Mattias Villani, and Minh-Ngoc Tran. Speeding up MCMC by efficient data subsampling. Journal of the American Statistical Association, 114(526):831–843, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Qu:2010:HEA

[QLL10] Annie Qu, Bruce G. Lindsay, and Lin Lu. Highly efficient aggregate unbiased estimating functions approach for correlated data with missing at random. Journal of the American Statistical Association, 105(489):194–204, March 2010. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Qin:2011:MLE

[QNLS11] Jing Qin, Jing Ning, Hao Liu, and Yu Shen. Maximum likelihood estimations and EM algorithms with length-biased data. Journal of the American Statistical Association, 106(496):1434–1449, December 2011. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Qin:2013:MLE

[QP13] Yichen Qin and Carey E. Priebe. Maximum L_{q}-likelihood estimation via the expectation-maximization algorithm: A robust estimation of mixture models. Journal of the American Statistical Association, 108(503):914–928, 2013. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Qiao:2010:WDW

[QZL^{+}10] Xingye Qiao, Hao Helen Zhang, Yufeng Liu, Michael J. Todd, and J. S. Marron. Weighted distance weighted discrimination and its asymptotic properties. Journal of the American Statistical Association, 105(489):401–414, March 2010. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Raftery:2017:CEL

[Raf17] Adrian E. Raftery. Comment: Extending the latent position model for networks. Journal of the American Statistical Association, 112(520):1531–1534, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ranganath:2018:CRM

[RB18] Rajesh Ranganath and David M. Blei. Correlated random measures. Journal of the American Statistical Association, 113(521): 417–430, 2018. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Reich:2013:NSM

[RBB13] Brian J. Reich, Dipankar Bandyopadhyay, and Howard D. Bondell. A nonparametric spatial model for periodontal data with nonrandom missingness. Journal of the American Statistical Association, 108(503):820–831, 2013. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ren:2017:BNO

[RBF^{+}17] Boyu Ren, Sergio Bacallado, Stefano Favaro, Susan Holmes, and Lorenzo Trippa. Bayesian nonparametric ordination for the analysis of microbial communities. Journal of the American Statistical Association, 112(520):1430–1442, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Rodrigues:2012:BEP

[RD12] Alexandre Rodrigues and Peter J. Diggle. Bayesian estimation and prediction for inhomogeneous spatiotemporal log-Gaussian Cox processes using low-rank models, with application to criminal surveillance. Journal of the American Statistical Association, 107(497):93–101, 2012. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Rodriguez:2010:LSB

[RDG10] Abel Rodríguez, David B. Dunson, and Alan E. Gelfand. Latent stick-breaking processes. Journal of the American Statistical Association, 105(490):647–659, June 2010. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Rubin-Delanchy:2019:MAM

[RDHL19] Patrick Rubin-Delanchy, Nicholas A. Heard, and Daniel J. Lawson. Meta-analysis of mid-p-values: Some new results based on the convex order. Journal of the American Statistical Association, 114(527):1105–1112, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ren:2010:DNB

[RDLC10] Lu Ren, David Dunson, Scott Lindroth, and Lawrence Carin. Dynamic nonparametric Bayesian models for analysis of music. Journal of the American Statistical Association, 105(490):458– 472, June 2010. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Reich:2011:BSQ

[RFD11] Brian J. Reich, Montserrat Fuentes, and David B. Dunson. Bayesian spatial quantile regression. Journal of the American Statistical Association, 106(493):6–20, March 2011. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Rockova:2014:EEA

[RG14] Veronika Rocková and Edward I. George. EMVS: The EM approach to Bayesian variable selection. Journal of the American Statistical Association, 109(506):828–846, 2014. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Reiss:2017:C

[RG17a] Philip T. Reiss and Jeff Goldsmith. Comment. Journal of the American Statistical Association, 112(517):161–164, 2017. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Rockova:2017:FBF

[RG17b] Veronika Rocková and Edward I. George. Fast Bayesian factor analysis via automatic rotations to sparsity. Journal of the American Statistical Association, 111(516):1608–1622, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Rockovxe:2018:SSL

[RG18] Veronika Rocková and Edward I. George. The spike-and-slab LASSO. Journal of the American Statistical Association, 113 (521):431–444, 2018. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Robbins:2013:IHD

[RGH13a] Michael W. Robbins, Sujit K. Ghosh, and Joshua D. Habiger. Imputation in high-dimensional economic data as applied to the agricultural resource management survey. Journal of the American Statistical Association, 108(501):81–95, 2013. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Rougier:2013:SOE

[RGH13b] Jonathan Rougier, Michael Goldstein, and Leanna House. Second-order exchangeability analysis for multimodel ensembles. Journal of the American Statistical Association, 108(503):852– 863, 2013. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Robbins:2016:GRC

[RGL16] Michael W. Robbins, Colin M. Gallagher, and Robert B. Lund. A general regression changepoint test for time series data. Journal of the American Statistical Association, 111(514):670–??, 2016. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Rizopoulos:2014:CDP

[RHCT14] Dimitris Rizopoulos, Laura A. Hatfield, Bradley P. Carlin, and Johanna J. M. Takkenberg. Combining dynamic predictions from joint models for longitudinal and time-to-event data using Bayesian model averaging. Journal of the American Statistical Association, 109(508):1385–1397, 2014. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Radchenko:2010:VSU

[RJ10] Peter Radchenko and Gareth M. James. Variable selection using adaptive nonlinear interaction structures in high dimensions. Journal of the American Statistical Association, 105(492):1541– 1553, December 2010. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ruth:2011:NTH

[RK11] David M. Ruth and Robert A. Koyak. Nonparametric tests for homogeneity based on non-bipartite matching. Journal of the American Statistical Association, 106(496):1615–1625, December 2011. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ren:2019:TFH

[RKFL19] Zhao Ren, Yongjian Kang, Yingying Fan, and Jinchi Lv. Tuningfree heterogeneous inference in massive networks. Journal of the American Statistical Association, 114(528):1908–1925, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Robbins:2011:CNA

[RLGL11] Michael W. Robbins, Robert B. Lund, Colin M. Gallagher, and QiQi Lu. Changepoints in the North Atlantic tropical cyclone record. Journal of the American Statistical Association, 106 (493):89–99, March 2011. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ryu:2013:SST

[RLM13] Duchwan Ryu, Faming Liang, and Bani K. Mallick. Sea surface temperature modeling using radial basis function networks with a dynamically weighted particle filter. Journal of the American Statistical Association, 108(501):111–123, 2013. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Rybin:2018:PRL

[RLP^{+}18] Denis Rybin, Robert Lew, Michael J. Pencina, Maurizio Fava, and Gheorghe Doros. Placebo response as a latent characteristic: Application to analysis of sequential parallel comparison design studies. Journal of the American Statistical Association, 113(524):1411–1430, 2018. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Rosenblum:2014:OTT

[RLY14] Michael Rosenblum, Han Liu, and En-Hsu Yen. Optimal tests of treatment effects for the overall population and two subpopulations in randomized trials, using sparse linear programming. Journal of the American Statistical Association, 109(507):1216– 1228, 2014. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Risk:2019:LNG

[RMR19] Benjamin B. Risk, David S. Matteson, and David Ruppert. Linear non-Gaussian component analysis via maximum likelihood. Journal of the American Statistical Association, 114(525):332– 343, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Reimherr:2016:EVC

[RN16] Matthew Reimherr and Dan Nicolae. Estimating variance components in functional linear models with applications to genetic heritability. Journal of the American Statistical Association, 111 (513):407–422, 2016. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Rockovxe:2018:PEV

[Roc18] Veronika Rocková. Particle EM for variable selection. Journal of the American Statistical Association, 113(524):1684–1697, 2018. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Rodriguez:2013:BBT

[Rod13] Robert N. Rodriguez. Building the big tent for statistics. Journal of the American Statistical Association, 108(501):1–6, 2013. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Rosenbaum:2010:DSE

[Ros10] Paul R. Rosenbaum. Design sensitivity and efficiency in observational studies. Journal of the American Statistical Association, 105(490):692–702, June 2010. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Rosenbaum:2011:SAE

[Ros11] Paul R. Rosenbaum. Some approximate evidence factors in observational studies. Journal of the American Statistical Association, 106(493):285–295, March 2011. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Rosenbaum:2014:WSS

[Ros14] Paul R. Rosenbaum. Weighted M-statistics with superior design sensitivity in matched observational studies with multiple controls. Journal of the American Statistical Association, 109(507): 1145–1158, 2014. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Rosenbaum:2015:BES

[Ros15a] Paul R. Rosenbaum. Bahadur efficiency of sensitivity analyses in observational studies. Journal of the American Statistical Association, 110(509):205–217, 2015. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Rosenbaum:2015:SCU

[Ros15b] Paul R. Rosenbaum. Some counterclaims undermine themselves in observational studies. Journal of the American Statistical Association, 110(512):1389–1398, 2015. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.tandfonline.com/doi/abs/10.1080/01621459.
2015.1054489.

Rosenblum:2017:C

[Ros17] Michael Rosenblum. Comment. Journal of the American Statistical Association, 111(516):1541–1542, 2017. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [CZK17a, CZK17b].

Roth:2018:C

[Rot18] Aaron Roth. Comment. Journal of the American Statistical Association, 113(521):208–211, 2018. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Risser:2019:SDM

[RPS19] Mark D. Risser, Christopher J. Paciorek, and Dáithí A. Stone. Spatially dependent multiple testing under model misspecification, with application to detection of anthropogenic influence on extreme climate events. Journal of the American Statistical Association, 114(525):61–78, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Radchenko:2015:IMS

[RQJ15] Peter Radchenko, Xinghao Qiao, and Gareth M. James. Index models for sparsely sampled functional data. Journal of the American Statistical Association, 110(510):824–836, 2015. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Rossell:2018:TBV

[RR18] David Rossell and Francisco J. Rubio. Tractable Bayesian variable selection: Beyond normality. Journal of the American Statistical Association, 113(524):1742–1758, 2018. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Richardson:2017:MER

[RRW17] Thomas S. Richardson, James M. Robins, and Linbo Wang. On modeling and estimation for the relative risk and risk difference. Journal of the American Statistical Association, 112(519):1121– 1130, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Racz:2010:BFM

[RS10] Michael J. Racz and J. Sedransk. Bayesian and frequentist methods for provider profiling using risk-adjusted assessments of medical outcomes. Journal of the American Statistical Association, 105(489):48–58, March 2010. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Roberts:2016:MTE

[RSA16] Margaret E. Roberts, Brandon M. Stewart, and Edoardo M. Airoldi. A model of text for experimentation in the social sciences. Journal of the American Statistical Association, 111(515): 988–1003, 2016. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Rashid:2014:SSS

[RSI14] Naim Rashid, Wei Sun, and Joseph G. Ibrahim. Some statistical strategies for DAE-seq data analysis: Variable selection and modeling dependencies among observations. Journal of the American Statistical Association, 109(505):78–94, 2014. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Robbins:2017:FSC

[RSK17] Michael W. Robbins, Jessica Saunders, and Beau Kilmer. A framework for synthetic control methods with high-dimensional, micro-level data: Evaluating a neighborhood-specific crime intervention. Journal of the American Statistical Association, 112 (517):109–126, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Rossell:2017:NPH

[RT17] David Rossell and Donatello Telesca. Nonlocal priors for highdimensional estimation. Journal of the American Statistical Association, 112(517):254–265, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Reid:2018:GFE

[RTT18] Stephen Reid, Jonathan Taylor, and Robert Tibshirani. A general framework for estimation and inference from clusters of features. Journal of the American Statistical Association, 113(521): 280–293, 2018. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Rao:2010:PEL

[RW10] J. N. K. Rao and Changbao Wu. Pseudo–empirical likelihood inference for multiple frame surveys. Journal of the American Statistical Association, 105(492):1494–1503, December 2010. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Rothe:2013:MTC

[RW13] Christoph Rothe and Dominik Wied. Misspecification testing in a class of conditional distributional models. Journal of the American Statistical Association, 108(501):314–324, 2013. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Raymer:2013:IME

[RWF^{+}13] James Raymer, Arkadiusz Wiśniowski, Jonathan J. Forster, Peter W. F. Smith, and Jakub Bijak. Integrated modeling of European migration. Journal of the American Statistical Association, 108(503):801–819, 2013. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Rosenthal:2014:SRM

[RWKS14] Michael Rosenthal, Wei Wu, Eric Klassen, and Anuj Srivastava. Spherical regression models using projective linear transformations. Journal of the American Statistical Association, 109(508): 1615–1624, 2014. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Rosen:2012:AAS

[RWS12] Ori Rosen, Sally Wood, and David S. Stoffer. AdaptSPEC: Adaptive spectral estimation for nonstationary time series. Journal of the American Statistical Association, 107(500):1575–1589, 2012. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Sadinle:2017:BEB

[Sad17] Mauricio Sadinle. Bayesian estimation of bipartite matchings for record linkage. Journal of the American Statistical Association, 112(518):600–612, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Sahu:2010:CCS

[Sah10] Sujit K. Sahu. Comment: Comment: Statistical dependence in stream networks. Journal of the American Statistical Association, 105(489):21–22, March 2010. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Samworth:2011:CAC

[Sam11] Richard J. Samworth. Comment: “Adaptive Confidence Intervals for the Test Error in Classification”. Journal of the American Statistical Association, 106(495):914–915, September 2011. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Sardy:2012:SBI

[Sar12] Sylvain Sardy. Smooth blockwise iterative thresholding: a smooth fixed point estimator based on the likelihood’s block gradient. Journal of the American Statistical Association, 107(498): 800–813, 2012. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Santacatterina:2018:OPW

[SB18] Michele Santacatterina and Matteo Bottai. Optimal probability weights for inference with constrained precision. Journal of the American Statistical Association, 113(523):983–991, 2018. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Schofield:2016:MBA

[SBG^{+}16] Matthew R. Schofield, Richard J. Barker, Andrew Gelman, Edward R. Cook, and Keith R. Briffa. A model-based approach to climate reconstruction using tree-ring data. Journal of the American Statistical Association, 111(513):93–106, 2016. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Sen:2012:FTC

[SC12] Bodhisattva Sen and Probal Chaudhuri. On fractile transformation of covariates in regression. Journal of the American Statistical Association, 107(497):349–361, 2012. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Sewell:2015:LSM

[SC15] Daniel K. Sewell and Yuguo Chen. Latent space models for dynamic networks. Journal of the American Statistical Association, 110(512):1646–1657, 2015. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/01621459.2014.988214.

Sarkar:2013:MTT

[SCG13] Sanat K. Sarkar, Jingjing Chen, and Wenge Guo. Multiple testing in a two-stage adaptive design with combination tests controlling FDR. Journal of the American Statistical Association, 108(504):1385–1401, 2013. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Schwartzman:2010:CCV

[Sch10] Armin Schwartzman. Comment: “Correlated z-Values and the Accuracy of Large-Scale Statistical Estimates”. Journal of the American Statistical Association, 105(491):1059–1063, September 2010. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Schweinberger:2011:ISD

[Sch11] Michael Schweinberger. Instability, sensitivity, and degeneracy of discrete exponential families. Journal of the American Statistical Association, 106(496):1361–1370, December 2011. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Schwartzman:2012:CFV

[Sch12] Armin Schwartzman. Comment: FDP vs FDR and the effect of conditioning. Journal of the American Statistical Association, 107(499):1039–1041, 2012. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Schenker:2015:WYI

[Sch15] Nathaniel Schenker. Why your involvement matters. Journal of the American Statistical Association, 110(509):1–5, 2015. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Small:2013:CDD

[SCHR13] Dylan S. Small, Jing Cheng, M. Elizabeth Halloran, and Paul R. Rosenbaum. Case definition and design sensitivity. Journal of the American Statistical Association, 108(504):1457–1468, 2013. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Strait:2019:ADU

[SCK19] Justin Strait, Oksana Chkrebtii, and Sebastian Kurtek. Automatic detection and uncertainty quantification of landmarks on elastic curves. Journal of the American Statistical Association, 114(527):1002–1017, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Sarkar:2018:BSMb

[SCM^{+}18] Abhra Sarkar, Jonathan Chabout, Joshua Jones Macopson, Erich D. Jarvis, and David B. Dunson. Bayesian semiparametric mixed effects Markov models with application to vocalization syntax. Journal of the American Statistical Association, 113(524):1515–1527, 2018. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Scarpa:2014:ESB

[SD14] Bruno Scarpa and David B. Dunson. Enriched stick-breaking processes for functional data. Journal of the American Statistical Association, 109(506):647–660, 2014. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Sarkar:2017:BNM

[SD17a] Abhra Sarkar and David B. Dunson. Bayesian nonparametric modeling of higher order Markov chains. Journal of the American Statistical Association, 111(516):1791–1803, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Schnurr:2017:TSB

[SD17b] Alexander Schnurr and Herold Dehling. Testing for structural breaks via ordinal pattern dependence. Journal of the American Statistical Association, 112(518):706–720, 2017. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Scealy:2015:RPC

[SdCG^{+}15] J. L. Scealy, Patrice de Caritat, Eric C. Grunsky, Michail T. Tsagris, and A. H. Welsh. Robust principal component analysis for power transformed compositional data. Journal of the American Statistical Association, 110(509):136–148, 2015. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Steingrimsson:2019:CUR

[SDS19] Jon Arni Steingrimsson, Liqun Diao, and Robert L. Strawderman. Censoring unbiased regression trees and ensembles. Journal of the American Statistical Association, 114(525):370–383, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Schwartzman:2010:GCE

[SDT10] Armin Schwartzman, Robert F. Dougherty, and Jonathan E. Taylor. Group comparison of eigenvalues and eigenvectors of diffusion tensors. Journal of the American Statistical Association, 105(490):588–599, June 2010. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Sun:2018:OPF

[SDWM18] Xiaoxiao Sun, Pang Du, Xiao Wang, and Ping Ma. Optimal penalized function-on-function regression under a reproducing kernel Hilbert space framework. Journal of the American Statistical Association, 113(524):1601–1611, 2018. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Stoner:2019:HFC

[SEdS19] Oliver Stoner, Theo Economou, and Gabriela Drummond Marques da Silva. A hierarchical framework for correcting underreporting in count data. Journal of the American Statistical Association, 114(528):1481–1492, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Sadinle:2013:GFS

[SF13] Mauricio Sadinle and Stephen E. Fienberg. A generalized Fellegi– Sunter framework for multiple record linkage with application to homicide record systems. Journal of the American Statistical Association, 108(502):385–397, 2013. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Scheipl:2012:SSP

[SFK12] Fabian Scheipl, Ludwig Fahrmeir, and Thomas Kneib. Spikeand-slab priors for function selection in structured additive regression models. Journal of the American Statistical Association, 107(500):1518–1532, 2012. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Swihart:2018:SMS

[SFM18] Bruce J. Swihart, Michael P. Fay, and Kazutoyo Miura. Statistical methods for standard membrane-feeding assays to measure transmission blocking or reducing activity in malaria. Journal of the American Statistical Association, 113(522):534–545, 2018. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Sloughter:2010:PWS

[SGR10] J. McLean Sloughter, Tilmann Gneiting, and Adrian E. Raftery. Probabilistic wind speed forecasting using ensembles and Bayesian model averaging. Journal of the American Statistical Association, 105(489):25–35, March 2010. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Stingo:2013:IBM

[SGVC13] Francesco C. Stingo, Michele Guindani, Marina Vannucci, and Vince D. Calhoun. An integrative Bayesian modeling approach to imaging genetics. Journal of the American Statistical Association, 108(503):876–891, 2013. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Shen:2010:GPT

[SH10] Xiaotong Shen and Hsin-Cheng Huang. Grouping pursuit through a regularization solution surface. Journal of the American Statistical Association, 105(490):727–739, June 2010. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Schennach:2013:NIS

[SH13] S. M. Schennach and Yingyao Hu. Nonparametric identification and semiparametric estimation of classical measurement error models without side information. Journal of the American Statistical Association, 108(501):177–186, 2013. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Shen:2015:ISA

[SH15] Juan Shen and Xuming He. Inference for subgroup analysis with a structured logistic-normal mixture model. Journal of the American Statistical Association, 110(509):303–312, 2015. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Shao:2010:DWB

[Sha10] Xiaofeng Shao. The dependent wild bootstrap. Journal of the American Statistical Association, 105(489):218–235, March 2010. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Shao:2015:SNT

[Sha15] Xiaofeng Shao. Self-normalization for time series: A review of recent developments. Journal of the American Statistical Association, 110(512):1797–1817, 2015. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/01621459.2015.
1050493.

Shedden:2011:CPV

[She11] Kerby Shedden. Comment: “Population Value Decomposition, a Framework for the Analysis of Image Populations”. Journal of the American Statistical Association, 106(495):796–797, September 2011. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Shedden:2013:C

[She13] Kerby Shedden. Comment. Journal of the American Statistical Association, 108(504):1162–1163, 2013. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [MJPW13b].

Steorts:2017:BAG

[SHF17] Rebecca C. Steorts, Rob Hall, and Stephen E. Fienberg. A Bayesian approach to graphical record linkage and deduplication. Journal of the American Statistical Association, 111(516): 1660–1672, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Swanson:2018:PIA

[SHM^{+}18] Sonja A. Swanson, Miguel A. Hernán, Matthew Miller, James M. Robins, and Thomas S. Richardson. Partial identification of the average treatment effect using instrumental variables: Review of methods for binary instruments, treatments, and outcomes. Journal of the American Statistical Association, 113(522):933– 947, 2018. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Saul:2019:DEU

[SHM19] Bradley C. Saul, Michael G. Hudgens, and Michael A. Mallin. Downstream effects of upstream causes. Journal of the American Statistical Association, 114(528):1493–1504, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Savchuk:2010:ICV

[SHS10] Olga Y. Savchuk, Jeffrey D. Hart, and Simon J. Sheather. Indirect cross-validation for density estimation. Journal of the American Statistical Association, 105(489):415–423, March 2010. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Sun:2017:NBR

[SHW17] Yifei Sun, Chiung-Yu Huang, and Mei-Cheng Wang. Nonparametric benefit-risk assessment using marker process in the presence of a terminal event. Journal of the American Statistical Association, 112(518):826–836, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Siddique:2013:PIT

[Sid13] Zahra Siddique. Partially identified treatment effects under imperfect compliance: The case of domestic violence. Journal of the American Statistical Association, 108(502):504–513, 2013. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Stroud:2014:BMF

[SJ14] Jonathan R. Stroud and Michael S. Johannes. Bayesian modeling and forecasting of 24-hour high-frequency volatility. Journal of the American Statistical Association, 109(508):1368–1384, 2014. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Shao:2019:BMC

[SJDT19] Stephane Shao, Pierre E. Jacob, Jie Ding, and Vahid Tarokh. Bayesian model comparison with the Hyvärinen score: Computation and consistency. Journal of the American Statistical Association, 114(528):1826–1837, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Smith:2012:ECM

[SK12] Michael S. Smith and Mohamad A. Khaled. Estimation of copula models with discrete margins via Bayesian data augmentation. Journal of the American Statistical Association, 107(497):290– 303, 2012. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Strait:2017:LCE

[SKBM17] Justin Strait, Sebastian Kurtek, Emily Bartha, and Steven N. MacEachern. Landmark-constrained elastic shape analysis of planar curves. Journal of the American Statistical Association, 112(518):521–533, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Scott:2015:FDR

[SKS^{+}15] James G. Scott, Ryan C. Kelly, Matthew A. Smith, Pengcheng Zhou, and Robert E. Kass. False discovery rate regression: An application to neural synchrony detection in primary visual cortex. Journal of the American Statistical Association, 110(510): 459–471, 2015. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Storelvmo:2011:CGW

[SL11] T. Storelvmo and T. Leirvik. Comment: “Global Warming and Local Dimming: The Statistical Evidence”. Journal of the American Statistical Association, 106(494):465–467, June 2011. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Sobel:2014:CIF

[SL14] Michael E. Sobel and Martin A. Lindquist. Causal inference for fMRI time series data with systematic errors of measurement in a balanced on/off study of social evaluative threat. Journal of the American Statistical Association, 109(507):967–976, 2014. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Song:2015:HDV

[SL15] Qifan Song and Faming Liang. High-dimensional variable selection with reciprocal L_{1}-regularization. Journal of the American Statistical Association, 110(512):1607–1620, 2015. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/01621459.2014.984812.

Sun:2019:DTC

[SL19] Will Wei Sun and Lexin Li. Dynamic tensor clustering. Journal of the American Statistical Association, 114(528):1894–1907, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Sun:2015:IDD

[SLC^{+}15] Wei Sun, Yufeng Liu, James J. Crowley, Ting-Huei Chen, Hua Zhou, Haitao Chu, Shunping Huang, Pei-Fen Kuan, Yuan Li, Darla Miller, Ginger Shaw, Yichao Wu, Vasyl Zhabotynsky, Leonard McMillan, Fei Zou, Patrick F. Sullivan, and Fernando Pardo-Manuel De Villena. IsoDOT detects differential RNAisoform expression/usage with respect to a categorical or continuous covariate with high sensitivity and specificity. Journal of the American Statistical Association, 110(511):975–986, 2015. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Schwartz:2011:BSA

[SLM11] Scott L. Schwartz, Fan Li, and Fabrizia Mealli. A Bayesian semiparametric approach to intermediate variables in causal inference. Journal of the American Statistical Association, 106(496): 1331–1344, December 2011. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Storlie:2015:CCM

[SLR^{+}15] Curtis B. Storlie, William A. Lane, Emily M. Ryan, James R. Gattiker, and David M. Higdon. Calibration of computational models with categorical parameters and correlated outputs via Bayesian smoothing spline ANOVA. Journal of the American Statistical Association, 110(509):68–82, 2015. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Shi:2018:MDF

[SLS18] Chengchun Shi, Wenbin Lu, and Rui Song. A massive data framework for M-estimators with cubic-rate. Journal of the American Statistical Association, 113(524):1698–1709, 2018. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

She:2016:ROC

[SLW16] Yiyuan She, Shijie Li, and Dapeng Wu. Robust orthogonal complement principal component analysis. Journal of the American Statistical Association, 111(514):763–??, 2016. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Sadinle:2019:LAS

[SLW19] Mauricio Sadinle, Jing Lei, and Larry Wasserman. Least ambiguous set-valued classifiers with bounded error levels. Journal of the American Statistical Association, 114(525):223–234, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Senturk:2010:FVC

[ŞM10] Damla Şentürk and Hans-Georg Müller. Functional varying coefficient models for longitudinal data. Journal of the American Statistical Association, 105(491):1256–1264, September 2010. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Sharma:2012:OST

[SM12a] Gaurav Sharma and Thomas Mathew. One-sided and two-sided tolerance intervals in general mixed and random effects models using small-sample asymptotics. Journal of the American Statistical Association, 107(497):258–267, 2012. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Sun:2012:MTC

[SM12b] Wenguang Sun and Alexander C. McLain. Multiple testing of composite null hypotheses in heteroscedastic models. Journal of the American Statistical Association, 107(498):673–687, 2012. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Sinha:2016:APO

[SM16] Samiran Sinha and Yanyuan Ma. Analysis of proportional odds models with censoring and errors-in-covariates. Journal of the American Statistical Association, 111(515):1301–1312, 2016. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Smith:2010:MLD

[SMAC10] Michael Smith, Aleksey Min, Carlos Almeida, and Claudia Czado. Modeling longitudinal data using a pair-copula decomposition of serial dependence. Journal of the American Statistical Association, 105(492):1467–1479, December 2010. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Smith:2010:CCH

[Smi10] Richard L. Smith. Comment: “Comment: Hierarchical Statistical Modeling for Paleoclimate Reconstruction”. Journal of the American Statistical Association, 105(491):905–910, September 2010. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Storlie:2013:BRA

[SMQ^{+}13] Curtis B. Storlie, Sarah E. Michalak, Heather M. Quinn, Andrew J. Dubois, Steven A. Wender, and David H. Dubois. A Bayesian reliability analysis of neutron-induced errors in high performance computing hardware. Journal of the American Statistical Association, 108(502):429–440, 2013. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Shirani-Mehr:2018:DBV

[SMRGG18] Houshmand Shirani-Mehr, David Rothschild, Sharad Goel, and Andrew Gelman. Disentangling bias and variance in election polls. Journal of the American Statistical Association, 113(522): 607–614, 2018. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Snavely:2019:BRR

[Sna19] Anna Snavely. Book review: Randomization, Masking, and Allocation Concealment. Journal of the American Statistical Association, 114(527):1423–1424, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

She:2011:ODU

[SO11] Yiyuan She and Art B. Owen. Outlier detection using nonconvex penalized regression. Journal of the American Statistical Association, 106(494):626–639, June 2011. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Schrxfder:2019:FFS

[SO19] Anna Louise Schröder and Hernando Ombao. FreSpeD: Frequency-specific change-point detection in epileptic seizure multi-channel EEG data. Journal of the American Statistical Association, 114(525):115–128, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Sobel:2012:DMB

[Sob12] Michael E. Sobel. Does marriage boost men’s wages?: Identification of treatment effects in fixed effects regression models for panel data. Journal of the American Statistical Association, 107 (498):521–529, 2012. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Sarkar:2018:BSMa

[SPC^{+}18] Abhra Sarkar, Debdeep Pati, Antik Chakraborty, Bani K. Mallick, and Raymond J. Carroll. Bayesian semiparametric multivariate density deconvolution. Journal of the American Statistical Association, 113(521):401–416, 2018. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Sun:2016:GQR

[SPHL16] Xiaoyan Sun, Limin Peng, Yijian Huang, and HuiChuan J. Lai. Generalizing quantile regression for counting processes with applications to recurrent events. Journal of the American Statistical Association, 111(513):145–156, 2016. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Spieker:2019:CPS

[Spi19] Andrew J. Spieker. Comment on penalized spline of propensity methods for treatment comparison by Zhou, Elliott, and Little. Journal of the American Statistical Association, 114(525):20–23, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [ZEL19a] and rejoinder [ZEL19b].

Satopaa:2017:MPF

[SPU17] Ville A. Satopää, Robin Pemantle, and Lyle H. Ungar. Modeling probability forecasts via information diversity. Journal of the American Statistical Association, 111(516):1623–1633, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Shen:2012:LBS

[SPZ12] Xiaotong Shen, Wei Pan, and Yunzhang Zhu. Likelihood-based selection and sharp parameter estimation. Journal of the American Statistical Association, 107(497):223–232, 2012. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Spirling:2010:IIV

[SQ10] Arthur Spirling and Kevin Quinn. Identifying intraparty voting blocs in the U.K. House of Commons. Journal of the American Statistical Association, 105(490):447–457, June 2010. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Sun:2016:SNN

[SQC16] Will Wei Sun, Xingye Qiao, and Guang Cheng. Stabilized nearest neighbor classifier and its statistical properties. Journal of the American Statistical Association, 111(515):1254–1265, 2016. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Storlie:2017:SMN

[SRR^{+}17] Curtis B. Storlie, Brian J. Reich, William N. Rust, Lawrence O. Ticknor, Amanda M. Bonnie, Andrew J. Montoya, and Sarah E. Michalak. Spatiotemporal modeling of node temperatures in supercomputers. Journal of the American Statistical Association, 112(517):92–108, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Shah:2015:C

[SS15] Rajen D. Shah and Richard J. Samworth. Comment. Journal of the American Statistical Association, 110(512):1439–1442, 2015. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/01621459.2015.1102142.

Steingrimsson:2017:ESA

[SS17] Jon Arni Steingrimsson and Robert L. Strawderman. Estimation in the semiparametric accelerated failure time model with missing covariates: Improving efficiency through augmentation. Journal of the American Statistical Association, 112(519):1221– 1235, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Stroud:2010:EKF

[SSL^{+}10] Jonathan R. Stroud, Michael L. Stein, Barry M. Lesht, David J. Schwab, and Dmitry Beletsky. An ensemble Kalman filter and smoother for satellite data assimilation. Journal of the American Statistical Association, 105(491):978–990, September 2010. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Sommerfeld:2018:CRS

[SSS18] Max Sommerfeld, Stephan Sain, and Armin Schwartzman. Confidence regions for spatial excursion sets from repeated random field observations, with an application to climate. Journal of the American Statistical Association, 113(523):1327–1340, 2018. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Sun:2012:JAL

[SSZL12] Liuquan Sun, Xinyuan Song, Jie Zhou, and Lei Liu. Joint analysis of longitudinal data with informative observation times and a dependent terminal event. Journal of the American Statistical Association, 107(498):688–700, 2012. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Simon:2015:PAT

[ST15] Noah Simon and Robert Tibshirani. A permutation approach to testing interactions for binary response by comparing correlations between classes. Journal of the American Statistical Association, 110(512):1707–1716, 2015. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/01621459.2014.993079.

Sun:2017:MCS

[ST17] Fasheng Sun and Boxin Tang. A method of constructing spacefilling orthogonal designs. Journal of the American Statistical Association, 112(518):683–689, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Sun:2018:IPW

[ST18] BaoLuo Sun and Eric J. Tchetgen Tchetgen. On inverse probability weighting for nonmonotone missing at random data. Journal of the American Statistical Association, 113(521):369–379, 2018. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Sussman:2012:CAS

[STFP12] Daniel L. Sussman, Minh Tang, Donniell E. Fishkind, and Carey E. Priebe. A consistent adjacency spectral embedding for stochastic blockmodel graphs. Journal of the American Statistical Association, 107(499):1119–1128, 2012. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Su:2019:BR

[Su19] Jing Su. Book review. Journal of the American Statistical Association, 114(526):948, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Sengupta:2016:SDB

[SVS16] Srijan Sengupta, Stanislav Volgushev, and Xiaofeng Shao. A subsampled double bootstrap for massive data. Journal of the American Statistical Association, 111(515):1222–1232, 2016. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Schoen:2017:TLO

[SVTG17] Eric D. Schoen, Nha Vo-Thanh, and Peter Goos. Two-level orthogonal screening designs with 24, 28, 32, and 36 runs. Journal of the American Statistical Association, 112(519):1354–1369, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Sun:2011:MTP

[SW11] Wenguang Sun and Zhi Wei. Multiple testing for pattern identification, with applications to microarray time-course experiments. Journal of the American Statistical Association, 106(493):73–88, March 2011. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Song:2014:PHM

[SW14] Xiao Song and Ching-Yun Wang. Proportional hazards model with covariate measurement error and instrumental variables. Journal of the American Statistical Association, 109(508):1636– 1646, 2014. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Scealy:2017:DME

[SW17a] J. L. Scealy and A. H. Welsh. A directional mixed effects model for compositional expenditure data. Journal of the American Statistical Association, 112(517):24–36, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Schennach:2017:SPM

[SW17b] Susanne M. Schennach and Daniel Wilhelm. A simple parametric model selection test. Journal of the American Statistical Association, 112(520):1663–1674, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Sun:2017:EUM

[SW17c] Yifei Sun and Mei-Cheng Wang. Evaluating utility measurement from recurrent marker processes in the presence of competing terminal events. Journal of the American Statistical Association, 112(518):745–756, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Scealy:2019:SMF

[SW19] J. L. Scealy and Andrew T. A. Wood. Scaled von Mises–Fisher distributions and regression models for paleomagnetic directional data. Journal of the American Statistical Association, 114(528): 1547–1560, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Steele:2016:LML

[SWCB16] Fiona Steele, Elizabeth Washbrook, Christopher Charlton, and William J. Browne. A longitudinal mixed logit model for estimation of push and pull effects in residential location choice. Journal of the American Statistical Association, 111(515):1061–1074, 2016. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

She:2018:GRE

[SWJ18] Yiyuan She, Zhifeng Wang, and He Jiang. Group regularized estimation under structural hierarchy. Journal of the American Statistical Association, 113(521):445–454, 2018. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Salter:2019:UQC

[SWSK19] James M. Salter, Daniel B. Williamson, John Scinocca, and Viatcheslav Kharin. Uncertainty quantification for computer models with spatial output using calibration-optimal bases. Journal of the American Statistical Association, 114(528):1800–1814, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Stefanski:2014:VSN

[SWW14] L. A. Stefanski, Yichao Wu, and Kyle White. Variable selection in nonparametric classification via measurement error model selection likelihoods. Journal of the American Statistical Association, 109(506):574–589, 2014. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Shephard:2017:CTA

[SY17] Neil Shephard and Justin J. Yang. Continuous time analysis of fleeting discrete price moves. Journal of the American Statistical Association, 112(519):1090–1106, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Shi:2018:PCC

[SY18] Peng Shi and Lu Yang. Pair copula constructions for insurance experience rating. Journal of the American Statistical Association, 113(521):122–133, 2018. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Shin:2019:BSE

[SYS^{+}19] Seung Jun Shin, Ying Yuan, Louise C. Strong, Jasmina Bojadzieva, and Wenyi Wang. Bayesian semiparametric estimation of cancer-specific age-at-onset penetrance with application to Li–Fraumeni Syndrome. Journal of the American Statistical Association, 114(526):541–552, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Sgouropoulos:2015:MDM

[SYY15] Nikolaos Sgouropoulos, Qiwei Yao, and Claudia Yastremiz. Matching a distribution by matching quantiles estimation. Journal of the American Statistical Association, 110(510):742–759, 2015. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Shao:2010:TCP

[SZ10] Xiaofeng Shao and Xianyang Zhang. Testing for change points in time series. Journal of the American Statistical Association, 105(491):1228–1240, September 2010. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Shao:2014:MDC

[SZ14] Xiaofeng Shao and Jingsi Zhang. Martingale difference correlation and its use in high-dimensional variable screening. Journal of the American Statistical Association, 109(507):1302–1318, 2014. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Schmertmann:2014:BFC

[SZGM14] Carl Schmertmann, Emilio Zagheni, Joshua R. Goldstein, and Mikko Myrskylä. Bayesian forecasting of cohort fertility. Journal of the American Statistical Association, 109(506):500–513, 2014. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Shi:2012:IRM

[SZI^{+}12] Xiaoyan Shi, Hongtu Zhu, Joseph G. Ibrahim, Faming Liang, Jeffrey Lieberman, and Martin Styner. Intrinsic regression models for medial representation of subcortical structures. Journal of the American Statistical Association, 107(497):12–23, 2012. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Sun:2015:SSP

[SZLI15] Qiang Sun, Hongtu Zhu, Yufeng Liu, and Joseph G. Ibrahim. SPReM: Sparse projection regression model for high-dimensional linear regression. Journal of the American Statistical Association, 110(509):289–302, 2015. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Taddy:2010:AMM

[Tad10] Matthew A. Taddy. Autoregressive mixture models for dynamic spatial Poisson processes: Application to tracking intensity of violent crime. Journal of the American Statistical Association, 105(492):1403–1417, December 2010. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Taddy:2013:MIR

[Tad13a] Matt Taddy. Multinomial inverse regression for text analysis. Journal of the American Statistical Association, 108(503):755– 770, 2013. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See comments [Gri13, Ble13, Tad13b].

Taddy:2013:RES

[Tad13b] Matt Taddy. Rejoinder: Efficiency and structure in MNIR. Journal of the American Statistical Association, 108(503):772–774, 2013. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [Tad13a].

Taddy:2017:C

[Tad17] Matt Taddy. Comment. Journal of the American Statistical Association, 111(516):1403–1405, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [AB17a, Air17].

Tian:2014:SME

[TAGT14] Lu Tian, Ash A. Alizadeh, Andrew J. Gentles, and Robert Tibshirani. A simple method for estimating interactions between a treatment and a large number of covariates. Journal of the American Statistical Association, 109(508):1517–1532, 2014. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Tan:2010:MNS

[Tan10] Zhiqiang Tan. Marginal and nested structural models using instrumental variables. Journal of the American Statistical Association, 105(489):157–169, March 2010. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Tansey:2017:MSD

[TARS17] Wesley Tansey, Alex Athey, Alex Reinhart, and James G. Scott. Multiscale spatial density smoothing: An application to largescale radiological survey and anomaly detection. Journal of the American Statistical Association, 112(519):1047–1063, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Tran:2017:C

[TB17] Dustin Tran and David M. Blei. Comment. Journal of the American Statistical Association, 112(517):156–158, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Toth:2011:BCR

[TE11] Daniell Toth and John L. Eltinge. Building consistent regression trees from complex sample data. Journal of the American Statistical Association, 106(496):1626–1636, December 2011. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Telesca:2012:MCC

[TEKM12] Donatello Telesca, Elena A. Erosheva, Derek A. Kreager, and Ross L. Matsueda. Modeling criminal careers as departures from a unimodal population age-crime curve: The case of marijuana use. Journal of the American Statistical Association, 107(500): 1427–1440, 2012. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Tang:2018:MRE

[TGHS18] Xueying Tang, Malay Ghosh, Neung Soo Ha, and Joseph Sedransk. Modeling random effects using global-local shrinkage priors in small area estimation. Journal of the American Statistical Association, 113(524):1476–1489, 2018. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Taddy:2011:DTL

[TGP11] Matthew A. Taddy, Robert B. Gramacy, and Nicholas G. Polson. Dynamic trees for learning and design. Journal of the American Statistical Association, 106(493):109–123, March 2011. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Tang:2015:RAC

[TGZ15] Xu Tang, Fah F. Gan, and Lingyun Zhang. Risk-adjusted cumulative sum charting procedure based on multiresponses. Journal of the American Statistical Association, 110(509):16–26, 2015. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Titsias:2016:SIH

[THY16] Michalis K. Titsias, Christopher C. Holmes, and Christopher Yau. Statistical inference in hidden Markov models using ksegment constraints. Journal of the American Statistical Association, 111(513):200–215, 2016. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Tansey:2018:FDR

[TKPS18] Wesley Tansey, Oluwasanmi Koyejo, Russell A. Poldrack, and James G. Scott. False discovery rate smoothing. Journal of the American Statistical Association, 113(523):1156–1171, 2018. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Tang:2016:GQL

[TLG16] Jin Tang, Yehua Li, and Yongtao Guan. Generalized quasilikelihood ratio tests for semiparametric analysis of covariance models in longitudinal data. Journal of the American Statistical Association, 111(514):736–??, 2016. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Telesca:2012:MPE

[TMK^{+}12] Donatello Telesca, Peter Müller, Steven M. Kornblau, Marc A. Suchard, and Yuan Ji. Modeling protein expression and protein signaling pathways. Journal of the American Statistical Association, 107(500):1372–1384, 2012. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Thomas:2018:RCI

[TMP18] Zachary M. Thomas, Steven N. MacEachern, and Mario Peruggia. Reconciling curvature and importance sampling based procedures for summarizing case influence in Bayesian models. Journal of the American Statistical Association, 113(524):1669–1683, 2018. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Thall:2014:OSD

[TNZM14] Peter F. Thall, Hoang Q. Nguyen, Sarah Zohar, and Pierre Maton. Optimizing sedative dose in preterm infants undergoing treatment for respiratory distress syndrome. Journal of the American Statistical Association, 109(507):931–943, 2014. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Tavakoli:2016:DLD

[TP16a] Shahin Tavakoli and Victor M. Panaretos. Detecting and localizing differences in functional time series dynamics: A case study in molecular biophysics. Journal of the American Statistical Association, 111(515):1020–1035, 2016. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Trippa:2016:C

[TP16b] Lorenzo Trippa and Giovanni Parmigiani. Comment. Journal of the American Statistical Association, 111(515):947–948, 2016. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Tavakoli:2019:RSM

[TPAC19a] Shahin Tavakoli, Davide Pigoli, John A. D. Aston, and John S. Coleman. Rejoinder for “A Spatial Modeling Approach for Linguistic Object Data: Analyzing Dialect Sound Variations Across Great Britain”. Journal of the American Statistical Association, 114(527):1103–1104, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Tavakoli:2019:SMA

[TPAC19b] Shahin Tavakoli, Davide Pigoli, John A. D. Aston, and John S. Coleman. A spatial modeling approach for linguistic object data: Analyzing dialect sound variations across Great Britain. Journal of the American Statistical Association, 114(527):1081–1096, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Tarpey:2010:OPL

[TPLG10] Thaddeus Tarpey, Eva Petkova, Yimeng Lu, and Usha Govindarajulu. Optimal partitioning for linear mixed effects models: Applications to identifying placebo responders. Journal of the American Statistical Association, 105(491):968–977, September 2010. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Tibshirani:2019:EOH

[TR19] Ryan J. Tibshirani and Saharon Rosset. Excess optimism: How biased is the apparent error of an estimator tuned by SURE? Journal of the American Statistical Association, 114(526):697– 712, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Trapani:2018:RSP

[Tra18] Lorenzo Trapani. A randomized sequential procedure to determine the number of factors. Journal of the American Statistical Association, 113(523):1341–1349, 2018. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Trindade:2019:BRL

[Tri19] A. Alexandre Trindade. Book review: Linear Models and the Relevant Distributions and Matrix Algebra, by David A. Harville. Boca Raton, FL: Chapman & Hall/CRC Press, 2018, xiii + 524 pp., $135.00 (H), ISBN: 978-1-13-857833-3. Journal of the American Statistical Association, 114(528):1928–1929, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Tsionas:2017:WWH

[Tsi17] Mike G. Tsionas. “when, where, and how” of efficiency estimation: Improved procedures for stochastic frontier modeling. Journal of the American Statistical Association, 112(519):948– 965, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Tsai:2010:CFM

[TT10] Henghsiu Tsai and Ruey S. Tsay. Constrained factor models. Journal of the American Statistical Association, 105(492): 1593–1605, December 2010. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Todorov:2012:IRL

[TT12] Viktor Todorov and George Tauchen. Inverse realized Laplace transforms for nonparametric volatility density estimation in jump-diffusions. Journal of the American Statistical Association, 107(498):622–635, 2012. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Tibshirani:2016:EPS

[TTLT16a] Ryan J. Tibshirani, Jonathan Taylor, Richard Lockhart, and Robert Tibshirani. Exact post-selection inference for sequential regression procedures. Journal of the American Statistical Association, 111(514):600–??, 2016. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Tibshirani:2016:R

[TTLT16b] Ryan J. Tibshirani, Jonathan Taylor, Richard Lockhart, and Robert Tibshirani. Rejoinder. Journal of the American Statistical Association, 111(514):618–??, 2016. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Tu:2017:C

[Tu17] Wanzhu Tu. Comment. Journal of the American Statistical Association, 112(517):158–161, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Tao:2011:LVM

[TWYZ11] Minjing Tao, Yazhen Wang, Qiwei Yao, and Jian Zou. Large volatility matrix inference via combining low-frequency and highfrequency approaches. Journal of the American Statistical Association, 106(495):1025–1040, September 2011. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Titsias:2017:HBS

[TY17] Michalis K. Titsias and Christopher Yau. The Hamming ball sampler. Journal of the American Statistical Association, 112 (520):1598–1611, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Tang:2019:STM

[TYY^{+}19] Xueying Tang, Yang Yang, Hong-Jie Yu, Qiao-Hong Liao, and Nikolay Bliznyuk. A spatio-temporal modeling framework for surveillance data of multiple infectious pathogens with small laboratory validation sets. Journal of the American Statistical Association, 114(528):1561–1573, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Tao:2015:ASD

[TZF^{+}15] Ran Tao, Donglin Zeng, Nora Franceschini, Kari E. North, Eric Boerwinkle, and Dan-Yu Lin. Analysis of sequence data under multivariate trait-dependent sampling. Journal of the American Statistical Association, 110(510):560–572, 2015. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Tao:2017:ESI

[TZL17] Ran Tao, Donglin Zeng, and Dan-Yu Lin. Efficient semiparametric inference under two-phase sampling, with applications to genetic association studies. Journal of the American Statistical Association, 112(520):1468–1476, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Uematsu:2017:TOR

[UL17] Kazuki Uematsu and Yoonkyung Lee. On theoretically optimal ranking functions in bipartite ranking. Journal of the American Statistical Association, 112(519):1311–1322, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Utts:2017:AS

[Utt17] Jessica Utts. Appreciating statistics. Journal of the American Statistical Association, 111(516):1373–1380, 2017. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Vats:2019:BRS

[Vat19] Dootika Vats. Book review: Simulation and the Monte Carlo Method, 3rd ed. Journal of the American Statistical Association, 114(527):1425, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ventz:2019:BUD

[VCBT19] Steffen Ventz, Matteo Cellamare, Sergio Bacallado, and Lorenzo Trippa. Bayesian uncertainty directed trial designs. Journal of the American Statistical Association, 114(527):962–974, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Volfovsky:2015:TND

[VH15] Alexander Volfovsky and Peter D. Hoff. Testing for nodal dependence in relational data matrices. Journal of the American Statistical Association, 110(511):1037–1046, 2015. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Vanderweele:2013:MSE

[VHJB13] Tyler J. Vanderweele, Guanglei Hong, Stephanie M. Jones, and Joshua L. Brown. Mediation and spillover effects in grouprandomized trials: A case study of the 4Rs educational intervention. Journal of the American Statistical Association, 108 (502):469–482, 2013. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Vandekar:2019:IHD

[VRS19] Simon N. Vandekar, Philip T. Reiss, and Russell T. Shinohara. Interpretable high-dimensional inference via score projection with an application in neuroimaging. Journal of the American Statistical Association, 114(526):820–830, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Vallejos:2015:OBS

[VS15] Catalina A. Vallejos and Mark F. J. Steel. Objective Bayesian survival analysis using shape mixtures of log-normal distributions. Journal of the American Statistical Association, 110(510): 697–710, 2015. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Vincent:2017:EPS

[VT17] Kyle Vincent and Steve Thompson. Estimating population size with link-tracing sampling. Journal of the American Statistical Association, 112(519):1286–1295, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Vermeulen:2015:BRD

[VV15] Karel Vermeulen and Stijn Vansteelandt. Bias-reduced doubly robust estimation. Journal of the American Statistical Association, 110(511):1024–1036, 2015. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

VanAelst:2011:REO

[VW11] Stefan Van Aelst and Gert Willems. Robust and efficient one-way MANOVA tests. Journal of the American Statistical Association, 106(494):706–718, June 2011. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Villa:2015:OAP

[VW15] C. Villa and S. G. Walker. An objective approach to prior mass functions for discrete parameter spaces. Journal of the American Statistical Association, 110(511):1072–1082, 2015. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wager:2018:EIH

[WA18] Stefan Wager and Susan Athey. Estimation and inference of heterogeneous treatment effects using random forests. Journal of the American Statistical Association, 113(523):1228–1242, 2018. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wickramasuriya:2019:OFR

[WAH19] Shanika L. Wickramasuriya, George Athanasopoulos, and Rob J. Hyndman. Optimal forecast reconciliation for hierarchical and grouped time series through trace minimization. Journal of the American Statistical Association, 114(526):804–819, 2019. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wang:2015:EOC

[Wan15] Weizhen Wang. Exact optimal confidence intervals for hypergeometric parameters. Journal of the American Statistical Association, 110(512):1491–1499, 2015. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/01621459.2014.966191.

Wand:2017:FAI

[Wan17a] M. P. Wand. Fast approximate inference for arbitrarily large semiparametric regression models via message passing. Journal of the American Statistical Association, 112(517):137–168, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wand:2017:R

[Wan17b] M. P. Wand. Rejoinder. Journal of the American Statistical Association, 112(517):166–168, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wasserman:2012:C

[Was12] Larry Wasserman. Comment. Journal of the American Statistical Association, 107(499):1035–1036, 2012. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wang:2019:BMCa

[WB19a] Yixin Wang and David M. Blei. The blessings of multiple causes. Journal of the American Statistical Association, 114(528):1574– 1596, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See comments [D’A19, AIP19, IJ19, OST19] and rejoinder [WB19b].

Wang:2019:BMCb

[WB19b] Yixin Wang and David M. Blei. The blessings of multiple causes: Rejoinder. Journal of the American Statistical Association, 114 (528):1616–1619, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [WB19a, IJ19, OST19].

Wang:2019:FCV

[WB19c] Yixin Wang and David M. Blei. Frequentist consistency of variational Bayes. Journal of the American Statistical Association, 114(527):1147–1161, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wang:2015:BPI

[WBCD15] Liangliang Wang, Alexandre Bouchard-Côté, and Arnaud Doucet. Bayesian phylogenetic inference using a combinatorial sequential Monte Carlo method. Journal of the American Statistical Association, 110(512):1362–1374, 2015. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/01621459.2015.1054487.

Wallace:2018:VSS

[WBG^{+}18] Meredith L. Wallace, Daniel J. Buysse, Anne Germain, Martica H. Hall, and Satish Iyengar. Variable selection for skewed model-based clustering: Application to the identification of novel sleep phenotypes. Journal of the American Statistical Association, 113(521):95–110, 2018. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wang:2018:QLE

[WC18] Chao Wang and Kung-Sik Chan. Quasi-likelihood estimation of a censored autoregressive model with exogenous variables. Journal of the American Statistical Association, 113(523):1135–1145, 2018. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wang:2013:AMA

[WCZ^{+}13] Yuanjia Wang, Huaihou Chen, Donglin Zeng, Christine Mauro, Naihua Duan, and M. Katherine Shear. Auxiliary markerassisted classification in the absence of class identifiers. Journal of the American Statistical Association, 108(502):553–565, 2013. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wheeler:2014:MHG

[WDP^{+}14] Matthew W. Wheeler, David B. Dunson, Sudha P. Pandalai, Brent A. Baker, and Amy H. Herring. Mechanistic hierarchical Gaussian processes. Journal of the American Statistical Association, 109(507):894–904, 2014. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wang:2010:BSM

[WDSL10] C. Wang, M. J. Daniels, D. O. Scharfstein, and S. Land. A Bayesian shrinkage model for incomplete longitudinal binary data with application to the breast cancer prevention trial. Journal of the American Statistical Association, 105(492):1333–1346, December 2010. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Westfall:2010:CCV

[Wes10] Peter H. Westfall. Comment: “Correlated z-Values and the Accuracy of Large-Scale Statistical Estimates”. Journal of the American Statistical Association, 105(491):1063–1066, September 2010. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wang:2012:MIR

[WF12] Huixia Judy Wang and Xingdong Feng. Multiple imputation for M-regression with censored covariates. Journal of the American Statistical Association, 107(497):194–204, 2012. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Weinstein:2013:SAC

[WFB13] Asaf Weinstein, William Fithian, and Yoav Benjamini. Selection adjusted confidence intervals with more power to determine the sign. Journal of the American Statistical Association, 108(501): 165–176, 2013. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Windmeijer:2019:ULI

[WFDS19] Frank Windmeijer, Helmut Farbmacher, Neil Davies, and George Davey Smith. On the use of the lasso for instrumental variables estimation with some invalid instruments. Journal of the American Statistical Association, 114(527):1339–1350, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wang:2012:CIF

[WfTQ12] Peng Wang, Guei feng Tsai, and Annie Qu. Conditional inference functions for mixed-effects models with unspecified randomeffects distribution. Journal of the American Statistical Association, 107(498):725–736, 2012. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wang:2018:LOP

[WFZ18] Yuanjia Wang, Haoda Fu, and Donglin Zeng. Learning optimal personalized treatment rules in consideration of benefit and risk:

With an application to treating type 2 diabetes patients with insulin therapies. Journal of the American Statistical Association, 113(521):1–13, 2018. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Woodard:2011:OMB

[WG11] Dawn B. Woodard and Moises Goldszmidt. Online modelbased clustering for crisis identification in distributed computing. Journal of the American Statistical Association, 106(493): 49–60, March 2011. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wang:2014:MSS

[WG14] Fangpo Wang and Alan E. Gelfand. Modeling space and spacetime directional data using projected Gaussian processes. Journal of the American Statistical Association, 109(508):1565–1580, 2014. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wierzbicki:2014:R

[WGDG14a] Michael R. Wierzbicki, Li-Bing Guo, Qing-Tao Du, and Wensheng Guo. Rejoinder. Journal of the American Statistical Association, 109(508):1353–1354, 2014. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wierzbicki:2014:SSN

[WGDG14b] Michael R. Wierzbicki, Li-Bing Guo, Qing-Tao Du, and Wensheng Guo. Sparse semiparametric nonlinear model with application to chromatographic fingerprints. Journal of the American Statistical Association, 109(508):1339–1349, 2014. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Warnick:2018:BAE

[WGE^{+}18] Ryan Warnick, Michele Guindani, Erik Erhardt, Elena Allen, Vince Calhoun, and Marina Vannucci. A Bayesian approach for estimating dynamic functional network connectivity in fMRI data. Journal of the American Statistical Association, 113(521): 134–151, 2018. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wang:2012:NEC

[WGM12] Yuanjia Wang, Tanya P. Garcia, and Yanyuan Ma. Nonparametric estimation for censored mixture data with application to the cooperative Huntington’s observational research trial. Journal of the American Statistical Association, 107(500):1324–1338, 2012. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wikle:2015:C

[WH15] Christopher K. Wikle and Scott H. Holan. Comment. Journal of the American Statistical Association, 110(511):901–903, 2015. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wu:2017:BHM

[WH17] Guohui Wu and Scott H. Holan. Bayesian hierarchical multipopulation multistate jolly-seber models with covariates: Application to the pallid sturgeon population assessment program. Journal of the American Statistical Association, 112(518):471– 483, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Willis:2019:CSP

[Wil19] Amy Willis. Confidence sets for phylogenetic trees. Journal of the American Statistical Association, 114(525):235–244, 2019. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wang:2013:RVS

[WJHZ13] Xueqin Wang, Yunlu Jiang, Mian Huang, and Heping Zhang. Robust variable selection with exponential squared loss. Journal of the American Statistical Association, 108(502):632–643, 2013. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wang:2013:HBA

[WJYJ13] Yueqing Wang, Xin Jiang, Bin Yu, and Ming Jiang. A hierarchical Bayesian approach for aerosol retrieval using MISR data. Journal of the American Statistical Association, 108(502):483– 493, 2013. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wei:2013:LSL

[WK13] Susan Wei and Michael R. Kosorok. Latent supervised learning. Journal of the American Statistical Association, 108(503):957– 970, 2013. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wiesenfarth:2012:DSI

[WKKS12] Manuel Wiesenfarth, Tatyana Krivobokova, Stephan Klasen, and Stefan Sperlich. Direct simultaneous inference in additive models and its application to model undernutrition. Journal of the American Statistical Association, 107(500):1286–1296, 2012. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wang:2013:EEC

[WL13] Huixia Judy Wang and Deyuan Li. Estimation of extreme conditional quantiles through power transformation. Journal of the American Statistical Association, 108(503):1062–1074, 2013. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wang:2012:EHC

[WLH12] Huixia Judy Wang, Deyuan Li, and Xuming He. Estimation of high conditional quantiles for heavy-tailed distributions. Journal of the American Statistical Association, 107(500):1453–1464, 2012. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Womack:2014:IIB

[WLNC14] Andrew J. Womack, Luis León-Novelo, and George Casella. Inference from intrinsic Bayes’ procedures under model selection and uncertainty. Journal of the American Statistical Association, 109(507):1040–1053, 2014. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wang:2017:URS

[WLS17] Xinlei Wang, Johan Lim, and Lynne Stokes. Using ranked set sampling with cluster randomized designs for improved inference on treatment effects. Journal of the American Statistical Association, 111(516):1576–1590, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wood:2017:GAM

[WLSA17] Simon N. Wood, Zheyuan Li, Gavin Shaddick, and Nicole H. Augustin. Generalized additive models for gigadata: Modeling the U.K. black smoke network daily data. Journal of the American Statistical Association, 112(519):1199–1210, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wu:2014:SAO

[WLXL14] Hulin Wu, Tao Lu, Hongqi Xue, and Hua Liang. Sparse additive ordinary differential equations for dynamic gene regulatory network modeling. Journal of the American Statistical Association, 109(506):700–716, 2014. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wang:2014:SET

[WLY^{+}14] Chao Wang, Heng Liu, Jian-Feng Yao, Richard A. Davis, and Wai Keung Li. Self-excited threshold Poisson autoregression. Journal of the American Statistical Association, 109(506):777– 787, 2014. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wong:2019:PLF

[WLZ19] Raymond K. W. Wong, Yehua Li, and Zhengyuan Zhu. Partially linear functional additive models for multivariate functional data. Journal of the American Statistical Association, 114 (525):406–418, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wang:2014:ELE

[WM14] Christina D. Wang and Per A. Mykland. The estimation of leverage effect with high-frequency data. Journal of the American Statistical Association, 109(505):197–215, 2014. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wang:2012:NRM

[WMA^{+}12] Yuan Wang, J. S. Marron, Burcu Aydin, Alim Ladha, Elizabeth Bullitt, and Haonan Wang. A nonparametric regression model with tree-structured response. Journal of the American Statistical Association, 107(500):1272–1285, 2012. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Weinstein:2018:GLE

[WMBZ18] Asaf Weinstein, Zhuang Ma, Lawrence D. Brown, and Cun-Hui Zhang. Group-linear empirical Bayes estimates for a heteroscedastic normal mean. Journal of the American Statistical Association, 113(522):698–710, 2018. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wallace:2017:C

[WMS17] Michael P. Wallace, Erica E. M. Moodie, and David A. Stephens. Comment. Journal of the American Statistical Association, 111 (516):1530–1534, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [CZK17a, CZK17b].

Wu:2015:SCS

[WMY15] Yuanshan Wu, Yanyuan Ma, and Guosheng Yin. Smoothed and corrected score approach to censored quantile regression with measurement errors. Journal of the American Statistical Association, 110(512):1670–1683, 2015. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/01621459.2014.989323.

Wei:2011:CAC

[WN11] Susan Wei and Andrew B. Nobel. Comment: “Adaptive Confidence Intervals for the Test Error in Classification”. Journal of the American Statistical Association, 106(495):931–936, September 2011. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wang:2019:APC

[WO19] Jingshu Wang and Art B. Owen. Admissibility in partial conjunction testing. Journal of the American Statistical Association, 114 (525):158–168, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wolfson:2011:EGM

[Wol11] Julian Wolfson. EEBoost: a general method for prediction and variable selection based on estimating equations. Journal of the American Statistical Association, 106(493):296–305, March 2011. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wood:2017:C

[Woo17] Simon N. Wood. Comment. Journal of the American Statistical Association, 112(517):164–166, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wang:2015:CDC

[WPH^{+}15] Xueqin Wang, Wenliang Pan, Wenhao Hu, Yuan Tian, and Heping Zhang. Conditional distance correlation. Journal of the American Statistical Association, 110(512):1726–1734, 2015. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/01621459.2014.993081.

Wang:2015:HDN

[WPL15] Lan Wang, Bo Peng, and Runze Li. A high-dimensional nonparametric multivariate test for mean vector. Journal of the American Statistical Association, 110(512):1658–1669, 2015. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/01621459.2014.988215.

Wood:2017:R

[WPS17a] Simon N. Wood, Natalya Pya, and Benjamin Säfken. Rejoinder. Journal of the American Statistical Association, 111(516):1573– 1575, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [GS17b, Kne17, WPS17b, Yee17].

Wood:2017:SPM

[WPS17b] Simon N. Wood, Natalya Pya, and Benjamin Säfken. Smoothing parameter and model selection for general smooth models. Journal of the American Statistical Association, 111(516):1548–1563, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See comments [Kne17, Yee17, GS17b] and rejoinder [WPS17a].

Wu:2019:PEV

[WQxYW19] Leqin Wu, Xing Qiu, Ya xiang Yuan, and Hulin Wu. Parameter estimation and variable selection for big systems of linear ordinary differential equations: a matrix-based approach. Journal of the American Statistical Association, 114(526):657–667, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wheldon:2013:RPP

[WRCG13] Mark C. Wheldon, Adrian E. Raftery, Samuel J. Clark, and Patrick Gerland. Reconstructing past populations with uncertainty from fragmentary data. Journal of the American Statistical Association, 108(501):96–110, 2013. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wang:2010:NRM

[WRL10] Lu Wang, Andrea Rotnitzky, and Xihong Lin. Nonparametric regression with missing outcomes using weighted kernel estimating equations. Journal of the American Statistical Association, 105(491):1135–1146, September 2010. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wang:2012:EVD

[WRL^{+}12a] Lu Wang, Andrea Rotnitzky, Xihong Lin, Randall E. Millikan, and Peter F. Thall. Evaluation of viable dynamic treatment regimes in a sequentially randomized trial of advanced prostate cancer. Journal of the American Statistical Association, 107 (498):493–508, 2012. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wang:2012:R

[WRL^{+}12b] Lu Wang, Andrea Rotnitzky, Xihong Lin, Randall E. Millikan, and Peter F. Thall. Rejoinder. Journal of the American Statistical Association, 107(498):518–520, 2012. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wang:2014:GGP

[WS14] Bo Wang and Jian Qing Shi. Generalized Gaussian process regression model for non-Gaussian functional data. Journal of the American Statistical Association, 109(507):1123–1133, 2014. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wakefield:2016:CGS

[WSG16] Jon Wakefield, Daniel Simpson, and Jessica Godwin. Comment: Getting into space with a weight problem. Journal of the American Statistical Association, 111(515):1111–1118, 2016. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wang:2014:C

[WSL14] Lan Wang, Ben Sherwood, and Runze Li. Comment. Journal of the American Statistical Association, 109(507):1007–1010, 2014. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wei:2016:QRS

[WSL^{+}16] Ying Wei, Xiaoyu Song, Mengling Liu, Iuliana Ionita-Laza, and Joan Reibman. Quantile regression in the secondary analysis of case-control data. Journal of the American Statistical Association, 111(513):344–354, 2016. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wang:2016:CUP

[WSSQ16] Junhui Wang, Xiaotong Shen, Yiwen Sun, and Annie Qu. Classification with unstructured predictors and an application to sentiment analysis. Journal of the American Statistical Association, 111(515):1242–1253, 2016. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

White:2017:VSK

[WSW17] Kyle R. White, Leonard A. Stefanski, and Yichao Wu. Variable selection in kernel regression using measurement error selection likelihoods. Journal of the American Statistical Association, 112(520):1587–1597, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wahl:2010:CCH

[WSWT10] Eugene Wahl, Christian Schoelzel, John Williams, and Seyitriza Tigrek. Comment: “Comment: Hierarchical Statistical Modeling for Paleoclimate Reconstruction”. Journal of the American Statistical Association, 105(491):900–905, September 2010. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Witten:2010:C

[WT10a] Daniela Witten and Robert Tibshirani. Correction. Journal of the American Statistical Association, 105(492):1637, December 2010. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Witten:2010:FFS

[WT10b] Daniela M. Witten and Robert Tibshirani. A framework for feature selection in clustering. Journal of the American Statistical Association, 105(490):713–726, June 2010. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wu:2013:NEC

[WT13] Colin O. Wu and Xin Tian. Nonparametric estimation of conditional distributions and rank-tracking probabilities with timevarying transformation models in longitudinal studies. Journal of the American Statistical Association, 108(503):971–982, 2013. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wang:2019:JIS

[WTM19] Yifei Wang, Daniel J. Tancredi, and Diana L. Miglioretti. Joint indirect standardization when only marginal distributions are observed in the index population. Journal of the American Statistical Association, 114(526):622–630, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wu:2011:CPV

[Wu11] Ying Nian Wu. Comment: “Population Value Decomposition, a Framework for the Analysis of Image Populations”. Journal of the American Statistical Association, 106(495):802–803, September 2011. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wu:2015:PFE

[Wu15] C. F. Jeff Wu. Post-Fisherian experimentation: From physical to virtual. Journal of the American Statistical Association, 110 (510):612–620, 2015. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wen:2015:ITB

[WW15] Kuangyu Wen and Ximing Wu. An improved transformationbased kernel estimator of densities on the unit interval. Journal of the American Statistical Association, 110(510):773–783, 2015. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wang:2012:QRA

[WWL12] Lan Wang, Yichao Wu, and Runze Li. Quantile regression for analyzing heterogeneity in ultra-high dimension. Journal of the American Statistical Association, 107(497):214–222, 2012. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wang:2015:WLE

[WX15] Tianhao Wang and Yingcun Xia. Whittle likelihood estimation of nonlinear autoregressive models with moving average residuals. Journal of the American Statistical Association, 110(511):1083– 1099, 2015. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wu:2013:CRQ

[WY13] Yuanshan Wu and Guosheng Yin. Cure rate quantile regression for censored data with a survival fraction. Journal of the American Statistical Association, 108(504):1517–1531, 2013. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wang:2019:CHT

[WY19] Shulei Wang and Ming Yuan. Combined hypothesis testing on graphs with applications to gene set enrichment analysis. Journal of the American Statistical Association, 114(527):1320–1338, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wang:2019:IBO

[WYS19] HaiYing Wang, Min Yang, and John Stufken. Information-based optimal subdata selection for big data linear regression. Journal of the American Statistical Association, 114(525):393–405, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wasserman:2010:SFD

[WZ10] Larry Wasserman and Shuheng Zhou. A statistical framework for differential privacy. Journal of the American Statistical Association, 105(489):375–389, March 2010. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wang:2017:CPM

[WZ17a] Tao Wang and Hongyu Zhao. Constructing predictive microbial signatures at multiple taxonomic levels. Journal of the American Statistical Association, 112(519):1022–1031, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wang:2017:GSI

[WZ17b] Xiao Wang and Hongtu Zhu. Generalized scalar-on-image regression models via total variation. Journal of the American Statistical Association, 112(519):1156–1168, 2017. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). For the the Alzheimer’s Disease Neuroimaging Initiative.

Wu:2010:RMF

[WZL10] Yichao Wu, Hao Helen Zhang, and Yufeng Liu. Robust modelfree multiclass probability estimation. Journal of the American Statistical Association, 105(489):424–436, March 2010. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wong:2018:EES

[WZL18] Kin Yau Wong, Donglin Zeng, and D. Y. Lin. Efficient estimation for semiparametric structural equation models with censored data. Journal of the American Statistical Association, 113 (522):893–905, 2018. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wong:2019:RST

[WZL19] Kin Yau Wong, Donglin Zeng, and D. Y. Lin. Robust score tests with missing data in genomics studies. Journal of the American Statistical Association, 114(528):1778–1786, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wang:2018:OSL

[WZM18] HaiYing Wang, Rong Zhu, and Ping Ma. Optimal subsampling for large sample logistic regression. Journal of the American Statistical Association, 113(522):829–844, 2018. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Wang:2018:QOT

[WZSS18] Lan Wang, Yu Zhou, Rui Song, and Ben Sherwood. Quantileoptimal treatment regimes. Journal of the American Statistical Association, 113(523):1243–1254, 2018. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Xiao:2014:FMS

[XAM^{+}14] Yongling Xiao, Michal Abrahamowicz, Erica E. M. Moodie, Rainer Weber, and James Young. Flexible marginal structural models for estimating the cumulative effect of a time-dependent treatment on the hazard: Reassessing the cardiovascular risks of didanosine treatment in the Swiss HIV cohort study. Journal of the American Statistical Association, 109(506):455–464, 2014. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Xi:2014:LCC

[XBZ^{+}14] Liqun Xi, Kristin Brogaard, Qingyang Zhang, Bruce Lindsay, Jonathan Widom, and Ji-Ping Wang. A locally convoluted cluster model for nucleosome positioning signals in chemical maps. Journal of the American Statistical Association, 109(505):48–62, 2014. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Xu:2014:SMU

[XC14] Chen Xu and Jiahua Chen. The sparse MLE for ultrahighdimensional feature screening. Journal of the American Statistical Association, 109(507):1257–1269, 2014. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Xia:2018:MTS

[XCC18] Yin Xia, Tianxi Cai, and T. Tony Cai. Multiple testing of submatrices of a precision matrix with applications to identification of between pathway interactions. Journal of the American Statistical Association, 113(521):328–339, 2018. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Xu:2017:JSC

[XCH^{+}17] Gongjun Xu, Sy Han Chiou, Chiung-Yu Huang, Mei-Cheng Wang, and Jun Yan. Joint scale-change models for recurrent events and failure time. Journal of the American Statistical Association, 112(518):794–805, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Xun:2013:PEP

[XCM^{+}13] Xiaolei Xun, Jiguo Cao, Bani Mallick, Arnab Maity, and Raymond J. Carroll. Parameter estimation of partial differential Equation models. Journal of the American Statistical Association, 108(503):1009–1020, 2013. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Xu:2015:SRL

[XCQ15] Jin Xu, Jiajie Chen, and Peter Z. G. Qian. Sequentially refined latin hypercube designs: Reusing every point. Journal of the American Statistical Association, 110(512):1696–1706, 2015. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/01621459.2014.993078.

Xu:2017:TRF

[XG17] Ganggang Xu and Marc G. Genton. Tukey g-and-h random fields. Journal of the American Statistical Association, 112(519): 1236–1249, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Xie:2012:SEH

[XKB12] Xianchao Xie, S. C. Kou, and Lawrence D. Brown. SURE estimates for a heteroscedastic hierarchical model. Journal of the American Statistical Association, 107(500):1465–1479, 2012. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Xie:2017:GAV

[XKBS17] Weiyi Xie, Sebastian Kurtek, Karthik Bharath, and Ying Sun. A geometric approach to visualization of variability in functional data. Journal of the American Statistical Association, 112(519): 979–993, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Xu:2018:NHF

[XLN18] Yuhang Xu, Yehua Li, and Dan Nettleton. Nested hierarchical functional data modeling and inference for the analysis of functional plant phenotypes. Journal of the American Statistical Association, 113(522):593–606, 2018. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Xu:2016:R

[XMWT16a] Yanxun Xu, Peter Müller, Abdus S. Wahed, and Peter Thall. Rejoinder. Journal of the American Statistical Association, 111 (515):948–950, 2016. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Xu:2016:BNE

[XMWT16b] Yanxun Xu, Peter Müller, Abdus S. Wahed, and Peter F. Thall. Bayesian nonparametric estimation for dynamic treatment regimes with sequential transition times. Journal of the American Statistical Association, 111(515):921–950, 2016. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Xu:2015:MBT

[XMY^{+}15] Yanxun Xu, Peter Müller, Yuan Yuan, Kamalakar Gulukota, and Yuan Ji. MAD Bayes for tumor heterogeneity-feature allocation with exponential family sampling. Journal of the American Statistical Association, 110(510):503–514, 2015. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Xue:2012:PDP

[XMZ12] Lingzhou Xue, Shiqian Ma, and Hui Zou. Positive-definite l_{1}- penalized estimation of large covariance matrices. Journal of the American Statistical Association, 107(500):1480–1491, 2012. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Xue:2010:CMS

[XQZ10] Lan Xue, Annie Qu, and Jianhui Zhou. Consistent model selection for marginal generalized additive model for correlated data. Journal of the American Statistical Association, 105(492): 1518–1530, December 2010. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Xu:2018:ILS

[XS18] Gongjun Xu and Zhuoran Shang. Identifying latent structures in restricted latent class models. Journal of the American Statistical Association, 113(523):1284–1295, 2018. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Xie:2011:CDU

[XSS11] Minge Xie, Kesar Singh, and William E. Strawderman. Confidence distributions and a unifying framework for meta-analysis. Journal of the American Statistical Association, 106(493):320– 333, March 2011. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Xu:2017:EIQ

[XSWH17] Gongjun Xu, Tony Sit, Lan Wang, and Chiung-Yu Huang. Estimation and inference of quantile regression for survival data under biased sampling. Journal of the American Statistical Association, 112(520):1571–1586, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Xiao:2014:BMM

[XTR^{+}14] Luo Xiao, Sally W. Thurston, David Ruppert, Tanzy M. T. Love, and Philip W. Davidson. Bayesian models for multiple outcomes in domains with application to the Seychelles Child Development Study. Journal of the American Statistical Association, 109(505): 1–10, 2014. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Xu:2019:SQL

[XWG19] Ganggang Xu, Rasmus Waagepetersen, and Yongtao Guan. Stochastic quasi-likelihood for case-control point pattern data. Journal of the American Statistical Association, 114(526):631– 644, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Xiao:2011:MTD

[XWK11] Guanghua Xiao, Xinlei Wang, and Arkady B. Khodursky. Modeling three-dimensional chromosome structures using gene expression data. Journal of the American Statistical Association, 106 (493):61–72, March 2011. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Xie:2019:SVS

[XXL^{+}19] Yimeng Xie, Li Xu, Jie Li, Xinwei Deng, Yili Hong, Korine Kolivras, and David N. Gaines. Spatial variable selection and an application to Virginia Lyme disease emergence. Journal of the American Statistical Association, 114(528):1466–1480, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Xing:2012:SCP

[XY12] Haipeng Xing and Zhiliang Ying. A semiparametric changepoint regression model for longitudinal observations. Journal of the American Statistical Association, 107(500):1625–1637, 2012. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Xia:2010:DRS

[XZX10] Yingcun Xia, Dixin Zhang, and Jinfeng Xu. Dimension reduction and semiparametric estimation of survival models. Journal of the American Statistical Association, 105(489):278–290, March 2010. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Yang:2019:BRF

[Yan19] Shu Yang. Book review: Flexible Imputation of Missing Data, 2nd ed. Journal of the American Statistical Association, 114 (527):1421, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Yang:2013:ODN

[YBT13] Min Yang, Stefanie Biedermann, and Elina Tang. On optimal designs for nonlinear models: A general and efficient algorithm. Journal of the American Statistical Association, 108(504):1411– 1420, 2013. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Yang:2011:CAC

[YC11] Yuhong Yang and Gang Cheng. Comment: “Adaptive Confidence Intervals for the Test Error in Classification”. Journal of the American Statistical Association, 106(495):924–931, September 2011. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Cheng:2013:LLR

[yCtW13] Ming yen Cheng and Hau tieng Wu. Local linear regression on manifolds and its geometric interpretation. Journal of the American Statistical Association, 108(504):1421–1434, 2013. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Yang:2016:BCT

[YD16] Yun Yang and David B. Dunson. Bayesian conditional tensor factorizations for high-dimensional classification. Journal of the American Statistical Association, 111(514):656–??, 2016. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Yu:2016:TPG

[YDZ16] Zhou Yu, Yuexiao Dong, and Li-Xing Zhu. Trace pursuit: A general framework for model-free variable selection. Journal of the American Statistical Association, 111(514):813–??, 2016. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Yee:2017:C

[Yee17] Thomas W. Yee. Comment. Journal of the American Statistical Association, 111(516):1565–1568, 2017. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [WPS17b, WPS17a].

Yang:2010:MCI

[YHD^{+}10] Yang Yang, M. Elizabeth Halloran, Michael J. Daniels, Ira M. Longini, Jr., Donald S. Burke, and Derek A. T. Cummings. Modeling competing infectious pathogens from a Bayesian perspective: Application to influenza studies with incomplete laboratory results. Journal of the American Statistical Association, 105(492):1310–1322, December 2010. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Huang:2012:CPL

[yHQ12] Chiung yu Huang and Jing Qin. Composite partial likelihood estimation under length-biased sampling, with application to a prevalent cohort study of dementia. Journal of the American Statistical Association, 107(499):946–957, 2012. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Yan:2019:SID

[YJFL19] Ting Yan, Binyan Jiang, Stephen E. Fienberg, and Chenlei Leng. Statistical inference in a directed network model with covariates. Journal of the American Statistical Association, 114(526):857– 868, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Yao:2015:BIM

[YKC^{+}15] Hui Yao, Sungduk Kim, Ming-Hui Chen, Joseph G. Ibrahim, Arvind K. Shah, and Jianxin Lin. Bayesian inference for multivariate meta-regression with a partially observed within-study sample covariance matrix. Journal of the American Statistical Association, 110(510):528–544, 2015. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Yu:2013:SSE

[YL13] Tao Yu and Pengfei Li. Spatial shrinkage estimation of diffusion tensors on diffusion-weighted imaging data. Journal of the American Statistical Association, 108(503):864–875, 2013. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Yang:2015:R

[YL15] Y. Claire Yang and Kenneth C. Land. Reply. Journal of the American Statistical Association, 110(509):457, 2015. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Yu:2016:SRI

[YL16] Guan Yu and Yufeng Liu. Sparse regression incorporating graphical structure among predictors. Journal of the American Statistical Association, 111(514):707–??, 2016. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Young:2019:IPW

[YLRH19] Jessica G. Young, Roger W. Logan, James M. Robins, and Miguel A. Hernán. Inverse probability weighted estimation of risk under representative interventions in observational studies. Journal of the American Statistical Association, 114(526):938– 947, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Yi:2015:FSM

[YMSC15] Grace Y. Yi, Yanyuan Ma, Donna Spiegelman, and Raymond J. Carroll. Functional and structural methods with mixed measurement error and misclassification in covariates. Journal of the American Statistical Association, 110(510):681–696, 2015. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Yao:2019:CIM

[YNLC19] Weixin Yao, Debmalya Nandy, Bruce G. Lindsay, and Francesca Chiaromonte. Covariate information matrix for sufficient dimension reduction. Journal of the American Statistical Association, 114(528):1752–1764, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Yang:2011:NBS

[YOD11] Hongxia Yang, Sean O’Brien, and David B. Dunson. Nonparametric Bayes stochastically ordered latent class models. Journal of the American Statistical Association, 106(495):807–817, September 2011. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Youngman:2019:GAM

[You19] Benjamin D. Youngman. Generalized additive models for exceedances of high thresholds with an application to return level estimation for U.S. wind gusts. Journal of the American Statistical Association, 114(528):1865–1879, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Yang:2017:WSD

[YP17] Qing Yang and Guangming Pan. Weighted statistic in detecting faint and sparse alternatives for high-dimensional covariance matrices. Journal of the American Statistical Association, 112 (517):188–200, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Yu:2018:BVS

[YPOR18] Cheng-Han Yu, Raquel Prado, Hernando Ombao, and Daniel Rowe. A Bayesian variable selection approach yields improved detection of brain activation from complex-valued fMRI. Journal of the American Statistical Association, 113(524):1395–1410, 2018. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Yu:2016:UIS

[YPQ^{+}16] Zhe Yu, Raquel Prado, Erin Burke Quinlan, Steven C. Cramer, and Hernando Ombao. Understanding the impact of stroke on brain motor function: A hierarchical Bayesian approach. Journal of the American Statistical Association, 111(514):549–??, 2016. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Yauck:2019:CRM

[YRR19] Mamadou Yauck, Louis-Paul Rivest, and Greg Rothman. Capture–recapture methods for data on the activation of applications on mobile phones. Journal of the American Statistical Association, 114(525):105–114, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Yang:2017:JEQ

[YT17] Yun Yang and Surya T. Tokdar. Joint estimation of quantile planes over arbitrary predictor spaces. Journal of the American Statistical Association, 112(519):1107–1120, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Yau:2015:EMR

[YTL15] Chun Yip Yau, Chong Man Tang, and Thomas C. M. Lee. Estimation of multiple-regime threshold autoregressive models with structural breaks. Journal of the American Statistical Association, 110(511):1175–1186, 2015. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Yu:2019:BRD

[Yu19] Kaixian Yu. Book review: Data Science Foundations: Geometry and Topology of Complex Hierarchic Systems and Big Data Analytics. Journal of the American Statistical Association, 114(527): 1420–1421, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Yuan:2011:REC

[YY11] Ying Yuan and Guosheng Yin. Robust EM continual reassessment method in oncology dose finding. Journal of the American Statistical Association, 106(495):818–831, September 2011. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Yan:2016:CFM

[YY16] Ying Yan and Grace Y. Yi. A class of functional methods for error-contaminated survival data under additive hazards models with replicate measurements. Journal of the American Statistical Association, 111(514):684–??, 2016. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Yun:2017:APF

[YYC17] Jonghyun Yun, Fan Yang, and Yuguo Chen. Augmented particle filters. Journal of the American Statistical Association, 112(517): 300–313, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ye:2018:SOI

[YYY18] Chenglong Ye, Yi Yang, and Yuhong Yang. Sparsity oriented importance learning for high-dimensional linear regression. Journal of the American Statistical Association, 113(524):1797–1812, 2018. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Yang:2019:DPS

[YZ19] Shu Yang and Donglin Zeng. Discussion of “Penalized Spline of Propensity Methods for Treatment Comparison” by Zhou, Elliott, and Little. Journal of the American Statistical Association, 114(525):30–32, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [ZEL19a, ZEL19b].

Yue:2016:PWM

[YZB^{+}16] Chen Yue, Vadim Zipunnikov, Pierre-Louis Bazin, Dzung Pham, Daniel Reich, Ciprian Crainiceanu, and Brian Caffo. Parameterization of white matter manifold-like structures using principal surfaces. Journal of the American Statistical Association, 111(515):1050–1060, 2016. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Yao:2018:ENS

[YZBE18] Zhigang Yao, Ye Zhang, Zhidong Bai, and William F. Eddy. Estimating the number of sources in magnetoencephalography using spiked population eigenvalues. Journal of the American Statistical Association, 113(522):505–518, 2018. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zhang:2016:BIC

[ZA16] Ning Zhang and Daniel W. Apley. Brownian integrated covariance functions for Gaussian process modeling: Sigmoidal versus localized basis functions. Journal of the American Statistical Association, 111(515):1182–1195, 2016. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zajonc:2012:BID

[Zaj12] Tristan Zajonc. Bayesian inference for dynamic treatment regimes: Mobility, equity, and efficiency in student tracking. Journal of the American Statistical Association, 107(497):80–92, 2012. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zhang:2013:PBA

[ZB13] Jin Zhang and Thomas M. Braun. A Phase I Bayesian adaptive design to simultaneously optimize dose and schedule assignments both between and within patients. Journal of the American Statistical Association, 108(503):892–901, 2013. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zhu:2018:LHT

[ZB18] Yinchu Zhu and Jelena Bradic. Linear hypothesis testing in dense high-dimensional linear models. Journal of the American Statistical Association, 113(524):1583–1600, 2018. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zhou:2015:BFB

[ZBHD15] Jing Zhou, Anirban Bhattacharya, Amy H. Herring, and David B. Dunson. Bayesian factorizations of big sparse tensors. Journal of the American Statistical Association, 110(512): 1562–1576, 2015. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/01621459.2014.983233.

Zhu:2011:RAF

[ZBM11] Hongxiao Zhu, Philip J. Brown, and Jeffrey S. Morris. Robust, adaptive functional regression in functional mixed model framework. Journal of the American Statistical Association, 106(495): 1167–1179, September 2011. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zhang:2016:FCM

[ZBZ^{+}16] Lin Zhang, Veerabhadran Baladandayuthapani, Hongxiao Zhu, Keith A. Baggerly, Tadeusz Majewski, Bogdan A. Czerniak, and Jeffrey S. Morris. Functional CAR models for large spatially correlated functional datasets. Journal of the American Statistical Association, 111(514):772–??, 2016. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zhong:2011:THD

[ZC11] Ping-Shou Zhong and Song Xi Chen. Tests for high-dimensional regression coefficients with factorial designs. Journal of the American Statistical Association, 106(493):260–274, March 2011. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zhang:2013:SCI

[ZC13] Jingfei Zhang and Yuguo Chen. Sampling for conditional inference on network data. Journal of the American Statistical Association, 108(504):1295–1307, 2013. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zhang:2017:FCM

[ZC17a] Jingfei Zhang and Jiguo Cao. Finding common modules in a time-varying network with application to the Drosophila Melanogaster gene regulation network. Journal of the American Statistical Association, 112(519):994–1008, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zhang:2017:SIH

[ZC17b] Xianyang Zhang and Guang Cheng. Simultaneous inference for high-dimensional linear models. Journal of the American Statistical Association, 112(518):757–768, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zhao:2017:SSS

[ZCC^{+}17] Sihai Dave Zhao, T. Tony Cai, Thomas P. Cappola, Kenneth B. Margulies, and Hongzhe Li. Sparse simultaneous signal detection for identifying genetically controlled disease genes. Journal of the American Statistical Association, 112(519):1032–1046, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zhao:2017:HFS

[ZCJ^{+}17] Yize Zhao, Matthias Chung, Brent A. Johnson, Carlos S. Moreno, and Qi Long. Hierarchical feature selection incorporating known and novel biological information: Identifying genomic features related to prostate cancer recurrence. Journal of the American Statistical Association, 111(516):1427–1439, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zhang:2011:LNA

[ZCL11] Hao Helen Zhang, Guang Cheng, and Yufeng Liu. Linear or nonlinear? Automatic structure discovery for partially linear models. Journal of the American Statistical Association, 106(495): 1099–1112, September 2011. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zhu:2013:LAB

[ZD13] Bin Zhu and David B. Dunson. Locally adaptive Bayes nonparametric regression via nested Gaussian processes. Journal of the American Statistical Association, 108(504):1445–1456, 2013.

CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zigler:2014:UPS

[ZD14] Corwin Matthew Zigler and Francesca Dominici. Uncertainty in propensity score estimation: Bayesian methods for variable selection and model-averaged causal effects. Journal of the American Statistical Association, 109(505):95–107, 2014. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zhang:2019:NBM

[ZDD19] Zhengwu Zhang, Maxime Descoteaux, and David B. Dunson. Nonparametric Bayes models of fiber curves connecting brain regions. Journal of the American Statistical Association, 114(528): 1505–1517, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zhigljavsky:2010:NAO

[ZDP10] Anatoly Zhigljavsky, Holger Dette, and Andrey Pepelyshev. A new approach to optimal design for linear models with correlated observations. Journal of the American Statistical Association, 105(491):1093–1103, September 2010. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zantedeschi:2011:PMF

[ZDP11] Daniel Zantedeschi, Paul Damien, and Nicholas G. Polson. Predictive macro-finance with dynamic partition models. Journal of the American Statistical Association, 106(494):427–439, June 2011. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zhou:2019:PSPa

[ZEL19a] Tingting Zhou, Michael R. Elliott, and Roderick J. A. Little. Penalized spline of propensity methods for treatment comparison. Journal of the American Statistical Association, 114(525):1–19, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See comments [Spi19, YZ19, PL19] and rejoinder [ZEL19b].

Zhou:2019:PSPb

[ZEL19b] Tingting Zhou, Michael R. Elliott, and Roderick J. A. Little. Penalized spline of propensity methods for treatment comparison:

Rejoinder. Journal of the American Statistical Association, 114 (525):35–38, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [ZEL19a, Spi19, YZ19, PL19].

Zhao:2018:FME

[ZEMD18] Shiwen Zhao, Barbara E. Engelhardt, Sayan Mukherjee, and David B. Dunson. Fast moment estimation for generalized latent Dirichlet models. Journal of the American Statistical Association, 113(524):1528–1540, 2018. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zeng:2013:Cb

[Zen13] Donglin Zeng. Comment. Journal of the American Statistical Association, 108(504):1257–1258, 2013. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [CCJ13a].

Zhu:2014:SVC

[ZFK14] Hongtu Zhu, Jianqing Fan, and Linglong Kong. Spatially varying coefficient model for neuroimaging data with jump discontinuities. Journal of the American Statistical Association, 109(507): 1084–1098, 2014. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zhou:2017:FFD

[ZFW17] Mingyuan Zhou, Stefano Favaro, and Stephen G. Walker. Frequency of frequencies distributions and size-dependent exchangeable random partitions. Journal of the American Statistical Association, 112(520):1623–1635, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zhou:2018:NDB

[ZG18] Quan Zhou and Yongtao Guan. On the null distribution of Bayes factors in linear regression. Journal of the American Statistical Association, 113(523):1362–1371, 2018. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zhou:2018:UFF

[ZH18] Haiming Zhou and Timothy Hanson. A unified framework for fitting Bayesian semiparametric models to arbitrarily censored survival data, including spatially referenced data. Journal of the American Statistical Association, 113(522):571–581, 2018. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zhang:2019:OSS

[ZH19] Anru Zhang and Rungang Han. Optimal sparse singular value decomposition for high-dimensional high-order data. Journal of the American Statistical Association, 114(528):1708–1725, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zhang:2005:AAD

[Zha05] Jin-Ting Zhang. Approximate and asymptotic distributions of chi-squared-type mixtures with applications. Journal of the American Statistical Association, 100(469):273–285, March 2005. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://miranda.asa.catchword.org/vl=
1515603/cl=66/nw=1/rpsv/cw/asa/01621459/v100n469/s31/
p273;   http://www.jstor.org/stable/27590537. See cor-

rection [Dui16].

Zhao:2011:PPD

[Zha11] Yihua Zhao. Posterior probability of discovery and expected rate of discovery for multiple hypothesis testing and high throughput assays. Journal of the American Statistical Association, 106 (495):984–996, September 2011. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zhang:2013:ECM

[Zha13a] Jian Zhang. Epistatic clustering: A model-based approach for identifying links between clusters. Journal of the American Statistical Association, 108(504):1366–1384, 2013. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zhang:2013:CHD

[Zha13b] Ting Zhang. Clustering high-dimensional time series based on parallelism. Journal of the American Statistical Association, 108 (502):577–588, 2013. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zhang:2019:BI

[Zha19a] Kai Zhang. BET on independence. Journal of the American Statistical Association, 114(528):1620–1637, 2019. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zhao:2019:SVP

[Zha19b] Qingyuan Zhao. On sensitivity value of pair-matched observational studies. Journal of the American Statistical Association, 114(526):713–722, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zhou:2010:RRM

[ZHM^{+}10] Lan Zhou, Jianhua Z. Huang, Josue G. Martinez, Arnab Maity, Veerabhadran Baladandayuthapani, and Raymond J. Carroll. Reduced rank mixed effects models for spatially correlated hierarchical functional data. Journal of the American Statistical Association, 105(489):390–400, March 2010. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zhou:2011:MDS

[Zho11] Qing Zhou. Multi-domain sampling with applications to structural inference of Bayesian networks. Journal of the American Statistical Association, 106(496):1317–1330, December 2011. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zhou:2013:HAR

[Zho13] Zhou Zhou. Heteroscedasticity and autocorrelation robust structural change detection. Journal of the American Statistical Association, 108(502):726–740, 2013. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zhou:2014:MCS

[Zho14] Qing Zhou. Monte Carlo simulation for lasso-type problems by estimator augmentation. Journal of the American Statistical Association, 109(508):1495–1516, 2014. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zhou:2019:BRR

[Zho19] Chen Zhou. Book review: Risk Theory: A Heavy Tail Approach. Journal of the American Statistical Association, 114(527):1424– 1425, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zhou:2017:SSM

[ZHS17] Qingning Zhou, Tao Hu, and Jianguo Sun. A sieve semiparametric maximum likelihood approach for regression analysis of bivariate interval-censored failure time data. Journal of the American Statistical Association, 112(518):664–672, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zhu:2012:NCA

[ZJZ12] Wensheng Zhu, Yuan Jiang, and Heping Zhang. Nonparametric covariate-adjusted association tests based on the generalized Kendall’s Tau. Journal of the American Statistical Association, 107(497):1–11, 2012. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zhu:2012:RIS

[ZK12] Ruoqing Zhu and Michael R. Kosorok. Recursively imputed survival trees. Journal of the American Statistical Association, 107 (497):331–340, 2012. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zubizarreta:2017:OMM

[ZK17] José R. Zubizarreta and Luke Keele. Optimal multilevel matching in clustered observational studies: A case study of the effectiveness of private schools under a large-scale voucher system. Journal of the American Statistical Association, 112(518):547– 560, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zhu:2014:BGL

[ZKLI14] Hongtu Zhu, Zakaria Khondker, Zhaohua Lu, and Joseph G. Ibrahim. Bayesian generalized low rank regression models for neuroimaging phenotypes and genetic markers. Journal of the American Statistical Association, 109(507):977–990, 2014. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zhang:2011:FAA

[ZL11] Yu Zhang and Jun S. Liu. Fast and accurate approximation to significance tests in genome-wide association studies. Journal of the American Statistical Association, 106(495):846–857, September 2011. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zeng:2014:EES

[ZL14] Donglin Zeng and D. Y. Lin. Efficient estimation of semiparametric transformation models for two-phase cohort studies. Journal of the American Statistical Association, 109(505):371–383, 2014. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zhang:2015:C

[ZL15a] Yichi Zhang and Eric B. Laber. Comment. Journal of the American Statistical Association, 110(512):1451–1454, 2015. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/01621459.2015.1106403.

Zhu:2015:LBI

[ZL15b] Ke Zhu and Shiqing Ling. LADE-based inference for ARMA models with unspecified and heavy-tailed heteroscedastic noises. Journal of the American Statistical Association, 110(510):784– 794, 2015. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zhang:2018:USN

[ZL18] Ting Zhang and Liliya Lavitas. Unsupervised self-normalized change-point testing for time series. Journal of the American Statistical Association, 113(522):637–648, 2018. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zhang:2018:IDT

[ZLDT18] Yichi Zhang, Eric B. Laber, Marie Davidian, and Anastasios A. Tsiatis. Interpretable dynamic treatment regimes. Journal of the American Statistical Association, 113(524):1541–1549, 2018. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zhou:2018:EEN

[ZLL18] Ling Zhou, Huazhen Lin, and Hua Liang. Efficient estimation of the nonparametric mean and covariance functions for longitudinal and sparse functional data. Journal of the American Statistical Association, 113(524):1550–1564, 2018. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zhu:2011:MFF

[ZLLZ11] Li-Ping Zhu, Lexin Li, Runze Li, and Li-Xing Zhu. Model-free feature screening for ultrahigh-dimensional data. Journal of the American Statistical Association, 106(496):1464–1475, December 2011. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zhao:2014:DUI

[ZLP^{+}14] Yinshan Zhao, David K. B. Li, A. John Petkau, Andrew Riddehough, and Anthony Traboulsee. Detection of unusual increases in MRI lesion counts in individual multiple sclerosis patients. Journal of the American Statistical Association, 109(505):119– 132, 2014. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zhang:2010:RPS

[ZLT10] Yiyun Zhang, Runze Li, and Chih-Ling Tsai. Regularization parameter selections via generalized information criterion. Journal of the American Statistical Association, 105(489):312–323, March 2010. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zhang:2010:ATM

[ZLW10] Heping Zhang, Ching-Ti Liu, and Xueqin Wang. An association test for multiple traits based on the generalized kendall’s tau. Journal of the American Statistical Association, 105(490):473– 481, June 2010. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zou:2017:CRA

[ZLWT17] Tao Zou, Wei Lan, Hansheng Wang, and Chih-Ling Tsai. Covariance regression analysis. Journal of the American Statistical Association, 112(517):266–281, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zhou:2013:TRA

[ZLZ13] Hua Zhou, Lexin Li, and Hongtu Zhu. Tensor regression with applications in neuroimaging data analysis. Journal of the American Statistical Association, 108(502):540–552, 2013. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zhou:2016:DBM

[ZMB^{+}16] Bo Zhou, David E. Moorman, Sam Behseta, Hernando Ombao, and Babak Shahbaba. A dynamic Bayesian model for characterizing cross-neuronal interactions during decision-making. Journal of the American Statistical Association, 111(514):459–??, 2016. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zhou:2017:RWL

[ZMHKK17] Xin Zhou, Nicole Mayer-Hamblett, Umer Khan, and Michael R. Kosorok. Residual weighted learning for estimating individualized treatment rules. Journal of the American Statistical Association, 112(517):169–187, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zhou:2015:STP

[ZMW^{+}15] Zhengyi Zhou, David S. Matteson, Dawn B. Woodard, Shane G. Henderson, and Athanasios C. Micheas. A spatio-temporal point process model for ambulance demand. Journal of the American Statistical Association, 110(509):6–15, 2015. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zubizarreta:2012:CEW

[ZNSR12] José R. Zubizarreta, Mark Neuman, Jeffrey H. Silber, and Paul R. Rosenbaum. Contrasting evidence within and between institutions that provide treatment in an observational study of alternate forms of anesthesia. Journal of the American Statistical Association, 107(499):901–915, 2012. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zhang:2013:TVA

[ZPlW13] Xiaoke Zhang, Byeong U. Park, and Jane ling Wang. Timevarying additive models for longitudinal data. Journal of the American Statistical Association, 108(503):983–998, 2013. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zhou:2016:PRC

[ZPS16] Mingyuan Zhou, Oscar Hernan Madrid Padilla, and James G. Scott. Priors for random count matrices derived from a family of negative binomial processes. Journal of the American Statistical Association, 111(515):1144–1156, 2016. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zhou:2012:IES

[ZQ12] Jianhui Zhou and Annie Qu. Informative estimation and selection of correlation structure for longitudinal data. Journal of the American Statistical Association, 107(498):701–710, 2012. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zhu:2015:CPS

[ZR15] Jian Zhu and Trivellore E. Raghunathan. Convergence properties of a sequential regression multiple imputation algorithm. Journal of the American Statistical Association, 110(511):1112–1124, 2015. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zhang:2019:ICE

[ZRY19] Rongmao Zhang, Peter Robinson, and Qiwei Yao. Identifying cointegration by eigenanalysis. Journal of the American Statistical Association, 114(526):916–927, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zhang:2015:TAB

[ZS15a] Jian Zhang and Li Su. Temporal autocorrelation-based beamforming with MEG neuroimaging data. Journal of the American Statistical Association, 110(512):1375–1388, 2015. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/01621459.2015.1054488.

Zhao:2015:SPL

[ZS15b] Jiwei Zhao and Jun Shao. Semiparametric pseudo-likelihoods in generalized linear models with nonignorable missing data. Journal of the American Statistical Association, 110(512):1577–1590, 2015. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/01621459.2014.983234.

Zoh:2018:PBT

[ZSCM18] Roger S. Zoh, Abhra Sarkar, Raymond J. Carroll, and Bani K. Mallick. A powerful Bayesian test for equality of means in high dimensions. Journal of the American Statistical Association, 113(524):1733–1741, 2018. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zhang:2011:USS

[ZSL^{+}11] Kai Zhang, Dylan S. Small, Scott Lorch, Sindhu Srinivas, and Paul R. Rosenbaum. Using split samples and evidence factors in an observational study of neonatal outcomes. Journal of the American Statistical Association, 106(494):511–524, June 2011. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zhu:2013:SGP

[ZSP13] Yunzhang Zhu, Xiaotong Shen, and Wei Pan. Simultaneous grouping pursuit and feature selection over an undirected graph. Journal of the American Statistical Association, 108(502):713– 725, 2013. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zhu:2014:SPM

[ZSP14] Yunzhang Zhu, Xiaotong Shen, and Wei Pan. Structural pursuit over multiple undirected graphs. Journal of the American Statistical Association, 109(508):1683–1696, 2014. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zhu:2017:MMW

[ZSPL17] Hongtu Zhu, Dan Shen, Xuewei Peng, and Leo Yufeng Liu. MW-PCR: Multiscale weighted principal component regression for high-dimensional prediction. Journal of the American Statistical Association, 112(519):1009–1021, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zhao:2018:CSO

[ZSR18] Qingyuan Zhao, Dylan S. Small, and Paul R. Rosenbaum. Crossscreening in observational studies that test many hypotheses. Journal of the American Statistical Association, 113(523):1070– 1084, 2018. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zhao:2019:MTW

[ZSS19] Qingyuan Zhao, Dylan S. Small, and Weijie Su. Multiple testing when many p-values are uniformly conservative, with application to testing qualitative interaction in educational interventions. Journal of the American Statistical Association, 114(527): 1291–1304, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zhou:2012:IRT

[ZST12] Qian M. Zhou, Peter X.-K. Song, and Mary E. Thompson. Information ratio test for model misspecification in quasi-likelihood inference. Journal of the American Statistical Association, 107 (497):205–213, 2012. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zhu:2016:PPS

[ZSY16] Yunzhang Zhu, Xiaotong Shen, and Changqing Ye. Personalized prediction and sparsity pursuit in latent factor models. Journal of the American Statistical Association, 111(513):241–252, 2016. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zheng:2012:GMC

[ZSZ12] Shurong Zheng, Ning-Zhong Shi, and Zhengjun Zhang. Generalized measures of correlation for asymmetry, nonlinearity, and beyond. Journal of the American Statistical Association, 107(499): 1239–1252, 2012. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zhao:2013:EST

[ZTC^{+}13] Lihui Zhao, Lu Tian, Tianxi Cai, Brian Claggett, and L. J. Wei. Effectively selecting a target population for a future comparative study. Journal of the American Statistical Association, 108(502): 527–539, 2013. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zhu:2011:SSM

[ZTS11] Bin Zhu, Jeremy M. G. Taylor, and Peter X.-K. Song. Semiparametric stochastic modeling of the rate function in longitudinal studies. Journal of the American Statistical Association, 106(496):1485–1495, December 2011. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zubizarreta:2012:UMI

[Zub12] José R. Zubizarreta. Using mixed integer programming for matching in an observational study of kidney failure after surgery. Journal of the American Statistical Association, 107 (500):1360–1371, 2012. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zubizarreta:2015:SWB

[Zub15] José R. Zubizarreta. Stable weights that balance covariates for estimation with incomplete outcome data. Journal of the American Statistical Association, 110(511):910–922, 2015. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zeng:2013:Ca

[ZW13] Donglin Zeng and Yuanjia Wang. Comment. Journal of the American Statistical Association, 108(504):1164, 2013. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [MJPW13b].

Zhou:2014:GPA

[ZW14] Hua Zhou and Yichao Wu. A generic path algorithm for regularized statistical estimation. Journal of the American Statistical Association, 109(506):686–699, 2014. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zhang:2015:DDM

[ZWL^{+}15] Tingting Zhang, Jingwei Wu, Fan Li, Brian Caffo, and Dana Boatman-Reich. A dynamic directional model for effective brain connectivity using electrocorticographic (ECoG) time series. Journal of the American Statistical Association, 110(509): 93–106, 2015. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zhang:2017:LMS

[ZWMC17] Xinyu Zhang, Haiying Wang, Yanyuan Ma, and Raymond J. Carroll. Linear model selection when covariates contain errors. Journal of the American Statistical Association, 112(520):1553– 1561, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zhang:2018:RRO

[ZWQ18] Chong Zhang, Wenbo Wang, and Xingye Qiao. On reject and refine options in multicategory classification. Journal of the American Statistical Association, 113(522):730–745, 2018. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zhu:2019:MTM

[ZWZZ19] Rong Zhu, Alan T. K. Wan, Xinyu Zhang, and Guohua Zou. A Mallows-type model averaging estimator for the varyingcoefficient partially linear model. Journal of the American Statistical Association, 114(526):882–892, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zhou:2014:SFF

[ZX14] Yong-Dao Zhou and Hongquan Xu. Space-filling fractional factorial designs. Journal of the American Statistical Association, 109(507):1134–1144, 2014. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zhou:2017:CDB

[ZX17] Yong-Dao Zhou and Hongquan Xu. Composite designs based on orthogonal arrays and definitive screening designs. Journal of the American Statistical Association, 112(520):1675–1683, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zheng:2014:SSC

[ZYH14] Shuzhuan Zheng, Lijian Yang, and Wolfgang K. Härdle. A smooth simultaneous confidence corridor for the mean of sparse functional data. Journal of the American Statistical Association, 109(506):661–673, 2014. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zhang:2017:OMA

[ZYZL17] Xinyu Zhang, Dalei Yu, Guohua Zou, and Hua Liang. Optimal model averaging estimation for generalized linear models and generalized linear mixed-effects models. Journal of the American Statistical Association, 111(516):1775–1790, 2017. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zhang:2018:C

[ZZ18] Anderson Y. Zhang and Harrison H. Zhou. Comment. Journal of the American Statistical Association, 113(521):201–203, 2018. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zhu:2010:DRR

[ZZF10] Li-Ping Zhu, Li-Xing Zhu, and Zheng-Hui Feng. Dimension reduction in regressions through cumulative slicing estimation. Journal of the American Statistical Association, 105(492):1455– 1466, December 2010. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zhu:2015:RLT

[ZZK15] Ruoqing Zhu, Donglin Zeng, and Michael R. Kosorok. Reinforcement learning trees. Journal of the American Statistical Association, 110(512):1770–1784, 2015. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/01621459.2015.
1036994.

Zhao:2015:NSL

[ZZLK15] Ying-Qi Zhao, Donglin Zeng, Eric B. Laber, and Michael R. Kosorok. New statistical learning methods for estimating optimal dynamic treatment regimes. Journal of the American Statistical Association, 110(510):583–598, 2015. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zhao:2012:EIT

[ZZRK12] Yingqi Zhao, Donglin Zeng, A. John Rush, and Michael R. Kosorok. Estimating individualized treatment rules using outcome weighted learning. Journal of the American Statistical Association, 107(499):1106–1118, 2012. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zhu:2019:PIV

[ZZS19] Wensheng Zhu, Donglin Zeng, and Rui Song. Proper inference for value function in high-dimensional Q-learning for dynamic treatment regimes. Journal of the American Statistical Association, 114(527):1404–1417, 2019. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Zhou:2012:LBE

[ZZWJ12] Qin Zhou, Changliang Zou, Zhaojun Wang, and Wei Jiang. Likelihood-based EWMA charts for monitoring Poisson count data with time-varying sample sizes. Journal of the American Statistical Association, 107(499):1049–1062, 2012. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).